A Bibliography of Papers in Lecture Notes in Computer Science (2012): Volumes 7125–7199

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 March 2018 Version 1.04

Title word cross-reference

(2, 2) [1342]. 1 [662]. 2 [158, 358, 60]. 2^{nd} [597]. 3 [1265, 1008, 665, 316, 663, 570, 394, 1009]. 3/4 [798]. 5 [947]. 6^{th} [845]. 8

[947]. ^{++} [435]. ◯^{R} [638]. ℓ_{1} [1189, 1190]. ℓ_{2} [1190]. GF(2) [808]. h [720]. k

[670, 201, 327, 784, 799, 332]. λ [1018, 309]. K [987]. EL [1120]. N [1053, 637]. P_{k} [660]. π [970]. Q^{2} [1377]. r [655].

-Anonymization [201]. -Approximation [798]. -Body [637]. -Calculi [309]. -Calculus [1018, 970]. -Center [662]. -Cliques [1053]. -Connectivity [784]. -D [1009]. -Free [660]. -Generation [332]. -Graphs [663]. -Grids [670]. -Median [662]. -Partization [655]. -Penalized [1189]. -Point [947]. -Random [720]. -Server [799]. -Trees [665].

10Gbps [29]. 10th [1399, 1400, 1394, 1453]. 12th [1446, 1455]. 13th [1395, 1443, 1444, 1412, 1454, 1432]. 14th [1413, 1454]. 15th [1393]. 16th [1425]. 18th [1442, 1456, 1397].

2011 [584, 571, 830]. 20th [1403, 1435]. 25th [1441].

3G [1205].

4*4 [860]. 4*4-Pattern [860]. 4th [1407, 1423, 1405, 1416, 1459, 1457, 1458, 1451].

5th [566, 562].

60th [1424]. 65nm [731]. 6th [639, 1436, 1445, 1430, 1421].

7th [833, 1417, 1434]. 7x7 [858].

8th [1406, 1439, 1460, 1426, 1427, 1418, 1414, 1435].

9th [1422, 1433, 1450, 1428].

A* [326]. Aachen [1436]. Abduction [1057]. Abductive [1128]. Able [296]. Abnormal [1210]. Abnormalities [1357]. Above [655]. ABox [1120]. Absence [165]. Abstract [1156, 137, 402, 503, 417, 1412]. Abstraction [1020]. Abstractions [987, 1090, 408]. Academic [1319]. ACARM [213]. ACARM-ng [213]. Accelerated [146, 851, 949, 954]. Accelerating [171]. Acceleration [758, 1389, 1380]. Accelerators [1004]. Acceptance [254]. Accepting [904]. Access [900, 209, 899, 415, 1201, 894, 1268, 601, 1255, 31, 294, 696, 211, 896, 182]. Accessible [222, 740]. Accessing [219]. Accident [910]. According [223]. Accounting [381]. Accumulation [177]. Accuracy [683]. accurate [995]. ACG [1432]. Achievements [1402]. Achieving [84, 901]. ACIIDS [1459, 1457, 1458]. Acquisition [831, 1075, 188, 1037]. across [24]. Act [882]. Action [231, 1186, 877, 356]. Actionable [203]. Actions [1193]. Active [1185, 1109, 1112, 1454, 1270]. ActiveBPEL [938]. Activity [121, 1263, 1262, 1297]. Actor [519]. Actor-Network [519]. Actualization [145]. Acyclic [1233, 789, 77]. Ad [524, 732, 535, 1333, 435, 674]. Adams [1281]. Adapt [258]. Adaptation [249, 841, 847]. Adaptive [1, 1357, 262, 353, 82, 1108, 250, 53, 1364, 922, 260, 246, 701, 1317]. ADB [740]. ADCONS [1401]. Adding [1361]. Address [1202]. Addresses [1267]. Adequacy [1043]. Adjunction [756]. Adjustable [1379]. Adjustment [489]. Administration [1313]. Admission [452]. Adopting [1293, 186]. Adoption [1348]. ADT [227, 330]. Advance [208]. Advanced [208, 643, 1401]. Advances [423, 1450, 1404, 1397, 1432]. Adversarial [1225]. Advertisements [437]. Advice [1096, 799]. Advocating [367]. Aesthetic [110, 1321]. Affect [844]. Affective [1221, 251, 349]. AFRAs [135]. Africa [1415]. African [468]. AFs [129]. against [1338, 793]. Age [706, 1408]. Agent [823, 875, 886, 589, 1225, 881, 1224, 1310, 1221, 1228, 1312, 258, 878, 1433, 264, 1311, 882, 1308]. Agent-Based [589, 1312]. Agents [392, 549, 1221, 1220, 514, 1227]. Aggregate [658, 388, 516]. Aggregating [661, 1325]. Aggregation [28, 775, 741]. Aggregations [931]. Aggressive [142]. Agreement [70]. Agreements [606, 608, 611]. AHDAM [946]. ahead [1197]. al [530]. Alarms [421]. Algebra [302, 113, 531]. Algebraic [238, 240, 1403, 1240]. Algebras [748]. Algorithm [405, 951, 786, 1089, 796, 68, 1108, 1360, 1361, 1340, 638, 326, 1359, 48, 1338, 142, 336, 585, 156, 1362, 195, 1332, 1363, 161, 51, 912, 570, 271, 436]. Algorithmic [755, 1032]. Algorithms [1159, 584, 788, 277, 223, 571, 353, 173, 572, 791, 826, 79, 783, 674, 1250, 1247, 434, 58, 1101, 1421, 799, 1184, 789, 793, 1428, 1181, 1132, 798]. Alignable [1064]. Aligning [1297, 1064]. Alignment [501]. Alignments [1062]. Alleviating [1234]. Allocation [185, 451, 604, 208]. Allocator [946, 634]. Alphabet [667]. Alternating [419]. Ambiguities [1049, 959]. Ambiguous [1186]. AMBTC [1342]. AMG [953]. Analog [1110]. Analyses [1234]. Analysing [292, 312]. Analysis [900, 987, 1264, 1140, 753, 25, 181, 1305, 978, 1315, 602, 223, 198, 616, 921, 1231, 35, 1170, 1238, 1201, 3, 314, 989, 1405, 1222, 864, 1242, 648, 766, 421, 1236, 1396, 1117, 975, 984, 614, 1245, 1440, 1451, 493, 1296, 1345, 1239, 297, 615, 1453, 341, 1292, 169, 1341, 1381]. Analytic [368]. Analytics [325]. Analyzer [1041]. Analyzing [31, 305, 917, 116]. Anaphor [1365]. Anchorage [351]. Android [740]. Angles [669]. Animation [393, 532]. Annealing [166]. Annihilation [14]. Annotated [1040, 1138]. Annotation [106, 1073, 112]. Annotations [986, 1039]. Announcement [1202]. Anomalous [11]. Anomaly [1209, 732, 1149, 91]. Anomaly-Based [1149]. Anonymity [524, 297]. Anonymization [201]. Anonymous [297]. Answering [1128, 1288]. Answers [512]. Ant [1358, 817]. Ant-Based [817]. Antecedent [1365]. Anti [1320]. Anti-Diabetic [1320]. Anticipation [1033]. Appeal [1350]. Application [690, 1243, 218, 208, 624, 409, 67, 1358, 377, 27, 644, 1322, 167, 1069, 612, 851, 1173, 992, 1247, 627, 1232, 594, 918, 1377, 467, 622, 702, 550, 347, 383, 1208, 676]. Application-Level [702]. Application-Replay [690]. Application/OS [624]. Applications [1393, 1024, 555, 560, 683, 1127, 1460, 1369, 222, 368, 628, 1445, 548, 34, 250, 292, 1257, 601, 950, 553, 1425, 542, 387, 114, 556, 325, 1398, 922, 544, 1427, 700, 609, 216, 617, 1131]. Applicative [1284]. Applied [1335, 470, 1241, 1399, 1400]. Applying [214, 978]. Approach [76, 253, 1259, 1351, 1354, 541, 644, 196, 264, 1128, 1052, 943, 425, 1053, 1005, 1337, 393, 388, 648, 1057, 304, 455, 1314, 174, 697, 1329, 1312, 917, 1239, 615, 356, 341, 1292, 470, 433]. Approaches [691, 215, 937, 1352, 1072, 346, 257, 90]. Approximability [652]. Approximate [1092, 1238]. Approximated [431]. Approximating [659, 865, 784]. Approximation [802, 787, 788, 783, 674, 797, 789, 793, 1192, 570, 798, 1428].

Approximations [320, 3]. April [1407]. Arbitrary [1256]. ARC [1460]. Architects [1290]. Architectural [233, 305]. Architecture [997, 549, 34, 881, 638, 1374, 1291, 1385, 1144, 908, 361, 772, 806, 1372, 978, 1441]. Architecture-Aware [1372]. Architectures [591, 993, 1460, 979, 547, 468, 1011, 637, 999]. Architecure [903]. Archives [274]. ARCS [1441]. Area [399, 348, 193, 1312]. Areas [891]. Argument [125, 427]. Argumentation [135, 124, 137, 134, 503, 248, 139, 132, 123, 133, 1398, 136, 130, 138, 128]. Arguments [126]. Arithmetic [983, 442]. armed [1182]. Armstrong [506]. Array [945, 568, 1373, 1384]. Arrays [1020, 667, 1384]. Art [853, 87, 706]. Articulation [923]. Artificial [1442, 1319, 883, 202]. AS-Level [1271]. Asian [1459, 1457, 1458]. ASKALON [558]. ASP [429, 428]. Aspects [92, 429, 1406, 1437, 160, 251, 1440, 1413]. Aspectual [1364]. Assamese [920]. assembled [223]. Assembler [692, 478]. Assessing [1199, 1312, 843]. Assessment [907, 454, 910, 1204, 847]. Assignment [659]. Assistance [965]. Assistant [404, 555]. Assisted [199]. Assistive [268]. Association [735, 561]. Associations [1366]. Associative [1331]. Assurance [1302]. Astrocytes [1114]. Asymmetric [946, 1358]. Asymmetries [1234]. Asymptotic [718]. Asynchronism [146]. Asynchronous [153, 927, 525, 420, 1114]. Asynchronously [406]. At-Most-Once [73]. Athens [1450]. Athletic [383]. Atlantic [1394]. Atlas [387]. Atmospheric [154]. Atomic [65, 1152]. Atomic-Set [1152]. Attack [690, 700, 741]. Attacking [315]. Attacks [875, 296, 773]. Attempto [962]. Attention [260, 1350]. Attestation [902, 297]. Attractor [824]. Attribute [201, 436, 1326]. Auckland [1424]. Auction [447]. Audio [909, 53, 52, 918]. Audit [775]. Auditory [913, 108, 918]. Augmented [253]. Augmenting [249, 359]. August [1419, 1420, 1446, 1415, 1410]. Austria [1451, 1397, 1454]. Authenticating [292, 60]. Authentication [530, 41, 895, 23, 51, 738]. Authorization [32, 1165]. Auto [173, 740, 617]. Auto-connecting [740]. Auto-tuning [173, 617]. Automata [1085, 1078, 1107, 1445, 1094, 1079, 291, 989, 418, 469, 498, 1102]. Automata-Based [1078]. Automated [831, 1019, 900, 402, 67, 1157]. Automatic [1, 1014, 1305, 618, 1193, 196, 113, 415, 1030, 1038, 1344, 1343, 109, 106, 1182, 697, 614, 1065, 362, 1051, 558, 341]. Automatically [1074, 1310, 258]. Automating [1371, 385]. Automation [215]. Automorphisms [1033]. Automotive [255, 997, 1294]. Autonomic [555, 549, 1364, 1313]. Autonomous [1178, 34, 314, 1327]. Availability [372, 629]. Aviv [1453]. Avoid [308]. Avoiding [288, 577]. Aware [899, 934, 881, 1321, 360, 687, 1372]. Awareness [736, 188].

B [47, 470]. B-Spline [47]. Back [42, 990, 1012, 1176, 1217, 1229, 1285, 100, 140, 244, 280, 299, 342, 379, 443, 459, 474, 595, 649, 688, 704, 745, 781, 814, 828, 887, 905, 924, 940, 956, 968, 980, 778]. Backbone [29, 727, 1256]. Backgammon [862]. Background [268].

Backing [503]. BackStreamDB [1256]. Backward [1021]. Bacterial [9]. Bag [108]. Bags [506]. Balance [298]. Balancing [943, 154, 635]. Ballot [775]. Bandits [1182]. Bands [352]. Bangla [351]. Bangladesh [1421]. Banking [777]. Barcelona [1398]. Barcode [1339]. Barcodes [60]. Barriers [66]. Base [1041, 1124]. Based [838, 771, 405, 587, 499, 317, 1243, 729, 575, 1208, 201, 920, 129, 825, 501, 1078, 1089, 441, 277, 1028, 1105, 397, 618, 1060, 1071, 972, 377, 899, 1335, 588, 589, 885, 69, 1301, 935, 873, 82, 399, 1128, 220, 625, 215, 608, 774, 694, 769, 1224, 826, 1036, 293, 601, 902, 760, 1340, 1053, 992, 989, 955, 1371, 1110, 1311, 1344, 1378, 109, 1359, 1222, 1386, 1338, 199, 55, 316, 1242, 768, 1310, 648, 640, 1334, 1121, 1342, 1300, 304, 1125, 1291, 1153, 172, 1331, 305, 149, 52, 1280, 1314, 1324, 390, 58, 985, 387, 738, 1373]. Based [530, 1115, 1150, 1364, 114, 1117, 450, 469, 1149, 984, 882, 565, 1228, 1329, 1312, 1339, 1044, 110, 346, 817, 189, 775, 1390, 514, 515, 1184, 915, 350, 362, 626, 22, 498, 1321, 457, 1332, 446, 679, 1244, 297, 395, 442, 57, 743, 275, 777, 484, 1237, 323, 332, 1356, 912, 995, 1341, 118, 271, 1068, 999, 1375, 194, 1405]. Bases [1310]. Basic [1113]. Basis [479, 305]. Batching [75]. Bayes [1059]. Bayesian [1, 1358, 1196, 1192, 860]. BD [527]. BD-TTCS [527]. BDI [880]. Be [724]. Been [942]. Behavior [381, 972, 846, 1340, 974, 1149]. Behavioral [319]. Behaviour [345, 382]. Beijing [1405]. Belgium [1406, 1429, 1427]. Belief [500, 191]. Bellman [1189]. Benchmark [374, 367, 366]. Benchmarking [371, 763]. Benchmarks [563, 645, 365]. Benefit [631]. Benefits [455]. Besicovitch [1091]. Best [214, 1174]. Better [283, 1300]. between [284, 1029, 608, 74, 721, 1366, 835, 1136]. Beware [1290]. Beyond [1424, 646, 849]. Bhubaneswar [1418, 1435]. Bialgebraic [239]. Bidirectionality [721]. Big [1200]. Bilinear [1188]. Bilingual [1063, 1066]. Bimodal [272]. Binary [409, 983, 667, 1387, 1157, 744]. Binary-Field [983]. Biobjective [802, 787]. Bioinformatics [597, 600]. Biological [598]. Biology [1099, 1100]. Biomedical [601]. Biomedicine [597]. Biometric [895]. BIP [748]. BipRank [1133]. Birthday [1424]. Bisimulation [931, 1193, 761]. Bisimulations [318]. bit [57]. BitTorrent [1262, 1204]. Bivariate [170]. Black [863, 115]. Black-Box [863, 115]. Blacklisting [681]. Blasting [391]. Blind [1243, 1231, 1240, 268]. Block [1231, 1250]. Blocking [144]. Blocks [945]. Bloom [22]. BloomCasting [22]. BlueGene [614]. BlueGene/P [614]. Blum [711]. Blunder [869]. Blurred [1232]. Blurred/ Noisy [1232]. Body [8, 637]. Boltzmann [954]. Boogie [486]. Boolean [1084, 504, 808]. Boosting [642, 1380, 912]. Booths [733]. Bootstrap [1246]. Bootstrapping [759]. Bordeaux [1419, 1420]. Border [276]. BOSCH [493]. Both [1374]. Bottleneck [659]. Bottom [138]. Bottom-Up [138]. Bound [657, 791, 800, 766, 1236]. Boundary [178]. Bounded [710, 487, 315]. Boundedly [1226]. Boundedness [130]. Bounding [654]. Bounds [1014, 1196, 718]. Box [1035, 863, 989, 1385, 115]. Box-Based [989]. BPEL [977, 976]. Brain [911]. Brainware [619]. Branch [1007]. Branching [765, 239]. Brand [1350]. Brands [1351]. breakthrough [867].

Bridge [102]. Bridging [1099]. Broadband [152]. Brown [1312]. Browser [730, 1166, 360]. Browser- [1166]. Browsers [283]. BSS [1242]. Bubble [399]. Budget [694, 452]. Budget-Deadline [452]. Bug [1166]. Build [1066]. Building [1040, 105, 226, 945, 1310, 1124, 1402]. Buildings [382]. Built [222, 54]. Burning [390]. Bus [779]. Business [368, 1348]. Buying [1354]. Bytecode [289]. Byzantine [70].

C [582, 282, 492, 483, 763, 179, 487, 478]. C-XSC [179]. CA [1448, 933]. Cache [1088, 144, 627]. Cache-Oblivious [1088]. Caching [1131]. Calculation [11, 489]. Calculi [309, 319]. Calculus [1018, 229, 970, 878, 884]. Calibration [686]. Calude [1424, 707]. CAM [1387]. Camera [684, 276, 279, 355, 275, 1405]. Camera-Based [275, 1405]. Camera-Captured [276, 279]. Caml [1277]. Camouflage [348]. Campus [1252]. Can [724, 301, 631, 651]. Candidate [1365]. Canonical [418, 1242]. Capabilities [208]. Capability [202]. Capture [65]. Captured [276, 279]. Capturing [1258]. Car [250, 928]. Card [34]. Cardinality [506]. Cards [690, 948, 34]. Carlo [854, 163, 1386, 857, 856, 852, 860]. Carrying [893]. Cascade [286]. Cascaded [812]. Case [135, 1304, 208, 993, 1315, 67, 223, 1301, 1166, 74, 635, 565, 1346, 1296, 243]. Case-Based [1301]. Cases [910]. Casl [232, 233]. Casl-Mdl [232]. CASP [428]. Cast [1163]. Casting [387]. Catalytic [1093]. Catastrophe [621]. Categorization [109, 1067, 1068]. Category [1060, 110, 742]. Category-Based [110]. Causal [1233, 1234]. Causality [1103, 1127]. Cavernous [147]. Cavernous/Fractured [147]. CBDAR [1405]. CBM [2]. CCPI [1419, 1420, 539]. CCS [320]. CCSL [841]. Cell [774, 591, 568, 152]. Cell-Based [774]. Cellular [1094, 1079, 469, 1267]. Center [662]. Centrality [679]. Centric [973, 609, 1372, 1271]. Cerebrovascular [910]. Certificates [893]. Certification [778]. Certified [477, 476]. Certifying [986, 988]. CGWS [1419, 1420]. CGWS2011 [551]. CH [366]. CH-benCHmark [366]. Chaff [651]. Chain [1371, 1222, 915]. Chaining [1083]. Chaitin [720]. Challenge [282, 1437, 645, 718, 928]. Challenges [472, 2, 694, 546, 1288, 534, 607, 104]. Change [611, 1271, 1292]. Changes [846]. Changing [811]. Channel [152, 8, 682]. Channels [286, 94]. Characteristic [760, 31]. Characteristics [1353, 1255, 1345]. Characterization [754, 982, 1410, 1203]. Characterize [1085]. Characterizing [1265]. Characters [351, 272]. Chargeback [449]. Charrelation [1244]. Checker [466, 494]. Checking [409, 1028, 1106, 554, 413, 758, 1112, 1227, 301, 1412, 487, 467]. Checkpoint [632, 630, 631]. Checkpoint/Restart [631]. Checkpoint/Rollback [632]. Chemical [1110, 1314, 559, 559]. Chemistry [208]. Chess [870]. China [1395, 1460, 1405, 1447, 1408]. Chinese [1135, 340, 1124]. Chip [946, 996, 998, 637, 566, 1011]. Chips [813]. Chirp [50]. Choice [1056]. Choose [1295]. Choquet [1324]. Church [1079]. CICLing [1443, 1444]. CIIP [87]. Ciphers [40]. Circle [525]. Circuit [574, 729, 302, 4, 1384].

Circuits [1084, 813, 812, 808, 807, 810, 809]. Circular [1104, 1030, 1236]. City [1394]. clairvoyant [794]. Classes [1137]. Classical [907]. Classification [1357, 1219, 1208, 992, 913, 451, 739, 112, 1330]. Classification-Based [992]. Classified [437]. Classifier [1331]. Classifiers [198, 1067]. Clermont [1439]. Clermont-Ferrand [1439]. Client [771, 451]. Client-Based [771]. Clinical [1075]. Cliques [1053]. Clock [841, 1110, 1113]. Clock-Free [1113]. Cloning [902]. CLOP [863]. Closed [715]. Closer [723]. Closure [504]. Cloud [639, 186, 557, 541, 549, 539, 376, 548, 543, 547, 371, 821, 648, 455, 542, 451, 458, 450, 544, 536, 447, 457, 446, 540, 449, 550, 629, 998, 145]. Cloud-Based [450, 446]. Clouds [644, 551, 543, 448, 640, 646, 185, 556, 620, 1414]. CLP [432]. Clue [195]. Cluster [458, 1336, 1333, 999]. Cluster-Based [999]. Clustering [1243, 201, 1335, 1212, 1069, 388, 1334, 421, 1329, 915, 1332, 1070]. Clusters [590, 586, 1385, 641, 620]. CMC [1446]. CMMR [1435]. CMOS [813]. CNL [1438, 960, 967]. Co [208, 373, 126]. Co-allocation [208]. Co-located [373]. Co-operative [126]. Coarse [166]. Code [282, 618, 844, 293, 1342, 479, 569, 1339, 1157]. Codeco [963]. CodeContracts [494]. Codes [718]. Coding [1339]. Coercion [1165]. Coercion-Resistant [1165]. Coexisting [642]. Cognition [923, 360]. Cognitive [1052, 1353, 1221, 910, 843]. Cographs [654]. Coinductive [1274]. Collaboration [247, 246]. Collaborative [1293, 545, 464, 248, 336, 338]. Collection [684, 694, 672]. Collective [381, 1452, 626]. Collector [690]. Collider [312]. Colonies [1106]. Colony [1358]. Color [657, 196, 1316]. Coloring [1096]. Colorization [113]. Colour [271]. Colour-Based [271]. Column [588]. Column-Based [588]. CoMA [1156]. Combination [1243, 339, 1059]. Combinations [1069]. Combinatorial [651]. Combining [238, 106, 1072, 327]. Come [678]. Comments [253, 177, 340]. commerce [1350]. Commitment [884]. Commitments [876, 877]. Commodity [1215, 1258]. Common [1087, 224, 949]. Communicating [406]. Communication [359, 993, 732, 94, 1355, 827, 1313, 882, 585, 577]. Communication-Avoiding [577]. Communications [971, 626]. Communities [27, 337, 454]. Community [734, 816]. Commuting [825]. Compact [312]. Companions [259]. Comparable [1062, 1063]. Comparative [1264, 1303]. Comparing [661, 491]. Comparison [1019, 1354, 731, 215, 1353, 1255, 864, 486, 197]. Compatibility [664, 307, 973, 663]. Compensation [48]. Compensations [230]. Competitive [796, 72, 801]. Compiler [759, 1373, 487, 1148]. Compilers [1275, 1276]. Complementary [1223]. Complements [507]. Complete [1093]. Completion [1204]. Complex [824, 370, 576, 1236, 150, 1240, 1239, 819, 1237, 268]. Complexity [804, 711, 1094, 712, 1080, 125, 1096, 1294, 716, 1015, 1016, 516, 1381]. Compliance [1289]. Component [552, 1231, 1238, 846, 1344, 305, 1236].

Component-Based [305]. Component-Hue-Difference-Based [1344]. Components [1293, 1250, 206]. Composing [307]. Composition [972, 885, 1324, 558, 191]. Compositional [754, 310, 466, 241]. Compositionality [699]. Compositions [190]. Compound [1062, 1061]. Comprehension [618]. Comprehensive [63]. Compress [1072]. Compressed [1214, 53]. Compression [590, 630, 199, 654, 51]. COMPSs [542]. Computability [718]. Computable [1084]. Computation [1429, 4, 949, 766, 1373, 1421, 159, 170, 1424]. Computational [1392, 636, 1094, 2, 215, 1050, 1373, 1452, 219, 1443, 1444]. Computational-Memory [1373]. Computations [579, 158, 146, 176, 8, 442, 433]. Compute [155, 600]. Computer [1424, 399, 5, 992, 1081, 1144, 114, 707, 357, 998, 722, 1432]. Computers [1110]. Computing [135, 1099, 591, 1395, 208, 584, 181, 1436, 1460, 551, 539, 125, 1277, 1446, 1441, 1399, 1400, 205, 179, 455, 1376, 1116, 1364, 1313, 808, 565, 536, 166, 1418, 593, 620, 562, 540, 1401, 169, 703, 639, 633, 1437]. ConArg [129]. Concept [1231, 479, 108]. Concepts [839, 290]. Concern [554]. Concerns [979]. Concrete [9]. Concurrency [226, 526, 1113, 1141]. Concurrent [432, 930, 307, 1153, 528, 446]. Condition [477, 491]. Conditional [1021, 137, 509, 939]. Conditions [397, 178, 1352, 9]. Conduction [178, 469]. Conference [1392, 1407, 1393, 1442, 1395, 1426, 1445, 757, 1443, 1444, 1446, 1441, 1399, 1400, 1416, 1448, 1449, 1412, 1409, 1425, 1410, 1459, 1457, 1458, 1447, 1422, 1418, 1456, 1397, 1454, 1453, 1401, 1432]. Confident [863]. Confidential [211]. Configurable [278]. Configuring [290, 1313]. Conflict [693, 554, 1033, 1152]. Conflicts [876]. Conformance [1156]. Confounders [1233]. Confused [690]. Congestion [786]. Congruences [931]. Conjecture [714]. Conjectures [712]. Conjunctive [1128]. Connect [868]. Connect-Four [868]. Connect6 [861]. connecting [740]. Connection [441]. Connections [1210]. Connectivity [74, 784, 679]. Connector [748]. Consecutive [47]. Consensus [826]. Consent [30]. Conservative [1095]. Considering [734, 634, 1292]. Consistency [3]. Consistent [555]. Constant [1084]. Constant-Depth [1084]. Constants [1375]. Constrained [453, 679, 452]. Constraint [129, 1090, 408, 397, 431, 1273]. Constraint-Based [129]. Constraint-Free [1273]. Constraints [1024, 1049, 841, 932, 664, 834, 790, 506, 305, 516]. Constructing [709, 1122]. Construction [1193, 69, 1065, 1039]. Constructions [1051]. Constructive [306]. Consuming [601]. Consumption [1014, 677]. Contacts [1206]. Content [87, 109, 189, 1321, 51, 59]. Content-Aware [1321]. Content-Based [109]. Context [247, 1086, 934, 1303, 188, 112]. Context-Free [1086]. Contexts [1086]. Contextual [348]. Contextualisation [263]. Continuous [13, 195, 62]. Continuous-Time [13]. Continuous-Tone [62]. Continuously [526]. Contract [412]. Contraction [1029]. Contracts [426, 932, 971]. Contrail [557]. Contrast [353, 1245]. Contrastive [1119]. Contravariant [318]. Contributions [365].

Control [919, 900, 95, 153, 899, 289, 212, 894, 294, 1110, 312, 419, 750, 696, 896, 1184, 1336, 743, 1009, 452, 872]. Controlled [965, 961, 962, 963, 964, 1438]. Controller [1378]. Controlling [959]. Convenience [821]. Conventional [813]. Convergence [1246]. Conversational [258]. Conversion [811]. convex [417]. Convolutional [272]. Coo [880]. Coo-BDI [880]. Cooperation [1225, 535]. Cooperative [624, 647]. Coordinated [632]. Coordination [520, 878]. Copenhagen [1434]. Coprocessors [1005]. Copy [59]. Core [1046, 983, 155, 638, 567, 251, 1382, 1379, 1215, 1258, 1003, 1002, 156]. Core-Periphery [1046]. CoreGRID [552, 551]. CoreGRID/ERCIM [551]. Cores [1001, 599]. Corpora [1062, 1063, 1065]. Corpus [1040, 1064, 1039, 1037]. Corpus-Driven [1037]. Correcting [718]. Correction [625]. Correctness [936, 761]. Correlates [911]. Correlation [213, 1242, 938]. Cortex [348]. Coruña [1445]. Cosine [1208]. COSMO [154]. COSMO-SPECS [154]. Cost [1082, 986, 869, 375, 1010, 376, 790, 686, 448, 114, 458, 1387, 274, 457, 1388]. Cost-Effective [1010, 458]. Cost-Efficient [375]. Costs [1294, 585]. COTS [1000]. Council [365]. Count [516]. Counter [1090, 408, 989, 764]. Counterfactual [865]. Counters [668, 564]. Countries [1361, 468, 1363]. Covariance [1241]. Covariant [318]. Covariant-Contravariant [318]. Cover [655, 667]. Coverability [764]. Coverage [80, 79, 1153]. Coverage-Enhancing [79]. Covering [1091]. Covers [507]. CPUs [955]. Cramér [1236]. Crank [178, 174]. Creating [615]. Creativity [1355, 1295]. Credential [33]. Cristian [1424]. Criteria [1069, 1353, 176, 128]. Critical [86, 9, 1396]. Critiques [1300]. Cross [730, 540]. Cross-Browser [730]. Cross-Platform [540]. Crossing [15]. Crossings [669]. Crowd [381]. Crowfoot [411]. Cryptanalysis [530]. Crypto [692]. Crypto-Processor [692]. Cryptographic [282, 896]. Cryptography [61, 60, 62]. Cryptyc [296]. CS [1360]. CS/PSO [1360]. CS1 [1282]. Csec [282]. CTMC [931]. CTMC-Level [931]. Cube [505]. CUDA [1105, 158]. Cultural [396]. Cumulant [164]. Curve [1386]. Custom [945]. Customisation [1386]. Cut [1032]. Cut-Introduction [1032]. Cutting [429]. Cutting-Edge [429]. CV [373]. Cyber [95, 313, 89, 476]. Cyber-Physical [313]. Cycles [15]. Cyclogram [145]. Cyprus [1414].

D [1008, 316, 1009, 158, 60, 570, 394]. Dafny [482]. DAG [575]. DAG-Based [575]. Daisyworld [822]. DALT [1433]. Data [523, 590, 226, 545, 402, 602, 921, 368, 583, 1205, 505, 694, 543, 1138, 1234, 430, 638, 388, 1242, 1374, 339, 58, 801, 973, 600, 211, 918, 1137, 241, 901, 1048, 631, 516, 741, 456, 1348, 216, 103, 161, 265, 1130, 1267, 672, 49, 188, 1126, 1136, 435]. Data-Centric [973]. Data-Collection [672]. Data-Intensive [216]. Database [1214, 376, 372, 1459, 1457, 1458, 915]. Databases [335, 506, 55, 508]. Datacenters [451]. Datalog [431]. Dataset [279, 278]. DCS [92]. Deadline [452]. Deadlock [1146]. Deblurring [1232].

Debugging [755]. Decapod [274]. December [1447, 1401, 1414]. Decentralized [265]. Decidability [319]. Deciding [481]. Decision [409, 377, 1354, 310, 256, 1353, 699, 521, 254, 203]. Decision-Making [1353]. Decisions [912]. Declarative [425, 463, 428, 423, 1150, 1413, 442, 1433]. Decoding [1059]. Decomposition [632, 52]. Decompositional [303]. Decompositions [1237]. Decreasing [1046]. Dedicated [1424]. Deduction [488]. Deductive [132]. Deduplication [640]. Defeasible [1022]. Defense [214, 700, 1396]. Defense-in-Depth [214]. Defining [1086, 1283]. Definition [516]. Definitions [1297]. Deflation [142]. Degree [19, 1353]. Degrees [713]. Delay [738]. Delays [1265]. Delegation [288, 872]. Deletion [1104]. Delft [1430]. Delhi [1443, 1444]. Deliberation [131]. Delivering [603]. Delivery [465]. Demonstrator [777]. Demuth [717]. Demystified [26]. Denmark [1434]. Density [1324]. Deontic [231]. Dependencies [1088, 505, 506, 508]. Dependency [1043, 285, 240, 1292]. Dependent [1227, 394]. Depending [1353]. Deployment [679]. Depth [1084, 214, 384]. Depth-Varying [384]. Derivation [738]. Derivational [1015]. Derivations [1030]. Derivatives [1092]. Derive [253]. Describing [434]. Description [1024, 547, 109, 1016, 1136]. Descriptions [1133, 1308]. Descriptors [350]. Design [579, 261, 560, 1007, 71, 977, 1111, 1364, 1313, 975, 701, 1001, 360, 809, 772, 806, 744]. Designed [215]. Designing [727, 255]. Desktop [592]. Detached [389]. Detail [386, 844]. Detect [296]. Detecting [752, 876, 1210, 532, 512, 1262, 340, 1152]. Detection [1209, 276, 693, 270, 83, 732, 198, 196, 625, 909, 1337, 119, 108, 175, 91, 1148, 77, 197, 1316, 54, 63]. Determinism [1145]. Deterministic [1000, 791, 72]. Developed [392]. Developing [458, 325, 468, 482]. Development [335, 691, 218, 921, 1225, 1294, 306, 304, 1403, 484, 835, 612]. Device [1010, 269]. Dewarping [277]. Dhaka [1421]. Diabetic [1320]. Diagnosis [1314, 461, 1317]. Diagnostic [1135]. Diagonalization [1250, 1249, 1239, 1237]. Diagrams [656, 657, 409, 1297, 1296]. Dialogue [131, 127]. Dialogues [139, 258]. Difference [3, 1190, 1344, 178, 862, 861]. Different [198, 1038, 1353, 1295, 363, 794]. Differential [284, 1362, 17]. Difficulties [1264]. Difficulty [295]. Diffie [295]. Digit [1381]. Digital [412, 901, 395, 391, 1394]. Digitization [274]. Digraphs [789]. Dimension [1367]. Dimensional [591, 178, 8, 822, 653]. Dimensionality [346, 315]. Direct [958, 297]. Directed [1183, 197]. Directional [79, 1341]. Directives [950]. Directly [1172]. Disabling [33]. Disambiguation [1052, 263, 1054]. Disaster [456]. Discourse [1057]. Discovering [1047]. Discovery [1233, 139, 1182, 698, 189, 192, 203, 646]. Discrete [591]. Discretization [10]. Discriminant [1341]. Disequations [1024]. Disjunction [742]. Display [155, 743]. Disposable [287]. Dissemination [84]. Dissimilarities [499]. Distance [531, 332]. Distances [661, 1071]. Distilleries [363]. Distortion [61]. Distributability [747]. Distributed [823, 157, 78, 485, 643, 230, 1402, 602, 927, 1319, 437, 1212, 163, 934, 1168, 1277, 881, 1224, 970, 79, 1310, 313, 149, 1256, 510, 520, 816, 1101, 534, 166, 647, 152, 1102, 1132, 577, 1395, 1418]. Distributed-Memory [157, 577]. Distributing [589]. Distribution [463, 1223, 456]. Distributional [1075, 1246]. Distributions [1094, 870]. District [341]. Disunification [1024]. Diversified [203]. Divisible [592]. Division [1375]. DL [497]. DL-Programs [497]. DNA [648, 1081]. DNS [1199]. Do [301, 114, 449]. Do-It-Yourself [114]. Doctoral [830]. Document [276, 279, 1405, 1058]. Document-Specific [1058]. Documentation [691]. Documents [1224, 1300, 334, 1122]. Does [844, 376]. Doing [1158]. DOM [1071]. Domain [831, 440, 1320, 979, 466, 398, 1373, 967, 1039, 738]. Domain-Specific [979, 1373]. Domains [148, 402, 756, 417, 1036, 24, 1002]. Donut [417]. Double [119]. dP [1107]. DRAM [624, 1008, 625]. Draw [868]. Drawing [670, 669, 665, 653]. Drawings [394]. Drift [48]. Driven [190, 1135, 185, 514, 1039, 835, 1037]. Driver [1384]. Driving [1377]. Drone [911]. Drugs [1320]. Drying [16]. DSLs [1140]. Dual [1025, 1385]. Dual-Priced [1025]. Dual-VT [1385]. Duality [1029]. Dudo [865]. Dumb [1304]. Duplex [971]. Durations [1025]. DWT [398]. Dyadic [818]. Dynamic [78, 946, 285, 1088, 397, 232, 1225, 943, 1125, 339, 154, 185, 530, 975, 556, 1004, 532, 816, 647, 1148, 994]. Dynamically [1313, 1312, 340]. Dynamics [121, 1199, 381, 537, 1158].

E-commerce [1350]. E-election [1174]. e-Infrastructure [209, 1402, 222, 607]. e-Infrastructures [611, 211]. E-Learning [1317]. E-Mobility [928]. e-Science [222, 542]. E-Voting [1449, 1175]. Early [142]. Earthquake [200]. ECC [297, 744]. ECC-Based [297]. ECC-Hardware [744]. Economic [644, 445]. Economics [1414]. Edge [429, 196, 199]. Editing [389]. Editors [963]. Education [833, 104, 383]. Educator [833]. Edutainment [1411]. Effect [351, 1296]. Effective [359, 1010, 30, 420, 1223, 334, 458, 1388]. Effectiveness [73, 329]. Effects [1276, 1175, 1186, 348]. Efficiency [1146, 329, 915, 1216]. Efficient [247, 1209, 1089, 190, 375, 180, 324, 417, 430, 1167, 1121, 953, 1376, 510, 1155, 367, 1165, 457, 1332, 1102, 145, 1309]. eGovernment [978]. eHealth [1315, 472]. Eigenvalue [1390]. Eindhoven [1423]. EinStein [855]. election [1174]. Elections [1170, 1173, 775]. Electrical [194, 911, 469]. Electromagnetics [636]. Electronic [317, 875, 1319, 473, 1287, 471, 363]. Element [636, 735, 242]. Elicitation [1299]. Eligibility [1191]. Elimination [1330]. Embeddability [666]. Embedded [440, 1383, 750, 1011, 701]. Embedding [53, 49, 653]. Embeddings [72]. Embracing [1435]. Emergable [59]. Emerging [104]. Emotion [358, 345]. Empathic [1221]. Empathy [259]. Emphasis [1352]. Empirical [124, 131, 842]. Employing [1382]. Empowering [914]. Emulator [1387]. Enabled [335, 636, 548]. Enabling [574, 623, 472, 902, 542, 600, 211]. Encoding [508, 46]. Encryption [41]. Encryptions [770]. Encyclopedia [1124]. End [694, 750, 1254]. End-Host [1254]. End-to-End [694, 750]. Energy [163, 677, 1376, 816, 533, 367, 687, 617]. Energy-Aware [687].

Energy-Efficient [367]. Enforcement [930, 927, 293, 695, 535]. Enforcing [472, 883]. Enfragmo [1017]. Engine [1387, 152]. Engineering [831, 1407, 1099, 691, 1304, 886, 961, 1431, 1415, 1295, 832, 200, 1456, 1423]. Engineers [1299]. English [962, 963, 1061]. Enhance [249]. Enhanced [933, 199, 367]. Enhancement [353, 272]. Enhancements [856]. Enhancing [549, 322, 757, 79, 619, 1130]. Enlarged [1059]. Enriched [253, 240]. Enrichment [1323]. Ensuring [78]. Enterprise [978]. Entire [222]. Entities [510, 1366, 1136]. Entropy [286, 509, 331]. Entry [752]. Environment [261, 218, 541, 222, 1225, 495, 427, 565, 150, 357, 1348]. Environmental [919, 1010]. Environments [92, 247, 83, 33, 645, 594, 544, 257, 1213, 1266, 452]. Epic [1275]. Episodic [1219]. Equation [441, 178, 954, 18]. Equation-Based [441]. Equations [1024]. Equilibria [929]. Equilibrium [528]. Equipped [683]. Equivalences [760, 319]. Erasure [35]. Error [194, 993, 628, 625, 718, 1273, 484]. Error-Correcting [718]. Errors [628, 627, 626, 702]. Ershov [1426]. Espoo [1393]. Essen [1456]. Essential [947]. ESSoS [1423]. Establish [1162]. Establishment [697]. Estimates [1246]. Estimating [1259]. Estimation [947, 1005, 1232, 594, 582, 1154]. Estonia [1449, 1425, 728]. Estonian [1173]. ESXi [563]. Etelsen [1403]. EU [901]. Euro [1419, 1420]. Euro-Par [1419, 1420]. European [1427, 1422, 1450, 572]. EuroPKI [1427]. Evacuation [382]. Evaluating [627, 1132]. Evaluation [563, 587, 124, 277, 636, 353, 497, 864, 1291, 645, 1061, 1410, 511, 1260, 1070, 360, 772, 484, 1309]. Evaluation-Function [864]. Evaluators [759]. Event [1074, 754, 780, 84, 370, 119, 77, 743, 884, 1073, 470]. Event-B [470]. Event-Based [743]. Events [1219, 966]. Ever [621]. Everyday [964]. Evolution [879, 34, 1362, 942, 837]. Evolutionary [1111]. Evolving [526, 553, 228]. EWRL [1450]. EWSN [1422]. Exact [931, 177, 1238, 16]. Exactly [18]. Examining [224]. Exascale [572, 622]. Exceptional [974]. Exceptions [938, 1195]. Exchange [1355, 516]. Exchanged [590]. Executable [610]. Executables [1141]. Executing [434, 433]. Execution [432, 1000, 491, 33, 988, 1141, 762]. Exercise [483]. Existence [929]. Exo [1104]. Exo-insertion [1104]. Expansion [164]. Experience [282, 757, 491, 494, 634]. Experiences [249, 842, 950, 466]. Experiencing [1210]. Experiment [2, 221, 1213, 312]. Experimental [76, 25, 628, 1368, 1204]. Experiments [859, 224, 1416, 827]. Expert [1359, 363]. Experts [967]. Explanation [255]. Exploiting [1206, 1374]. Exploration [1178, 1007, 488, 182]. Explore [1368]. Exploring [1351, 1056, 744]. Explosion [308]. Exponential [758]. Exponents [11]. Exposing [1271]. Expression [358, 1117, 1044, 362]. Expressions [1092, 233, 808]. Expressive [385]. Expressiveness [504]. Extendable [211]. Extended [1193, 1114, 1366, 62]. Extending [638, 372, 740, 152, 685]. Extensible [757]. Extension [586]. Extensional [294]. Extensions [1087, 134, 760, 221]. Extensive [1056]. External [1140, 250]. Extortion [890]. Extracting [328]. Extraction [1074, 1071, 1205, 1038, 1063]. Extremal [149, 1390]. Extreme [630]. Eye [1351, 1354, 114, 45, 260]. Eye-Tracking [1351, 1354, 260]. Eyetracking [1369].

Fabric [626]. Fabric-Based [626]. Face [355, 195, 350, 1316, 347]. Facial [358, 362, 1316]. Facilitating [71]. Facing [1437]. Factor [735]. Failed [1287]. Failure [1296]. Failures [1141]. Fairness [898]. Faithful [105]. False [523]. Families [715]. Family [224]. Far [942]. Fare [489]. Farthest [656, 657]. Farthest-Color [657]. FAST [1406, 148, 1088, 177, 796, 40, 172, 1067, 51, 995]. Faster [797]. FastICA [1246]. Fat [525, 668]. Father [707]. Fault [624, 71, 630, 1314, 565, 679, 994]. Fault-Tolerant [679]. FD [507]. FD4 [154]. FDTD [159]. Feasibility [1170, 493]. Feature [838, 1364, 739, 1187, 1056, 350, 356, 1341]. Feature-Based [838]. Features [277, 1337, 736, 1316, 63, 1330]. February [1423, 1424, 1441, 1422, 1421, 1418]. Federal [1170]. Federated [546]. Federations [557]. Feedback [1291]. Feedforward [200]. Femtocell [25]. Ferrand [1439]. fetching [186]. Few [790, 8]. Few-Body [8]. FF [731]. FHIES [1415]. Field [10, 983, 744, 1381]. File [186, 893, 457, 360]. Files [58]. Filling [1072]. Filter [1121, 922, 22, 54]. Filtering [523, 1121, 336, 338]. Financial [1279, 98, 1386, 1228, 1356]. Finding [1026, 80, 484]. Fine [377, 567]. Fine-Grained [377, 567]. Fingerprinting [1087, 730]. Finite [1085, 1026, 636, 1107, 3, 178, 7, 1381]. Finland [1393]. Fire [382]. Firewall [1259]. First [231, 842, 139, 962, 1409, 1415, 174, 1398, 942]. First-Order [231, 962]. Fits [889]. Five [15]. Fixed [802, 1255, 1016, 865]. Fixed-Access [1255]. Fixed-Parameter [1016]. Fixed-Strategy [865]. Flat [1072]. Flaws [532]. FLD [398]. Flexible [1215, 1057, 61, 211, 510, 274]. Flickr [1346]. Flight [682]. Floating [19]. Floorplanning [1372]. Flow [930, 1209, 285, 413, 1123, 289, 419, 695, 77, 939, 794]. Flows [368]. Fluid [1100, 172]. Flux [1388]. Fly [285, 686]. FM [1439]. Focused [328]. FoIKS [1417]. Foisting [675]. Folding [224, 1274]. Folksonomies [264, 322]. Folkview [264]. Follows [462]. Fontainebleau [1446]. FOPARA [1440]. Forecasting [1324, 1356]. Foreground [892]. Forensics [45, 1394]. Forgetting [1022]. Form [462, 1327]. Formal [464, 1439, 232, 1079, 1311, 1112, 461, 1398, 882, 476, 478, 297, 874, 1406]. Formalisation [696]. Formalization [1045, 985, 469, 493]. Formalized [490]. Formalizing [515, 877]. Formants [923]. Formation [525, 822]. Formats [435]. Forming [1278]. Forms [395]. Formulas [1182, 511]. Formulation [1127, 1335]. Forwarding [68, 535]. Foundation [232, 1456]. Foundational [1440]. Foundations [1227, 205, 498, 1415, 1417]. Four [868]. Fourier [164]. Fourth [612, 633]. FPGA [947, 1373, 1390]. FPGA-Based [1373, 1390]. FPGAs [1005, 1385, 1382, 1381]. Fragile [56, 51]. Fragmentation [1207]. Fragmenting [793]. Frame [276, 1039]. Framework [987, 213, 623, 181, 589, 32, 628, 1386, 763, 387, 133, 976, 450, 1377, 1260, 128, 874]. Frameworks [135, 124, 575, 134, 235, 503, 123, 219, 847]. France [1419, 1420, 1439, 1446]. Francisco [1448]. Free [1086, 948, 660, 163, 376, 1113, 663, 1273, 720, 57, 1147]. Free-View [57]. French [604]. Frequency [1110, 918, 916]. Fresco [400]. Friendly [219]. Front [21, 43, 538, 596, 650, 671, 689, 705, 726, 746, 767, 782, 803, 815, 829, 850, 871, 888, 906, 925, 941, 957, 969, 981, 991, 1013, 1034, 1055, 64, 1076, 1097, 1118, 1139, 1160, 1177, 1198, 1218, 1230, 1251, 1272, 1286, 1307, 1328, 1349, 1370, 1391, 85, 101, 122, 141, 162, 183, 204, 225, 245, 266, 281, 300, 321, 343, 364, 380, 401, 422, 444, 460, 475, 496, 517]. FRP [439]. FRSM [1435]. FSEN [1407]. Fully [1168, 362]. Function [462, 1188, 864, 1192]. Functional [986, 426, 1279, 1027, 588, 1322, 505, 846, 506, 6, 1282, 1455, 1290, 243, 424, 481, 805, 554]. Functionals [165]. Functions [1084, 866, 165, 1245, 1184, 242, 1388, 47]. Functory [1277]. Fundamental [916]. Fundamentals [1407]. Funds [317]. Fusion [1284, 119]. Future [607, 891, 1345]. Fuzzy [1357, 1320, 353, 215, 1069, 358, 1329, 346, 593, 1326].

G [1265]. GA [1356]. Game [859, 965, 866, 455]. Game-Theoretical [455]. Game-Tree [866]. Games [929, 392, 851, 1057, 528, 1282, 857, 1432]. Gaming [104]. Gap [1090, 408, 102]. Gap-Order [1090, 408]. Garbage [690]. Gas [686, 96, 687]. Gate [811, 1384]. Gateway [220]. Gaussian [1233]. Gaze [116]. GECON [1414]. Gene [1117]. General [1019, 1247]. Generalisations [719]. Generalization [202]. Generalized [655, 18, 1141, 1241]. Generated [1074, 735]. Generating [1275, 569, 323]. Generation [574, 213, 1215, 1319, 571, 368, 375, 113, 1030, 1371, 491, 385, 1334, 1125, 243, 332]. Generator [477, 39]. Generators [759]. Generic [1003, 1129, 837, 470]. Genetic [713, 1359, 585]. GeniUS [1129]. Genre [109, 913]. Gent [1429]. Geo [322]. Geo-folksonomies [322]. Geographical [1211]. Geolocating [1267]. Geometric [386, 955]. German [1170, 1271]. Germany [1436, 1441, 1417, 1403, 1456, 1428]. Gestures [357]. Gets [690]. GHz [1266]. gigabit [1258]. GiN [1283]. Girona [1404]. GIS [325]. Given [508]. Global [153, 1436, 1360, 1361, 866]. GMRES [590]. Go [854, 853, 869, 858, 102, 860]. Goal [1183, 874]. Goal-Directed [1183]. Gödel [719]. Going [849]. Gone [678]. Governance [607]. Governmental [727]. GPS [683]. GPS-Equipped [683]. GPU [951, 590, 591, 636, 146, 173, 949, 172, 586, 387, 1004, 569, 954, 580]. GPU-Accelerated [954]. GPU-Based [172]. GPUs [618, 1005, 950, 955, 578, 168]. GR [420]. Gradient [1194, 1184]. Grain [223]. Grained [377, 567]. Graining [166]. Grammar [1050]. Grammars [1086, 436, 963]. Granger [1234]. Granger-Causal [1234]. Graph [951, 229, 236, 1036, 107, 1033, 238, 434, 793, 1132, 653, 1053]. Graph-Based [1036, 1053]. Graphemes [1046]. Graphic [948]. Graphical [1283]. Graphics [171, 114, 170]. Graphs [664, 669, 504, 15, 1223, 663, 1152, 654, 315]. Gratification [1347]. Greece [1450]. Greedy [951, 570]. Grid [606, 552, 1402, 212, 445, 558, 604, 635, 206, 603, 772, 561, 209, 222, 205, 585, 217]. Grids [670, 551, 160, 207, 620, 605, 1414]. GridSpace [221]. GridSpace2 [223, 222, 224, 211]. Group [256, 640, 1249, 593]. Group-Based [640]. Groupoids [1095]. Groups [852]. Growing [1252]. Growth [765]. Guarantee [655]. Guarantees [750]. Guidance [344]. Guide [779]. Guided [82]. Guidelines [293, 308].

H [374, 378]. H.264 [48, 1338]. H.264/AVC [48, 1338]. HA-OSCAR [629]. Hadron [312]. Half [971]. Half-Duplex [971]. Hamming [1342]. Hand [357, 195]. Handling [502, 1186, 904, 537]. Handoff [680]. Handover [738]. Hands [1270]. Hands-on [1270]. Handshake [768]. Handshaking [438]. Handwriting [361]. Hangzhou [1447]. Hardness [666]. Hardware [1258, 1000, 1389, 744]. Harmonic [578]. Hartmanis [714]. Hash [1041]. Haskell [1278, 762, 434, 1280, 424]. Havannah [856]. Head [371, 269, 910, 1085, 1107]. Head-Mounted [269]. healing [9]. Health [1415, 465, 468, 104]. Healthcare [464, 463, 466, 534]. Heaps [668]. Heart [952, 469]. Heat [20, 178]. Height [984]. Helios [1166]. Hellman [295]. Herbrand [1026]. Heritage [396]. Heterogeneous [587, 584, 588, 235, 592, 953, 1376, 635, 1377, 593, 1204]. HeteroPar [1419, 1420, 584]. Hets [235]. Heuristic [823, 1300, 570, 999]. Hex [869]. Hexagonal [223]. HiBB [1419, 1420, 597]. Hidden [729, 1247]. Hiding [234, 1338, 58, 896, 49, 54]. Hierarchical [1040, 532, 631, 915, 637]. Hierarchies [105, 1383]. Hierarchy [1125, 627, 843]. High [639, 1279, 1215, 683, 597, 1302, 571, 526, 1294, 633, 945, 950, 678, 179, 569, 1387, 565, 1390, 75, 620, 562, 400, 580, 629, 49, 169]. High-Accuracy [683]. High-Level [1294, 950, 569]. High-Performance [639, 571, 945, 1387, 565, 580]. High-Throughput [75]. Higher [411, 982, 559]. Higher-Order [411, 982]. Highly [218, 566, 601, 638, 1167, 154]. Hijacker [779]. Hillebrand [223]. Hindi [1061]. Hints [260]. Histogram [1332]. History [303, 324]. Hitting [788]. HMI [250]. HMM [915]. Hoc [524, 732, 535, 1333, 435, 674]. Holistic [501]. Home [1261, 1254]. Homogeneous [946]. Homogenizing [601]. Homonyms [70]. Hong [1395, 1460]. HOP [680]. Hopper [1312]. Hospital [743]. Host [31, 24, 1254]. Hosts [29]. HPC [612, 563, 583, 625, 945, 621, 645, 641, 1002]. HPC-Techniques [583]. HPCVirt [1419, 1420, 562]. HPPC [1419, 1420, 566]. HPSS [1419, 1420, 571]. HTPR* [324]. HTPR*-Tree [324]. Hue [1344]. Human [1019, 1351, 384, 357, 1058, 1345, 1350, 345, 382]. Hunting [702]. Hybrid [76, 1007, 591, 840, 1360, 1361, 1005, 294, 448, 952, 808]. Hydrodynamics [11]. Hypergraph [388]. Hypervisors [478]. Hyponym [1037]. Hypothesis [1059].

i* [1298]. I-Voting [1173]. I/O [573, 642]. I2P [1203]. iatmon [1269]. IBM [614]. ICA [1361, 1244]. ICA/PSO [1361]. ICDCIT [1418]. ICDCN [1395].

Iceland [1399, 1400]. ICondensation [195]. ICT [1355, 928]. ID [530]. ID-Based [530]. Idea [222, 698]. Identical [792]. Identification [1151, 1044, 1365, 1051, 1042, 912, 31]. Identify [1048]. Identifying [253, 1206, 116, 1137]. Identities [24]. Identity [771, 1347, 1449, 738]. Identity-Based [738]. IFIP [1430]. II [1437, 1420, 1444, 1400, 1282, 1458]. III [1459]. ILU [157]. Im [74]. Image [105, 194, 277, 1344, 1343, 199, 55, 106, 569, 1232, 278, 1329, 350, 1350, 1321, 51, 169, 54, 112, 63, 272]. Image-Level [106]. Images [591, 276, 279, 270, 267, 593, 116, 1316, 62]. Imaging [1004]. Immune [674]. Immunity [1146]. Impact [632, 146, 504, 147, 626, 1350, 1348, 910, 1292, 843]. Impairment [910]. Impedance [194]. Impediments [1289]. Implementation [223, 283, 180, 947, 1167, 330, 696, 882, 378, 39, 1390, 217, 1388]. Implementations [40, 945, 1146, 480, 744]. Implemented [492, 127]. Implementing [35, 694]. Implicit [1089, 1301]. Importance [605]. Imprecise [1308]. ImPrESS [228]. Improve [1036, 827, 1054, 1044, 1213]. Improved [523, 800, 195, 1184, 46]. Improvement [992, 530]. Improvements [148]. Improving [1242, 984, 862, 1058, 1216, 118, 1330]. Impulses [469]. In-Car [250]. INAM [616]. Incident [26]. Incompleteness [719]. Inconsistency [499]. Inconsistency-Tolerance [499]. Inconsistent [1119, 512]. Incorporated [202]. Increasing [1046, 736, 1179]. Incremental [485, 840, 58, 435]. Independence [756]. Independent [1238, 1361, 1236, 698, 1245, 1239]. Independently [1169]. Independently-Verifiable [1169]. Index [324]. Indexing [873, 430]. India [1443, 1444, 1409, 1418, 1401, 1435, 534, 360]. Indian [270, 909, 907, 195, 910]. Indicators [59]. Indistinguishable [770]. Indo [1409]. Indo-Japan [1409]. Indoor [193, 1266, 323]. Inductive [1274]. Industrial [188]. Industry [98, 1355, 365, 928, 835]. Inequality [431]. Inference [1014, 1357, 353, 415, 1179, 509, 667, 435]. Inferring [418]. Infested [1312]. Infimum [709]. InfiniBand [616]. Infinite [900, 135, 402, 868]. Influence [1234, 1355]. Influences [608]. Informal [396]. Informatics [1426, 1415]. Information [684, 284, 1077, 930, 691, 285, 1071, 724, 1347, 413, 1038, 88, 1338, 96, 26, 695, 28, 176, 770, 1355, 1396, 86, 1417, 468, 1459, 1457, 1458, 512, 1346, 904, 349, 1408, 939, 723, 1393, 1425]. Infrastructure [187, 209, 1402, 222, 212, 88, 210, 96, 205, 1396, 86, 476, 607, 217, 145]. Infrastructures [215, 546, 611, 676, 211, 544, 1427]. Infusion [470]. Inherit [1345]. Initial [1019, 826, 149, 827]. Initiated [672]. Initiative [604, 572]. Initiatives [539]. Injecting [628]. Injection [293, 565]. Injuries [910]. Inlining [285, 737]. Innovative [1319, 395]. Inpainting [1343]. Inputs [651]. Inquiry [127]. Ins [1171]. Insertion [660, 1104]. Insider [296, 773]. Inspection [1157]. Inspiration [445]. Inspired [812]. Installation [27]. Instance [1326]. Institution [236, 241]. Institutions [875, 883]. Instructions [1299]. Instrument [912]. Instrumental [145]. Instrumentation [1155]. Insulin [470]. Integer [407, 1375]. Integers [790]. Integral [1324]. Integrals [164]. Integrated [12, 464, 218, 776, 638, 544, 478, 874]. Integrating [1347, 248, 976, 206]. Integration [1323, 250, 1061, 903, 1308, 1318]. Integrationon [1326]. Integrity [901]. Intel [638, 998]. Intelligence [1442, 1319, 368, 1409, 1452, 1348]. Intelligent [186, 1443, 1444, 1310, 1459, 1457, 1458]. Intended [1163]. Intensive [600, 216]. Intentions [1350]. Inter [405, 1194, 738]. Inter-domain [738]. Inter-procedural [405]. Interacting [1138]. Interaction [357, 772, 1180]. Interactions [354, 974, 874]. Interactive [155, 389, 1224, 736, 1368, 395, 275, 914]. Interconnect [641]. Interconnection [522]. Interface [255, 226, 601]. Interfacing [813]. Interference [533]. Intermediate [478, 486]. Intermediate-Language [478]. Internal [1140, 819]. International [1392, 1407, 1423, 1406, 845, 1442, 1395, 1436, 1439, 842, 1460, 1426, 1429, 1445, 1424, 1443, 1444, 1446, 1441, 1405, 1399, 1400, 1416, 1448, 1449, 1430, 1412, 1415, 1417, 1434, 1398, 1403, 1447, 1440, 1455, 1451, 1421, 1418, 1456, 1438, 1413, 1433, 1397, 1394, 1428, 1454, 1453, 1401, 1414, 1408, 1435, 1432]. Internet [1161, 1418, 1163, 1201, 1268, 890, 534, 891, 1165, 1211, 1162, 728, 1271]. Interoperability [879, 1123, 832]. Interplay [74]. Interpolants [1020, 414]. Interpolation [405, 199, 984, 635]. Interpolation-Based [405, 984]. Interpretation [234, 437, 417, 1412]. Interpretations [1015]. Interpreter [559]. Interpreting [960]. Interruption [230]. Interval [78, 180, 178, 358, 985, 174, 175, 1329]. Interval-Based [985]. Intonation [921]. Intra [1194]. Introducing [374, 1175, 736]. Introduction [1032]. intrusive [1003]. Invariance [716, 165]. Invariant [573, 355]. Invariants [1030]. Invasive [998]. Inventory [212]. Inverse [516]. Inversion [143]. Investigating [1253]. Investigation [164]. Invited [1178, 1424, 1179, 1433, 1180, 1181]. Involution [40]. Involve [967]. Involvement [1354, 1353]. Involving [242]. IP [1267, 1270]. IPC [1335]. IPC-Based [1335]. IPM [1407]. IPv6 [527, 1253]. IR [487, 54]. Iran [1407]. IRIW [55]. Irregular [148, 12]. Irrelevant [1330]. Isabelle [1027, 495]. Isabelle/Circus [495]. Island [1438]. Islands [996]. ISO [691]. Isolating [770, 1145]. Isolation [373, 195]. Isomorphism [1084]. ISP [1202]. Israel [1453]. Issue [1379]. Issue-Slots [1379]. Issues [1225, 96, 536, 254, 470]. Italian [978]. Italy [1422, 1438]. iTask [1283]. Item [118, 338]. Iteration [865]. Iterative [146, 922]. IUPR [279]. IWDW [1394]. IWSOS [1430]. IX [1433].

Jacobi [158]. January [1395, 1416, 1412, 1409, 1413, 1397]. Japan [1409]. Java [1156, 752, 690, 289, 34, 694, 311, 149]. JavaScript [38]. JIST [1447]. Job [208, 867]. Job-Level [867]. Jobs [792]. Johannesburg [1415]. Join [15, 599, 332]. Joins [431]. Joint [1250, 1447, 1239, 1237]. Journals [707]. JPEG [46, 51]. July [1392, 1404, 1426, 1429, 1398, 1403]. June [1436, 1426, 1399, 1400, 1434]. Justification [125]. JVM [930]. JVM-like [930].

Kansas [438]. Kaohsiung [1459, 1457, 1458]. Kazantsev [11]. Kendall [661]. Kennedy [436]. Kernel [945, 108]. Kernels [569, 1184]. Key [1225, 738, 532, 1427, 41]. Keyed [703]. Keys [675]. Keyword [334]. Kiel [1417]. Klagenfurt [1397]. KNN [1340]. Knowledge [1299, 831, 499, 1041, 1322, 1224, 1135, 1310, 1223, 1417, 520, 203, 1124]. Knowledge-Based [499, 1224]. Knowledge-Driven [1135]. Kolkata [1409]. Kong [1395, 1460]. Korean [1355]. Kraichnan [11]. Kriging [161].

Labelling [137]. Labels [742]. Laboratory [223, 222, 224, 211]. Lambda [233]. Lanaguage [586]. Lanczos [1390, 577]. Land [1287]. Landscape [866, 365]. Language [831, 1024, 1049, 1035, 930, 359, 959, 965, 961, 1038, 221, 1283, 1373, 1367, 559, 964, 195, 1438, 478, 884, 1037, 805, 1445]. Language-Specific [1038]. Languages [1077, 440, 441, 944, 935, 1079, 1095, 504, 1038, 882, 1054, 1413, 1433, 486, 877, 1102, 847]. Large [573, 171, 65, 591, 921, 375, 583, 669, 389, 1361, 55, 648, 952, 28, 510, 1124, 312]. Large-Scale [573, 375, 583, 55, 648, 952]. Largest [621]. Laser [1384]. LATA [1445]. Latencies [66]. Latent [1233, 1453]. Lattice [1331, 184]. Lattices [223]. LAV [484]. Lava [438]. Layer [1168, 133, 107, 49]. Layered [292]. Lazy [1020, 416]. Leader [340]. Leadership [1345]. Leaf [789]. Leakage [284, 286]. Learning [253, 1185, 261, 396, 1196, 248, 1188, 715, 1190, 864, 1179, 1300, 1186, 1194, 785, 108, 1228, 1187, 1183, 862, 513, 1450, 257, 1191, 1181, 860, 1317, 861, 1126, 1068]. LearnPADS [435]. Least [1190, 1191]. Least-Squares [1191]. Legal [965]. Legislation [901]. Legislative [959]. Length [1207, 679, 722]. Length-Constrained [679]. Lesná [1392]. Leuven [1406, 1427]. Level [522, 606, 931, 610, 71, 844, 608, 1294, 950, 765, 611, 106, 569, 192, 609, 867, 562, 1211, 910, 702, 217, 605, 994, 1326, 1271, 995, 659, 1386, 627]. Levels [1353]. Lexical [1041, 1075, 1054, 1056]. Lexicon [1063, 1066]. Lexicons [1074]. LHCb [181]. Libraries [752, 572]. Library [560, 1275, 1319, 36, 625, 1277, 1129, 582]. Lifecycles [222]. Lifetime [685]. Lifted [509]. Lifting [66, 52]. Light [983, 355, 1051, 966]. Light-Camera [355]. Lightweight [485, 878]. Like [232, 930]. Limited [525, 592]. Limiter [1388]. Limits [812, 143]. Line [1299, 598, 657, 606, 80, 665, 736, 849, 1341, 394, 30]. Line-Sets [665]. Linear [590, 171, 1233, 1164, 239, 1232, 667, 314]. Lines [838, 483, 1327, 197]. Lingua [1115]. Linguistic [1047, 1056, 1048]. Linguistics [1443, 1444]. Link [29, 681, 1137, 1266, 1009]. Linked [288, 1138, 1130, 1126, 1136]. Linking [479, 1137]. Liquid [403]. List [587, 402, 795, 1042]. List-Manipulating [402]. Lists [242, 481]. Live [1130]. Lizards [1340]. LLBMC [487]. LLVM [1148]. LNN [811]. Load [943, 154]. Loads [592, 453]. Local [1088, 863, 1323, 288, 107, 1254]. Locality [795]. Localization [267, 77]. Locally [955]. Located [793, 373]. Location [25, 479, 897]. Locations [53, 328]. Locking [311]. Locks [581]. Locomotion [385]. Locomotor [381]. LODDO [1136]. Log [1130]. Logarithmic [1085]. Logging [780]. Logic [1017, 1022, 426, 302, 1028, 231, 983, 628, 215, 438, 760, 294, 313, 509, 1385, 1016, 498, 806, 1442]. Logical [235, 1164, 702]. Logics [1024, 894]. Long [230]. Long-Running [230]. Longest [1087, 949]. Look [1197, 723, 1270]. Look-ahead [1197]. Loop [11, 766]. Loops [407]. Loss [1184]. Low [1007, 694, 686, 678, 738, 114, 8, 1387, 274, 999]. Low-Budget [694]. Low-Cost [686, 114, 1387]. Low-Dimensional [8]. Low-Overhead [999]. Lower [1196, 791, 800]. LPAR [1442]. LPAR-18 [1442]. LPC [779]. LQ [1237]. LSB [47]. LSD [404]. LTL [1150]. LTL-Based [1150]. LU [1237]. LVA [1453]. LVA/ICA [1453].

M [1324]. M-Density [1324]. MAC [643]. Mach [352]. Machine [253, 628, 354, 1409, 453, 785, 1061, 565, 1065, 1058, 559]. Machines [1104, 642, 1156, 804, 791, 1100, 621, 641, 647]. Macro [478]. Macro-Assembler [478]. Made [1166]. Madrid [1440, 1455]. Magnetic [11]. MAHT [1041]. Main [1041, 599]. Mainstream [568]. Maintainability [844]. Maintenance [1223]. Major [270]. Make [424]. Making [1353, 521, 254, 1327]. Mal [1297]. Mal-activity [1297]. Malay [1365]. Malicious [29, 1263, 1149]. Malt [363]. Malthusian [621]. Malware [739]. Man [354]. Man-Machine [354]. Manage [543]. Management [771, 854, 997, 691, 786, 545, 208, 602, 1297, 644, 610, 554, 1294, 776, 445, 337, 821, 567, 611, 465, 1434, 556, 1004, 532, 816, 467, 91, 609, 217, 605, 1336, 1408]. Managing [222, 1301, 1224, 603]. Manipulating [402, 695]. Manipuri [1044]. Many [155, 638, 567, 1382]. Many-Core [155, 567, 1382]. Manycore [946, 1383]. Mapping [149, 996, 1377, 550]. MapReduce [1122, 703]. Maps [386, 1211]. Marathon [1152]. March [1442, 1460, 1445, 1443, 1444, 1441, 1430, 1417, 1459, 1457, 1458, 1451, 1456, 1454, 1453, 1435]. Marettimo [1438]. Marked [756]. Market [447, 452]. Market-Oriented [452]. Marketplace [450]. Markov [310, 1247, 915]. Markovian [931]. Marriage [821, 488]. Mars [145]. Mash [261]. Mash-Up [261]. Mashups [37]. Masking [1207, 918]. Massively [392]. Matching [951, 652, 192, 47, 1122]. Matchings [1089, 797]. Matchmaking [1128]. Mathematica [4]. Mathematical [12, 712, 14, 7, 9, 1392]. Mathematics [1019, 725]. Matrices [579, 1249, 1241]. Matrix [588, 180, 947, 955, 1015, 46, 1239, 1244, 580]. Matrix-Based [955]. Matrix-Vector [580]. Matroids [715]. Matter [42, 990, 1012, 1176, 1217, 1229, 1285, 100, 140, 244, 280, 299, 342, 379, 443, 459, 474, 595, 649, 688, 704, 745, 781, 814, 828, 887, 905, 924, 940, 956, 968, 980, 21, 43, 538, 596, 650, 671, 689, 705, 726, 746, 767, 782, 803, 815, 829, 850, 871, 888, 906, 925, 941, 957, 969, 981, 991, 1013, 1034, 1055, 64, 1076, 1097, 1118, 1139, 1160, 1177, 1198, 1218, 1230, 1251, 1272, 1286, 1307, 1328, 1349, 1370, 1391, 85, 101, 122, 141, 162, 183, 204]. Matter [225, 245, 266, 281, 300, 321, 343, 364, 380, 401, 422, 444, 460, 475, 496, 517]. Matting [1344]. Maude [243]. Max [787, 798]. Maximal [1089, 658].

Maximization [802]. Maximum [660, 509, 789]. May [1440, 1455, 1433]. MCMC [1389]. MCT [810]. MCTS [859, 858, 855]. MDEA [1362]. MDGS [1419, 1420, 603]. Mdl [232]. MDP [1185]. MDPs [1193]. Me [1166, 694]. ME-Based [694]. Meaning [103]. Means [1295]. Measure [712, 824, 372, 1136]. Measurement [399, 1201, 1410, 17, 1216, 819, 1454]. Measurements [1261, 1268, 1256]. Measures [711, 1162]. Measuring [1077, 366, 370, 331]. Mechanism [680, 824, 696, 211, 41]. Mechanisms [631, 701, 1317]. Media [102, 147, 339, 361]. Median [662]. MediaWiki [961]. Medical [473, 1004, 461, 471]. Medicine [1135]. Medium [274]. Meets [45]. Member [1355]. Membrane [1103, 1098, 1111, 1446]. Membranes [1109, 1112, 1116]. Memorability [735]. Memorial [1426]. Memory [157, 66, 1041, 1008, 1378, 592, 913, 640, 1373, 1383, 599, 152, 577]. Mental [733]. Merging [1029]. Mérida [1442]. Mesh [560, 389]. Meshes [955]. Message [68, 74]. Meta [838, 858]. Meta-MCTS [858]. Meta-modeling [838]. Metacognitive [251]. Metadata [1213]. Metamodels [832]. Metaphorical [1367]. Metaphysics [725]. Metasystem [765]. Meter [909]. Method [1243, 201, 277, 636, 158, 223, 399, 1119, 947, 1036, 1063, 1343, 178, 1242, 1121, 738, 174, 1339, 50, 1390, 1067, 202, 197, 275, 1333, 54, 577, 1066, 1068]. Methodological [429]. Methodology [1003, 1117, 995]. Methods [754, 1439, 734, 955, 6, 1194, 461, 637, 1009]. MetNet [145]. Metric [1028, 593]. Metrics [656, 1193, 366, 1153]. MetroNet [825]. Metropolitan [825]. Microcontroller [1157]. Microworld [248]. Mid [1126]. Mid-Ontology [1126]. Middleware [209, 218, 208, 206, 1336]. Migration [1379, 583, 23]. Min [286]. Min-Entropy [286]. Mind [354, 1367]. Mine [391]. Minimal [1082, 1109, 1038]. Minimization [865]. Minimum [1059, 533]. Mining [1047, 1075, 638, 1048, 103, 203, 1130, 561, 751, 703]. Mismatches [1293, 835]. Mission [145]. Mitigation [24]. Mixed [369, 366, 178, 999]. Mixnet [775]. MMCP [1392]. MMM [1397]. Mobile [518, 524, 1369, 83, 549, 680, 694, 1224, 1255, 892, 1380, 535, 777]. Mobile-Access [1255]. Mobility [1205, 928]. Modal [318, 226, 1025, 120]. Modalities [894]. Mode [36, 1296]. Model [11, 247, 317, 552, 171, 464, 825, 1000, 409, 1028, 1060, 377, 234, 554, 413, 1322, 758, 4, 1053, 466, 1112, 913, 1227, 448, 301, 1310, 304, 358, 952, 238, 1412, 1324, 1221, 487, 467, 537, 17, 152, 159, 1326, 145, 1366, 1356, 835, 843]. Model-Based [317, 304]. Model-Driven [835]. Modeling [1392, 1404, 12, 1264, 129, 833, 441, 834, 264, 375, 1129, 1100, 425, 316, 7, 154, 1397, 967, 489, 884, 1192, 995, 838, 979]. Modelling [1017, 839, 886, 842, 232, 840, 463, 14, 259, 975, 241, 90, 382]. MODELS [1431, 249, 253, 1185, 352, 845, 1026, 584, 1003, 749, 67, 1233, 588, 844, 846, 445, 393, 1311, 106, 1247, 1396, 1150, 469, 251, 1228, 698, 110, 1183, 200, 18, 837, 847, 1309, 9, 794, 833, 830, 1431]. Modern [44]. Modes [753]. MoDeVVa [848]. Modifiability [1290]. Modified [1362]. Modular [32, 732, 1276]. Modularity [1106]. Modulation [916]. Module [1364, 1384]. Modules [1110, 243]. Modulo [1083]. Moisture [20, 16]. Molecules [65]. MOLTO [964]. Monadic [1024]. Monitor [1028]. Monitor-Based [1028]. Monitoring [1156, 606, 1159, 121, 1142, 1269, 927, 541, 616, 1010, 215, 846, 1257, 1256, 1149, 544, 1451]. Monitors [285, 737, 467]. Monotonic [764]. Monte [854, 163, 1386, 857, 856, 852, 860]. Moral [1031]. Morpho [151]. Morphological [1041]. Morphology [223]. Mortality [1180]. mOSAIC [548, 609]. mOSAIC-Enabled [548]. Moses [1061]. Most [73]. Motifs [224, 818]. Motion [1338, 362]. Motivation [222]. Mounted [269]. Move [852]. Movements [116]. Moving [324, 1343, 323]. MP [1117]. MPI [943, 594, 998]. MPU [1387]. Mstar [522]. MT [1060]. Much [724, 301, 1072]. Multi [823, 839, 1379, 1085, 875, 1215, 1258, 886, 1003, 409, 636, 397, 410, 840, 1107, 264, 659, 979, 477, 1311, 326, 762, 1386, 311, 1227, 107, 267, 1310, 765, 176, 586, 676, 627, 1182, 1116, 8, 882, 857, 521, 1002, 156, 682, 195, 599, 136, 456, 1308, 57, 1326, 128, 49, 120, 794, 1145]. Multi-Agent [875, 886, 1310, 264, 1311, 882, 1308]. Multi-agents [1227]. Multi-application [676]. Multi-armed [1182]. Multi-attribute [1326]. Multi-bit [57]. Multi-channel [8, 682]. Multi-core [1379, 1003, 1002, 156]. Multi-cores [599]. Multi-criteria [176, 128]. Multi-execution [762]. Multi-GPU [636]. Multi-head [1085, 1107]. Multi-layer [107, 49]. Multi-level [659, 1386, 627]. Multi-membranes [1116]. Multi-modal [120]. Multi-modeling [979]. Multi-node [586]. Multi-oriented [267]. Multi-Paradigm [839]. Multi-player [857]. Multi-processor [794]. Multi-prover [477]. Multi-result [765]. Multi-robot [521]. Multi-Rooted [409]. Multi-samples [397]. Multi-script [267]. Multi-site [456]. Multi-sorted [136]. Multi-Target [326]. Multi-threaded [410, 311, 1145]. Multi-valued [1326]. Multi-view [840]. Multiadaptive [1216]. Multiagent [1222, 1309]. Multiagent-System [1309]. Multicast [22]. Multicore [153, 955, 1437, 529, 1011, 570]. Multidimensional [330, 1250]. Multifractal [819]. Multigrid [955]. Multilayer [200]. Multilevel [157]. Multilingual [958]. Multimedia [119, 1397, 103, 104]. Multimodal [1305, 108]. Multiparty [927, 974]. Multiplayer [392]. Multiple [65, 554, 445, 1337, 108, 1054, 996, 1339, 156, 168, 203, 1325]. Multiplication [588, 180, 580]. Multipliers [1381]. Multiprocess [69]. Multiprocessors [157, 996]. Multirole [926]. Multiscale [147]. Multiverse [725]. Multiword [1044]. Munich [1441]. Muon [312]. Music [909, 913, 1334, 915, 1435, 914]. Musical [912]. Musically [914]. Mutual [902, 1206]. My [227, 1158]. Myths [1306].

Name [319]. Named [1136]. Names [288]. Nanotechnology [220, 1081]. Naproche [960]. Narrative [395]. Nash [929, 528]. Nation [1271]. Nation-Centric [1271]. National [1402, 205, 604, 468]. Natural [1049, 1035, 965, 270, 961, 269, 278, 1438, 271, 272]. Navigation [82]. Navigational [504]. NCVW [810]. Near [1155, 1056]. Near-Optimal [1155]. Near-Synonym [1056]. Nearest [1341, 1330]. Nearly [73, 678].

Negotiation [256, 451]. Negotiations [446]. Neighbor [1323, 1330]. Neighbors [336]. Neoclassical [1062]. NEOCR [278]. ness [331]. Nested [936, 1190]. Net [747]. Netherlands [1423, 1430, 1432]. Nets [748, 1093]. Network [522, 684, 78, 753, 683, 1358, 1347, 163, 616, 80, 1261, 1257, 326, 1206, 674, 1252, 182, 519, 682, 200, 1216, 818, 265, 202, 637, 188, 117, 272, 1011]. Network-on-Chip [637, 1011]. Networked [824, 313]. Networking [76, 1395, 423, 1401]. Networks [523, 524, 780, 82, 1265, 72, 681, 1123, 680, 677, 1255, 79, 673, 1263, 28, 676, 327, 903, 533, 817, 535, 1422, 820, 513, 1262, 741, 679, 687, 77, 1267, 675, 1333]. Neural [1105, 1115, 1114, 908, 200, 202, 272]. Neuro [1357, 353]. Neuro-Fuzzy [353]. Neurodynamical [13]. News [513, 340, 120, 111]. Next [574, 213, 1319, 571, 368, 102]. Next-Generation [574, 571]. ng [213]. NGI PL [214]. Nicht [855]. Nicolson [178, 174]. Night [198]. Nightclubs [914]. No [404]. NoCs [316, 1009]. Node [441, 683, 685, 586]. Node-Based [441]. Nodes [155, 1206, 864]. Noise [276, 593]. Noisy [863, 1064, 651, 1330]. Non [1003, 1233, 554, 298, 1094, 417, 314, 898, 1352, 421, 892, 1290, 1237, 794]. Non-clairvoyant [794]. Non-convex [417]. Non-Functional [1290, 554]. Non-Gaussian [1233]. Non-intrusive [1003]. Non-linear [314]. Non-Orthogonal [1237]. Non-Repudiation [898, 298]. Non-standards [892]. Non-statistical [421]. Non-stressed [1352]. Non-uniform [1094]. Nonlinear [143]. Nonmonotonic [497]. Nonparametric [1337]. Nonstandardized [1064]. Nordic [1393, 1425]. NordSec [1393, 1425]. Norm [881]. Norm-Aware [881]. Normal [386]. Normalisation [309]. Normalization [369]. Norms [883]. NORT [1149]. North [910]. Norwegian [1265, 1161, 1162]. Nose [363]. Notation [232, 963]. Note [711]. Notes [13, 908]. Notified [26]. Notions [895]. Noun [1074, 1061]. NOVA [465]. Novel [1278, 1337, 393, 1324, 58, 193, 1339, 361, 51]. November [1432]. Novosibirsk [1426]. Number [802, 724, 867, 654]. Numbers [15]. Numerical [19, 20, 572, 160, 164]. NY [1394].

O [573, 642]. OBDD [1089]. OBDD-Based [1089]. Obfuscated [50]. Object [441, 1343, 107, 390]. Object-Oriented [441]. Objective [1277, 907, 866]. Objectives [576, 528]. Objects [432, 324, 307, 116, 323, 439]. Oblivious [1088]. Observables [17]. Occasion [1424]. October [1393, 1431, 1425, 1394]. Octree [386]. Off [284]. Offered [221]. OFLOPS [1260]. Oil [96]. OLAP [369, 516]. OLTP [369]. Omics [602]. On-Line [598, 606, 736, 30]. On-Policy [1184]. On-the-Fly [285, 686]. On/Off [1009]. Onboard [145]. Once [73, 893]. One [1269, 178, 301, 1274, 336, 707]. One-Dimensional [178]. One-Way [1269]. Onion [729]. Online [392, 786, 1351, 1235, 220, 791, 1096, 800, 1247, 556, 1183, 1346, 799, 1428, 265]. OnlineMin [796]. Only [1095, 942]. Ontological [1127]. Ontologies [958, 1119, 1123, 558, 1120, 515]. Ontology [501, 1320, 1323, 514, 1039, 1326, 1122, 1126]. Ontology-Driven [514, 1039]. OODBMS [526]. OP2 [560]. Open [1293, 853, 763, 468, 775, 1260, 809, 1136]. Open-Audit [775]. OpenFlow [1260]. OpenMP [994]. Operable [222]. Operating [1109, 478]. Operation [95]. Operational [1134, 976, 17]. Operations [192, 207]. operative [126]. Operators [1029, 1226, 17]. Opinion [340, 341]. OPM [1127]. Opportunistic [672]. Opportunities [248]. Optical [12, 1168]. Optical-Scan [1168]. Optically [1384]. Optimal [190, 681, 73, 1250, 811, 1155, 865, 456, 77, 1042, 703]. Optimally [786]. OPTIMIS [543]. Optimisation [501]. Optimistic [298]. Optimization [1358, 863, 368, 167, 1360, 1361, 790, 866, 165, 149, 150, 521, 582, 807, 1132, 651, 112]. Optimizations [673, 1146, 613]. Optimize [334]. Optimized [737, 1197, 908]. Optimizing [431, 336]. Option [1270]. Options [1194, 1195]. Oracle [849]. Orchestrating [935, 937]. Orchestration [32]. Orchestrations [976]. Order [1046, 1090, 408, 231, 411, 982, 962, 760, 15, 176, 1045, 811, 559, 63]. Ordered [581]. Ordering [436, 697]. Orders [661]. Organisation [885]. Organisation-Based [885]. Organisations [1300]. Organised [823, 820]. Organization [1046, 824, 818]. Organizations [1008, 1224]. Organized [816, 819]. Organizing [826, 822, 1430]. Oriented [838, 441, 979, 1257, 210, 428, 882, 1309, 452, 267]. Oriya [922]. Ornamentation [907]. Orthogonal [1237]. Orthonormal [165]. OSCAR [629]. Oscillation [720]. Oscillation-Free [720]. Other [1221]. Our [494]. Outerplanar [663]. Over-Decomposition [632]. Overhead [999]. Overheads [573]. Overlap [1136]. Overlapping [175]. Overlay [78]. Overview [1, 86].

P [1104, 1105, 1106, 1108, 1109, 1110, 1112, 1113, 1115, 1101, 614, 1114, 1102]. P-Lingua [1115]. P2P [551, 502, 821]. PA [1416, 1412, 1413]. PACHI [853]. Packet [1258, 1207, 535]. Packing [800, 783]. PACUE [634]. Padding [1207]. PADL [1413]. Page [276, 277, 1069]. PageRank [529]. Pages [328]. Paging [796]. Pain [467]. Paint [399]. Pair [121]. Pair-Wise [121]. Pairs [1232]. Pairwise [664, 663]. Palovca [434]. PAM [1454]. Panel [363]. Paper [468]. Paperless [1169]. Papers [1419, 1420, 610, 1399, 1400, 1392, 1407, 1404, 1393, 1406, 1436, 1439, 1426, 1429, 1424, 1446, 1405, 1448, 1449, 1431, 1425, 1415, 1434, 1398, 1403, 1410, 1440, 1455, 1427, 1438, 1433, 1450, 1394, 1428, 1401, 1414, 1408, 1435, 1432]. PaperUI [273]. Paphos [1414]. PARA [1399, 1400]. Paraconsistent [1220]. Paradigm [839, 1257, 891]. Paradigms [1437]. ParaForming [1278]. Parallel [303, 148, 151, 1419, 1420, 590, 574, 153, 218, 584, 1278, 588, 146, 566, 638, 955, 1399, 1400, 142, 568, 1437, 1006, 1389, 175, 1101, 150, 1065, 942, 1120, 954, 615, 161, 170, 1067, 654, 1064, 582]. Parallelism [1109, 567, 1374, 1147]. Parallelizability [1102]. Parallelization [157, 952, 529]. Parallelizing [598]. Parameter [863, 1016]. Parameterised [989]. Parameterized [788, 1080, 1016, 315]. Parametric [1151, 1356]. Pareto [176]. Pareto-set [176]. Parliamentary [1173]. Parser [1043]. Part [1419, 1420, 1443, 1444, 776, 1399, 1400, 1459, 1457, 1458]. Partial [318, 661, 324, 759, 760, 506, 770, 1388, 1372]. Partially [1232]. Participate [914]. Partitioning [588, 576, 999]. Partization [655]. Parts [501]. Pass [813, 54]. Pass-Transistor [813]. PASSERINE [41]. Passive [1454]. Password [769]. Password-Based [769]. Passwords [735, 473]. Past [778]. Patch [438]. Path [167, 1121, 717, 164]. Path-Based [1121]. Paths [1096, 15]. Pattern [1219, 1378, 1364, 1313, 822, 200, 967, 860]. Patterns [579, 1047, 381, 1202, 1354, 1205, 548, 31, 1048, 867]. Paxos [75]. Payload [56]. PCA [398]. PCMs [1008]. PDE [3]. PDEs [575]. PDF [58]. PDGF [378]. Peak [453]. Pedagogic [258]. pedagogical [261]. Pedestrian [198]. Pedophile [1262]. Peer [593]. Peers [897]. Penalization [1190]. Penalized [1189]. AVC [48, 1338]. Circus [495]. ERCIM [551]. Fold [761]. Fractured [147]. ICA [1453]. Noisy [1232]. Off [1009]. OS [624]. Pilar [486]. PSO [1360, 1361]. Restart [631]. Rollback [632]. TPCE [375]. TPCH [375]. uTP [1204]. PENG [966]. Penultimate [765]. People [1347, 1295, 193, 914]. PEPR [194]. Perception [348, 1409]. Perceptual [919, 352, 351]. Perceptually [361]. Perfect [758]. perfectness [74]. Performance [639, 563, 587, 1279, 560, 1259, 1215, 277, 636, 597, 377, 944, 588, 571, 198, 366, 608, 612, 370, 1210, 1005, 633, 945, 673, 619, 1352, 678, 179, 827, 1387, 365, 1410, 594, 565, 1380, 1044, 614, 862, 1002, 620, 562, 564, 580, 615, 1266, 49, 169, 1309, 1241]. Performance-Based [377]. Performances [870]. Periphery [1046]. PerMIn [1409]. Permission [288, 894, 1147]. Permissions [415]. Permutative [1018]. Persian [1051]. Personal [250, 735, 903]. Personalising [261]. Personalization [249, 257]. Personalized [27, 250]. Persons [268]. Perspective [552, 795, 1144, 1271, 843]. Perspectives [205, 1426]. Persuasion [127]. Pervasive [478]. Petri [747, 748, 1093]. Petrochemical [96]. PGAS [586]. Phantom [194]. Phase [20, 16, 721]. Philadelphia [1416, 1412, 1413]. Phonemes [920]. Photo [45]. Photographs [110]. Phrasal [1060]. Phrase [1060]. Phrase-Based [1060]. Phrasebook [964]. Physical [1079, 313, 383]. Physically [390]. Physics [181, 1424, 8]. Physiological [1353, 1352]. Picture [45]. Piece [866]. PIGA [643]. PIGA-Virt [643]. Pilot [1064]. Pinch [904]. Pipelining [1374, 75]. Pixel [197]. Pixels [47]. PKCS#11 [290]. PKI [776, 777]. PKI-Based [777]. PL [1402, 209, 222, 205, 217]. PL-Grid [1402, 209, 222, 205, 217]. Place [923, 322]. Placement [453]. Plagiarizing [700]. Plan [873]. Planar [314, 665]. Plane [658]. Planning [1127, 885, 1057, 456, 452]. Plans [873, 126]. Plant [1312]. Platform [1003, 206, 540, 484, 221]. Platforms [65, 1215, 584, 902, 1377, 561]. Play [855, 865]. Player [851, 857]. Players [1200]. Playing [1182]. Playout [857]. PLCA [1235]. Plosives [923]. Plurals [960]. POI [326]. Point [19, 666, 947, 393]. Point-Set [666]. Pointers [479]. Points [752, 351, 658].

Poisson [148]. Policies [642, 900, 930, 32, 35, 289, 451, 1197, 38, 939]. Policy [693, 437, 1194, 896, 1184]. Polygonal [160]. Polyhedra [1078]. Polynomial [982, 1015, 170]. PoP [1211]. Population [501]. Population-Based [501]. Porting [950]. POS [1043, 1044]. Position [468]. Positional [1046]. Positioning [683]. Positions [669]. Post [1104, 203]. Power [997, 786, 1007, 95, 1179, 678, 1309, 995]. Power-State-Based [995]. PPMC [1378]. Prêt [1170, 1171]. Practical [429, 194, 28, 963, 1440, 1180, 1413]. Practice [1174, 1187, 835]. Practices [214, 30, 395]. Pragmatic [424]. Pray [1261]. Pre [186]. Pre-fetching [186]. Precise [751]. Precision [19, 400]. Preconditioner [954]. Preconditioners [157]. Predict [9]. Predictable [453]. Predicting [1141]. Prediction [908]. Predictive [324, 963, 1228, 1183]. Predictors [1007]. Preemption [791]. Preferences [502, 128]. Presence [1233, 512, 1042]. Preserved [199]. Preserving [81, 741, 772]. President [1166]. Prespecified [1270]. Prevent [54]. Prevention [293]. Preview [1305, 1301, 1291]. Priced [1082, 1025]. Pricing [1226, 816]. Primary [534]. Prime [744]. Principles [1079, 1001]. Printed [685]. Prioritizing [1291]. Priority [1280]. PRISM [1180]. Privacy [284, 933, 81, 283, 472, 734, 736, 30, 536, 897, 741, 772, 773]. Privacy-CA [933]. Privacy-Preserving [772]. Private [28]. Proactive [255]. Probabilistic [409, 982, 795, 937, 306, 992, 509, 123, 912]. Probability [1243]. Probability-Based [1243]. Probe [1261, 350]. Probing [875, 1270]. Problem [1082, 660, 1358, 659, 308, 1294, 73, 764, 176, 1247, 428, 1390, 799, 516, 789, 159, 315]. Problem-Oriented [428]. Problems [1017, 802, 786, 712, 445, 652, 1359, 783, 576, 1310, 1368, 784, 662]. procedural [405]. Procedure [699, 258]. Procedures [335, 519]. Proceedings [1443, 1444, 1459, 1457, 1458, 1423, 1442, 1395, 1460, 1445, 1441, 1416, 1430, 1412, 1409, 1417, 1447, 1451, 1422, 1421, 1418, 1456, 1413, 1397, 1454, 1453]. Process [753, 978, 495, 319, 14, 224, 1314, 1150, 467]. Processes [303, 931, 825, 87, 310, 695, 241]. Processing [1419, 1420, 1035, 566, 1443, 1444, 370, 388, 1121, 1006, 569, 922, 365, 534, 170, 349, 188, 333, 1435]. Processor [1379, 692, 634, 567, 1001, 806, 794]. Processors [171, 153, 588, 792, 529, 1380, 1002, 156, 570]. Product [838, 1299, 1354, 1353, 849]. Production [515]. Productivity [944, 612, 619, 1058]. Products [177]. Professional [207]. Profile [1127, 628, 537, 615]. Profiles [1334]. Profiling [1003, 1210, 1377, 363]. Profitability [1348]. Program [853, 153, 485, 618, 712, 483, 761, 430, 855, 1158, 1158]. Program-Size [712]. Programmable [1378]. Programmer [1158]. Programming [151, 426, 948, 1088, 1442, 886, 440, 944, 571, 293, 950, 568, 882, 1455, 424]. Programs [986, 1156, 402, 1278, 1027, 410, 411, 69, 71, 497, 306, 149, 487, 482, 1145]. Progress [882]. Project [545, 228]. Projected [1189]. Projection [194]. Projections [573, 709, 507]. Projects [539, 1294]. Proof [234, 893, 404, 867]. Proof-Carrying [893]. Proofs [1019, 1023, 290, 404, 722]. Propagation [194, 431, 147, 1346, 265]. Propagation-based [194]. ProPer [1419, 1420, 612]. Properties [1098, 1151, 352, 1159, 320, 1045, 242]. Proposal [1288]. Proposed [1335, 1365, 1288]. Proposition [302]. Propositional [511]. Propositions [395]. Prospective [878]. Prospects [817]. Protect [214]. Protecting [38]. Protection [643, 734, 94, 1396, 86]. Protein [224]. Proteins [5]. Protocol [65, 632, 694, 461, 39, 93, 741, 672, 1161]. Protocols [555, 412, 298, 1164, 292, 898, 480, 773]. Prototype [985]. Prototyping [547]. Provenance [749]. prover [477]. Provided [116]. Provider [734, 454]. Providing [1299, 1256]. Proving [761, 1227, 242]. Provisioning [210, 216]. Pruning [1359, 202]. Pseudo [17, 1241]. Pseudo-covariance [1241]. Pseudo-Differential [17]. PSI [1426]. Psycho [261, 353]. Psycho-pedagogical [261]. Psycho-Visual [353]. Psychoacoustical [918]. Public [41, 743, 145, 341, 1427]. PUF [731]. Pump [470]. Pupil [114]. Pupil-Based [114]. Purchase [1350]. Pushdown [1082]. Puzzles [24]. Python [36].

Q [1188, 228]. Q-ImPrESS [228]. QAlign [1063]. QCA [1363]. QoS [541, 644, 190, 216]. QosCosGrid [218, 208]. QR [142, 1339]. Quad [387, 1363]. Quad-Tree [387]. Quadrature [1, 19]. Qualitative [121, 128]. Quality [80, 322, 992, 1306, 199, 1291, 1456]. Quantification [582]. Quantitative [1103, 223, 824]. Quantization [356]. Quantum [208, 812, 1238, 1079, 4, 7, 1367, 808, 17]. Quantum-Inspired [812]. Quasi [50, 995]. Quasi-accurate [995]. Quasi-Chirp [50]. Quaternion [113]. Qubit [1238]. Queries [78, 1302, 326, 1130]. Query [81, 1128, 324, 504, 388, 1121, 339, 327, 1130, 1132, 333]. Querying [1023]. ‘QUESO’ [582]. Questions [258]. Queues [1280].

Race [867, 1148, 1147]. Race-Free [1147]. Radio [674]. Raga [908]. Ragas [910]. Random [708, 182, 720]. Randomized [796]. Randomness [710, 713, 717]. Range [658]. Ranging [682]. Ranking [1133, 1182, 1366, 118]. RANSAC [947]. Rao [1236]. Rapid [547]. Ratification [894]. Ratings [27]. Rational [1226]. Raven [182, 182]. Ray [387]. RC [1429]. RDF [1133]. RDPC [197]. Re [885, 31, 1263, 738, 54]. Re-authentication [738]. Re-identification [31]. Re-recording [54]. Re-wiring [1263]. Reachability [1082, 314, 528, 315]. Reaching [706]. Reactive [413, 884]. Read [581]. Read-Write [581]. Reading [275, 268]. Real [249, 1159, 1199, 1208, 724, 1235, 714, 307, 1256, 329, 1006, 556, 511, 493, 701, 439, 347]. Real-Time [1159, 1208, 1235, 307, 1256, 1006, 556, 493, 701]. Real-Valued [511]. Real-World [1199, 439, 347]. Realistic [381, 677]. Reality [1318, 383]. Really [376]. Reals [708]. Reasoning [303, 986, 1442, 402, 1320, 1301, 310, 1119, 962, 1031, 506, 1311, 1227, 1016, 1220, 241, 1120, 874]. Recall [758].

Receiver [672]. Receiver-Initiated [672]. Recharge [287]. Recognition [351, 1219, 920, 355, 1405, 107, 358, 278, 200, 195, 362, 345, 356, 1073, 347]. Recognize [1095]. Recognizer [920]. Recognizing [269, 272]. Recommendation [1320, 336, 537, 254, 118, 338, 117]. Recommendations [255]. Recommender [254, 252]. Reconfigurable [1374, 191, 1389, 1380, 1377, 1384, 1460]. Reconfiguration [1371, 1388, 1372]. Reconfiguration-Centric [1372]. Reconfigurations [975]. Reconfigured [807]. Reconstruction [194, 419]. Record [473]. recording [54]. Records [471]. Recovering [386]. Recovery [624, 456, 910, 56]. Recurrence [984]. Recursive [1191]. Reduced [654]. Reducing [573, 66, 505, 626]. Reduction [804, 171, 739, 346]. Redundancy [994, 1384]. Reduplicated [1044]. Refactoring [1278]. References [309]. Refined [955]. Refinement [307, 316, 480]. Refinement-Based [316]. Refinements [318]. Reflection [54]. REFSQ [1456]. Regard [215]. Regions [1147]. Register [418]. Registry [1287]. Regression [410, 1359, 1356]. Regret [865]. Regular [1092, 1095, 668]. Regularization [194]. Regularized [1190]. Reinforcement [1196, 1188, 1179, 1187, 1450, 1181]. Related [791, 1242]. Relatedness [1136]. Relating [318]. Relation [1274, 897, 819, 356]. Relational [509]. Relations [121, 939]. Relatively [156]. Relaxed [581]. Relaxing [664]. Relay [679]. Relevant [253, 1053, 1137]. Reliability [937, 1213]. Reliable [84, 680, 971, 74]. Remote [902, 895]. Removal [276]. Remove [593]. Reoptimization [660]. Replay [690]. Replica [83, 821]. Replication [187]. Report [491]. reported [1355]. Reporting [26]. Reports [1431]. Representation [1069, 569, 1180]. Representational [1179]. Representations [1078, 918, 156, 349]. Represented [1223]. Representing [356]. Repudiation [898, 298]. Reputation [926, 337]. Request [768, 1288]. Request-Based-Revealing [768]. Requirement [1292]. Requirements [1299, 1293, 691, 1304, 1305, 1297, 1302, 1301, 256, 1294, 30, 1300, 1291, 1295, 1289, 1290, 493, 514, 550, 847, 1456]. Requirements-Compliance [1289]. Rerouting [679]. Research [76, 1305, 1301, 220, 1340, 1291, 1435, 1068]. Researching [468]. Reservation [208]. Reservoirs [147]. Residential [1268]. Residual [1189]. Resilience [1419, 1420, 628, 622, 620]. Resiliency [620]. Resilient [993]. Resistant [1165, 51]. Resizing [1321]. Resolution [1293, 554, 1113, 510]. Resolving [1049]. Resource [823, 1014, 440, 547, 445, 985, 185, 976, 604, 556, 1004, 1440, 514, 130, 1333, 999]. Resource-Safe [440]. Resources [208, 322, 219, 447, 550]. Respiratory [1357]. Response [1368]. Responses [1353]. Restart [577]. RESTful [187]. restoration [51]. Restoring [813]. result [765]. Resultants [170]. Results [237, 826]. Resumption [247]. Retrieval [102, 55, 350, 120]. Retrospect [817]. Retrospective [883, 881, 878, 884]. Reuse [1333]. Revealing [500, 768]. Reverse [447]. Reversible [1085, 804, 813, 812, 807, 809, 806, 49, 805, 1429]. Review [883]. Revised [1392, 1419, 1420, 1407, 1404, 1393, 1406, 1436, 1439, 1426, 1429, 1424, 1446, 1405, 1399, 1400, 1448, 1449, 1431, 1425, 1415, 1434, 1398, 1403, 1410, 1440, 1455, 1427, 1438, 1433, 1450, 1394, 1428, 1401, 1414, 1408, 1435, 1432]. Revising [500]. Revision [67, 132, 191]. Revisited [708, 436, 792, 895, 1281, 47]. Revisiting [852]. RevKit [809]. Revocable [893]. Rewrite [1021, 1015]. Rewriting [229, 320]. Reykjavík [1399, 1400]. RFID [287, 903, 773]. Rice [1312]. Rich [325]. Rigid [1343]. Rigidity [1046]. Ring [729]. RISC [692]. Risk [899, 1297, 298, 776, 1059, 454, 889]. Risk-Aware [899]. RL [1180]. RMWE [1044]. RNN [920]. Road [326, 327, 820]. Roadpricing [774]. robot [521]. Robotic [259]. Robotics [1345]. Robots [525, 1345, 1327, 1325]. Robust [78, 1196, 1010, 947, 48, 52, 156, 1181]. Robustness [790, 826, 992]. Role [1248, 899, 923, 901]. Role-Based [899]. Roles [1345]. Rollback [555]. Rooted [409]. Rooting [740]. Rotatable [79]. Round [662]. Round-Trip [662]. Router [801]. Routing [1202, 1385, 820]. RP [1121]. RP-Filter [1121]. Rule [1094, 346, 514, 515, 197]. Rule-Based [346, 514, 515]. Rules [1106, 921, 215, 561, 344]. Run [1379, 753, 376, 1371]. Run-Time [1379, 753, 1371]. run.time [845]. Running [230]. Runtime [1142, 846, 695, 737, 1143, 1150, 1144, 1149, 1155, 1154, 1448]. Runtimes [722]. Rural [360]. Russia [1426]. RV [1448].

S [1424]. Saarbrücken [1428]. SADmote [1010]. Safe [440, 463, 311, 1280, 38]. Safety [1302, 471, 188]. Saga [457]. Salesman [1358]. Saliency [1337]. Salt [904]. samples [397]. Sampling [1209, 1216]. San [1448]. Sane [38]. Sanitizable [901]. SAP [293]. SARA [212]. SAT [798]. Satisfying [508]. Saturation [1153]. Saturation-Based [1153]. SCADA [92, 1264, 94, 93]. Scaffolding [248]. Scalability [1202, 536]. Scalable [616, 154, 1383, 614, 599, 1382]. ScalableCore [1382]. Scale [573, 65, 921, 375, 583, 1361, 630, 55, 648, 952, 1124]. Scaling [352, 1179, 169]. Scan [1168]. Scanner [400]. Scanning [1387]. ScanTraffic [684]. Scenario [369]. Scenarios [255]. Scene [270, 267, 111, 271, 272]. Scenes [269, 394]. Schedulability [307]. Schedule [585]. Scheduling [587, 642, 948, 791, 608, 592, 783, 792, 553, 801, 785, 1004, 647, 570, 794]. Scheme [10, 1379, 523, 56, 527, 48, 768, 1342, 52, 61, 530, 46, 49, 62]. Schemes [532, 901, 297, 1073]. Schrödinger [18]. Science [1392, 222, 1424, 220, 542, 1081, 707]. Scientific [579, 1402, 571, 548, 1399, 1400, 1376, 617]. Scoping [319]. Scotch [363]. Scott [756]. Scout [613]. Screen [54]. Scribble [113]. script [267]. Scripting [221]. Scripts [270]. SDM [546]. Seamless [209]. Search [1017, 823, 929, 854, 866, 1033, 1153, 334, 1197, 857, 599, 915, 867, 856, 852, 344]. Search-Based [1153]. Searching [224, 329]. Seattle [1410]. Second [1448, 176, 892, 1440, 1438]. Second-Order [176]. Secrecy [699]. Secret [768, 1342]. Secrets [500]. Sector [96]. Sector-Specific [96]. Sectors [88]. Secure [1393, 1423, 285, 780, 40, 936, 287, 34, 694, 293, 210, 762, 1425, 211, 903].

Securing [59]. Security [92, 518, 1393, 214, 1406, 691, 524, 933, 1297, 283, 37, 212, 979, 776, 1166, 735, 26, 695, 1425, 892, 480, 891, 701, 22, 895, 1401, 1434]. See [1290]. Segment [106, 433]. Segment-Level [106]. Segment-Swapping [433]. Segmentation [160, 1329, 271]. Segments [657]. Seismic [147]. Selected [1392, 1419, 1420, 1407, 1404, 1393, 1406, 1436, 1439, 1426, 1424, 1446, 1405, 1399, 1400, 1448, 1449, 1431, 1425, 1415, 1434, 1398, 1403, 1410, 1440, 1455, 1427, 1433, 1450, 1394, 1428, 1401, 1414, 1408, 1435, 1432]. Selection [190, 681, 53, 1130]. Selective [132, 265]. Self [823, 786, 56, 223, 824, 826, 1355, 38, 816, 820, 822, 818, 819, 51, 9, 1430]. Self-assembled [223]. Self-healing [9]. Self-Organised [823, 820]. Self-Organization [824, 818]. Self-Organized [816, 819]. Self-Organizing [826, 822, 1430]. Self-recovery [56]. Self-reported [1355]. Self-restoration [51]. Self-tuning [786]. Semantic [105, 841, 262, 1319, 1301, 263, 1123, 1129, 1053, 547, 1220, 1447, 189, 192, 1366, 323, 1131, 1136, 961]. Semantic-Based [323]. Semantically [253, 1298, 1138]. Semantics [105, 441, 1134, 509, 239, 976, 882, 404, 478, 1039]. Semi [1235, 558, 1064, 341, 51, 1356]. Semi-automatic [558, 341]. Semi-fragile [51]. Semi-parallel [1064]. Semi-parametric [1356]. Semi-supervised [1235]. Semicomputable [708]. Semiconductors [954]. Sense [1052, 1065]. SenShare [676]. Sensing [81, 676, 687, 345]. Sensitive [201, 1142]. Sensor [523, 780, 683, 83, 80, 82, 681, 680, 677, 79, 673, 676, 1314, 533, 682, 1422, 741, 679, 687, 685, 675, 1333, 188]. Sensors [686, 79, 903]. Sensory [363]. Sentences [967]. Sentiment [341]. Separate [651]. Separation [1248, 1243, 1231, 1235, 1250, 490, 1240, 1453, 1241]. September [1419, 1420, 1406, 1439, 1405, 1448, 1449, 1410, 1427, 1438, 1450, 1428, 1408]. Sequence [1047, 1060, 1257, 648, 428, 1296, 599]. Sequence-Oriented [1257]. Sequences [1046, 720]. Sequential [931, 1048]. Serial [1381]. Serializability [1152]. Serialization [1325]. Series [1214, 1228, 908, 654, 1356]. Series-Parallel [654]. Serious [104]. Server [563, 799]. Service [838, 606, 932, 972, 885, 610, 190, 608, 979, 291, 734, 992, 736, 30, 611, 738, 191, 1313, 740, 189, 454, 192, 207, 609, 457, 217, 605]. Service-Level [606]. Service-Oriented [838, 979]. Services [187, 729, 1439, 1319, 541, 190, 937, 98, 970, 465, 1313, 973, 976, 450, 536, 189, 1427, 206, 603, 449, 1414]. Session [927, 23]. Sessions [926, 1212, 31]. Set [1024, 788, 235, 795, 666, 508, 1152, 332, 176]. Sets [715, 665, 1242, 938, 1056, 481]. Setting [743]. Settings [694]. Setup [355]. Shallow [168]. Shaping [365]. Share [545, 1347]. Shared [451, 152, 1213]. Shares [60]. Sharing [229, 1342, 490]. Shiksha [361]. Shogi [866]. Shopping [1351]. Short [302, 1070, 54]. Short-Circuit [302]. Shortcut [1284]. Should [1137]. Shrine [227]. SI [373]. SI-CV [373]. Sibling [864]. Side [1175]. Side-Effects [1175]. SIFT [277, 350]. Sign [195]. Signal [1453]. Signaling [1205]. Signature [901]. Signatures [769, 240, 111]. Significant [361]. Signing [412]. SIMD [1374, 613]. SIMD-Optimizations [613].

Similar [329]. Similarity [1208, 1075, 599, 332, 338]. Simple [507, 1111]. Simpler [797, 798]. Simulated [166]. Simulating [622, 382]. Simulation [574, 825, 1003, 155, 4, 1100, 826, 677, 224, 1222, 147, 390, 582, 1292, 391, 383]. Simulation-Based [826]. Simulations [318, 598, 20, 381, 589, 163, 5, 1386, 172, 1228, 168]. Simulator [1105, 1115, 1382]. Simultaneous [1249]. Singing [907, 916]. Single [787, 1168, 326, 1380, 206, 998]. Single-Chip [998]. Single-Source [326]. Singular [52]. Sinhala [1050]. Sireum [486]. Sireum/Pilar [486]. Site [1334, 456]. Sites [1347]. Sixty [706]. Size [712, 505, 674]. Skeletonization [393]. Sketch [350]. Skew [374, 1249]. Skew-Symmetric [1249]. Skin [389]. Skin-Detached [389]. Skiplists [1280]. Skype [1206]. SLA [185, 451]. SLA- [185]. SLAM [1247]. SLCA [334]. Sleepy [1340]. Slicing [1021, 1273]. SLM [546]. Slope [336]. Slot [1072]. Slots [1379]. Slovakia [1392]. Small [184, 274, 1375]. Smart [684, 1027, 680, 641, 345, 34]. Smart-HOP [680]. Smartphone [732, 700]. Smoothers [955]. Smoothing [1356]. SMPSuperscalar [943]. SMT [484]. SMT-Based [484]. SNA [1408]. Snap [68]. Snap-Stabilizing [68]. Snapshot [373, 498]. Snapshotable [490]. SOA [228, 585]. SOA-Grid [585]. Social [262, 1029, 1347, 263, 757, 1129, 102, 337, 1334, 259, 339, 531, 1346, 513, 265, 252, 118, 117]. Socially [253]. Societies [881]. Society [1345]. Soft [628, 627, 1006, 626]. Software [1299, 831, 1407, 1423, 575, 886, 833, 842, 834, 571, 572, 625, 425, 1306, 312, 403, 1416, 1431, 1291, 1153, 150, 832, 1456, 835]. Software-Based [625]. Solenoid [312]. Solid [390]. Solution [377, 933, 858, 16, 1365, 954, 274, 120]. Solutions [802, 1288, 1240, 928]. Solvable [18]. Solve [534]. Solved [868]. Solver [148, 590, 984]. Solvers [572]. Solving [1017, 1024, 129, 445, 178, 73, 1310, 764, 428, 867]. Some [282, 802, 719]. Sood [530]. sorted [136]. Sound [919, 1167, 421, 917, 699, 1435]. Source [1293, 853, 1243, 618, 1231, 1235, 844, 326, 613, 1240, 809, 1130]. Source-to-Source [613]. Sources [1248, 339, 1137, 904]. South [1415]. Space [1085, 1007, 1091, 308, 1059, 744]. Spaced [161]. Spain [1404, 1445, 1398, 1440, 1455]. Spanish [1073]. Spanning [789]. SPARQL [1134, 1121, 1132]. Sparse [590, 579, 635, 1192]. Spatial [1099, 121, 388, 356, 112]. Spatial-Temporal [356]. Spatio [822]. Spatio-temporal [822]. Specific [440, 979, 1038, 96, 1373, 1058]. Specification [1049, 437, 32, 934, 495, 308, 238, 532, 442]. Specifications [426, 237, 233, 320, 420, 751]. Specified [669]. Specifying [883, 849]. SPECS [154]. Spectra [819]. Spectral [1344, 160]. Spectrum [53, 239, 57]. Spectrum-Based [57]. Specular [54]. Speech [921, 922, 882, 1066, 1435]. Speech-Act [882]. Speech-to-Speech [1066]. Speed [1268, 739]. SPH [172]. Spherical [578]. Spiking [1105, 1115, 1114]. Spirits [363]. Splicing [63]. Spline [47]. Splitting [124, 414]. Sports [104]. Spread [53, 57]. Sprite [384]. SQL [755]. Squares [1190, 1191]. SRAM [731]. SRL [1371]. SRL-Based [1371]. SRT [541]. SRT-15 [541]. Stabilizing [68]. Stable [134, 652, 797]. Staging [631]. Standard [365]. standards [892]. Stará [1392]. State [1156, 900, 853, 873, 308, 1154, 995]. State-Based [873]. Stateful [1259]. Statement [1294]. States [966]. Static [930, 753, 711, 212, 421, 494, 996]. Stations [849]. Statistical [1028, 917, 582, 1058, 63, 421]. Statistics [598, 1207]. Status [125, 734]. Statutory [901]. Stealing [1006, 675]. Stearns [714]. Steganalysis [44, 1338, 47]. Steganography [44, 46]. Stem [1042]. Stencil [158, 1373]. Step [139, 516, 49]. Steps [1205, 942]. STM [1434]. Stochastic [598, 6, 785]. Stochastically [1219, 793]. Stochastically-Located [793]. Stock [87]. Storage [211, 816, 457, 216]. Store [411]. Stored [335]. Stories [111]. Story [120]. Storytelling [395]. Straight [197]. Strategic [929, 937]. Strategies [1226, 1006, 535, 700]. Strategy [214, 230, 87, 1335, 776, 865, 1308]. Stream [1257, 648, 330, 1006]. Stream-Based [648]. Streaming [159]. Streamlined [1011]. Strengthening [473]. Stressed [1352]. Strings [667]. Strip [800]. Strong [1090, 309, 533, 1241]. Strongly [796]. Structural [192]. Structure [1358, 224, 240]. Structured [430, 1072, 974, 801, 331]. Structured-ness [331]. Structures [754]. Student [260, 258]. Studied [5]. Studies [842]. Study [194, 131, 993, 1315, 67, 223, 826, 1166, 1368, 1303, 329, 635, 536, 1064, 493, 1346, 1296, 637, 1246, 63]. Stylistics [1048]. Sub [1208]. Sub-application [1208]. Subgraph [660]. Subject [20]. Subjectivity [889]. Sublist [481]. Subsequence [949]. Subset [784]. Subspace [573, 1245, 1067, 1246]. Subterfuge [288]. Subtractive [1329]. Subtyping [1274]. Successful [449]. Suffix [1042]. Suitable [1303]. Sum [501]. Summarizing [1133]. Summary [839, 845, 842, 848]. SUP [1030]. Supercapacitors [685]. Supercompilation [764, 765]. Supervised [1068, 1235]. Supply [1222]. Support [554, 324, 256, 30, 611, 246]. Supporting [831, 247, 691, 840, 1300, 339, 1296]. Surathkal [1401]. Surface [389]. Survey [283, 37]. Suspect [512]. SVM [1044]. Swapping [433]. Swarm [827]. Swarms [827, 1204]. Switch [1385, 1260]. SyFi [1259]. Symbiotic [115]. Symbolic [432, 1078, 491, 766, 442]. Symmetric [1249, 1390]. Symmetry [723]. Symplectic [1249]. Symposia [1431]. Symposium [1423, 833, 1436, 1460, 830, 1415, 1417, 1455, 1413, 1435]. Synchronizability [406]. Synchronization [840, 1108, 581]. Synchronous [1248, 753, 70]. Synchrony [74]. Synonym [1056]. Syntactic [1049, 1060]. Syntax [1298]. Synthesis [919, 397, 921, 416, 384, 420, 916, 1180]. Synthesizable [1001]. Synthesizing [412]. Synthetic [1100]. System [1017, 823, 1163, 1357, 186, 131, 926, 623, 1315, 1105, 396, 353, 921, 264, 355, 212, 757, 215, 1224, 902, 1135, 1111, 1310, 1125, 154, 16, 1256, 556, 893, 1339, 1059, 127, 362, 562, 457, 345, 1382, 489, 1309, 995, 577, 338, 1108]. System-Level [562, 995]. Systematic [1259]. SystemC [1003, 71]. Systems [1082, 1103, 1104, 1098, 587, 499, 317, 1393, 78, 1423, 997, 406, 691, 1025, 171, 875, 95, 1142, 1090, 408, 440, 643, 1315, 1174, 502, 232, 1426, 1302, 934, 13, 81, 70, 413, 756, 1109, 758, 824, 3, 370, 1005, 314, 992, 1441, 989, 294, 6, 1311, 592, 1112, 1113, 473, 313, 1368, 420, 764, 7, 953, 1430, 305, 1425, 372, 1415, 1376, 1417, 1383, 1115, 827, 750, 1364, 1313, 1006, 1015, 468, 635, 150, 614, 1114, 1459, 1457, 1458, 817, 1346, 647, 91, 515, 254, 878, 701, 367, 478, 446, 819, 252, 743, 1414, 315]. Systems [1066, 1101]. Systolic [1373].

Table [1375]. Table-Based [1375]. Tables [1041]. Tabling [430]. TAFA [1398]. Tag [1334, 118]. Tag-Based [1334]. Taggers [1043]. Tagging [1044, 1065, 1346]. Tags [263, 116]. Taint [36, 35]. Taipei [1433]. Taiwan [1459, 1457, 1458, 1433]. Taken [942]. Taking [87]. Talk [1178, 1179, 1180, 1181]. Tallinn [1449, 1425]. Taming [182]. Tanpura [911]. Tape [804]. Target [326]. Targeting [465]. Task [1379, 1352, 1004, 697, 996, 994]. Task-Scheduling [1004]. Tasks [200]. Tau [661]. Taxonomy [1038]. TB [1040]. TBA [294]. TC [1430]. TCG [933]. TCP [1210, 673, 1204]. TCP/uTP [1204]. Teaching [834, 1143, 404, 835, 361]. Team [1355, 246]. Team-Member [1355]. Teams [1225, 521]. Technical [1402, 1162]. Technique [1209, 485, 1167, 199, 473]. Techniques [25, 1278, 1315, 1320, 173, 283, 583, 1403, 1155, 700, 1048, 967, 169]. Technologies [1437, 1433]. Technology [1393, 1335, 731, 1425, 1355, 1410, 1447, 534, 1418, 104, 849, 347, 1318]. Tehran [1407]. Tel [1453]. Tele [361]. Tele-teaching [361]. Telecommunication [1226]. Telecommunications [97]. Tempering [1389]. Template [1299]. Templates [807]. Temporal [1151, 1028, 1190, 1311, 1072, 862, 361, 356, 861, 822]. Temporally [1193]. Termination [407, 1090]. Terms [1158]. Terrain [397, 387]. Test [1157, 243]. Test-Case [243]. Test-Trace [1157]. Testing [1084, 317, 1027, 1369, 1153, 182, 849]. Tests [242]. Text [1075, 270, 1443, 1444, 267, 269, 278, 1070, 1067, 268, 271, 1068]. Texts [1064, 1073]. Texture [397]. TFP [1455]. TGC [1436]. Their [1092, 819]. Theorem [1227, 719]. Theoretic [234]. Theoretical [1424, 455]. Theorie [1398]. Theories [1021, 1347, 1416, 395]. Theory [10, 226, 1445, 1079, 354, 1117]. Therapeutic [910]. There [724, 45]. Thermal [567]. Thesis [1079]. Thick [577]. Thick-Restart [577]. Things [1295, 424]. Think [301]. Thinking [1035]. Third [1429, 1449, 1410, 63]. Thoughts [892]. Thread [1280, 1380]. Thread-Safe [1280]. threaded [410, 311, 1145]. Threat [793]. Threats [89]. Three [1043, 58, 133, 653]. Three-Dimensional [653]. Three-Layer [133]. Threshold [196, 681, 1332]. Throughput [1390, 75]. Thus [942]. Tickets [287]. Tight [657, 1196]. Tilburg [1432]. Tiled [996]. Time [1379, 804, 656, 854, 1159, 1025, 1208, 753, 1142, 198, 982, 13, 375, 1214, 1235, 714, 601, 1371, 307, 1227, 678, 453, 974, 1256, 329, 239, 1006, 556, 667, 1155, 918, 1228, 908, 682, 493, 988, 701, 1204, 1356, 794]. Time-Constrained [453]. Time-Consuming [601]. Time-Dependent [1227]. Time-Frequency [918]. Time-Sensitive [1142]. Time-Triggered [1155]. Timed [1082, 1077, 237, 754, 134, 756, 760, 528]. Timely [84]. Timestamp [1270]. Timing [987, 989, 470]. TLB [492]. TMA [1451]. Together [1099, 889]. Tolerance [499, 71, 630, 994]. Tolerant [679]. Tomography [194]. Tone [62]. Tool [771, 129, 235, 616, 977, 1283, 465, 1061, 615]. Toolkit [216, 809]. Tools [839, 429, 1304, 359, 584, 1460, 571, 612, 1416, 30, 763]. Top [222, 327]. Top- [327]. Topics [1410]. Topologies [68]. Topology [1091, 533, 1309]. TPC [365, 1410, 374, 372, 378]. TPC-E [372]. TPC-H [374, 378]. TPCC [375]. TPCC/TPCE/TPCH [375]. TPCTC [1410]. TPM [902]. Trace [1021, 1140, 1302, 1157]. Traceability [1303]. Traces [1191, 1141]. Tracings [917]. Tracker [114, 115]. Tracking [29, 730, 1351, 1354, 260, 193, 195]. Tractability [1016]. Trade [284, 818]. Trade-Off [284]. Tradeoff [678]. Trading [450]. Traditional [1135]. Traffic [684, 1264, 1208, 1215, 1269, 1201, 1255, 24, 1252, 1256, 1254, 1253, 1451]. Train [849, 489]. Training [104, 383]. Trajectories [329, 323]. Transaction [69, 71, 365]. Transaction-Based [69]. Transactional [66, 936]. Transactions [230, 936, 1452, 1411]. Transcendence [714]. Transfer [20, 317, 1353, 1068]. Transfinite [719]. Transform [591, 578, 52, 50]. Transformation [618, 236, 430, 238, 1241]. Transformations [761, 304]. Transformator [613]. Transforming [676, 810]. Transistor [813]. Transition [20, 1025, 756, 16, 721]. Transitions [765]. Transitive [504]. Translation [1061, 1065, 1058, 1066]. Translator [527, 1061]. Transmission [74]. Transparency [1162]. Transparent [66, 1298]. Transport [954]. Transportation [99]. Transversal [693]. Trapped [433]. Travel [656]. Traveling [1358]. Traversal [1132]. Tree [854, 78, 68, 324, 866, 695, 387, 1197, 857, 789, 856, 442, 852]. Tree-Based [442]. Treebank [1040]. Trees [670, 1053, 665, 490, 1281, 203, 654]. Trends [1225, 44, 1403, 1455]. Trento [1422]. Triadic [818]. Triangle [663]. Triangle-Free [663]. Triggered [1155]. Trilinos [574]. Trip [404, 662]. Triple [1121, 1384]. Trivitas [1172]. Troubleshooting [1212]. Trust [1406, 1168, 1353, 890, 892, 1434, 697, 901, 891, 889, 1162, 703]. Trusted [33, 897]. Trustworthy [27, 1436]. tsdb [1214]. TSP [787]. TTCS [527]. Tunable [625]. Tuning [863, 619, 75, 786, 173, 617]. Tunneling [167, 391]. Tuple [935]. Tuple-Based [935]. Turing [804, 1093, 1079]. Turkish [1037]. Tutorial [1080]. Two [522, 11, 591, 709, 40, 270, 13, 483, 215, 851, 1242, 15, 682, 822, 1366, 49]. Two-Dimensional [591, 822]. Two-Loop [11]. Two-Player [851]. Two-Step [49]. Two-Way [682]. Type [293, 178, 358, 1329, 1273]. Type- [358]. Type-2 [1329]. Type-Based [293]. Typed [692]. Typeful [958]. Types [403]. Typing [983].

UCHPC [1419, 1420, 633]. UCRL [1178]. UCT [851]. UDP [1252]. Ultra [1007]. UMAP [1404]. UML [67, 232, 844]. UML-Like [232]. Un-Alignable [1064]. Uncertain [336]. Uncertainty [1031, 582, 889]. Unconstrained [120, 111]. UnConventional [633]. Uncorrelating [1241]. Uncoupling [1238]. Uncovering [1200]. Undercutting [503]. Understandability [843]. Understanding [623, 970, 89, 1221, 260, 1346, 870]. Unevenly [161]. Unfold [761]. Unfold/ Fold [761]. UNICORE [209]. Unidirectionality [721]. Unification [1083]. Unified [1194]. Uniform [497, 792, 1094]. Unifying [338]. Unilaterally [292]. Unitary [165]. Units [170]. Universal [771, 69, 507, 665]. Universality [716]. Unmasking [1252]. Unordered [670]. Unreliable [937, 971]. Unsolicited [24]. Unstructured [560]. Unsupervised [1243]. Untrusted [904]. Unwinding [939]. Update [795, 1134]. Updates [58]. UPnP [1261]. Upon [621]. Upscaling [1312]. Urban [81]. Urdu [1040]. URDU.KON [1040]. URDU.KON-TB [1040]. USA [1416, 1448, 1412, 1410, 1413]. Usability [1369, 1368]. Usable [518]. Usage [985, 1203, 617]. Use [429, 208, 27, 248, 583, 250, 893, 964, 59]. Use-Once [893]. Used [743]. User [249, 253, 1404, 730, 1129, 734, 1255, 31, 634, 210, 736, 30, 1334, 1125, 251, 219, 206, 609, 537, 116, 338]. User-Friendly [219]. User-Oriented [210]. Users [1334, 360]. Uses [1347]. Using [1357, 1047, 121, 209, 1208, 920, 948, 875, 218, 841, 1278, 276, 277, 1041, 1075, 1193, 1320, 834, 1369, 1212, 1301, 35, 1261, 113, 288, 1038, 505, 438, 1053, 547, 761, 445, 430, 1337, 945, 735, 955, 1344, 913, 199, 179, 172, 176, 479, 1194, 336, 827, 191, 461, 487, 1232, 1389, 667, 532, 1387, 816, 1228, 1312, 1054, 908, 189, 1064, 807, 350, 593, 361, 1213, 687, 877, 967, 994, 1388, 47, 489, 1067, 197, 356, 49, 60, 54, 333, 363, 1122, 1136, 1270]. Utility [284].

V1 [348]. Valid [78]. Validating [489]. Validation [1314, 778, 698]. Value [52, 1192]. Valued [511, 1326]. Values [866]. Variability [838, 836, 837]. Variable [56, 1453]. Variables [819]. Variants [142, 862, 1102]. Various [206]. VARY [836]. Varying [384]. VDM [489]. Vector [1338, 1239, 580]. Vectors [1208]. Vehicle [772]. Vehicle-to-Grid [772]. Vehicular [1205]. Velocity [82]. Venezuela [1442]. Verb [1051]. Verbalization [958]. Verifiable [1169, 289]. Verification [405, 492, 1163, 406, 886, 25, 408, 485, 410, 320, 934, 977, 495, 483, 477, 1112, 403, 1167, 491, 763, 1448, 479, 1143, 985, 1150, 1144, 490, 1155, 476, 478, 486, 1154, 1309, 1412]. Verified [179, 482, 1416]. Verifier [411]. Verifying [282, 1172, 483, 480, 514]. VerisoftXT [478]. Version [174]. Versions [373]. versus [1140, 1107, 1255]. Vertex [660, 655]. Vertical [1009]. Vertices [669]. Very [995]. VHPC [1419, 1420, 639]. VI [1452]. via [1087, 730, 1319, 36, 885, 414, 1138, 108, 1380, 1130, 315]. Viability [630]. Video [384, 109, 48, 1282, 108, 57, 59, 1318, 111, 117]. Vienna [1451, 1454]. View [1367, 1254, 57, 1271, 394, 840]. View-Dependent [394]. Viewpoint [1125]. Views [755, 333]. VII [1411]. Violations [1152]. Virt [643]. Virtual [642, 78, 643, 396, 223, 222, 628, 72, 224, 453, 458, 211, 641, 565, 647, 357, 1318, 383, 1122]. Virtualization [639, 492, 377, 1002, 562]. Virtualized [210, 640, 645]. Virtualizing [564, 439]. Virus [1387]. Visibility [525]. Vision [399, 543]. Visual [353, 1298, 61, 348, 325, 1350, 60, 344, 62, 112]. Visualization [247, 155, 1303, 387, 182, 743]. Visualizations [615]. VMCAI [1412]. VMware [563]. Vocabulary [1133, 112]. Voice [1368, 916].

Volatility [1292]. Voltage [996]. Volunteer [594]. Voronoi [656, 657, 859]. Vote [1165]. VoteID [1449]. Voter [1167, 1170, 1171]. Voters [1172]. Votes [1172]. Voting [1163, 1169, 1164, 733, 1168, 1161, 1175, 1173, 1166, 1165, 1162, 728, 1449]. Vowels [923]. VPH [545]. VPH-Share [545]. vs [428, 1065, 744]. VSTTE [1416]. VT [1385]. Vulnerabilities [89, 93]. Vulnerability [95, 627, 698].

WA [1410]. WADT [1403]. WAIM [1408]. WALCOM [1421, 1421]. Walkthrough [388]. Wall [155]. WAN [84]. WAOA [1428]. War [476]. Warehouse [1257]. Warren [436]. Wasn’t [26]. Water [91, 168]. Watermark [56]. Watermarking [56, 48, 55, 53, 52, 50, 1394, 57, 51]. Watermarks [59]. Waveguides [12]. Wavelength [54]. Wavelet [591, 199, 52]. Waves [147]. Way [1269, 193, 682]. WCET [766]. Weak [931, 1234, 473]. Weakly [1068]. Weakly-Supervised [1068]. Weather [155]. Web [1447, 187, 730, 1439, 262, 1319, 222, 885, 1212, 190, 283, 37, 220, 1129, 1069, 1200, 601, 776, 337, 1166, 31, 191, 973, 976, 325, 1220, 189, 192, 252, 1408, 1131, 328]. Web-Accessible [222]. Web-Age [1408]. Web-Based [220, 601]. Web-Security [1166]. Webpages [1071]. Website [1350]. Weighted [660, 1028, 528, 428, 794]. Weighted-Sequence [428]. Wellington [1431]. WGIM [1408]. Whale [405]. Wheat [651]. Where [102]. Which [1303]. Whisky [363]. Whiteboard [275]. Whitening [1248]. Whose [678]. Why3 [483]. Widgets [1138]. Wiener [922]. WiFi [740]. WiFi-Accessible [740]. Wiki [1124]. Wikipedia [1038]. Wiktionary [1054]. Will [1044, 1345]. Window [520]. Windows [1149]. Wireless [523, 993, 683, 83, 82, 681, 680, 677, 673, 533, 817, 535, 741, 679, 687, 1266, 1009, 1422]. wiring [1263]. Wise [121]. within [1225]. without [500, 791, 668, 61, 295]. Word [1052, 1045]. WordNet [1036, 1065, 189]. Words [107, 108]. Work [1006, 1254]. Workbench [1222]. Workflow [121, 1311, 619, 465, 975, 467, 452]. Workflows [900, 464, 1283, 463, 558, 600, 585, 169]. Working [795, 1456, 966]. Workload [369, 376, 366, 635, 999]. Workloads [375]. Workshop [639, 839, 1406, 845, 597, 1439, 842, 551, 1429, 539, 1424, 566, 612, 633, 1405, 1430, 848, 1434, 1398, 1403, 1440, 1451, 1427, 1421, 1438, 1433, 1450, 620, 562, 1394, 1428, 1414]. Workshops [1419, 1420, 1404, 1431, 1408]. World [249, 1199, 184, 818, 439, 347]. Wrappers [38]. Write [581, 1171]. Write-Ins [1171]. Writing [1300]. Written [1046]. WS [1439]. WS-FM [1439]. WSD [1053]. WTCS [1424]. Wuhan [1408]. Würfelt [855].

x86 [1141]. XACML [696]. XcalableMP [586]. XIO [334]. XIO-SLCA [334]. XML [335, 334, 498, 331, 332, 333]. XMLDM [1408]. XSC [179].

Years [880]. Yourself [114]. YouTube [253].

Zealand [1424, 1431].

References

Adam:2012:BAA

[1] Gheorghe Adam and Sanda Adam. Bayesian automatic adaptive quadrature: An overview. Lecture Notes in Computer Science, 7125:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_1/.

Friese:2012:CCC

[2] Volker Friese. Computational challenges for the CBM experiment. Lecture Notes in Computer Science, 7125:17–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28212-6_2/.

Gerdt:2012:CAF

[3] Vladimir P. Gerdt. Consistency analysis of finite difference approximations to PDE systems. Lecture Notes in Computer Science, 7125:28– 42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_3/.

Gerdt:2012:CMQ

[4] Vladimir P. Gerdt and Alexander N. Prokopenya. The circuit model of quantum computation and its simulation with Mathematica. Lecture Notes in Computer Science, 7125:43–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28212-6_4/.

Hansmann:2012:PSC

[5] Ulrich H. E. Hansmann. Proteins studied by computer simulations. Lecture Notes in Computer Science, 7125:56–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28212-6_5/.

Honkonen:2012:FMS

[6] Juha Honkonen. Functional methods in stochastic systems. Lecture Notes in Computer Science, 7125:66–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28212-6_6/.

Kornyak:2012:MMF

[7] Vladimir V. Kornyak. Mathematical modeling of finite quantum systems. Lecture Notes in Computer Science, 7125:79–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28212-6_7/.

Melezhik:2012:MCC

[8] Vladimir S. Melezhik. Multi-channel computations in low-dimensional few-body physics. Lecture Notes in Computer Science, 7125:94–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_8/.

Zemskov:2012:MMP

[9] Serguey V. Zemskov, Henk M. Jonkers, and Fred J. Vermolen. Mathematical models to predict the critical conditions for bacterial selfhealing of concrete. Lecture Notes in Computer Science, 7125:108–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_9/.

Acatrinei:2012:NDS

[10] Ciprian Sorin Acatrinei. A new discretization scheme in field theory. Lecture Notes in Computer Science, 7125:122–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28212-6_10/.

Antonov:2012:TLC

[11] Nikolay V. Antonov and Nikolay M. Gulitskiy. Two-loop calculation of the anomalous exponents in the Kazantsev–Kraichnan model of magnetic hydrodynamics. Lecture Notes in Computer Science, 7125:128– 135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_11/.

Ayryan:2012:MMI

[12] Edik A. Ayryan, Alexander A. Egorov, Leonid A. Sevastyanov, and Konstantin P. Lovetskiy. Mathematical modeling of irregular integrated optical waveguides. Lecture Notes in Computer Science, 7125:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_12/.

Dano:2012:TNC

[13] Ivan Daňo. Two notes on continuous-time neurodynamical systems. Lecture Notes in Computer Science, 7125:148–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28212-6_13/.

Hnatic:2012:MMA

[14] Michal Hnatič, Juha Honkonen, and Tomáš Lučivjanský. On the mathematical modelling of the annihilation process. Lecture Notes in Computer Science, 7125:154–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28212-6_14/.

Klesc:2012:CNJ

[15] Marián Klešč and Štefan Schrötter. The crossing numbers of join of paths and cycles with two graphs of order five. Lecture Notes in Computer Science, 7125:160–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28212-6_15/.

Litavcova:2012:ESM

[16] Eva Litavcová, Miron Pavluš, Ján Seman, and Ibrohim Sarhadov. Exact solution of a moisture drying system with phase transition. Lecture Notes in Computer Science, 7125:168–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28212-6_16/.

Sevastyanov:2012:PDO

[17] Leonid Sevastyanov, Alexander Zorin, and Alexander Gorbachev. Pseudo-differential operators in an operational model of the quantum measurement of observables. Lecture Notes in Computer Science, 7125:174–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28212-6_17/.

Suzko:2012:ESM

[18] Alina Suzko and Elena Velicheva. Exactly solvable models for the generalized Schrödinger equation. Lecture Notes in Computer Science, 7125:182– 188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_18/.

Adam:2012:FPD

[19] Sanda Adam and Gheorghe Adam. Floating point degree of precision in numerical quadrature. Lecture Notes in Computer Science, 7125:189– 194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_19/.

Amirkhanov:2012:NSH

[20] Ilkizar V. Amirkhanov, Taisia P. Puzynina, Igor V. Puzynin, and Ibrohim Sarhadov. Numerical simulations of heat and moisture transfer subject to the phase transition. Lecture Notes in Computer Science, 7125:195– 200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28212-6_20/.

Anonymous:2012:FMa

[21] Anonymous. Front matter. Lecture Notes in Computer Science, 7125: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28212-6/1.

Sarela:2012:BSB

[22] Mikko Särelä, Christian Esteve Rothenberg, András Zahemszky, Pekka Nikander, and Jörg Ott. BloomCasting: Security in Bloom filter based multicast. Lecture Notes in Computer Science, 7127:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27937-9_1/.

Suoranta:2012:ASM

[23] Sanna Suoranta, Jani Heikkinen, and Pekka Silvekoski. Authentication session migration. Lecture Notes in Computer Science, 7127:17– 32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27937-9_2/.

Komu:2012:MUT

[24] Miika Komu, Sasu Tarkoma, and Andrey Lukyanenko. Mitigation of unsolicited traffic across domains with host identities and puzzles. Lecture Notes in Computer Science, 7127:33–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27937-9_3/.

Borgaonkar:2012:EAF

[25] Ravishankar Borgaonkar, Kevin Redon, and Jean-Pierre Seifert. Experimental analysis of the femtocell location verification techniques. Lecture Notes in Computer Science, 7127:49–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27937-9_4/.

Koivunen:2012:WWN

[26] Erka Koivunen. “why wasn’t I notified?”: Information security incident reporting demystified. Lecture Notes in Computer Science, 7127: 55–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27937-9_5/.

Chia:2012:URP

[27] Pern Hui Chia, Andreas P. Heiner, and N. Asokan. Use of ratings from personalized communities for trustworthy application installation. Lecture Notes in Computer Science, 7127:71–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27937-9_6/.

Kreitz:2012:PPI

[28] Gunnar Kreitz, Mads Dam, and Douglas Wikström. Practical private information aggregation in large networks. Lecture Notes in Computer Science, 7127:89–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27937-9_7/.

Almgren:2012:TMH

[29] Magnus Almgren and Wolfgang John. Tracking malicious hosts on a 10Gbps backbone link. Lecture Notes in Computer Science, 7127:104– 120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27937-9_8/.

Kani-Zabihi:2012:SUR

[30] Elahe Kani-Zabihi and Lizzie Coles-Kemp. Service users’ requirements for tools to support effective on-line privacy and consent practices. Lecture Notes in Computer Science, 7127:121–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27937-9_9/.

Herrmann:2012:ACH

[31] Dominik Herrmann, Christoph Gerber, Christian Banse, and Hannes Federrath. Analyzing characteristic host access patterns for reidentification of Web user sessions. Lecture Notes in Computer Science, 7127:136–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27937-9_10/.

Crampton:2012:FMS

[32] Jason Crampton and Michael Huth. A framework for the modular specification and orchestration of authorization policies. Lecture Notes in Computer Science, 7127:155–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27937-9_11/.

Kostiainen:2012:CDT

[33] Kari Kostiainen, N. Asokan, and Jan-Erik Ekberg. Credential disabling from trusted execution environments. Lecture Notes in Computer Science, 7127:171–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27937-9_12/.

Gadyatskaya:2012:JCA

[34] Olga Gadyatskaya, Fabio Massacci, Federica Paci, and Sergey Stankevich. Java card architecture for autonomous yet secure evolution of Smart Cards applications. Lecture Notes in Computer Science, 7127:187– 192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27937-9_13/.

DelTedesco:2012:IEP

[35] Filippo Del Tedesco, Alejandro Russo, and David Sands. Implementing erasure policies using Taint analysis. Lecture Notes in Computer Science, 7127:193–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27937-9_14/.

Conti:2012:TMP

[36] Juan José Conti and Alejandro Russo. A Taint mode for Python via a library. Lecture Notes in Computer Science, 7127:210–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27937-9_15/.

DeRyck:2012:SWM

[37] Philippe De Ryck, Maarten Decat, Lieven Desmet, Frank Piessens, and Wouter Joosen. Security of Web mashups: a survey. Lecture Notes in Computer Science, 7127:223–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27937-9_16/.

Magazinius:2012:SWS

[38] Jonas Magazinius, Phu H. Phung, and David Sands. Safe wrappers and sane policies for self protecting JavaScript. Lecture Notes in Computer Science, 7127:239–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27937-9_17/.

Quaresma:2012:PIG

[39] Jose Quaresma and Christian W. Probst. Protocol implementation generator. Lecture Notes in Computer Science, 7127:256–268, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27937-9_18/.

Brumley:2012:SFI

[40] Billy Bob Brumley. Secure and fast implementations of two involution ciphers. Lecture Notes in Computer Science, 7127:269–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27937-9_19/.

Saarinen:2012:PPK

[41] Markku-Juhani O. Saarinen. The PASSERINE public key encryption and authentication mechanism. Lecture Notes in Computer Science, 7127:283–288, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27937-9_20/.

Anonymous:2012:BMa

[42] Anonymous. Back matter. Lecture Notes in Computer Science, 7127: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27937-9/1.

Anonymous:2012:FMb

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27937-9/1.

Fridrich:2012:MTS

[44] Jessica Fridrich. Modern trends in steganography and steganalysis. Lecture Notes in Computer Science, 7128:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-32205-1_1.

Memon:2012:PFT

[45] Nasir Memon. Photo forensics — there is more to a picture than meets the eye. Lecture Notes in Computer Science, 7128:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-32205-1_2.

Sachnev:2012:IME

[46] Vasily Sachnev and Hyoung Joong Kim. An improved matrix encoding scheme for JPEG steganography. Lecture Notes in Computer Science, 7128:3–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32205-1_3/.

Tan:2012:SLM

[47] Shunquan Tan. Steganalysis of LSB matching revisited for consecutive pixels using B-spline functions. Lecture Notes in Computer Science, 7128: 16–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32205-1_4/.

Jiang:2012:DCA

[48] Xinghao Jiang, Tanfeng Sun, Yue Zhou, and Yun Q. Shi. A drift compensation algorithm for H.264/AVC video robust watermarking scheme. Lecture Notes in Computer Science, 7128:30–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32205-1_5/.

Wang:2012:HPM

[49] Junxiang Wang, Jiangqun Ni, and Jinwei Pan. A high performance multilayer reversible data hiding scheme using two-step embedding. Lecture Notes in Computer Science, 7128:42–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32205-1_6/.

Ohzeki:2012:NWM

[50] Kazuo Ohzeki, YuanYu Wei, Yutaka Hirakawa, and Kiyotsugu Sato. A new watermarking method with obfuscated quasi-chirp transform. Lecture Notes in Computer Science, 7128:57–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32205-1_7/.

Wang:2012:NFS

[51] Hui Wang, Anthony T. S. Ho, and Xi Zhao. A novel fast self-restoration semi-fragile watermarking algorithm for image content authentication resistant to JPEG compression. Lecture Notes in Computer Science, 7128: 72–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32205-1_8/.

Lei:2012:RAW

[52] Baiying Lei, Ing Yann Soon, and Zhen Li. A robust audio watermarking scheme based on lifting wavelet transform and singular value decomposition. Lecture Notes in Computer Science, 7128:86–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32205-1_9/.

Koz:2012:ASE

[53] Alper Koz and Claude Delpha. Adaptive selection of embedding locations for spread spectrum watermarking of compressed audio. Lecture Notes in Computer Science, 7128:97–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32205-1_10/.

Yamada:2012:IHM

[54] Takayuki Yamada, Seiichi Gohshi, and Isao Echizen. IR hiding: Method to prevent re-recording screen image built in short wavelength pass filter detection method using specular reflection. Lecture Notes in Computer Science, 7128:111–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32205-1_11/.

Jun:2012:IIR

[55] Jong Yun Jun, Kunho Kim, Jae-Pil Heo, and Sung eui Yoon. IRIW: Image retrieval based image watermarking for large-scale image databases. Lecture Notes in Computer Science, 7128:126–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32205-1_12/.

Chen:2012:SRF

[56] Fan Chen, Hongjie He, Yaoran Huo, and Hongxia Wang. Self-recovery fragile watermarking scheme with variable watermark payload. Lecture Notes in Computer Science, 7128:142–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32205-1_13/.

Tian:2012:SSB

[57] Huawei Tian, Zheng Wang, Yao Zhao, Rongrong Ni, and Lunming Qin. Spread spectrum-based multi-bit watermarking for free-view video. Lecture Notes in Computer Science, 7128:156–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32205-1_14/.

Liu:2012:TNA

[58] Hongmei Liu, Lei Li, Jian Li, and Jiwu Huang. Three novel algorithms for hiding data in PDF files based on incremental updates. Lecture Notes in Computer Science, 7128:167–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32205-1_15/.

Yamada:2012:UEW

[59] Takaaki Yamada and Yoshiyasu Takahashi. Use of “emergable watermarks” as copy indicators for securing video content. Lecture Notes in Computer Science, 7128:181–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32205-1_16/.

Weir:2012:AVC

[60] Jonathan Weir and WeiQi Yan. Authenticating visual cryptography shares using 2D barcodes. Lecture Notes in Computer Science, 7128:196– 210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32205-1_17/.

Liu:2012:FVC

[61] Feng Liu, Teng Guo, ChuanKun Wu, and Ching-Nung Yang. Flexible visual cryptography scheme without distortion. Lecture Notes in Computer Science, 7128:211–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32205-1_18/.

Yamaguchi:2012:EVC

[62] Yasushi Yamaguchi. An extended visual cryptography scheme for continuous-tone images. Lecture Notes in Computer Science, 7128:228– 242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32205-1_19/.

Zhao:2012:CST

[63] Xudong Zhao, Shilin Wang, Shenghong Li, and Jianhua Li. A comprehensive study on third order statistical features for image splicing detection. Lecture Notes in Computer Science, 7128:243–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32205-1_20/.

Anonymous:2012:FMc

[64] Anonymous. Front matter. Lecture Notes in Computer Science, 7128: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32205-1/1.

Bertier:2012:PAC

[65] Marin Bertier, Marko Obrovac, and Cédric Tedeschi. A protocol for the atomic capture of multiple molecules on large scale platforms. Lecture Notes in Computer Science, 7129:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-25959-3_1/.

Bieniusa:2012:LBR

[66] Annette Bieniusa and Thomas Fuhrmann. Lifting the barriers — reducing latencies with transparent transactional memory. Lecture Notes in Computer Science, 7129:16–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25959-3_2/.

Chen:2012:AAR

[67] Jingshu Chen and Sandeep Kulkarni. Application of automated revision for UML models: a case study. Lecture Notes in Computer Science, 7129:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_3/.

Cournier:2012:SSM

[68] Alain Cournier, Swan Dubois, and Anissa Lamani. Snap-stabilizing message forwarding algorithm on tree topologies. Lecture Notes in Computer Science, 7129:46–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_4/.

Crain:2012:TUC

[69] Tyler Crain, Damien Imbs, and Michel Raynal. Towards a universal construction for transaction-based multiprocess programs. Lecture Notes in Computer Science, 7129:61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25959-3_5/.

Delporte-Gallet:2012:BAH

[70] Carole Delporte-Gallet and Hugues Fauconnier. Byzantine agreement with homonyms in synchronous systems. Lecture Notes in Computer Science, 7129:76–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_6/.

Ebnenasir:2012:FDF

[71] Ali Ebnenasir and Reza Hajisheykhi. Facilitating the design of fault tolerance in transaction level SystemC programs. Lecture Notes in Computer Science, 7129:91–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_7/.

Even:2012:CDE

[72] Guy Even, Moti Medina, and Gregor Schaffrath. Competitive and deterministic embeddings of virtual networks. Lecture Notes in Computer Science, 7129:106–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_8/.

Kentros:2012:SMO

[73] Sotirios Kentros and Aggelos Kiayias. Solving the at-most-once problem with nearly optimal effectiveness. Lecture Notes in Computer Science, 7129:122–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_9/.

Mehta:2012:IBI

[74] Abhinav Mehta, Shashank Agrawal, and Kannan Srinathan. Interplay between (im)perfectness, synchrony and connectivity: The case of reliable message transmission. Lecture Notes in Computer Science, 7129:138–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_10/.

Santos:2012:TPH

[75] Nuno Santos and André Schiper. Tuning Paxos for high-throughput with batching and pipelining. Lecture Notes in Computer Science, 7129:153– 167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25959-3_11/.

Abidi:2012:HAE

[76] Amine Abidi, Sonia Mettali Gammar, and Farouk Kamoun. Hybrid approach for experimental networking research. Lecture Notes in Computer Science, 7129:168–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_12/.

Suresh:2012:TOE

[77] Mahima Agumbe Suresh, Radu Stoleru, and Ron Denton. Towards optimal event detection and localization in acyclic flow networks. Lecture Notes in Computer Science, 7129:179–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-25959-3_13/.

Baldoni:2012:VTR

[78] Roberto Baldoni, Silvia Bonomi, and Adriano Cerocchi. Virtual tree: a robust overlay network for ensuring interval valid queries in dynamic distributed systems. Lecture Notes in Computer Science, 7129:197–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25959-3_14/.

Hsu:2012:DCE

[79] Yin-Chung Hsu, Yen-Ting Chen, and Chiu-Kuo Liang. Distributed coverage-enhancing algorithms in directional sensor networks with rotatable sensors. Lecture Notes in Computer Science, 7129:201–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25959-3_15/.

Dash:2012:FQL

[80] Dinesh Dash, Arijit Bishnu, and Arobinda Gupta. Finding the quality of line coverage of a sensor network. Lecture Notes in Computer Science, 7129:214–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_16/.

DeCristofaro:2012:PQP

[81] Emiliano De Cristofaro and Roberto Di Pietro. Preserving query privacy in urban sensing systems. Lecture Notes in Computer Science, 7129:218– 233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-25959-3_17/.

Deshpande:2012:AVB

[82] Sarang Deshpande and Krishna M. Sivalingam. Adaptive velocity based guided navigation in wireless sensor networks. Lecture Notes in Computer Science, 7129:234–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_18/.

Conti:2012:WSR

[83] Mauro Conti, Roberto Di Pietro, and Angelo Spognardi. Wireless sensor replica detection in mobile environments. Lecture Notes in Computer Science, 7129:249–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-25959-3_19/.

Esposito:2012:ART

[84] Christian Esposito, Stefano Russo, and Roberto Beraldi. Achieving reliable and timely event dissemination over WAN. Lecture Notes in Computer Science, 7129:265–280, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-25959-3_20/.

Anonymous:2012:FMd

[85] Anonymous. Front matter. Lecture Notes in Computer Science, 7129: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-25959-3/1.

Lopez:2012:OCI

[86] Javier Lopez, Roberto Setola, and Stephen D. Wolthusen. Overview of critical information infrastructure protection. Lecture Notes in Computer Science, 7130:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28920-0_1/.

Cavelty:2012:ACS

[87] Myriam Dunn Cavelty and Manuel Suter. The art of CIIP strategy: Taking stock of content and processes. Lecture Notes in Computer Science, 7130:15–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28920-0_2/.

Glorioso:2012:ISI

[88] Andrea Glorioso and Andrea Servida. Infrastructure sectors and the information infrastructure. Lecture Notes in Computer Science, 7130:39– 51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28920-0_3/.

Luiijf:2012:UCT

[89] Eric Luiijf. Understanding cyber threats and vulnerabilities. Lecture Notes in Computer Science, 7130:52–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28920-0_4/.

Svendsen:2012:MA

[90] Nils Kalstad Svendsen and Stephen D. Wolthusen. Modelling approaches. Lecture Notes in Computer Science, 7130:68–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28920-0_5/.

Raciti:2012:ADW

[91] Massimiliano Raciti, Jordi Cucurull, and Simin Nadjm-Tehrani. Anomaly detection in water management systems. Lecture Notes in Computer Science, 7130:98–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28920-0_6/.

Alcaraz:2012:SAS

[92] Cristina Alcaraz, Gerardo Fernandez, and Fernando Carvajal. Security aspects of SCADA and DCS environments. Lecture Notes in Computer Science, 7130:120–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28920-0_7/.

Rrushi:2012:SPV

[93] Julian L. Rrushi. SCADA protocol vulnerabilities. Lecture Notes in Computer Science, 7130:150–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28920-0_8/.

Khelil:2012:PSC

[94] Abdelmajid Khelil, Daniel Germanus, and Neeraj Suri. Protection of SCADA communication channels. Lecture Notes in Computer Science, 7130:177–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28920-0_9/.

Bompard:2012:CVP

[95] Ettore Bompard, Paolo Cuccia, Marcelo Masera, and Igor Nai Fovino. Cyber vulnerability in power systems operation and control. Lecture Notes in Computer Science, 7130:197–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28920-0_10/.

Johnsen:2012:SSI

[96] Stig O. Johnsen, Andreas Aas, and Ying Qian. Sector-specific information infrastructure issues in the oil, gas, and petrochemical sector. Lecture Notes in Computer Science, 7130:235–279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28920-0_11/.

Goldman:2012:T

[97] Stuart Goldman and Huseyin Uzunalioglu. Telecommunications. Lecture Notes in Computer Science, 7130:280–300, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28920-0_12/.

Hammerli:2012:FSI

[98] Bernhard Hämmerli. Financial services industry. Lecture Notes in Computer Science, 7130:301–329, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28920-0_13/.

Hartong:2012:T

[99] Mark Hartong, Rajn Goel, and Duminda Wijesekera. Transportation. Lecture Notes in Computer Science, 7130:330–355, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28920-0_14/.

Anonymous:2012:BMb

[100] Anonymous. Back matter. Lecture Notes in Computer Science, 7130: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28920-0/1.

Anonymous:2012:FMe

[101] Anonymous. Front matter. Lecture Notes in Computer Science, 7130: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28920-0/1.

Hanjalic:2012:NGB

[102] Alan Hanjalic. A new gap to bridge: Where to go next in social media retrieval? Lecture Notes in Computer Science, 7131:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-27355-1_1.

Smith:2012:MMD

[103] John R. Smith. Mining multimedia data for meaning. Lecture Notes in Computer Science, 7131:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-27355-1_2.

Steinmetz:2012:CSG

[104] Ralf Steinmetz and Stefan Göbel. Challenges in serious gaming as emerging multimedia technology for education, training, sports and health. Lecture Notes in Computer Science, 7131:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-27355-1_3.

Bannour:2012:BSH

[105] Hichem Bannour and Céline Hudelot. Building semantic hierarchies faithful to image semantics. Lecture Notes in Computer Science, 7131:

4–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27355-1_4/.

Kuettel:2012:CIL

[106] Daniel Kuettel, Matthieu Guillaumin, and Vittorio Ferrari. Combining image-level and segment-level models for automatic annotation. Lecture Notes in Computer Science, 7131:16–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27355-1_5/.

Karaman:2012:MLL

[107] Svebor Karaman, Jenny Benois-Pineau, Rémi Mégret, and Aurélie Bugeau. Multi-layer local graph words for object recognition. Lecture Notes in Computer Science, 7131:29–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27355-1_6/.

Muhling:2012:MVC

[108] Markus Mühling, Ralph Ewerth, Jun Zhou, and Bernd Freisleben. Multimodal video concept detection via bag of auditory words and multiple kernel learning. Lecture Notes in Computer Science, 7131:40–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27355-1_7/.

Ionescu:2012:CBV

[109] Bogdan Ionescu, Klaus Seyerlehner, Christoph Rasche, Constantin Vertan, and Patrick Lambert. Content-based video description for automatic video genre categorization. Lecture Notes in Computer Science, 7131:51– 62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27355-1_8/.

Obrador:2012:TCB

[110] Pere Obrador, Michele A. Saad, Poonam Suryanarayan, and Nuria Oliver. Towards category-based aesthetic models of photographs. Lecture Notes in Computer Science, 7131:63–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27355-1_9/.

Younessian:2012:SSU

[111] Ehsan Younessian and Deepu Rajan. Scene signatures for unconstrained news video stories. Lecture Notes in Computer Science, 7131:77–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27355-1_10/.

Yang:2012:VVO

[112] Zhiguo Yang, Yuxin Peng, and Jianguo Xiao. Visual vocabulary optimization with spatial context for image annotation and classification. Lecture Notes in Computer Science, 7131:89–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27355-1_11/.

Ding:2012:CUQ

[113] Xiaowei Ding, Yi Xu, Lei Deng, and Xiaokang Yang. Colorization using quaternion algebra with automatic scribble generation. Lecture Notes in Computer Science, 7131:103–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27355-1_12/.

Mantiuk:2012:DIY

[114] Radosław Mantiuk, Michał Kowalik, Adam Nowosielski, and Bartosz Bazyluk. Do-it-yourself eye tracker: Low-cost pupil-based eye tracker for computer graphics applications. Lecture Notes in Computer Science, 7131:115–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27355-1_13/.

Zhang:2012:SBB

[115] Longfei Zhang, Yue Gao, Alexander Hauptmann, Rongrong Ji, Gangyi Ding, and Boaz Super. Symbiotic black-box tracker. Lecture Notes in Computer Science, 7131:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27355-1_14/.

Walber:2012:IOI

[116] Tina Walber, Ansgar Scherp, and Steffen Staab. Identifying objects in images from analyzing the users’ gaze movements for provided tags. Lecture Notes in Computer Science, 7131:138–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27355-1_15/.

Zhao:2012:VRS

[117] Xiaojian Zhao, Jin Yuan, Richang Hong, Meng Wang, Zhoujun Li, and Tat-Seng Chua. On video recommendation over social network. Lecture Notes in Computer Science, 7131:149–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27355-1_16/.

Yoshida:2012:IIR

[118] Taiga Yoshida, Go Irie, Takashi Satou, Akira Kojima, and Suguru Higashino. Improving item recommendation based on social tag ranking. Lecture Notes in Computer Science, 7131:161–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27355-1_17/.

Lan:2012:DFM

[119] Zhen zhong Lan, Lei Bao, Shoou-I Yu, Wei Liu, and Alexander G. Hauptmann. Double fusion for multimedia event detection. Lecture Notes in Computer Science, 7131:173–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27355-1_18/.

Younessian:2012:MMS

[120] Ehsan Younessian and Deepu Rajan. Multi-modal solution for unconstrained news story retrieval. Lecture Notes in Computer Science, 7131:186–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27355-1_19/.

Behera:2012:WAM

[121] Ardhendu Behera, Anthony G. Cohn, and David C. Hogg. Workflow activity monitoring using dynamics of pair-wise qualitative spatial relations. Lecture Notes in Computer Science, 7131:196–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27355-1_20/.

Anonymous:2012:FMf

[122] Anonymous. Front matter. Lecture Notes in Computer Science, 7131: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27355-1/1.

Li:2012:PAF

[123] Hengfei Li, Nir Oren, and Timothy J. Norman. Probabilistic argumentation frameworks. Lecture Notes in Computer Science, 7132:1– 16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29184-5_1/.

Baumann:2012:SAF

[124] Ringo Baumann, Gerhard Brewka, and Renata Wong. Splitting argumentation frameworks: An empirical evaluation. Lecture Notes in Computer Science, 7132:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29184-5_2/.

Dvorak:2012:CCJ

[125] Wolfgang Dvořák. On the complexity of computing the justification status of an argument. Lecture Notes in Computer Science, 7132:32– 49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29184-5_3/.

Medellin-Gasque:2012:ACO

[126] Rolando Medellin-Gasque, Katie Atkinson, Peter McBurney, and Trevor Bench-Capon. Arguments over co-operative plans. Lecture Notes in Computer Science, 7132:50–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29184-5_4/.

Riley:2012:IDS

[127] Luke Riley, Katie Atkinson, Terry Payne, and Elizabeth Black. An implemented dialogue system for inquiry and persuasion. Lecture Notes in Computer Science, 7132:67–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29184-5_5/.

Visser:2012:AFQ

[128] Wietske Visser, Koen V. Hindriks, and Catholijn M. Jonker. An argumentation framework for qualitative multi-criteria preferences. Lecture Notes in Computer Science, 7132:85–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29184-5_6/.

Bistarelli:2012:MSA

[129] Stefano Bistarelli and Francesco Santini. Modeling and solving AFs with a constraint-based tool: ConArg. Lecture Notes in Computer Science, 7132:99–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29184-5_7/.

Rotstein:2012:RBA

[130] Nicolás D. Rotstein, Nir Oren, and Timothy J. Norman. Resource boundedness and argumentation. Lecture Notes in Computer Science, 7132:117–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29184-5_8/.

Black:2012:ESD

[131] Elizabeth Black and Katie Bentley. An empirical study of a deliberation dialogue system. Lecture Notes in Computer Science, 7132:132– 146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29184-5_9/.

Krumpelmann:2012:SRD

[132] Patrick Krümpelmann, Matthias Thimm, Marcelo A. Falappa, and Alejandro J. García. Selective revision by deductive argumentation. Lecture Notes in Computer Science, 7132:147–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29184-5_10/.

Maio:2012:TLA

[133] Paulo Maio and Nuno Silva. A three-layer argumentation framework. Lecture Notes in Computer Science, 7132:163–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29184-5_11/.

Cobo:2012:SET

[134] Maria Laura Cobo, Diego C. Martinez, and Guillermo R. Simari. Stable extensions in timed argumentation frameworks. Lecture Notes in Computer Science, 7132:181–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29184-5_12/.

Baroni:2012:CIA

[135] Pietro Baroni, Federico Cerutti, Paul E. Dunne, and Massimiliano Giacomin. Computing with infinite argumentation frameworks: The case of AFRAs. Lecture Notes in Computer Science, 7132:197–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29184-5_13/.

Rienstra:2012:MSA

[136] Tjitze Rienstra, Alan Perotti, Serena Villata, Dov M. Gabbay, and Leendert van der Torre. Multi-sorted argumentation. Lecture Notes in Computer Science, 7132:215–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29184-5_14/.

Boella:2012:CLA

[137] Guido Boella, Dov M. Gabbay, Alan Perotti, Leendert van der Torre, and Serena Villata. Conditional labelling for abstract argumentation. Lecture Notes in Computer Science, 7132:232–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29184-5_15/.

Toni:2012:BA

[138] Francesca Toni and Paolo Torroni. Bottom-up argumentation. Lecture Notes in Computer Science, 7132:249–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29184-5_16/.

Fan:2012:FST

[139] Xiuyi Fan and Francesca Toni. A first step towards argumentation dialogues for discovery. Lecture Notes in Computer Science, 7132:263– 279, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29184-5_17/.

Anonymous:2012:BMc

[140] Anonymous. Back matter. Lecture Notes in Computer Science, 7132: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29184-5/1.

Anonymous:2012:FMg

[141] Anonymous. Front matter. Lecture Notes in Computer Science, 7132: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29184-5/1.

Kaagstrom:2012:AED

[142] Bo Kågström, Daniel Kressner, and Meiyue Shao. On aggressive early deflation in parallel variants of the QR algorithm. Lecture Notes in Computer Science, 7133:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28151-8_1/.

Mosegaard:2012:LNI

[143] Klaus Mosegaard. Limits to nonlinear inversion. Lecture Notes in Computer Science, 7133:11–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28151-8_2/.

Gustavson:2012:CB

[144] Fred G. Gustavson. Cache blocking. Lecture Notes in Computer Science, 7133:22–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28151-8_3/.

Vazquez-Poletti:2012:MEO

[145] Jose Luis Vázquez-Poletti, Gonzalo Barderas, Ignacio M. Llorente, and Pilar Romero. A model for efficient onboard actualization of an instrumental cyclogram for the Mars MetNet Mission on a public cloud infrastructure. Lecture Notes in Computer Science, 7133:33–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28151-8_4/.

Contassot-Vivier:2012:IAG

[146] Sylvain Contassot-Vivier, Thomas Jost, and Stéphane Vialle. Impact of asynchronism on GPU accelerated parallel iterative computations. Lecture Notes in Computer Science, 7133:43–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28151-8_5/.

Kostin:2012:SSW

[147] Victor Kostin, Vadim Lisitsa, Galina Reshetova, and Vladimir Tcheverda. Simulation of seismic waves propagation in multiscale media: Impact of cavernous/fractured reservoirs. Lecture Notes in Computer Science, 7133:54–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28151-8_6/.

Adelmann:2012:IFP

[148] Andreas Adelmann, Peter Arbenz, and Yves Ineichen. Improvements of a fast parallel Poisson solver on irregular domains. Lecture Notes in Computer Science, 7133:65–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28151-8_7/.

Laskowski:2012:DJP

[149] Eryk Laskowski, Marek Tudruj, Ivanoe De Falco, Umberto Scafuri, and Ernesto Tarantino. Distributed Java programs initial mapping based on extremal optimization. Lecture Notes in Computer Science, 7133:75– 85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28151-8_8/.

Niewiadomska-Szynkiewicz:2012:SEP

[150] Ewa Niewiadomska-Szynkiewicz and Michal Marks. Software environment for parallel optimization of complex systems. Lecture Notes in Computer Science, 7133:86–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28151-8_9/.

Agnarsson:2012:PPM

[151] Snorri Agnarsson. Parallel programming in morpho. Lecture Notes in Computer Science, 7133:97–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28151-8_10/.

Skovhede:2012:EDS

[152] Kenneth Skovhede, Morten N. Larsen, and Brian Vinter. Extending distributed shared memory for the Cell Broadband Engine to a channel model. Lecture Notes in Computer Science, 7133:108–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28151-8_11/.

Borkowski:2012:GAP

[153] Janusz Borkowski, Marek Tudruj, Adam Smyk, and Damian Kopanski. Global asynchronous parallel program control for multicore processors. Lecture Notes in Computer Science, 7133:119–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28151-8_12/.

Lieber:2012:HSD

[154] Matthias Lieber, Verena Grützun, Ralf Wolke, Matthias S. Müller, and Wolfgang E. Nagel. Highly scalable dynamic load balancing in the atmospheric modeling system COSMO-SPECS+FD4. Lecture Notes in Computer Science, 7133:131–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28151-8_13/.

Fjukstad:2012:IWS

[155] Bård Fjukstad, Tor-Magne Stien Hagen, Daniel Stødle, and Phuong Hoai Ha. Interactive weather simulation and visualization on a display wall with many-core compute nodes. Lecture Notes in Computer Science, 7133:142–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28151-8_14/.

Petschow:2012:AMR

[156] Matthias Petschow and Paolo Bientinesi. The algorithm of multiple relatively robust representations for multi-core processors. Lecture Notes in Computer Science, 7133:152–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28151-8_15/.

Aliaga:2012:PMI

[157] José I. Aliaga, Matthias Bollhöfer, Alberto F. Martín, and Enrique S. Quintana-Ortí. Parallelization of multilevel ILU preconditioners on distributed-memory multiprocessors. Lecture Notes in Computer Science, 7133:162–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28151-8_16/.

Cecilia:2012:CSC

[158] José María Cecilia, José Manuel García, and Manuel Ujaldón. CUDA 2D stencil computations for the Jacobi method. Lecture Notes in Computer Science, 7133:173–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28151-8_17/.

Smyk:2012:SMC

[159] Adam Smyk and Marek Tudruj. Streaming model computation of the FDTD problem. Lecture Notes in Computer Science, 7133:184–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28151-8_18/.

Matsekh:2012:NAS

[160] Anna Matsekh, Alexei Skurikhin, Lakshman Prasad, and Edward Rosten. Numerical aspects of spectral segmentation on polygonal grids. Lecture Notes in Computer Science, 7133:193–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28151-8_19/.

Strzelczyk:2012:PKA

[161] Jacek Strzelczyk and Stanislawa Porzycka. Parallel kriging algorithm for unevenly spaced data. Lecture Notes in Computer Science, 7133:204– 212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28151-8_20/.

Anonymous:2012:FMh

[162] Anonymous. Front matter. Lecture Notes in Computer Science, 7133: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28151-8/1.

Czapla:2012:FEM

[163] Luke Czapla, Alexey Siretskiy, John Grime, and Malek O. Khan. Free energy Monte Carlo simulations on a distributed network. Lecture Notes in Computer Science, 7134:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28145-7_1/.

Plattner:2012:NIC

[164] Nuria Plattner, Sharif Kunikeev, David L. Freeman, and Jimmie D. Doll. Numerical investigation of the cumulant expansion for Fourier path integrals. Lecture Notes in Computer Science, 7134:13–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28145-7_2/.

Klupfel:2012:OFO

[165] Peter Klüpfel, Simon Klüpfel, Kiril Tsemekhman, and Hannes Jónsson. Optimization of functionals of orthonormal functions in the absence of unitary invariance. Lecture Notes in Computer Science, 7134:23–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28145-7_3/.

Pedersen:2012:SAC

[166] Andreas Pedersen, Jean-Claude Berthet, and Hannes Jónsson. Simulated annealing with coarse graining and distributed computing. Lecture Notes in Computer Science, 7134:34–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28145-7_4/.

Einarsdottir:2012:POA

[167] Dóróthea M. Einarsdóttir, Andri Arnaldsson, Finnbogi Óskarsson, and Hannes Jónsson. Path optimization with application to tunneling. Lecture Notes in Computer Science, 7134:45–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28145-7_5/.

Saetra:2012:SWS

[168] Martin Lilleeng Sætra and André Rigland Brodtkorb. Shallow water simulations on multiple GPUs. Lecture Notes in Computer Science, 7134: 56–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28145-7_6/.

Widener:2012:HPC

[169] Patrick M. Widener, Tahsin Kurc, Wenjin Chen, Fusheng Wang, Lin Yang, and Jun Hu. High performance computing techniques for scaling image analysis workflows. Lecture Notes in Computer Science, 7134:67– 77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28145-7_7/.

Stussak:2012:PCB

[170] Christian Stussak and Peter Schenzel. Parallel computation of bivariate polynomial resultants on graphics processing units. Lecture Notes in Computer Science, 7134:78–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28145-7_8/.

Benner:2012:AMR

[171] Peter Benner, Pablo Ezzatti, Daniel Kressner, Enrique S. Quintana-Ortí, and Alfredo Remón. Accelerating model reduction of large linear systems with graphics processors. Lecture Notes in Computer Science, 7134:88– 97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28145-7_9/.

Krog:2012:FGB

[172] Øystein E. Krog and Anne C. Elster. Fast GPU-based fluid simulations using SPH. Lecture Notes in Computer Science, 7134:98–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28145-7_10/.

Davidson:2012:TTA

[173] Andrew Davidson and John Owens. Toward techniques for auto-tuning GPU algorithms. Lecture Notes in Computer Science, 7134:110–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28145-7_11/.

Marciniak:2012:IVC

[174] Andrzej Marciniak. An interval version of the Crank–Nicolson method — the first approach. Lecture Notes in Computer Science, 7134:120– 126, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28145-7_12/.

Nehmeier:2012:PDI

[175] Marco Nehmeier, Stefan Siegel, and Jürgen Wolff von Gudenberg. Parallel detection of interval overlapping. Lecture Notes in Computer Science, 7134:127–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28145-7_13/.

Kubica:2012:USO

[176] Bartłomiej Jacek Kubica and Adam Woźniak. Using the second-order information in Pareto-set computations of a multi-criteria problem. Lecture Notes in Computer Science, 7134:137–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28145-7_14/.

Bohlender:2012:CFE

[177] Gerd Bohlender and Ulrich Kulisch. Comments on fast and exact accumulation of products. Lecture Notes in Computer Science, 7134:148– 156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28145-7_15/.

Jankowska:2012:IFD

[178] Małgorzata A. Jankowska. An interval finite difference method of Crank– Nicolson type for solving the one-dimensional heat conduction equation with mixed boundary conditions. Lecture Notes in Computer Science, 7134:157–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28145-7_16/.

Kramer:2012:UCX

[179] Walter Krämer, Michael Zimmer, and Werner Hofschuster. Using C-XSC for high performance verified computing. Lecture Notes in Computer Science, 7134:168–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28145-7_17/.

Diep:2012:EII

[180] Nguyen Hong Diep. Efficient implementation of interval matrix multiplication. Lecture Notes in Computer Science, 7134:179–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28145-7_18/.

Britsch:2012:CFP

[181] Markward Britsch. The computing framework for physics analysis at LHCb. Lecture Notes in Computer Science, 7134:189–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28145-7_19/.

Neukirchen:2012:TRT

[182] Helmut Neukirchen. Taming the raven — testing the random access, visualization and exploration network RAVEN. Lecture Notes in Computer Science, 7134:196–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28145-7_20/.

Anonymous:2012:FMi

[183] Anonymous. Front matter. Lecture Notes in Computer Science, 7134: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28145-7/1.

Parakh:2012:NSW

[184] Abhishek Parakh and Subhash Kak. A new small world lattice. Lecture Notes in Computer Science, 7135:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29280-4_1/.

M:2012:SDD

[185] Siddesh G. M. and Srinivasa K. G. SLA- driven dynamic resource allocation on clouds. Lecture Notes in Computer Science, 7135:9–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29280-4_2/.

Balasubramaniyan:2012:ICS

[186] Jagadeesh Balasubramaniyan and Sumalatha Ramachandran. An intelligent cloud system adopting file pre-fetching. Lecture Notes in Computer Science, 7135:19–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29280-4_3/.

Bazydlo:2012:RIR

[187] Marcin Bazydło, Szymon Francuzik, Cezary Sobaniec, and Dariusz Wawrzyniak. Replication infrastructure for RESTful Web services. Lecture Notes in Computer Science, 7135:28–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29280-4_4/.

Yang:2012:CAA

[188] Joon-Mo Yang, Jun-Yong Park, and Ryum-Duck Oh. Context awareness acquisition for safety sensor data processing on industrial sensor network. Lecture Notes in Computer Science, 7135:38–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29280-4_5/.

Paulraj:2012:CBS

[189] D. Paulraj and S. Swamynathan. Content based service discovery in semantic Web services using WordNet. Lecture Notes in Computer Science, 7135:48–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29280-4_6/.

DMello:2012:OSW

[190] Demian Antony D’Mello, Manasa Shetty, Akshatha Prabhu, Akshaya, and Adithi Shenoy. Optimal selection of Web services for efficient QoS driven Web service compositions. Lecture Notes in Computer Science, 7135:57–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29280-4_7/.

Mallayya:2012:RWS

[191] Deivamani Mallayya and Baskaran Ramachandran. Reconfigurable Web service composition using belief revision. Lecture Notes in Computer Science, 7135:67–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29280-4_8/.

PonHarshavardhanan:2012:SWS

[192] PonHarshavardhanan, J. Akilandeswari, and K. A. Anjali Krishna. Semantic Web service discovery with structural level matching of operations. Lecture Notes in Computer Science, 7135:77–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29280-4_9/.

Narang:2012:NWT

[193] Aditya Narang, Sagar Prakash Joglekar, and Karthikeyan Balaji Dhanapal. A novel way of tracking people in an indoor area. Lecture Notes in Computer Science, 7135:85–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29280-4_10/.

Bera:2012:IRE

[194] Tushar Kanti Bera, Samir Kumar Biswas, K. Rajan, and J. Nagaraju. Image reconstruction in electrical impedance tomography (EIT) with projection error propagation-based regularization (PEPR): a practical phantom study. Lecture Notes in Computer Science, 7135:95–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29280-4_11/.

Rekha:2012:IHT

[195] J. Rekha, J. Bhattacharya, and S. Majumder. Improved hand tracking and isolation from face by ICondensation multi clue algorithm for continuous Indian sign language recognition. Lecture Notes in Computer Science, 7135:106–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29280-4_12/.

D:2012:ACE

[196] Arpitha M. D., Megha P. Arakeri, and G. Ram Mohan Reddy. An approach for color edge detection with automatic threshold detection. Lecture Notes in Computer Science, 7135:117–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29280-4_13/.

V:2012:DSL

[197] Anand T. V., Madhu S. Nair, and Rao Tatavarti. Detection of straight lines using rule directed pixel comparison (RDPC) method. Lecture Notes in Computer Science, 7135:125–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29280-4_14/.

Cyriac:2012:PAP

[198] Praveen Cyriac and Philomina Simon. Performance analysis of pedestrian detection at night time with different classifiers. Lecture Notes in Computer Science, 7135:135–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29280-4_15/.

Joseph:2012:EQP

[199] A. Benjamin Joseph and Baskaran Ramachandran. Enhanced quality preserved image compression technique using edge assisted wavelet based interpolation. Lecture Notes in Computer Science, 7135:146–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29280-4_16/.

Reddy:2012:MFN

[200] T. Ashwini Reddy, K. Renuka Devi, and Suryakanth V. Gangashetty. Multilayer feedforward neural network models for pattern recognition tasks in earthquake engineering. Lecture Notes in Computer Science, 7135:154–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29280-4_17/.

Bhaladhare:2012:SAB

[201] Pawan Bhaladhare and Devesh Jinwala. A sensitive attribute based clustering method for k-anonymization. Lecture Notes in Computer Science, 7135:163–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29280-4_18/.

Urolagin:2012:GCA

[202] Siddhaling Urolagin, Prema K. V., and N. V. Subba Reddy. Generalization capability of artificial neural network incorporated with pruning method. Lecture Notes in Computer Science, 7135:171–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29280-4_19/.

Subramani:2012:PMD

[203] Sudha Subramani and Sathiyabhama Balasubramaniam. Post mining of diversified multiple decision trees for actionable knowledge discovery. Lecture Notes in Computer Science, 7135:179–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29280-4_20/.

Anonymous:2012:FMj

[204] Anonymous. Front matter. Lecture Notes in Computer Science, 7135: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29280-4/1.

Kitowski:2012:PGF

[205] Jacek Kitowski and Michał Turała. PL-Grid: Foundations and perspectives of national computing infrastructure. Lecture Notes in Computer Science, 7136:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_1/.

Radecki:2012:IVG

[206] Marcin Radecki and Tadeusz Szymocha. Integrating various grid middleware components and user services into a single platform. Lecture Notes in Computer Science, 7136:15–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28267-6_2/.

Radecki:2012:TPS

[207] Marcin Radecki and Tomasz Szepieniec. Towards professional service operations in grids. Lecture Notes in Computer Science, 7136:27–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28267-6_3/.

Bosak:2012:NCQ

[208] Bartosz Bosak, Jacek Komasa, and Piotr Kopta. New capabilities in QosCosGrid middleware for advanced job management, advance reservation and co-allocation of computing resources — quantum chemistry application use case. Lecture Notes in Computer Science, 7136:40–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28267-6_4/.

Benedyczak:2012:SAP

[209] Krzysztof Benedyczak and Marcin Stolarek. Seamless access to the PL-Grid e-infrastructure using UNICORE middleware. Lecture Notes in Computer Science, 7136:56–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28267-6_5/.

Jarzab:2012:UOP

[210] Marcin Jarzab and Jacek Kosiński. User-oriented provisioning of secure virtualized infrastructure. Lecture Notes in Computer Science, 7136:73– 88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28267-6_6/.

Meizner:2012:FEM

[211] Jan Meizner, Eryk Ciepiela, and Piotr Nowakowski. Flexible and extendable mechanism enabling secure access to e-infrastructures and storage of confidential data for the GridSpace2 Virtual Laboratory. Lecture Notes in Computer Science, 7136:89–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28267-6_7/.

Frankowski:2012:SSI

[212] Gerard Frankowski. SARA — system for inventory and static security control in a grid infrastructure. Lecture Notes in Computer Science, 7136:102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_8/.

Balcerek:2012:ANN

[213] Bartłomiej Balcerek and Bartosz Szurgot. ACARM-ng: Next generation correlation framework. Lecture Notes in Computer Science, 7136:114– 127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28267-6_9/.

Balcerek:2012:SBP

[214] Bartłomiej Balcerek and Gerard Frankowski. Security best practices: Applying defense-in-depth strategy to protect the NGI PL. Lecture Notes in Computer Science, 7136:128–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28267-6_10/.

Funika:2012:ASM

[215] Włodzimierz Funika and Filip Szura. Automation of system monitoring based on fuzzy logic or rules; comparison of two designed approaches with regard to computational infrastructures. Lecture Notes in Computer Science, 7136:142–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_11/.

Slota:2012:TSQ

[216] Renata Słota and Dariusz Król. A toolkit for storage QoS provisioning for data-intensive applications. Lecture Notes in Computer Science, 7136:157–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_12/.

Szepieniec:2012:ISLa

[217] Tomasz Szepieniec and Małgorzata Tomanek. Implementation of service level management in PL-Grid infrastructure. Lecture Notes in Computer Science, 7136:171–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_13/.

Bosak:2012:HIE

[218] Bartosz Bosak and Jan Konczak. Highly integrated environment for parallel application development using QosCosGrid middleware. Lecture Notes in Computer Science, 7136:182–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28267-6_14/.

Palak:2012:UFF

[219] Bartek Palak and Paweł Wolniewicz. User-friendly frameworks for accessing computational resources. Lecture Notes in Computer Science, 7136:191–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_15/.

Dziubecki:2012:OWB

[220] Piotr Dziubecki and Piotr Grabowski. Online Web-based science gateway for Nanotechnology research. Lecture Notes in Computer Science, 7136:205–216, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_16/.

Harezlak:2012:SLE

[221] Daniel Hareżlak and Marek Kasztelnik. Scripting language extensions offered by the GridSpace Experiment Platform. Lecture Notes in Computer Science, 7136:217–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_17/.

Ciepiela:2012:MEL

[222] Eryk Ciepiela and Piotr Nowakowski. Managing entire lifecycles of e-science applications in the GridSpace2 Virtual Laboratory — from motivation through idea to operable Web-accessible environment built on top of PL-Grid e-infrastructure. Lecture Notes in Computer Science, 7136:228–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28267-6_18/.

Ciepiela:2012:GVL

[223] Eryk Ciepiela and Leszek Zaraska. GridSpace2 Virtual Laboratory case study: Implementation of algorithms for quantitative analysis of grain morphology in self-assembled hexagonal lattices according to the Hillebrand method. Lecture Notes in Computer Science, 7136:240–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28267-6_19/.

Jadczyk:2012:EPF

[224] Tomasz Jadczyk and Maciej Malawski. Examining protein folding process simulation and searching for common structure motifs in a protein family as experiments in the GridSpace2 Virtual Laboratory. Lecture Notes in Computer Science, 7136:252–264, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28267-6_20/.

Anonymous:2012:FMk

[225] Anonymous. Front matter. Lecture Notes in Computer Science, 7136: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28267-6/1.

Bauer:2012:BMI

[226] Sebastian S. Bauer, Rolf Hennicker, and Martin Wirsing. Building a modal interface theory for concurrency and data. Lecture Notes in Computer Science, 7137:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28412-0_1/.

Ehrich:2012:MAS

[227] Hans-Dieter Ehrich. My ADT shrine. Lecture Notes in Computer Science, 7137:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28412-0_2/.

Kofron:2012:ESQ

[228] Jan Kofroň and František Plášil. Evolving SOA in the Q-ImPrESS project. Lecture Notes in Computer Science, 7137:25–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/accesspage/chapter/10.1007/978-3-642-28412-
0_3.

Baldan:2012:SGR

[229] Paolo Baldan and Clara Bertolissi. Sharing in the graph rewriting calculus. Lecture Notes in Computer Science, 7137:27–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28412-0_4/.

Bruni:2012:NSD

[230] Roberto Bruni, Anne Kersten, Ivan Lanese, and Giorgio Spagnolo. A new strategy for distributed compensations with interruption in longrunning transactions. Lecture Notes in Computer Science, 7137:42–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28412-0_5/.

Castro:2012:TFO

[231] Pablo F. Castro and Tom S. E. Maibaum. Towards a first-order Deontic action logic. Lecture Notes in Computer Science, 7137:61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28412-0_6/.

Choppy:2012:CMM

[232] Christine Choppy and Gianna Reggio. Casl-Mdl, modelling dynamic systems with a formal foundation and a UML-like notation. Lecture Notes in Computer Science, 7137:76–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28412-0_7/.

Codescu:2012:LEC

[233] Mihai Codescu. Lambda expressions in Casl architectural specifications. Lecture Notes in Computer Science, 7137:98–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28412-0_8/.

Codescu:2012:PTI

[234] Mihai Codescu, Fulya Horozal, Michael Kohlhase, Till Mossakowski, and Florian Rabe. A proof theoretic interpretation of model theoretic hiding. Lecture Notes in Computer Science, 7137:118–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28412-0_9/.

Codescu:2012:TLF

[235] Mihai Codescu, Fulya Horozal, Michael Kohlhase, Till Mossakowski, and Florian Rabe. Towards logical frameworks in the heterogeneous tool set hets. Lecture Notes in Computer Science, 7137:139–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28412-0_10/.

Corradini:2012:IGT

[236] Andrea Corradini, Fabio Gadducci, and Leila Ribeiro. An institution for graph transformation. Lecture Notes in Computer Science, 7137:160– 174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28412-0_11/.

Bourke:2012:NRT

[237] Timothy Bourke, Alexandre David, Kim G. Larsen, Axel Legay, Didier Lime, and Ulrik Nyman. New results on timed specifications. Lecture Notes in Computer Science, 7137:175–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28412-0_12/.

Kreowski:2012:CGT

[238] Hans-Jörg Kreowski, Sabine Kuske, and Caroline von Totth. Combining graph transformation and algebraic specification into model transformation. Lecture Notes in Computer Science, 7137:193–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28412-0_13/.

Maldonado:2012:TBS

[239] Ana Paula Maldonado, Luís Monteiro, and Markus Roggenbach. Towards bialgebraic semantics for the linear time — branching time spectrum. Lecture Notes in Computer Science, 7137:209–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28412-0_14/.

Marczynski:2012:ASE

[240] Grzegorz Marczyński. Algebraic signatures enriched by dependency structure. Lecture Notes in Computer Science, 7137:226–250, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-28412-0_
15/.

OReilly:2012:CMR

[241] Liam O’Reilly, Till Mossakowski, and Markus Roggenbach. Compositional modelling and reasoning in an institution for processes and data. Lecture Notes in Computer Science, 7137:251–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28412-0_16/.

Seidel:2012:PPA

[242] Daniel Seidel and Janis Voigtländer. Proving properties about functions on lists involving element tests. Lecture Notes in Computer Science, 7137:270–286, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28412-0_17/.

Riesco:2012:TCG

[243] Adrián Riesco. Test-case generation for Maude functional modules. Lecture Notes in Computer Science, 7137:287–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28412-0_18/.

Anonymous:2012:BMd

[244] Anonymous. Back matter. Lecture Notes in Computer Science, 7137: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28412-0/1.

Anonymous:2012:FMl

[245] Anonymous. Front matter. Lecture Notes in Computer Science, 7137: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28412-0/1.

Paramythis:2012:AST

[246] Alexandros Paramythis, Lydia Lau, Stavros Demetriadis, and Manolis Tzagarakis. Adaptive support for team collaboration. Lecture Notes in Computer Science, 7138:1–4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28509-7_1/.

Ardissono:2012:VMS

[247] Liliana Ardissono, Gianni Bosio, and Marino Segnan. A visualization model supporting an efficient context resumption in collaboration environments. Lecture Notes in Computer Science, 7138:5–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28509-7_2/.

Dragon:2012:SCL

[248] Toby Dragon, Bruce M. McLaren, Manolis Mavrikis, and Eirini Geraniou. Scaffolding collaborative learning opportunities: Integrating microworld use and argumentation. Lecture Notes in Computer Science, 7138:18–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28509-7_3/.

Abel:2012:AUM

[249] Fabian Abel, Vania Dimitrova, Eelco Herder, and Geert-Jan Houben. Augmenting user models with real world experiences to enhance personalization and adaptation. Lecture Notes in Computer Science, 7138: 31–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28509-7_4/.

Garzon:2012:PAC

[250] Sandro Rodriguez Garzon and Mark Poguntke. The personal adaptive in-car HMI: Integration of external applications for personalized use. Lecture Notes in Computer Science, 7138:35–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28509-7_5/.

Moore:2012:CAA

[251] Adam Moore, Victoria Macarthur, and Owen Conlan. Core aspects of affective metacognitive user models. Lecture Notes in Computer Science, 7138:47–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28509-7_6/.

Tiroshi:2012:RSS

[252] Amit Tiroshi, Tsvi Kuflik, Judy Kay, and Bob Kummerfeld. Recommender systems and the social Web. Lecture Notes in Computer Science, 7138:60–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28509-7_7/.

Ammari:2012:IRY

[253] Ahmad Ammari, Vania Dimitrova, and Dimoklis Despotakis. Identifying relevant YouTube comments to derive socially augmented user models: a semantically enriched machine learning approach. Lecture Notes in Computer Science, 7138:71–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28509-7_8/.

Ricci:2012:DMR

[254] Francesco Ricci, Giovanni Semeraro, Marco de Gemmis, and Pasquale Lops. Decision making and recommendation acceptance issues in recommender systems. Lecture Notes in Computer Science, 7138:86–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28509-7_9/.

Bader:2012:DEI

[255] Roland Bader, Wolfgang Woerndl, Andreas Karitnig, and Gerhard Leitner. Designing an explanation interface for proactive recommendations in automotive scenarios. Lecture Notes in Computer Science, 7138:92– 104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28509-7_10/.

Felfernig:2012:GDS

[256] Alexander Felfernig, Christoph Zehentner, Gerald Ninaus, Harald Grabner, and Walid Maalej. Group decision support for requirements negotiation. Lecture Notes in Computer Science, 7138:105–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28509-7_11/.

Santos:2012:PAL

[257] Olga C. Santos, Milos Kravcik, and Diana Pérez-Marín. Personalization approaches in learning environments. Lecture Notes in Computer Science, 7138:117–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28509-7_12/.

Redondo-Hernandez:2012:PAA

[258] Alberto Redondo-Hernández and Diana Pérez-Marín. A procedure to automatically adapt questions in student–pedagogic conversational agent dialogues. Lecture Notes in Computer Science, 7138:122–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28509-7_13/.

Leite:2012:MES

[259] Iolanda Leite, André Pereira, Ginevra Castellano, Samuel Mascarenhas, and Carlos Martinho. Modelling empathy in social robotic companions. Lecture Notes in Computer Science, 7138:135–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28509-7_14/.

Muir:2012:USA

[260] Mary Muir and Cristina Conati. Understanding student attention to adaptive hints with eye-tracking. Lecture Notes in Computer Science, 7138:148–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28509-7_15/.

Berthold:2012:PPM

[261] Marcel Berthold, Pablo Lachmann, Alexander Nussbaumer, and Sergei Pachtchenko. Psycho-pedagogical mash-up design for personalising the learning environment. Lecture Notes in Computer Science, 7138:161– 175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28509-7_16/.

Cena:2012:SAS

[262] Federica Cena, Antonina Dattolo, Ernesto William De Luca, Pasquale Lops, and Till Plumbaum. Semantic adaptive social Web. Lecture Notes in Computer Science, 7138:176–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28509-7_17/.

Fernandez-Tobias:2012:SDC

[263] Ignacio Fernández-Tobías, Iván Cantador, and Alejandro Bellogín. Semantic disambiguation and contextualisation of social tags. Lecture Notes in Computer Science, 7138:181–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28509-7_18/.

Dattolo:2012:FMA

[264] Antonina Dattolo and Emanuela Pitassi. Folkview: a multi-agent system approach to modeling folksonomies. Lecture Notes in Computer Science, 7138:198–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28509-7_19/.

Tandukar:2012:SPS

[265] Udeep Tandukar and Julita Vassileva. Selective propagation of social data in decentralized online social network. Lecture Notes in Computer Science, 7138:213–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28509-7_20/.

Anonymous:2012:FMm

[266] Anonymous. Front matter. Lecture Notes in Computer Science, 7138: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28509-7/1.

Kasar:2012:MSM

[267] Thotreingam Kasar and Angarai G. Ramakrishnan. Multi-script and multi-oriented text localization from scene images. Lecture Notes in Computer Science, 7139:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29364-1_1/.

Yi:2012:ATR

[268] Chucai Yi and Yingli Tian. Assistive text reading from complex background for blind persons. Lecture Notes in Computer Science, 7139:15– 28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29364-1_2/.

Merino-Gracia:2012:HMD

[269] Carlos Merino-Gracia, Karel Lenc, and Majid Mirmehdi. A headmounted device for recognizing text in natural scenes. Lecture Notes in Computer Science, 7139:29–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29364-1_3/.

Chowdhury:2012:TDT

[270] Aruni Roy Chowdhury, Ujjwal Bhattacharya, and Swapan K. Parui. Text detection of two major Indian scripts in natural scene images. Lecture Notes in Computer Science, 7139:42–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29364-1_4/.

Zeng:2012:ACB

[271] Chao Zeng, Wenjing Jia, and Xiangjian He. An algorithm for colourbased natural scene text segmentation. Lecture Notes in Computer Science, 7139:58–68, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29364-1_5/.

Zhu:2012:RNS

[272] Yuanping Zhu, Jun Sun, and Satoshi Naoi. Recognizing natural scene characters by convolutional neural network and bimodal image enhancement. Lecture Notes in Computer Science, 7139:69–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29364-1_6/.

Liu:2012:P

[273] Qiong Liu and Chunyuan Liao. PaperUI. Lecture Notes in Computer Science, 7139:83–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29364-1_7/.

Shafait:2012:DFL

[274] Faisal Shafait, Michael Patrick Cutter, Joost van Beusekom, and Syed Saqib Bukhari. Decapod: a flexible, low cost digitization solution for small and medium archives. Lecture Notes in Computer Science, 7139:101–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29364-1_8/.

Vajda:2012:MCB

[275] Szilárd Vajda, Leonard Rothacker, and Gernot A. Fink. A method for camera-based interactive whiteboard reading. Lecture Notes in Computer Science, 7139:112–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29364-1_9/.

Bukhari:2012:BNR

[276] Syed Saqib Bukhari, Faisal Shafait, and Thomas M. Breuel. Border noise removal of camera-captured document images using page frame detection. Lecture Notes in Computer Science, 7139:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29364-1_10/.

Bukhari:2012:IBP

[277] Syed Saqib Bukhari, Faisal Shafait, and Thomas M. Breuel. An image based performance evaluation method for page dewarping algorithms using SIFT features. Lecture Notes in Computer Science, 7139:138–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29364-1_11/.

Nagy:2012:NCD

[278] Robert Nagy, Anders Dicker, and Klaus Meyer-Wegener. NEOCR: a configurable dataset for natural image text recognition. Lecture Notes in Computer Science, 7139:150–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29364-1_12/.

Bukhari:2012:IDC

[279] Syed Saqib Bukhari, Faisal Shafait, and Thomas M. Breuel. The IUPR dataset of camera-captured document images. Lecture Notes in Computer Science, 7139:164–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29364-1_13/.

Anonymous:2012:BMe

[280] Anonymous. Back matter. Lecture Notes in Computer Science, 7139: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29364-1/1.

Anonymous:2012:FMn

[281] Anonymous. Front matter. Lecture Notes in Computer Science, 7139: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29364-1/1.

Aizatulin:2012:VCC

[282] Mihhail Aizatulin and François Dupressoir. Verifying cryptographic code in C: Some experience and the Csec challenge. Lecture Notes in Computer Science, 7140:1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29420-4_1/.

DeGroef:2012:BSP

[283] Willem De Groef, Dominique Devriese, and Frank Piessens. Better security and privacy for Web browsers: a survey of techniques, and a new implementation. Lecture Notes in Computer Science, 7140:21–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29420-4_2/.

Alvim:2012:DPT

[284] Mário S. Alvim and Miguel E. Andrés. Differential privacy: On the tradeoff between utility and information leakage. Lecture Notes in Computer Science, 7140:39–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29420-4_3/.

Bello:2012:FID

[285] Luciano Bello and Eduardo Bonelli. On-the-fly inlining of dynamic dependency monitors for secure information flow. Lecture Notes in Computer Science, 7140:55–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29420-4_4/.

Espinoza:2012:MEL

[286] Barbara Espinoza and Geoffrey Smith. Min-entropy leakage of channels in cascade. Lecture Notes in Computer Science, 7140:70–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29420-4_5/.

Focardi:2012:SRD

[287] Riccardo Focardi and Flaminia L. Luccio. Secure recharge of disposable RFID tickets. Lecture Notes in Computer Science, 7140:85–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29420-4_6/.

Foley:2012:ADS

[288] Simon N. Foley and Samane Abdi. Avoiding delegation subterfuge using linked local permission names. Lecture Notes in Computer Science, 7140:100–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29420-4_7/.

Fontaine:2012:VCF

[289] Arnaud Fontaine, Samuel Hym, and Isabelle Simplot-Ryl. Verifiable control flow policies for Java bytecode. Lecture Notes in Computer Science, 7140:115–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29420-4_8/.

Froschle:2012:CPC

[290] Sibylle Fröschle and Nils Sommer. Concepts and proofs for configuring PKCS#11. Lecture Notes in Computer Science, 7140:131–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29420-4_9/.

Gay:2012:SA

[291] Richard Gay, Heiko Mantel, and Barbara Sprick. Service automata. Lecture Notes in Computer Science, 7140:148–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29420-4_10/.

Gibson-Robinson:2012:AAL

[292] Thomas Gibson-Robinson and Gavin Lowe. Analysing applications layered on unilaterally authenticating protocols. Lecture Notes in Computer Science, 7140:164–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29420-4_11/.

Grabowski:2012:TBE

[293] Robert Grabowski, Martin Hofmann, and Keqin Li. Type-based enforcement of secure programming guidelines — code injection prevention at SAP. Lecture Notes in Computer Science, 7140:182–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29420-4_12/.

Hinrichs:2012:THL

[294] Timothy L. Hinrichs and William C. Garrison III. TBA: a hybrid of logic and extensional access control systems. Lecture Notes in Computer Science, 7140:198–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29420-4_13/.

Modersheim:2012:DHD

[295] Sebastian Mödersheim. Diffie–Hellman without difficulty. Lecture Notes in Computer Science, 7140:214–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29420-4_14/.

Sattarzadeh:2012:CAD

[296] Behnam Sattarzadeh and Mehran S. Fallah. Is cryptyc able to detect insider attacks? Lecture Notes in Computer Science, 7140:230–244, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29420-4_15/.

Smyth:2012:FAA

[297] Ben Smyth, Mark Ryan, and Liqun Chen. Formal analysis of anonymity in ECC-based direct anonymous attestation schemes. Lecture Notes in Computer Science, 7140:245–262, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29420-4_16/.

Dashti:2012:RBO

[298] Mohammad Torabi Dashti and Jan Cederquist. Risk balance in optimistic non-repudiation protocols. Lecture Notes in Computer Science, 7140:263–277, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29420-4_17/.

Anonymous:2012:BMf

[299] Anonymous. Back matter. Lecture Notes in Computer Science, 7140: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29420-4/1.

Anonymous:2012:FMo

[300] Anonymous. Front matter. Lecture Notes in Computer Science, 7140: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29420-4/1.

Katoen:2012:MCO

[301] Joost-Pieter Katoen. Model checking: One can do much more than you think! Lecture Notes in Computer Science, 7141:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29320-7_1/.

Bergstra:2012:PAS

[302] Jan A. Bergstra and Alban Ponse. Proposition algebra and short-circuit logic. Lecture Notes in Computer Science, 7141:15–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29320-7_2/.

Aceto:2012:DRA

[303] Luca Aceto, Arnar Birgisson, and Anna Ingólfsdóttir. Decompositional reasoning about the history of parallel processes. Lecture Notes in Computer Science, 7141:32–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29320-7_3/.

Kolahdouz-Rahimi:2012:MBD

[304] Shekoufeh Kolahdouz-Rahimi and Kevin Lano. A model-based development approach for model transformations. Lecture Notes in Computer Science, 7141:48–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29320-7_4/.

Lambertz:2012:ACB

[305] Christian Lambertz and Mila Majster-Cederbaum. Analyzing componentbased systems on the basis of architectural constraints. Lecture Notes in Computer Science, 7141:64–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29320-7_5/.

Haghighi:2012:CDP

[306] Hassan Haghighi and Mohammad Mahdi Javanmard. Constructive development of probabilistic programs. Lecture Notes in Computer Science, 7141:80–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29320-7_6/.

Jaghoori:2012:CRT

[307] Mohammad Mahdi Jaghoori. Composing real-time concurrent objects refinement, compatibility and schedulability. Lecture Notes in Computer Science, 7141:96–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29320-7_7/.

Groote:2012:SGA

[308] Jan Friso Groote and Tim W. D. M. Kouters. Specification guidelines to avoid the state space explosion problem. Lecture Notes in Computer Science, 7141:112–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29320-7_8/.

Demangeon:2012:SNC

[309] Romain Demangeon and Daniel Hirschkoff. Strong normalisation in λcalculi with references. Lecture Notes in Computer Science, 7141:128– 142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29320-7_9/.

Deng:2012:CRM

[310] Yuxin Deng and Matthew Hennessy. Compositional reasoning for Markov decision processes. Lecture Notes in Computer Science, 7141:143–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29320-7_10/.

Johnsen:2012:SLM

[311] Einar Broch Johnsen, Thi Mai Thuong Tran, and Olaf Owe. Safe locking for multi-threaded Java. Lecture Notes in Computer Science, 7141:158– 173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29320-7_11/.

Hwong:2012:ACS

[312] Yi-Ling Hwong and Vincent J. J. Kusters. Analysing the control software of the Compact Muon Solenoid Experiment at the Large Hadron Collider. Lecture Notes in Computer Science, 7141:174–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29320-7_12/.

Kim:2012:DLN

[313] Minyoung Kim, Mark-Oliver Stehr, and Carolyn Talcott. A distributed logic for networked cyber-physical systems. Lecture Notes in Computer Science, 7141:190–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29320-7_13/.

Hansen:2012:RAN

[314] Hallstein Asheim Hansen and Gerardo Schneider. Reachability analysis of non-linear planar autonomous systems. Lecture Notes in Computer Science, 7141:206–220, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29320-7_14/.

Yang:2012:ADP

[315] Qiusong Yang, Bei Zhang, Jian Zhai, and Mingshu Li. Attacking the dimensionality problem of parameterized systems via bounded reachability graphs. Lecture Notes in Computer Science, 7141:221–235, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29320-7_15/.

Kamali:2012:RBM

[316] Maryam Kamali, Luigia Petre, and Kaisa Sere. Refinement-based modeling of 3D NoCs. Lecture Notes in Computer Science, 7141:236–252, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29320-7_16/.

Asaadi:2012:TMB

[317] Hamid Reza Asaadi and Ramtin Khosravi. Towards model-based testing of electronic funds transfer systems. Lecture Notes in Computer Science, 7141:253–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29320-7_17/.

Aceto:2012:RMR

[318] Luca Aceto and Ignacio Fábregas. Relating modal refinements, covariantcontravariant simulations and partial bisimulations. Lecture Notes in Computer Science, 7141:268–283, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29320-7_18/.

He:2012:DBE

[319] Chaodong He, Yuxi Fu, and Hongfei Fu. Decidability of behavioral equivalences in process calculi with name scoping. Lecture Notes in Computer Science, 7141:284–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29320-7_19/.

Courbis:2012:RAP

[320] Roméo Courbis. Rewriting approximations for properties verification over CCS specifications. Lecture Notes in Computer Science, 7141:299– 315, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29320-7_20/.

Anonymous:2012:FMp

[321] Anonymous. Front matter. Lecture Notes in Computer Science, 7141: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29320-7/1.

ElGindy:2012:EQP

[322] Ehab ElGindy and Alia Abdelmoty. Enhancing the quality of place resources in geo-folksonomies. Lecture Notes in Computer Science, 7142: 1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28635-3_1/.

Wang:2012:GSB

[323] Huaishuai Wang, Peiquan Jin, Lei Zhao, Lanlan Zhang, and Lihua Yue. Generating semantic-based trajectories for indoor moving objects. Lecture Notes in Computer Science, 7142:13–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28635-3_2/.

Fang:2012:HTE

[324] Ying Fang, Jiaheng Cao, Junzhou Wang, Yuwei Peng, and Wei Song. HTPR*-tree: An efficient index for moving objects to support predictive query and partial history query. Lecture Notes in Computer Science, 7142:26–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28635-3_3/.

Meyers:2012:DRW

[325] Michael Meyers and Bruce A. Ralston. Developing rich Web GIS applications for visual analytics. Lecture Notes in Computer Science, 7142: 40–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28635-3_4/.

Htoo:2012:SSM

[326] Htoo Htoo, Yutaka Ohsawa, and Noboru Sonehara. Single-source multitarget A* algorithm for POI queries on road network. Lecture Notes in Computer Science, 7142:51–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28635-3_5/.

Liu:2012:CTQ

[327] Weimo Liu, Yinan Jing, Kunjie Chen, and Weiwei Sun. Combining top-k query in road networks. Lecture Notes in Computer Science, 7142:63– 75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28635-3_6/.

Zhang:2012:EFL

[328] Qingqing Zhang, Peiquan Jin, Sheng Lin, and Lihua Yue. Extracting focused locations for Web pages. Lecture Notes in Computer Science, 7142: 76–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28635-3_7/.

Ma:2012:SST

[329] Yuchi Ma, Chunyan Qu, Tingting Liu, Ning Yang, and Changjie Tang. Searching similar trajectories in real time: An effectiveness and efficiency study. Lecture Notes in Computer Science, 7142:90–102, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28635-3_8/.

Krizka:2012:MIS

[330] Filip Křižka, Michal Krátký, Radim Bača, and Peter Chovanec. Multidimensional implementation of stream ADT. Lecture Notes in Computer Science, 7142:103–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28635-3_9/.

Tang:2012:MXS

[331] Ruiming Tang, Huayu Wu, and Stéphane Bressan. Measuring XML structured-ness with entropy. Lecture Notes in Computer Science, 7142:113–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28635-3_10/.

Wang:2012:SJX

[332] Yue Wang, Hongzhi Wang, Yang Wang, and Hong Gao. Similarity join on XML based on k-generation set distance. Lecture Notes in Computer Science, 7142:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28635-3_11/.

Yao:2012:XQP

[333] Caiyun Yao, Jiaheng Lu, Wei Wang, and Xiaofang Zhou. XML query processing using views. Lecture Notes in Computer Science, 7142:136– 139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28635-3_12/.

Li:2012:XSO

[334] Xia Li, Zhanhuai Li, PeiYing Wang, Qun Chen, Lijun Zhang, and Ning Li. XIO-SLCA: Optimize SLCA for effective keyword search in XML documents. Lecture Notes in Computer Science, 7142:140–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28635-3_13/.

Alahmari:2012:DXS

[335] Fahad Alahmari and Eric Pardede. The development of XML stored procedures in XML enabled databases. Lecture Notes in Computer Science, 7142:150–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28635-3_14/.

Li:2012:SOC

[336] Jingjiao Li, Limei Sun, and Jiao Wang. A slope one collaborative filtering recommendation algorithm using uncertain neighbors optimizing. Lecture Notes in Computer Science, 7142:160–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28635-3_15/.

He:2012:SRM

[337] Di He, Zhiyong Peng, Liang Hong, and Yu Zhang. A social reputation management for Web communities. Lecture Notes in Computer Science, 7142:167–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28635-3_16/.

Zhang:2012:CFR

[338] Dongzhan Zhang and Chao Xu. A collaborative filtering recommendation system by unifying user similarity and item similarity. Lecture Notes in Computer Science, 7142:175–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28635-3_17/.

Li:2012:SQD

[339] Rongrong Li, Weixiang Zhai, and Zhiyong Peng. Supporting query over dynamic combination of data sources for social media. Lecture Notes in Computer Science, 7142:185–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28635-3_18/.

Song:2012:DOL

[340] Kaisong Song, Daling Wang, Shi Feng, and Ge Yu. Detecting opinion leader dynamically in Chinese news comments. Lecture Notes in Computer Science, 7142:197–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28635-3_19/.

Wang:2012:ASA

[341] Daling Wang, Shi Feng, Chao Yan, and Ge Yu. An approach of semiautomatic public sentiment analysis for opinion and district. Lecture Notes in Computer Science, 7142:210–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28635-3_20/.

Anonymous:2012:BMg

[342] Anonymous. Back matter. Lecture Notes in Computer Science, 7142: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28635-3/1.

Anonymous:2012:FMq

[343] Anonymous. Front matter. Lecture Notes in Computer Science, 7142: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28635-3/1.

Wolfe:2012:RGV

[344] Jeremy M. Wolfe. The rules of guidance in visual search. Lecture Notes in Computer Science, 7143:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27387-2_1/.

Suryadevara:2012:SSS

[345] N. K. Suryadevara, T. Quazi, and Subhas C. Mukhopadhyay. Smart sensing system for human emotion and behaviour recognition. Lecture Notes in Computer Science, 7143:11–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27387-2_2/.

Pal:2012:FRB

[346] Nikhil Ranjan Pal. Fuzzy rule-based approaches to dimensionality reduction. Lecture Notes in Computer Science, 7143:23–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27387-2_3/.

Yamaguchi:2012:FRT

[347] Osamu Yamaguchi. Face recognition technology and its real-world application. Lecture Notes in Computer Science, 7143:28–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27387-2_4/.

Mandal:2012:CEV

[348] Atanendu Sekhar Mandal. Contextual effects in the visual cortex area 1 (V1) and camouflage perception. Lecture Notes in Computer Science, 7143:35–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27387-2_5/.

Sugu:2012:AIP

[349] Dana Sugu and Amita Chatterjee. Affective information processing and representations. Lecture Notes in Computer Science, 7143:42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27387-2_6/.

S:2012:FIR

[350] Rakesh S., Kailash Atal, Ashish Arora, Pulak Purkait, and Bhabatosh Chanda. Face image retrieval based on probe sketch using SIFT feature descriptors. Lecture Notes in Computer Science, 7143:50–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27387-2_7/.

Bandyopadhyay:2012:EPA

[351] Asok Bandyopadhyay, Bhaswati Mukherjee, and Bidyut Baran Chaudhuri. Effect of perceptual anchorage points on recognition of Bangla characters. Lecture Notes in Computer Science, 7143:58–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27387-2_8/.

Bakshi:2012:SPM

[352] Ashish Bakshi and Kuntal Ghosh. Scaling properties of Mach bands and perceptual models. Lecture Notes in Computer Science, 7143:66– 74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27387-2_9/.

Das:2012:PVE

[353] Apurba Das and Suparna Parua. Psycho-visual evaluation of contrast enhancement algorithms by adaptive neuro-fuzzy inference system. Lecture Notes in Computer Science, 7143:75–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27387-2_10/.

Hamada:2012:TMM

[354] Fumito Hamada and Edson T. Miyamoto. Theory of mind in manmachine interactions. Lecture Notes in Computer Science, 7143:84–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27387-2_11/.

Dauthal:2012:FRS

[355] Naman Dauthal, Surya Prakash, and Phalguni Gupta. Face recognition system invariant to light-camera setup. Lecture Notes in Computer Science, 7143:90–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27387-2_12/.

Vishwakarma:2012:RFQ

[356] Sarvesh Vishwakarma and Anupam Agrawal. Representing feature quantization approach using spatial-temporal relation for action recognition. Lecture Notes in Computer Science, 7143:98–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27387-2_13/.

Rautaray:2012:HCI

[357] Siddharth S. Rautaray, Anand Kumar, and Anupam Agrawal. Human computer interaction with hand gestures in virtual environment. Lecture Notes in Computer Science, 7143:106–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27387-2_14/.

Konar:2012:ITF

[358] Amit Konar, Aruna Chakraborty, Anisha Halder, Rajshree Mandal, and Ramadoss Janarthanan. Interval type-2 fuzzy model for emotion recognition from facial expression. Lecture Notes in Computer Science, 7143:114–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27387-2_15/.

Bondale:2012:ALT

[359] Nandini Bondale and Gaurav Gupta. Augmenting language tools for effective communication. Lecture Notes in Computer Science, 7143:122– 128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27387-2_16/.

Sinha:2012:DEC

[360] Debmalya Sinha and Anupam Basu. Design and evaluation of a cognition aware file browser for users in rural India. Lecture Notes in Computer Science, 7143:129–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27387-2_17/.

Singhal:2012:SNA

[361] Amit Singhal, Santanu Chaudhury, and Sumantra Dutta Roy. Shiksha: a novel architecture for tele-teaching using handwriting as a perceptually significant temporal media. Lecture Notes in Computer Science, 7143:137–144, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27387-2_18/.

Saha:2012:NMB

[362] Chandrani Saha, Washef Ahmed, and Soma Mitra. A new motion based fully automatic facial expression recognition system. Lecture Notes in Computer Science, 7143:145–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27387-2_19/.

Yoshida:2012:PSM

[363] Koichi Yoshida, Emiko Ishikawa, Maltesh Joshi, Hervé Lechat, and Fatma Ayouni. Profiling Scotch malt whisky spirits from different distilleries using an electronic nose and an expert sensory panel. Lecture Notes in Computer Science, 7143:155–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27387-2_20/.

Anonymous:2012:FMr

[364] Anonymous. Front matter. Lecture Notes in Computer Science, 7143: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27387-2/1.

Nambiar:2012:SLI

[365] Raghunath Nambiar and Nicholas Wakou. Shaping the landscape of industry standard benchmarks: Contributions of the Transaction Processing Performance Council (TPC). Lecture Notes in Computer Science, 7144:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32627-1_1/.

Funke:2012:MMP

[366] Florian Funke and Alfons Kemper. Metrics for measuring the performance of the mixed workload CH-benCHmark. Lecture Notes in Computer Science, 7144:10–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32627-1_2/.

Schall:2012:TEB

[367] Daniel Schall, Volker Hoefner, and Manuel Kern. Towards an enhanced benchmark advocating energy-efficient systems. Lecture Notes in Computer Science, 7144:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32627-1_3/.

Dayal:2012:OAD

[368] Umeshwar Dayal and Kevin Wilkinson. Optimization of analytic data flows for next generation business intelligence applications. Lecture Notes in Computer Science, 7144:46–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32627-1_4/.

Bog:2012:NMO

[369] Anja Bog, Kai Sachs, and Alexander Zeier. Normalization in a mixed OLTP and OLAP workload scenario. Lecture Notes in Computer Science, 7144:67–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32627-1_5/.

Grabs:2012:MPC

[370] Torsten Grabs. Measuring performance of complex event processing systems. Lecture Notes in Computer Science, 7144:83–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32627-1_6/.

Huppler:2012:BYH

[371] Karl Huppler. Benchmarking with your head in the cloud. Lecture Notes in Computer Science, 7144:97–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32627-1_7/.

Li:2012:ETM

[372] Yantao Li and Charles Levine. Extending TPC-E to measure availability in database systems. Lecture Notes in Computer Science, 7144:111– 122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32627-1_8/.

Gottstein:2012:SCS

[373] Robert Gottstein and Ilia Petrov. SI-CV: Snapshot isolation with colocated versions. Lecture Notes in Computer Science, 7144:123–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32627-1_9/.

Crolotte:2012:IST

[374] Alain Crolotte and Ahmad Ghazal. Introducing skew into the TPC-H benchmark. Lecture Notes in Computer Science, 7144:137–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32627-1_10/.

Delimitrou:2012:TCE

[375] Christina Delimitrou and Sriram Sankar. Time and cost-efficient modeling and generation of large-scale TPCC/TPCE/TPCH workloads. Lecture Notes in Computer Science, 7144:146–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32627-1_11/.

Floratou:2012:WFR

[376] Avrilia Floratou and Jignesh M. Patel. When free is not really free: What does it cost to run a database workload in the cloud? Lecture Notes in Computer Science, 7144:163–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32627-1_12/.

Chen:2012:FGP

[377] Jianhai Chen, Dawei Huang, and Bei Wang. A fine-grained performancebased decision model for virtualization application solution. Lecture Notes in Computer Science, 7144:180–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32627-1_13/.

Poess:2012:PIT

[378] Meikel Poess, Tilmann Rabl, and Michael Frank. A PDGF implementation for TPC-H. Lecture Notes in Computer Science, 7144:196–212, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32627-1_14/.

Anonymous:2012:BMh

[379] Anonymous. Back matter. Lecture Notes in Computer Science, 7144: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32627-1/1.

Anonymous:2012:FMs

[380] Anonymous. Front matter. Lecture Notes in Computer Science, 7144: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32627-1/1.

Bonneaud:2012:APC

[381] Stéphane Bonneaud, Kevin Rio, Pierre Chevaillier, and William H. Warren. Accounting for patterns of collective behavior in crowd locomotor dynamics for realistic simulations. Lecture Notes in Computer Science, 7145:1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_1/.

Valentin:2012:HBM

[382] Julien Valentin, Florent Coudret, Eric Gouardères, and Wilfrid Lefer. Human behaviour modelling for simulating evacuation of buildings on fire. Lecture Notes in Computer Science, 7145:12–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29050-3_2/.

Zhang:2012:ASV

[383] Lin Zhang and Qing Liu. Application of simulation and virtual reality to physical education and athletic training. Lecture Notes in Computer Science, 7145:24–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_3/.

Hua:2012:DVH

[384] Wei Hua, Wenzhuo Yang, Zilong Dong, and Guofeng Zhang. Depthvarying human video Sprite synthesis. Lecture Notes in Computer Science, 7145:34–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_4/.

Kim:2012:AEL

[385] Yejin Kim and Michael Neff. Automating expressive locomotion generation. Lecture Notes in Computer Science, 7145:48–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29050-3_5/.

Fan:2012:RGD

[386] Wenshan Fan, Bin Wang, Bin Chan, Jean-Claude Paul, and Jiaguang Sun. Recovering geometric detail by octree normal maps. Lecture Notes in Computer Science, 7145:62–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29050-3_6/.

Luo:2012:QTA

[387] Jianxin Luo, Guiqiang Ni, Ping Cui, Jinsong Jiang, Yifeng Duan, and Guyu Hu. Quad-tree Atlas ray casting: a GPU based framework for terrain visualization and its applications. Lecture Notes in Computer Science, 7145:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_7/.

Hung:2012:CSD

[388] Shao-Shin Hung, Chih Ming Chiu, Tsou Tsun Fu, Jung-Tsung Chen, and Derchian Tsaih. Clustering spatial data for aggregate query processing in walkthrough: a hypergraph approach. Lecture Notes in Computer Science, 7145:86–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_8/.

Gao:2012:SDS

[389] Yujian Gao, Aimin Hao, and Qinping Zhao. Skin-detached surface for interactive large mesh editing. Lecture Notes in Computer Science, 7145:99–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_9/.

Liu:2012:PBS

[390] Shiguang Liu, Tai An, Zheng Gong, and Ichiro Hagiwara. Physically based simulation of solid objects’ burning. Lecture Notes in Computer Science, 7145:110–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_10/.

Wang:2012:TBS

[391] Chao Wang, Yu Wu, Tingting Zhu, Hongbo Li, and Mingliang Xu. Tunneling blasting simulation for digital mine. Lecture Notes in Computer Science, 7145:121–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_11/.

Aranda:2012:MMO

[392] Gustavo Aranda, Tomas Trescak, Marc Esteva, Inmaculada Rodriguez, and Carlos Carrascosa. Massively multiplayer online games developed with agents. Lecture Notes in Computer Science, 7145:129–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29050-3_12/.

He:2012:NSA

[393] Zhiying He, Xiaohui Liang, and Qinping Zhao. A novel skeletonization and animation approach for point models. Lecture Notes in Computer Science, 7145:139–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_13/.

Zhang:2012:VDL

[394] Xin Zhang, Zi’ang Ding, Chuan Zhu, Wei Chen, and Qunsheng Peng. View-dependent line drawings for 3D scenes. Lecture Notes in Computer Science, 7145:151–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_14/.

Szilas:2012:PIF

[395] Nicolas Szilas, Monica Axelrad, and Urs Richle. Propositions for innovative forms of digital interactive storytelling based on narrative theories and practices. Lecture Notes in Computer Science, 7145:161–179, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29050-3_15/.

Chen:2012:VIL

[396] Shengnan Chen, Zhigeng Pan, and Mingmin Zhang. A virtual informal learning system for cultural heritage. Lecture Notes in Computer Science, 7145:180–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29050-3_16/.

Cai:2012:MST

[397] Xingquan Cai, Jie Li, Haiyan Sun, and Jinhong Li. Multi-samples texture synthesis for dynamic terrain based on constraint conditions. Lecture Notes in Computer Science, 7145:188–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29050-3_17/.

Liu:2012:PFD

[398] Zhicheng Liu and Zhijun Fang. PCA and FLD in DWT domain. Lecture Notes in Computer Science, 7145:197–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29050-3_18/.

Dong:2012:AMM

[399] Shidu Dong, Xiaofan Yang, Huaqiu Wang, Xiang Liu, and Hengyang Liu. Area measurement method of paint bubble based on computer vision. Lecture Notes in Computer Science, 7145:205–213, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29050-3_19/.

Shi:2012:HPF

[400] Xifan Shi, Xianghua Chen, and Tiefeng Cai. A high precision fresco scanner. Lecture Notes in Computer Science, 7145:214–221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29050-3_20/.

Anonymous:2012:FMt

[401] Anonymous. Front matter. Lecture Notes in Computer Science, 7145: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29050-3/1.

Bouajjani:2012:ADA

[402] Ahmed Bouajjani and Cezara Drăgoi. Abstract domains for automated reasoning about list-manipulating programs with infinite data. Lecture Notes in Computer Science, 7148:1–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27940-9_1/.

Jhala:2012:SVL

[403] Ranjit Jhala. Software verification with liquid types. Lecture Notes in Computer Science, 7148:23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-27940-9_2.

Nipkow:2012:TSP

[404] Tobias Nipkow. Teaching semantics with a proof assistant: No more LSD trip proofs. Lecture Notes in Computer Science, 7148:24–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27940-9_3/.

Albarghouthi:2012:WIB

[405] Aws Albarghouthi and Arie Gurfinkel. Whale: An interpolation-based algorithm for inter-procedural verification. Lecture Notes in Computer Science, 7148:39–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27940-9_4/.

Basu:2012:SVA

[406] Samik Basu, Tevfik Bultan, and Meriem Ouederni. Synchronizability for verification of asynchronously communicating systems. Lecture Notes in Computer Science, 7148:56–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27940-9_5/.

Ben-Amram:2012:TIL

[407] Amir M. Ben-Amram and Samir Genaim. On the termination of integer loops. Lecture Notes in Computer Science, 7148:72–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27940-9_6/.

Bozzelli:2012:VGO

[408] Laura Bozzelli and Sophie Pinchinat. Verification of gap-order constraint abstractions of counter systems. Lecture Notes in Computer Science, 7148:88–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27940-9_7/.

Bugaychenko:2012:AMR

[409] Dmitry Bugaychenko. On application of multi-rooted binary decision diagrams to probabilistic model checking. Lecture Notes in Computer Science, 7148:104–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27940-9_8/.

Chaki:2012:RVM

[410] Sagar Chaki, Arie Gurfinkel, and Ofer Strichman. Regression verification for multi-threaded programs. Lecture Notes in Computer Science, 7148:119–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27940-9_9/.

Charlton:2012:CVH

[411] Nathaniel Charlton and Ben Horsfall. Crowfoot: a verifier for higherorder store programs. Lecture Notes in Computer Science, 7148:136– 151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27940-9_10/.

Chatterjee:2012:SPD

[412] Krishnendu Chatterjee and Vishwanath Raman. Synthesizing protocols for digital contract signing. Lecture Notes in Computer Science, 7148:152–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27940-9_11/.

Dimitrova:2012:MCI

[413] Rayna Dimitrova and Bernd Finkbeiner. Model checking information flow in reactive systems. Lecture Notes in Computer Science, 7148:169– 185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27940-9_12/.

Ermis:2012:SI

[414] Evren Ermis and Jochen Hoenicke. Splitting via interpolants. Lecture Notes in Computer Science, 7148:186–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27940-9_13/.

Ferrara:2012:AIA

[415] Pietro Ferrara and Peter Müller. Automatic inference of access permissions. Lecture Notes in Computer Science, 7148:202–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27940-9_14/.

Finkbeiner:2012:LS

[416] Bernd Finkbeiner and Swen Jacobs. Lazy synthesis. Lecture Notes in Computer Science, 7148:219–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27940-9_15/.

Ghorbal:2012:DDE

[417] Khalil Ghorbal and Franjo Ivančić. Donut domains: Efficient non-convex domains for abstract interpretation. Lecture Notes in Computer Science, 7148:235–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27940-9_16/.

Howar:2012:ICR

[418] Falk Howar, Bernhard Steffen, and Bengt Jonsson. Inferring canonical register automata. Lecture Notes in Computer Science, 7148:251– 266, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27940-9_17/.

Kinder:2012:ACF

[419] Johannes Kinder and Dmitry Kravchenko. Alternating control flow reconstruction. Lecture Notes in Computer Science, 7148:267–282, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27940-9_18/.

Klein:2012:ESA

[420] Uri Klein, Nir Piterman, and Amir Pnueli. Effective synthesis of asynchronous systems from GR(1) specifications. Lecture Notes in Computer Science, 7148:283–298, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27940-9_19/.

Lee:2012:SNS

[421] Woosuk Lee, Wonchan Lee, and Kwangkeun Yi. Sound non-statistical clustering of static analysis alarms. Lecture Notes in Computer Science, 7148:299–314, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27940-9_20/.

Anonymous:2012:FMu

[422] Anonymous. Front matter. Lecture Notes in Computer Science, 7148: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27940-9/1.

Loo:2012:RAD

[423] Boon Thau Loo, Harjot Gill, and Changbin Liu. Recent advances in declarative networking. Lecture Notes in Computer Science, 7149:1– 16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27694-1_1/.

Stewart:2012:MTN

[424] Don Stewart. Make things now! pragmatic functional programming in Haskell. Lecture Notes in Computer Science, 7149:17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-27694-1_2.

Goldberg:2012:DAS

[425] Mayer Goldberg and Guy Wiener. A declarative approach for software modeling. Lecture Notes in Computer Science, 7149:18–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27694-1_3/.

Antoy:2012:CSF

[426] Sergio Antoy and Michael Hanus. Contracts and specifications for functional logic programming. Lecture Notes in Computer Science, 7149:33– 47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27694-1_4/.

Martins:2012:EA

[427] Pedro M. Martins and Julie A. McCann. The environment as an argument. Lecture Notes in Computer Science, 7149:48–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27694-1_5/.

Lierler:2012:WSP

[428] Yuliya Lierler and Shaden Smith. Weighted-sequence problem: ASP vs CASP and declarative vs problem-oriented solving. Lecture Notes in Computer Science, 7149:63–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27694-1_6/.

Balduccini:2012:PMA

[429] Marcello Balduccini and Yulia Lierler. Practical and methodological aspects of the use of cutting-edge ASP tools. Lecture Notes in Computer Science, 7149:78–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27694-1_7/.

Have:2012:ETS

[430] Christian Theil Have and Henning Christiansen. Efficient tabling of structured data using indexing and program transformation. Lecture Notes in Computer Science, 7149:93–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27694-1_8/.

Campagna:2012:OIJ

[431] Dario Campagna and Beata Sarna-Starosta. Optimizing inequality joins in datalog with approximated constraint propagation. Lecture Notes in Computer Science, 7149:108–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27694-1_9/.

Albert:2012:SEC

[432] Elvira Albert and Puri Arenas. Symbolic execution of concurrent objects in CLP. Lecture Notes in Computer Science, 7149:123–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27694-1_10/.

deGuzman:2012:SSA

[433] Pablo Chico de Guzmán and Amadeo Casas. A segment-swapping approach for executing trapped computations. Lecture Notes in Computer Science, 7149:138–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27694-1_11/.

Lesniak:2012:PDE

[434] Michael Lesniak. Palovca: Describing and executing graph algorithms in Haskell. Lecture Notes in Computer Science, 7149:153–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27694-1_12/.

Zhu:2012:LII

[435] Kenny Q. Zhu, Kathleen Fisher, and David Walker. LearnPADS^{++}: Incremental inference of ad hoc data formats. Lecture Notes in Computer Science, 7149:168–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27694-1_13/.

Bransen:2012:KWA

[436] Jeroen Bransen and Arie Middelkoop. The Kennedy–Warren Algorithm revisited: Ordering attribute grammars. Lecture Notes in Computer Science, 7149:183–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27694-1_14/.

Coleman:2012:DPS

[437] Nicholas Coleman. Distributed policy specification and interpretation with classified advertisements. Lecture Notes in Computer Science, 7149:198–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27694-1_15/.

Gill:2012:HKL

[438] Andy Gill and Bowe Neuenschwander. Handshaking in Kansas Lava using patch logic. Lecture Notes in Computer Science, 7149:212–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27694-1_16/.

Winograd-Cort:2012:VRW

[439] Daniel Winograd-Cort, Hai Liu, and Paul Hudak. Virtualizing realworld objects in FRP. Lecture Notes in Computer Science, 7149:227– 241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27694-1_17/.

Brady:2012:RSS

[440] Edwin Brady and Kevin Hammond. Resource-safe systems programming with embedded domain specific languages. Lecture Notes in Computer Science, 7149:242–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27694-1_18/.

Broman:2012:NBC

[441] David Broman and Henrik Nilsson. Node-based connection semantics for equation-based object-oriented modeling languages. Lecture Notes in Computer Science, 7149:258–272, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27694-1_19/.

Tarau:2012:DST

[442] Paul Tarau. A declarative specification of tree-based symbolic arithmetic computations. Lecture Notes in Computer Science, 7149:273–288, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27694-1_20/.

Anonymous:2012:BMi

[443] Anonymous. Back matter. Lecture Notes in Computer Science, 7149: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-27694-1/1.

Anonymous:2012:FMv

[444] Anonymous. Front matter. Lecture Notes in Computer Science, 7149: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27694-1/1.

Haque:2012:ISG

[445] Aminul Haque, Saadat M. Alhashmi, and Rajendran Parthiban. An inspiration for solving grid resource management problems using multiple economic models. Lecture Notes in Computer Science, 7150:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28675-9_1/.

Siebenhaar:2012:CNC

[446] Melanie Siebenhaar, The An Binh Nguyen, Ulrich Lampe, Dieter Schuller, and Ralf Steinmetz. Concurrent negotiations in cloud-based systems. Lecture Notes in Computer Science, 7150:17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28675-9_2/.

Roovers:2012:RAM

[447] Joris Roovers, Kurt Vanmechelen, and Jan Broeckhove. A reverse auction market for cloud resources. Lecture Notes in Computer Science, 7150:32– 45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28675-9_3/.

Kashef:2012:CMH

[448] Mohammad Mahdi Kashef and Jörn Altmann. A cost model for hybrid clouds. Lecture Notes in Computer Science, 7150:46–60, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28675-9_4/.

Stefanov:2012:HDS

[449] Hristo Stefanov, Slinger Jansen, Ronald Batenburg, Eugene van Heusden, and Ravi Khadka. How to do successful chargeback for cloud services. Lecture Notes in Computer Science, 7150:61–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28675-9_5/.

Menychtas:2012:MFT

[450] Andreas Menychtas, Sergio Garcia Gomez, Andrea Giessmann, and Anna Gatzioura. A marketplace framework for trading cloud-based services. Lecture Notes in Computer Science, 7150:76–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28675-9_6/.

Macias:2012:CCP

[451] Mario Macías and Jordi Guitart. Client classification policies for SLA negotiation and allocation in shared cloud datacenters. Lecture Notes in Computer Science, 7150:90–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28675-9_7/.

Zheng:2012:BDC

[452] Wei Zheng and Rizos Sakellariou. Budget-deadline constrained workflow planning for admission control in market-oriented environments. Lecture Notes in Computer Science, 7150:105–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28675-9_8/.

Li:2012:VMP

[453] Wubin Li, Johan Tordsson, and Erik Elmroth. Virtual machine placement for predictable and time-constrained peak loads. Lecture Notes in Computer Science, 7150:120–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28675-9_9/.

Petri:2012:RAS

[454] Ioan Petri, Omer F. Rana, Yacine Regzui, and Gheorghe Cosmin Silaghi. Risk assessment in service provider communities. Lecture Notes in Computer Science, 7150:135–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28675-9_10/.

Kunsemoller:2012:GTA

[455] Jörn Künsemöller and Holger Karl. A game-theoretical approach to the benefits of cloud computing. Lecture Notes in Computer Science, 7150:148–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28675-9_11/.

Sengupta:2012:POM

[456] Shubhashis Sengupta and K. M. Annervaz. Planning for optimal multisite data distribution for disaster recovery. Lecture Notes in Computer Science, 7150:161–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28675-9_12/.

Shi:2012:SCE

[457] Wei Shi, Dapeng Ju, and Dongsheng Wang. Saga: a cost efficient file system based on cloud storage service. Lecture Notes in Computer Science, 7150:173–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28675-9_13/.

McGough:2012:DCE

[458] A. Stephen McGough. Developing a cost-effective virtual cluster on the cloud. Lecture Notes in Computer Science, 7150:185–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28675-9_14/.

Anonymous:2012:BMj

[459] Anonymous. Back matter. Lecture Notes in Computer Science, 7150: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28675-9/1.

Anonymous:2012:FMw

[460] Anonymous. Front matter. Lecture Notes in Computer Science, 7150: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28675-9/1.

Mery:2012:MPD

[461] Dominique Méry and Neeraj Kumar Singh. Medical protocol diagnosis using formal methods. Lecture Notes in Computer Science, 7151: 1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32355-3_1/.

Davies:2012:FFF

[462] Jim Davies, Jeremy Gibbons, Radu Calinescu, Charles Crichton, Steve Harris, and Andrew Tsui. Form follows function. Lecture Notes in Computer Science, 7151:21–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32355-3_2/.

Hildebrandt:2012:DMS

[463] Thomas Hildebrandt, Raghava Rao Mukkamala, and Tijs Slaats. Declarative modelling and safe distribution of healthcare workflows. Lecture Notes in Computer Science, 7151:39–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32355-3_3/.

Bertolini:2012:TFI

[464] Cristiano Bertolini, Martin Schäf, and Volker Stolz. Towards a formal integrated model of collaborative healthcare workflows. Lecture Notes in Computer Science, 7151:57–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32355-3_4/.

MacCaull:2012:NWW

[465] Wendy MacCaull and Fazle Rabbi. NOVA workflow: a workflow management tool targeting health services delivery. Lecture Notes in Computer Science, 7151:75–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32355-3_5/.

Hooman:2012:ECM

[466] Jozef Hooman, Robert Huis in ’t Veld, and Mathijs Schuts. Experiences with a compositional model checker in the healthcare domain. Lecture Notes in Computer Science, 7151:93–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32355-3_6/.

Rabbi:2012:MCW

[467] Fazle Rabbi, Ahmed Shah Mashiyat, and Wendy MacCaull. Model checking workflow monitors and its application to a pain management process. Lecture Notes in Computer Science, 7151:111–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32355-3_7/.

Moodley:2012:PPR

[468] Deshendran Moodley, Anban W. Pillay, and Christopher J. Seebregts. Position paper: Researching and developing open architectures for national health information systems in developing African countries. Lecture Notes in Computer Science, 7151:129–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32355-3_8/.

Mery:2012:FHM

[469] Dominique Méry and Neeraj Kumar Singh. Formalization of heart models based on the conduction of electrical impulses and cellular automata. Lecture Notes in Computer Science, 7151:140–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32355-3_9/.

Xu:2012:EBA

[470] Hao Xu and Tom Maibaum. An event-B approach to timing issues applied to the generic insulin infusion pump. Lecture Notes in Computer Science, 7151:160–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32355-3_10/.

Weber-Jahnke:2012:SEM

[471] Jens H. Weber-Jahnke and Fieran Mason-Blakley. On the safety of electronic medical records. Lecture Notes in Computer Science, 7151:177– 194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32355-3_11/.

Dong:2012:CEE

[472] Naipeng Dong, Hugo Jonker, and Jun Pang. Challenges in eHealth: From enabling to enforcing privacy. Lecture Notes in Computer Science, 7151:195–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32355-3_12/.

Kalyango:2012:TSW

[473] Samuel Tusubira Kalyango and Gilbert Maiga. A technique for strengthening weak passwords in electronic medical record systems. Lecture Notes in Computer Science, 7151:207–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32355-3_13/.

Anonymous:2012:BMk

[474] Anonymous. Back matter. Lecture Notes in Computer Science, 7151: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32355-3/1.

Anonymous:2012:FMx

[475] Anonymous. Front matter. Lecture Notes in Computer Science, 7151: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32355-3/1.

Paul:2012:CWF

[476] Wolfgang Paul. Cyber war, formal verification and certified infrastructure. Lecture Notes in Computer Science, 7152:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer.com/accesspage/chapter/10.1007/978-3-642-27705-
4_1.

Herms:2012:CMP

[477] Paolo Herms, Claude Marché, and Benjamin Monate. A certified multiprover verification condition generator. Lecture Notes in Computer Science, 7152:2–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27705-4_2/.

Schmaltz:2012:ISI

[478] Sabine Schmaltz and Andrey Shadrin. Integrated semantics of intermediate-language C and macro-assembler for pervasive formal verification of operating systems and hypervisors from VerisoftXT. Lecture Notes in Computer Science, 7152:18–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27705-4_3/.

Kulczycki:2012:LLC

[479] Gregory Kulczycki and Hampton Smith. The location linking concept: a basis for verification of code using pointers. Lecture Notes in Computer Science, 7152:34–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27705-4_4/.

Polikarpova:2012:VIS

[480] Nadia Polikarpova and Michał Moskal. Verifying implementations of security protocols by refinement. Lecture Notes in Computer Science, 7152: 50–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27705-4_5/.

Wies:2012:DFL

[481] Thomas Wies, Marco Muñiz, and Viktor Kuncak. Deciding functional lists with sublist sets. Lecture Notes in Computer Science, 7152:66– 81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27705-4_6/.

Rustan:2012:DVP

[482] K. Rustan and M. Leino. Developing verified programs with dafny. Lecture Notes in Computer Science, 7152:82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-27705-4_7.

Filliatre:2012:VTL

[483] Jean-Christophe Filliâtre. Verifying two lines of C with Why3: An exercise in program verification. Lecture Notes in Computer Science, 7152: 83–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27705-4_8/.

Vujosevic-Janicic:2012:DEL

[484] Milena Vujošević-Janičić and Viktor Kuncak. Development and evaluation of LAV: An SMT-based error finding platform. Lecture Notes in Computer Science, 7152:98–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27705-4_9/.

Brain:2012:LTD

[485] Martin Brain and Florian Schanda. A lightweight technique for distributed and incremental program verification. Lecture Notes in Computer Science, 7152:114–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27705-4_10/.

Segal:2012:CIV

[486] Loren Segal and Patrice Chalin. A comparison of intermediate verification languages: Boogie and Sireum/Pilar. Lecture Notes in Computer Science, 7152:130–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27705-4_11/.

Merz:2012:LBM

[487] Florian Merz and Stephan Falke. LLBMC: Bounded model checking of C and C++ programs using a compiler IR. Lecture Notes in Computer Science, 7152:146–161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27705-4_12/.

Majumdar:2012:MED

[488] Rupak Majumdar. The marriage of exploration and deduction. Lecture Notes in Computer Science, 7152:162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-27705-4_13.

Tang:2012:MVT

[489] Nguyen Van Tang and Daisuke Souma. Modeling and validating the train fare calculation and adjustment system using VDM++. Lecture Notes in Computer Science, 7152:163–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27705-4_14/.

Mehnert:2012:FVS

[490] Hannes Mehnert and Filip Sieczkowski. Formalized verification of snapshotable trees: Separation and sharing. Lecture Notes in Computer Science, 7152:179–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27705-4_15/.

Kassios:2012:CVC

[491] Ioannis T. Kassios and Peter Müller. Comparing verification condition generation with symbolic execution: An experience report. Lecture Notes in Computer Science, 7152:196–208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27705-4_16/.

Alkassar:2012:VTV

[492] Eyad Alkassar, Ernie Cohen, and Mikhail Kovalev. Verification of TLB virtualization implemented in C. Lecture Notes in Computer Science, 7152:209–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27705-4_17/.

Post:2012:FAR

[493] Amalinda Post and Jochen Hoenicke. Formalization and analysis of realtime requirements: a feasibility study at BOSCH. Lecture Notes in Computer Science, 7152:225–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27705-4_18/.

Logozzo:2012:OEC

[494] Francesco Logozzo. Our experience with the CodeContracts static checker. Lecture Notes in Computer Science, 7152:241–242, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/accesspage/chapter/10.1007/978-3-642-
27705-4_19.

Feliachi:2012:ICP

[495] Abderrahmane Feliachi and Marie-Claude Gaudel. Isabelle/Circus: a process specification and verification environment. Lecture Notes in Computer Science, 7152:243–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27705-4_20/.

Anonymous:2012:FMy

[496] Anonymous. Front matter. Lecture Notes in Computer Science, 7152: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27705-4/1.

Eiter:2012:UEN

[497] Thomas Eiter, Thomas Krennwallner, Patrik Schneider, and Guohui Xiao. Uniform evaluation of nonmonotonic DL-programs. Lecture Notes in Computer Science, 7153:1–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28472-4_1/.

Schwentick:2012:FXB

[498] Thomas Schwentick. Foundations of XML based on logic and automata: a snapshot. Lecture Notes in Computer Science, 7153:23–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28472-4_2/.

Arieli:2012:ITK

[499] Ofer Arieli and Anna Zamansky. Inconsistency-tolerance in knowledgebased systems by dissimilarities. Lecture Notes in Computer Science, 7153:34–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28472-4_3/.

Biskup:2012:RBR

[500] Joachim Biskup and Cornelia Tadros. Revising belief without revealing secrets. Lecture Notes in Computer Science, 7153:51–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28472-4_4/.

Bock:2012:MTS

[501] Jürgen Bock, Sebastian Rudolph, and Michael Mutter. More than the sum of its parts — holistic ontology alignment by population-based optimisation. Lecture Notes in Computer Science, 7153:71–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28472-4_5/.

Caroprese:2012:HPP

[502] Luciano Caroprese and Ester Zumpano. Handling preferences in P2P systems. Lecture Notes in Computer Science, 7153:91–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28472-4_6/.

Cohen:2012:BUA

[503] Andrea Cohen, Alejandro J. García, and Guillermo R. Simari. Backing and undercutting in abstract argumentation frameworks. Lecture Notes in Computer Science, 7153:107–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28472-4_7/.

Fletcher:2012:ITC

[504] George H. L. Fletcher, Marc Gyssens, Dirk Leinders, Jan Van den Bussche, and Dirk Van Gucht. The impact of transitive closure on the Boolean expressiveness of navigational query languages on graphs. Lecture Notes in Computer Science, 7153:124–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28472-4_8/.

Garnaud:2012:UFD

[505] Eve Garnaud, Sofian Maabout, and Mohamed Mosbah. Using functional dependencies for reducing the size of a data cube. Lecture Notes in Computer Science, 7153:144–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28472-4_9/.

Hartmann:2012:ADR

[506] Sven Hartmann, Henning Köhler, Sebastian Link, and Bernhard Thalheim. Armstrong databases and reasoning for functional dependencies and cardinality constraints over partial bags. Lecture Notes in Computer Science, 7153:164–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28472-4_10/.

Hegner:2012:FCU

[507] Stephen J. Hegner. FD covers and universal complements of simple projections. Lecture Notes in Computer Science, 7153:184–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28472-4_11/.

Katona:2012:EDS

[508] Gyula O. H. Katona and Krisztián Tichler. Encoding databases satisfying a given set of dependencies. Lecture Notes in Computer Science, 7153:203–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28472-4_12/.

Kramer:2012:LIR

[509] Annika Krämer and Christoph Beierle. On lifted inference for a relational probabilistic conditional logic with maximum entropy semantics. Lecture Notes in Computer Science, 7153:224–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28472-4_13/.

Molnar:2012:FED

[510] András J. Molnár, András A. Benczúr, and Csaba István Sidló. Flexible and efficient distributed resolution of large entities. Lecture Notes in Computer Science, 7153:244–263, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28472-4_14/.

Perovic:2012:RVE

[511] Aleksandar Perović, Dragan Doder, and Zoran Ognjanović. On realvalued evaluation of propositional formulas. Lecture Notes in Computer Science, 7153:264–277, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28472-4_15/.

Pivert:2012:DSA

[512] Olivier Pivert and Henri Prade. Detecting suspect answers in the presence of inconsistent information. Lecture Notes in Computer Science, 7153:278–297, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28472-4_16/.

Rajagopalan:2012:LNS

[513] Krishnan Rajagopalan, Venkatesh Srinivasan, and Alex Thomo. Learning the news in social networks. Lecture Notes in Computer Science, 7153:298–311, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28472-4_17/.

Rakib:2012:VRR

[514] Abdur Rakib, Rokan Uddin Faruqui, and Wendy MacCaull. Verifying resource requirements for ontology-driven rule-based agents. Lecture Notes in Computer Science, 7153:312–331, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28472-4_18/.

Rezk:2012:FPS

[515] Martín Rezk and Michael Kifer. Formalizing production systems with rule-based ontologies. Lecture Notes in Computer Science, 7153:332– 351, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28472-4_19/.

Sacca:2012:CCI

[516] Domenico Saccà, Edoardo Serra, and Antonella Guzzo. Count constraints and the inverse OLAP problem: Definition, complexity and a step toward aggregate data exchange. Lecture Notes in Computer Science, 7153:352–369, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28472-4_20/.

Anonymous:2012:FMz

[517] Anonymous. Front matter. Lecture Notes in Computer Science, 7153: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28472-4/1.

Asokan:2012:UMS

[518] N. Asokan and Cynthia Kuo. Usable mobile security. Lecture Notes in Computer Science, 7154:1–6, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28073-3_1/.

Pavlovic:2012:ANP

[519] Dusko Pavlovic and Catherine Meadows. Actor-network procedures. Lecture Notes in Computer Science, 7154:7–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28073-3_2/.

Moses:2012:KWD

[520] Yoram Moses. Knowledge as a window into distributed coordination. Lecture Notes in Computer Science, 7154:27–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28073-3_3/.

Parker:2012:DMO

[521] Lynne E. Parker. Decision making as optimization in multi-robot teams. Lecture Notes in Computer Science, 7154:35–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28073-3_4/.

Adhikari:2012:MNT

[522] Nibedita Adhikari and C. R. Tripathy. Mstar: a new two level interconnection network. Lecture Notes in Computer Science, 7154:50–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28073-3_5/.

Anudeep:2012:ISF

[523] C. Anudeep and Manik Lal Das. An improved scheme for false data filtering in wireless sensor networks. Lecture Notes in Computer Science, 7154:62–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28073-3_6/.

Bolla:2012:ASM

[524] Jhansi Vazram Bolla and Valli Kumari Vatsavayi. Anonymity and security in mobile ad hoc networks. Lecture Notes in Computer Science, 7154: 71–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28073-3_7/.

Dutta:2012:CFA

[525] Ayan Dutta and Sruti Gan Chaudhuri. Circle formation by asynchronous fat robots with limited visibility. Lecture Notes in Computer Science, 7154:83–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28073-3_8/.

Geetha:2012:HCC

[526] V. Geetha and N. Sreenath. High concurrency for continuously evolving OODBMS. Lecture Notes in Computer Science, 7154:94–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28073-3_9/.

Hanumanthappa:2012:NSI

[527] J. Hanumanthappa and D. H. Manjaiah. A new scheme for IPv6 BD-TTCS translator. Lecture Notes in Computer Science, 7154:106–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28073-3_10/.

Krishna:2012:NEW

[528] Shankara Narayanan Krishna. Nash equilibrium in weighted concurrent timed games with reachability objectives. Lecture Notes in Computer Science, 7154:117–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28073-3_11/.

Kumar:2012:PPM

[529] Tarun Kumar, Parikshit Sondhi, and Ankush Mittal. Parallelization of PageRank on multicore processors. Lecture Notes in Computer Science, 7154:129–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28073-3_12/.

Ma:2012:CIS

[530] Chun-Guang Ma, Ding Wang, and Qi-Ming Zhang. Cryptanalysis and improvement of Sood et al.’s dynamic ID-based authentication scheme. Lecture Notes in Computer Science, 7154:141–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28073-3_13/.

Mohanty:2012:ASD

[531] Hrushikesha Mohanty. An algebra of social distance. Lecture Notes in Computer Science, 7154:153–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28073-3_14/.

Mundra:2012:DFD

[532] Anil Mundra, Anish Mathuria, and Manik Lal Das. Detecting flaws in dynamic hierarchical key management schemes using specification animation. Lecture Notes in Computer Science, 7154:166–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28073-3_15/.

Panda:2012:SME

[533] Bhawani S. Panda and D. Pushparaj Shetty. Strong minimum energy minimum interference topology in wireless sensor networks. Lecture Notes in Computer Science, 7154:177–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28073-3_16/.

Pande:2012:DPI

[534] Arun Pande and Sanjay Kimbahune. Distributed processing and Internet technology to solve challenges of primary healthcare in India. Lecture Notes in Computer Science, 7154:188–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28073-3_17/.

Patel:2012:PFS

[535] Nidhi Patel and Sanjay Srivastava. Packet forwarding strategies for cooperation enforcement in mobile ad hoc wireless networks. Lecture Notes in Computer Science, 7154:200–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28073-3_18/.

Patibandla:2012:SSS

[536] R. S. M. Lakshmi Patibandla. A study on scalability of services and privacy issues in cloud computing. Lecture Notes in Computer Science, 7154:212–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28073-3_19/.

Rana:2012:RMH

[537] Chhavi Rana and Sanjay Kumar Jain. A recommendation model for handling dynamics in user profile. Lecture Notes in Computer Science, 7154:231–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28073-3_20/.

Anonymous:2012:FMba

[538] Anonymous. Front matter. Lecture Notes in Computer Science, 7154: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28073-3/1.

DiMartino:2012:CWC

[539] Beniamino Di Martino and Dana Petcu. CCPI 2011: Workshop on Cloud Computing Projects and Initiatives. Lecture Notes in Computer Science, 7155:1–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29737-3_1/.

Slawinska:2012:TCP

[540] Magdalena Slawinska, Jaroslaw Slawinski, and Vaidy Sunderam. Towards cross-platform cloud computing. Lecture Notes in Computer Science, 7155:5–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29737-3_2/.

Cicotti:2012:QMC

[541] Giuseppe Cicotti, Luigi Coppolino, Rosario Cristaldi, Salvatore D’Antonio, and Luigi Romano. QoS monitoring in a cloud services environment: The SRT-15 approach. Lecture Notes in Computer Science, 7155:15–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29737-3_3/.

Lezzi:2012:ESA

[542] Daniele Lezzi, Roger Rafanell, Abel Carrión, Ignacio Blanquer Espert, and Vicente Hernández. Enabling e-science applications on the cloud with COMPSs. Lecture Notes in Computer Science, 7155:25–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29737-3_4/.

Gogouvitis:2012:OVC

[543] Spyridon V. Gogouvitis, George Kousiouris, George Vafiadis, and Elliot K. Kolodner. OPTIMIS and VISION cloud: How to manage data in clouds. Lecture Notes in Computer Science, 7155:35–44, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29737-3_5/.

Palmieri:2012:IMI

[544] Roberto Palmieri, Pierangelo di Sanzo, Francesco Quaglia, and Paolo Romano. Integrated monitoring of infrastructures and applications in cloud environments. Lecture Notes in Computer Science, 7155:45–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29737-3_6/.

Benkner:2012:TCD

[545] Siegfried Benkner, Jesus Bisbal, Gerhard Engelbrecht, Rod D. Hose, and Yuriy Kaniovskyi. Towards collaborative data management in the VPH-share project. Lecture Notes in Computer Science, 7155:54–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29737-3_7/.

Heikkurinen:2012:SSC

[546] Matti Heikkurinen and Owen Appleton. SLM and SDM challenges in federated infrastructures. Lecture Notes in Computer Science, 7155:64– 72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29737-3_8/.

Haitof:2012:RPA

[547] Houssam Haitof. Rapid prototyping of architectures on the cloud using semantic resource description. Lecture Notes in Computer Science, 7155: 73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29737-3_9/.

Fortis:2012:CPM

[548] Teodor-Florin Fortiş, Gorka Esnal Lopez, Imanol Padillo Cruz, Gábor Ferschl, and Tamás Máhr. Cloud patterns for mOSAIC-Enabled scientific applications. Lecture Notes in Computer Science, 7155:83–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29737-3_10/.

Cuomo:2012:EAC

[549] A. Cuomo, M. Rak, S. Venticinque, and U. Villano. Enhancing an autonomic cloud architecture with mobile agents. Lecture Notes in Computer Science, 7155:94–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29737-3_11/.

Sun:2012:MAR

[550] Yih Leong Sun, Terence Harmer, Alan Stewart, and Peter Wright. Mapping application requirements to cloud resources. Lecture Notes in Computer Science, 7155:104–112, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29737-3_12/.

Danelutto:2012:CEW

[551] Marco Danelutto, Frédéric Desprez, Vladimir Getov, and Wolfgang Ziegler. CoreGRID/ERCIM Workshop on Grids, Clouds and P2P Computing — CGWS2011. Lecture Notes in Computer Science, 7155:113– 114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-29737-3_13.

Baude:2012:PCG

[552] Françoise Baude. A perspective on the CoreGRID grid component model. Lecture Notes in Computer Science, 7155:115–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-29737-3_14.

Klein:2012:TSE

[553] Cristian Klein and Christian Pérez. Towards scheduling evolving applications. Lecture Notes in Computer Science, 7155:117–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29737-3_15/.

Danelutto:2012:MCS

[554] Marco Danelutto, P. Kilpatrick, C. Montangero, and L. Semini. Model checking support for conflict resolution in multiple non-functional concern management. Lecture Notes in Computer Science, 7155:128–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29737-3_16/.

Bertolli:2012:CRP

[555] Carlo Bertolli, Gabriele Mencagli, and Marco Vanneschi. Consistent rollback protocols for autonomic ASSISTANT applications. Lecture Notes in Computer Science, 7155:139–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29737-3_17/.

Meilander:2012:DRM

[556] Dominik Meiländer, Alexander Ploss, Frank Glinka, and Sergei Gorlatch. A dynamic resource management system for real-time online applications on clouds. Lecture Notes in Computer Science, 7155:149– 158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29737-3_18/.

Carlini:2012:CFC

[557] Emanuele Carlini, Massimo Coppola, Patrizio Dazzi, Laura Ricci, and Giacomo Righetti. Cloud federations in contrail. Lecture Notes in Computer Science, 7155:159–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29737-3_19/.

Malik:2012:SAC

[558] Muhammad Junaid Malik, Thomas Fahringer, and Radu Prodan. Semiautomatic composition of ontologies for ASKALON grid workflows. Lecture Notes in Computer Science, 7155:169–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29737-3_20/;   http:
//link.springer.com/content/pdf/10.1007/978-3-642-29737-3_20.

Rajcsanyi:2012:CMI

[559] Vilmos Rajcsányi and Zsolt Németh. The Chemical Machine: an interpreter for the higher order chemical language. Lecture Notes in Computer Science, 7155:181–190, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_21.

Bertolli:2012:DPO

[560] Carlo Bertolli, Adam Betts, Gihan Mudalige, Mike Giles, and Paul Kelly. Design and performance of the OP2 library for unstructured mesh applications. Lecture Notes in Computer Science, 7155:191–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29737-3_22.

Tlili:2012:MAR

[561] Raja Tlili and Yahya Slimani. Mining association rules on grid platforms. Lecture Notes in Computer Science, 7155:201–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_23.

Scott:2012:WSL

[562] Stephen L. Scott, Geoffroy Vallée, and Thomas Naughton. 5th Workshop on System-Level Virtualization for High Performance Computing (HPCVirt 2011). Lecture Notes in Computer Science, 7155:211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29737-3_24.

Ali:2012:PEH

[563] Qasim Ali, Vladimir Kiriansky, Josh Simons, and Puneet Zaroo. Performance evaluation of HPC benchmarks on VMware’s ESXi server. Lecture Notes in Computer Science, 7155:213–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29737-3_25.

Serebrin:2012:VPC

[564] Benjamin Serebrin and Daniel Hecht. Virtualizing performance counters. Lecture Notes in Computer Science, 7155:223–233, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_26.

Naughton:2012:CVM

[565] Thomas Naughton, Geoffroy Vallée, Christian Engelmann, and Stephen L. Scott. A case for virtual machine based fault injection in a highperformance computing environment. Lecture Notes in Computer Science, 7155:234–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-29737-3_27.

Forsell:2012:HWH

[566] Martti Forsell and Jesper Larsson Träff. HPPC 2010: 5th Workshop on Highly Parallel Processing on a Chip. Lecture Notes in Computer Science, 7155:245–247, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29737-3_28.

Keceli:2012:TMM

[567] Fuat Keceli, Tali Moreshet, and Uzi Vishkin. Thermal management of a many-core processor under fine-grained parallelism. Lecture Notes in Computer Science, 7155:249–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_29.

Keir:2012:MPA

[568] Paul Keir, Paul W. Cockshott, and Andrew Richards. Mainstream parallel array programming on Cell. Lecture Notes in Computer Science, 7155:260–269, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29737-3_30.

Membarth:2012:GGC

[569] Richard Membarth, Anton Lokhmotov, and Jürgen Teich. Generating GPU code from a high-level representation for image processing kernels. Lecture Notes in Computer Science, 7155:270–280, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_31.

Xu:2012:GHA

[570] Thomas Canhao Xu, Pasi Liljeberg, and Hannu Tenhunen. A greedy heuristic approximation scheduling algorithm for 3D multicore processors. Lecture Notes in Computer Science, 7155:281–291, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_32.

Corsaro:2012:APT

[571] Stefania Corsaro, Pasqua D’Ambra, and Francesca Perla. Algorithms and programming tools for next-generation high-performance scientific software HPSS 2011. Lecture Notes in Computer Science, 7155:293–294, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29737-3_33.

Duff:2012:EES

[572] Iain S. Duff. European Exascale Software Initiative: Numerical libraries, solvers and algorithms. Lecture Notes in Computer Science, 7155:295–304, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29737-3_34.

Aktulga:2012:ROL

[573] Hasan Metin Aktulga, Chao Yang, Ümit V. Çatalyürek, Pieter Maris, and James P. Vary. On reducing I/O overheads in large-scale invariant subspace projections. Lecture Notes in Computer Science, 7155:305–314, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29737-3_35.

Baker:2012:ENG

[574] Chris Baker, Erik Boman, Mike Heroux, Eric Keiter, Siva Rajamanickam, and Rich Schiek. Enabling next-generation parallel circuit simulation with Trilinos. Lecture Notes in Computer Science, 7155:315–323, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29737-3_36.

Berzins:2012:DBS

[575] Martin Berzins, Qingyu Meng, John Schmidt, and James C. Sutherland. DAG-based software frameworks for PDEs. Lecture Notes in Computer Science, 7155:324–333, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_37.

Kaya:2012:PPC

[576] Kamer Kaya, François-Henry Rouet, and Bora Uçar. On partitioning problems with complex objectives. Lecture Notes in Computer Science, 7155:334–344, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29737-3_38.

Yamazaki:2012:CAT

[577] Ichitaro Yamazaki and Kesheng Wu. A communication-avoiding thickrestart Lanczos method on a distributed-memory system. Lecture Notes in Computer Science, 7155:345–354, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_39.

Hupca:2012:SHT

[578] Ioan Ovidiu Hupca, Joel Falcou, Laura Grigori, and Radek Stompor. Spherical harmonic transform with GPUs. Lecture Notes in Computer Science, 7155:355–366, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/content/
pdf/10.1007/978-3-642-29737-3_40.

Barbieri:2012:DPS

[579] Davide Barbieri, Valeria Cardellini, Salvatore Filippone, and Damian Rouson. Design patterns for scientific computations on sparse matrices. Lecture Notes in Computer Science, 7155:367–376, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_41.

Sorensen:2012:HPM

[580] Hans Henrik Brandenborg Sørensen. High-performance matrix-vector multiplication on the GPU. Lecture Notes in Computer Science, 7155: 377–386, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29737-3_42.

Gustedt:2012:RSO

[581] Jens Gustedt and Emmanuel Jeanvoine. Relaxed synchronization with ordered read-write locks. Lecture Notes in Computer Science, 7155: 387–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29737-3_43.

Prudencio:2012:PCS

[582] Ernesto E. Prudencio and Karl W. Schulz. The Parallel C++ Statistical Library ‘QUESO’: Quantification of uncertainty for estimation, simulation and optimization. Lecture Notes in Computer Science, 7155:398–407, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29737-3_44.

Dunnweber:2012:UHT

[583] Jan Dünnweber, Valentin Mihaylov, René Glettler, Volker Maiborn, and Holger Wolff. Use of HPC-techniques for large-scale data migration. Lecture Notes in Computer Science, 7155:408–415, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29737-3_45.

Bosilca:2012:AMT

[584] George Bosilca. Algorithms, models and tools for parallel computing on heterogeneous platforms (HeteroPar 2011). Lecture Notes in Computer Science, 7155:417, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_46.

Nicod:2012:GAC

[585] Jean-Marc Nicod, Laurent Philippe, and Lamiel Toch. A genetic algorithm with communication costs to schedule workflows on a SOA-Grid. Lecture Notes in Computer Science, 7155:419–428, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_47.

Lee:2012:EXP

[586] Jinpil Lee, Minh Tuan Tran, Tetsuya Odajima, Taisuke Boku, and Mitsuhisa Sato. An extension of XcalableMP PGAS lanaguage for multinode GPU clusters. Lecture Notes in Computer Science, 7155:429–439, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29737-3_48.

Arabnejad:2012:PEL

[587] Hamid Arabnejad and Jorge G. Barbosa. Performance evaluation of list based scheduling on heterogeneous systems. Lecture Notes in Computer Science, 7155:440–449, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_49.

Clarke:2012:CBM

[588] David Clarke, Alexey Lastovetsky, and Vladimir Rychkov. Column-based matrix partitioning for parallel matrix multiplication on heterogeneous processors based on functional performance models. Lecture Notes in Computer Science, 7155:450–459, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_50.

Cordasco:2012:FDA

[589] Gennaro Cordasco, Rosario De Chiara, Ada Mancuso, Dario Mazzeo, and Vittorio Scarano. A framework for distributing agent-based simulations. Lecture Notes in Computer Science, 7155:460–470, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_51.

Bahi:2012:PSL

[590] Jacques M. Bahi, Raphaël Couturier, and Lilia Ziane Khodja. Parallel sparse linear solver GMRES for GPU clusters with compression of exchanged data. Lecture Notes in Computer Science, 7155:471–480, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29737-3_52.

Blazewicz:2012:TDD

[591] Marek Błażewicz, Miłosz Ciżnicki, Piotr Kopta, Krzysztof Kurowski, and Paweł Lichocki. Two-dimensional discrete wavelet transform on large images for hybrid computing architectures: GPU and CELL. Lecture Notes in Computer Science, 7155:481–490, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_53.

Ilic:2012:SDL

[592] Aleksandar Ilic and Leonel Sousa. Scheduling divisible loads on heterogeneous desktop systems with limited memory. Lecture Notes in Computer Science, 7155:491–501, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29737-3_54.

Sanchez:2012:PGF

[593] Ma. Guadalupe Sánchez, Vicente Vidal, and Jordi Bataller. Peer group and fuzzy metric to remove noise in images using heterogeneous computing. Lecture Notes in Computer Science, 7155:502–510, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_55.

Nandagudi:2012:EMA

[594] Girish Nandagudi, Jaspal Subhlok, Edgar Gabriel, and Judit Gimenez. Estimation of MPI application performance on volunteer environments. Lecture Notes in Computer Science, 7155:511–520, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29737-3_56.

Anonymous:2012:BMl

[595] Anonymous. Back matter. Lecture Notes in Computer Science, 7155: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29737-3/1.

Anonymous:2012:FMbb

[596] Anonymous. Front matter. Lecture Notes in Computer Science, 7155: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29737-3/1.

Cannataro:2012:HWH

[597] Mario Cannataro. HiBB 2011: 2^{nd} Workshop on High Performance Bioinformatics and Biomedicine. Lecture Notes in Computer Science, 7156:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-29740-3_1.

Aldinucci:2012:PLS

[598] Marco Aldinucci, Mario Coppo, Ferruccio Damiani, Maurizio Drocco, and Eva Sciacca. On parallelizing on-line statistics for stochastic biological simulations. Lecture Notes in Computer Science, 7156:3–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29740-3_2/.

Rheinlander:2012:SSS

[599] Astrid Rheinländer and Ulf Leser. Scalable sequence similarity search and join in main memory on multi-cores. Lecture Notes in Computer Science, 7156:13–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29740-3_3/.

Mehta:2012:EDC

[600] Gaurang Mehta, Ewa Deelman, James A. Knowles, Ting Chen, Ying Wang, and Jens Vöckler. Enabling data and compute intensive workflows in bioinformatics. Lecture Notes in Computer Science, 7156:23– 32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29740-3_4/.

Grasso:2012:HAH

[601] Luigi Grasso, Nuria Medina-Medina, Rosana Montes-Soldado, and María M. Abad-Grau. Homogenizing access to highly time-consuming biomedical applications through a Web-based interface. Lecture Notes in Computer Science, 7156:33–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29740-3_5/.

Cannataro:2012:DMA

[602] Mario Cannataro and Pietro Hiram Guzzi. Distributed management and analysis of omics data. Lecture Notes in Computer Science, 7156: 43–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29740-3_6/.

Schaaf:2012:MDG

[603] Thomas Schaaf, Adam S. Z. Belloum, Owen Appleton, and Joan Serrat-Fernández. Managing and delivering grid services (MDGS). Lecture Notes in Computer Science, 7156:53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-29740-3_7.

Mathieu:2012:RAF

[604] Gilles Mathieu and Hélène Cordier. Resource allocation for the French national grid initiative. Lecture Notes in Computer Science, 7156:55– 63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29740-3_8/.

Szepieniec:2012:ISLb

[605] Tomasz Szepieniec, Joanna Kocot, Thomas Schaaf, Owen Appleton, and Matti Heikkurinen. On importance of service level management in grids. Lecture Notes in Computer Science, 7156:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29740-3_9/.

Balis:2012:LMS

[606] Bartosz Balis, Renata Slota, Jacek Kitowski, and Marian Bubak. Online monitoring of service-level agreements in the grid. Lecture Notes in Computer Science, 7156:76–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29740-3_10/.

Petcu:2012:CFI

[607] Dana Petcu. Challenges of future e-infrastructure governance. Lecture Notes in Computer Science, 7156:86–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29740-3_11/.

Galizia:2012:IBP

[608] Antonella Galizia, Alfonso Quarati, Michael Schiffers, and Mark Yampolskiy. Influences between performance based scheduling and service level agreements. Lecture Notes in Computer Science, 7156:96–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29740-3_12/.

Rak:2012:UCS

[609] Massimiliano Rak, Rocco Aversa, Salvatore Venticinque, and Beniamino Di Martino. User centric service level management in mOSAIC applications. Lecture Notes in Computer Science, 7156:106–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29740-3_13/.

Cushing:2012:SLM

[610] Reginald Cushing, Spiros Koulouzis, Rudolf Strijkers, Adam S. Z. Belloum, and Marian Bubak. Service level management for executable papers. Lecture Notes in Computer Science, 7156:116–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29740-3_14/.

Knittl:2012:CMI

[611] Silvia Knittl, Thomas Schaaf, and Ilya Saverchenko. Change management in e-infrastructures to support service level agreements. Lecture Notes in Computer Science, 7156:124–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29740-3_15/.

Gerndt:2012:PFW

[612] Michael Gerndt. PROPER 2011: Fourth Workshop on Productivity and Performance Tools for HPC Application Development. Lecture Notes in Computer Science, 7156:135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29740-3_16.

Krzikalla:2012:SSS

[613] Olaf Krzikalla, Kim Feldhoff, Ralph Müller-Pfefferkorn, and Wolfgang E. Nagel. Scout: a source-to-source transformator for SIMD-optimizations. Lecture Notes in Computer Science, 7156:137–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29740-3_17/.

Oleynik:2012:SAP

[614] Yury Oleynik and Michael Gerndt. Scalable automatic performance analysis on IBM BlueGene/P systems. Lecture Notes in Computer Science, 7156:146–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29740-3_18/.

Spear:2012:ACP

[615] Wyatt Spear, Allen D. Malony, Chee Wai Lee, Scott Biersdorff, and Sameer Shende. An approach to creating performance visualizations in a parallel profile analysis tool. Lecture Notes in Computer Science, 7156:156–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29740-3_19/.

Dandapanthula:2012:ISI

[616] N. Dandapanthula, H. Subramoni, J. Vienne, K. Kandalla, S. Sur, and Dhabaleswar K. Panda. INAM — a scalable InfiniBand network analysis and monitoring tool. Lecture Notes in Computer Science, 7156:166– 177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29740-3_20/.

Tiwari:2012:ATE

[617] Ananta Tiwari, Michael A. Laurenzano, Laura Carrington, and Allan Snavely. Auto-tuning for energy usage in scientific applications. Lecture Notes in Computer Science, 7156:178–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29740-3_21.

Cantiello:2012:ASC

[618] Pasquale Cantiello and Beniamino Di Martino. Automatic source code transformation for GPUs based on program comprehension. Lecture Notes in Computer Science, 7156:188–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29740-3_22.

Iwainsky:2012:EBP

[619] Christian Iwainsky, Ralph Altenfeld, Dieter an Mey, and Christian Bischof. Enhancing brainware productivity through a performance tuning workflow. Lecture Notes in Computer Science, 7156:198–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29740-3_23.

Scott:2012:WRH

[620] Stephen L. Scott and Chokchai (Box) Leangsuksun. Workshop on resiliency in high performance computing (resilience) in clusters, clouds, and grids. Lecture Notes in Computer Science, 7156:209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29740-3_24.

Kovatch:2012:MCU

[621] Patricia Kovatch, Matthew Ezell, and Ryan Braby. The Malthusian Catastrophe is upon us! Are the largest HPC machines ever up? Lecture Notes in Computer Science, 7156:211–220, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29740-3_25.

Riesen:2012:SAR

[622] Rolf Riesen, Kurt B. Ferreira, Maria Ruiz Varela, Michela Taufer, and Arun Rodrigues. Simulating application resilience at Exascale. Lecture Notes in Computer Science, 7156:221–230, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29740-3_26.

Brandt:2012:FES

[623] J. Brandt, F. Chen, A. Gentile, Chokchai (Box) Leangsuksun, J. Mayo, P. Pebay, and D. Roe. Framework for enabling system understanding. Lecture Notes in Computer Science, 7156:231–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29740-3_27.

Bridges:2012:CAD

[624] Patrick G. Bridges, Mark Hoemmen, Kurt B. Ferreira, Michael A. Heroux, and Philip Soltero. Cooperative application/OS DRAM fault recovery. Lecture Notes in Computer Science, 7156:241–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29740-3_28.

Fiala:2012:TSB

[625] David Fiala, Kurt B. Ferreira, Frank Mueller, and Christian Engelmann. A tunable, software-based DRAM error detection and correction library for HPC. Lecture Notes in Computer Science, 7156:251–261, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29740-3_29.

Sancho:2012:RIS

[626] José Carlos Sancho, Ana Jokanovic, and Jesus Labarta. Reducing the impact of soft errors on fabric-based collective communications. Lecture Notes in Computer Science, 7156:262–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29740-3_30.

Ma:2012:EAV

[627] Zhe Ma, Trevor Carlson, Wim Heirman, and Lieven Eeckhout. Evaluating application vulnerability to soft errors in multi-level cache hierarchy. Lecture Notes in Computer Science, 7156:272–281, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29740-3_31.

DeBardeleben:2012:EFI

[628] Nathan DeBardeleben, Sean Blanchard, Qiang Guan, Ziming Zhang, and Song Fu. Experimental framework for injecting logic errors in a virtual machine to profile applications for soft error resilience. Lecture Notes in Computer Science, 7156:282–291, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29740-3_32.

Thanakornworakij:2012:HAC

[629] Thanadech Thanakornworakij, Rajan Sharma, Blaine Scroggs, and Chokchai (Box) Leangsuksun. High availability on cloud with HA-OSCAR. Lecture Notes in Computer Science, 7156:292–301, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29740-3_33.

Ibtesham:2012:VCC

[630] Dewan Ibtesham, Dorian Arnold, Kurt B. Ferreira, and Patrick G. Bridges. On the viability of checkpoint compression for extreme scale fault tolerance. Lecture Notes in Computer Science, 7156:302–311, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29740-3_34.

Rajachandrasekar:2012:CCR

[631] Raghunath Rajachandrasekar, Xiangyong Ouyang, Xavier Besseron, and Vilobh Meshram. Can checkpoint/restart mechanisms benefit from hierarchical data staging? Lecture Notes in Computer Science, 7156:312–321, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29740-3_35.

Besseron:2012:IDC

[632] Xavier Besseron and Thierry Gautier. Impact of over-decomposition on coordinated checkpoint/rollback protocol. Lecture Notes in Computer Science, 7156:322–332, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29740-3_36.

Hast:2012:UFW

[633] Anders Hast, Josef Weidendorfer, and Jan-Philipp Weiss. UCHPC 2011: Fourth Workshop on UnConventional High Performance Computing. Lecture Notes in Computer Science, 7156:333–334, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29740-3_37.

Horikawa:2012:PPA

[634] Tetsuro Horikawa, Michio Honda, Jin Nakazawa, Kazunori Takashio, and Hideyuki Tokuda. PACUE: Processor Allocator Considering User Experience. Lecture Notes in Computer Science, 7156:335–344, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29740-3_38.

Murarasu:2012:WBH

[635] Alin Muraraşu, Josef Weidendorfer, and Arndt Bode. Workload balancing on heterogeneous systems: a case study of sparse grid interpolation. Lecture Notes in Computer Science, 7156:345–354, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29740-3_39.

Cabel:2012:PEM

[636] Tristan Cabel, Joseph Charles, and Stéphane Lanteri. Performance evaluation of a multi-GPU enabled finite element method for computational electromagnetics. Lecture Notes in Computer Science, 7156:355–364, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29740-3_40.

Xu:2012:SHB

[637] Thomas Canhao Xu, Pasi Liljeberg, and Hannu Tenhunen. Study of hierarchical N-body methods for network-on-chip architectures. Lecture Notes in Computer Science, 7156:365–374, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29740-3_41.

Heinecke:2012:EHP

[638] Alexander Heinecke, Michael Klemm, Dirk Pflüger, Arndt Bode, and Hans-Joachim Bungartz. Extending a highly parallel data mining algorithm to the Intel◯^{R} many integrated core architecture. Lecture Notes in Computer Science, 7156:375–384, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29740-3_42.

Alexander:2012:VWV

[639] Michael Alexander and Gianluigi Zanetti. VHPC 2011: 6th Workshop on Virtualization in High-Performance Cloud Computing. Lecture Notes in Computer Science, 7156:385, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29740-3_43.

Kim:2012:GBM

[640] Sangwook Kim, Hwanju Kim, and Joonwon Lee. Group-based memory deduplication for virtualized clouds. Lecture Notes in Computer Science, 7156:387–397, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29740-3_44.

Nanos:2012:SHI

[641] Anastassios Nanos, Nikos Nikoleris, Stratos Psomadakis, and Elisavet Kozyri. A smart HPC interconnect for clusters of virtual machines. Lecture Notes in Computer Science, 7156:398–406, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29740-3_45.

Aragiorgis:2012:CSP

[642] Dimitris Aragiorgis, Anastassios Nanos, and Nectarios Koziris. Coexisting scheduling policies boosting I/O virtual machines. Lecture Notes in Computer Science, 7156:407–415, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29740-3_46.

Briffaut:2012:PVA

[643] J. Briffaut, E. Lefebvre, J. Rouzaud-Cornabas, and C. Toinard. PIGA-Virt: an advanced distributed MAC protection of virtual systems. Lecture Notes in Computer Science, 7156:416–425, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/content/pdf/10.1007/978-3-642-29740-3_47.

Costache:2012:EAA

[644] Stefania Costache, Nikos Parlavantzas, Christine Morin, and Samuel Kortas. An economic approach for application QoS management in clouds. Lecture Notes in Computer Science, 7156:426–435, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/978-3-642-29740-3_48.

Luszczek:2012:EHC

[645] Piotr Luszczek, Eric Meek, Shirley Moore, Dan Terpstra, and Vincent M. Weaver. Evaluation of the HPC challenge benchmarks in virtualized environments. Lecture Notes in Computer Science, 7156:436–445, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29740-3_49.

Lebre:2012:DBC

[646] Adrien Lèbre, Paolo Anedda, Massimo Gaggero, and Flavien Quesnel. DISCOVERY, beyond the clouds. Lecture Notes in Computer Science, 7156:446–456, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-

3349 (electronic). URL http://link.springer.com/content/pdf/10.
1007/978-3-642-29740-3_50.

Quesnel:2012:CDS

[647] Flavien Quesnel and Adrien Lèbre. Cooperative dynamic scheduling of virtual machines in distributed systems. Lecture Notes in Computer Science, 7156:457–466, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

content/pdf/10.1007/978-3-642-29740-3_51.

Kienzler:2012:LSD

[648] Romeo Kienzler, Rémy Bruggmann, Anand Ranganathan, and Nesime Tatbul. Large-scale DNA sequence analysis in the cloud: a streambased approach. Lecture Notes in Computer Science, 7156:467–476, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer.com/content/pdf/10.1007/978-3-642-
29740-3_52.

Anonymous:2012:BMm

[649] Anonymous. Back matter. Lecture Notes in Computer Science, 7156: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29740-3/1.

Anonymous:2012:FMbc

[650] Anonymous. Front matter. Lecture Notes in Computer Science, 7156: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29740-3/1.

Widmayer:2012:CON

[651] Peter Widmayer. Combinatorial optimization with noisy inputs: How can we separate the wheat from the chaff? Lecture Notes in Computer Science, 7157:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-28076-4_1.

Iwama:2012:ASM

[652] Kazuo Iwama. Approximability of stable matching problems. Lecture Notes in Computer Science, 7157:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-28076-4_2.

Whitesides:2012:TDG

[653] Sue Whitesides. On three-dimensional graph drawing and embedding. Lecture Notes in Computer Science, 7157:3–4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-28076-4_3.

Uno:2012:BNR

[654] Takeaki Uno, Ryuhei Uehara, and Shin ichi Nakano. Bounding the number of reduced trees, cographs, and series-parallel graphs by compression. Lecture Notes in Computer Science, 7157:5–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28076-4_4/.

Krithika:2012:GAG

[655] R. Krithika and N. S. Narayanaswamy. Generalized above guarantee vertex cover and r-partization. Lecture Notes in Computer Science, 7157: 17–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28076-4_5/.

Bae:2012:FVD

[656] Sang Won Bae and Kyung-Yong Chwa. Farthest Voronoi diagrams under travel time metrics. Lecture Notes in Computer Science, 7157:28– 39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28076-4_6/.

Bae:2012:TBF

[657] Sang Won Bae. Tight bound for farthest-color Voronoi diagrams of line segments. Lecture Notes in Computer Science, 7157:40–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28076-4_7/.

Das:2012:RAM

[658] Ananda Swarup Das, Prosenjit Gupta, Anil Kishore Kalavagattu, and Jatin Agarwal. Range aggregate maximal points in the plane. Lecture Notes in Computer Science, 7157:52–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28076-4_8/.

Dokka:2012:AML

[659] Trivikram Dokka, Anastasia Kouvela, and Frits C. R. Spieksma. Approximating the multi-level bottleneck assignment problem. Lecture Notes in Computer Science, 7157:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28076-4_9/.

Boria:2012:RMW

[660] Nicolas Boria, Jérôme Monnot, and Vangelis Th. Paschos. Reoptimization of the maximum weighted P_{k}-free subgraph problem under vertex insertion. Lecture Notes in Computer Science, 7157:76–87, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28076-4_10/.

Brandenburg:2012:CAP

[661] Franz J. Brandenburg, Andreas Gleißner, and Andreas Hofmeier. Comparing and aggregating partial orders with Kendall Tau distances. Lecture Notes in Computer Science, 7157:88–99, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28076-4_11/.

Wang:2012:RTC

[662] Biing-Feng Wang, Jhih-Hong Ye, and Pei-Jung Chen. On the roundtrip 1-center and 1-median problems. Lecture Notes in Computer Science, 7157:100–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28076-4_12/.

Salma:2012:TFO

[663] Sammi Abida Salma and Md. Saidur Rahman. Triangle-free outerplanar 3-graphs are pairwise compatibility graphs. Lecture Notes in Computer Science, 7157:112–123, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28076-4_13/.

Calamoneri:2012:RCP

[664] Tiziana Calamoneri, Rossella Petreschi, and Blerina Sinaimeri. On relaxing the constraints in pairwise compatibility graphs. Lecture Notes in Computer Science, 7157:124–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28076-4_14/.

Hossain:2012:ULS

[665] Md. Iqbal Hossain, Debajyoti Mondal, Md. Saidur Rahman, and Sammi Abida Salma. Universal line-sets for drawing planar 3-trees. Lecture Notes in Computer Science, 7157:136–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28076-4_15/.

Durocher:2012:HPS

[666] Stephane Durocher and Debajyoti Mondal. On the hardness of pointset embeddability. Lecture Notes in Computer Science, 7157:148–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28076-4_16/.

Moosa:2012:LTI

[667] Tanaeem M. Moosa, Sumaiya Nazeen, M. Sohel Rahman, and Rezwana Reaz. Linear time inference of strings from cover arrays using a binary alphabet. Lecture Notes in Computer Science, 7157:160–172, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28076-4_17/.

Elmasry:2012:FHR

[668] Amr Elmasry and Jyrki Katajainen. Fat heaps without regular counters. Lecture Notes in Computer Science, 7157:173–185, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28076-4_18/.

Fink:2012:DGV

[669] Martin Fink, Jan-Henrik Haunert, Tamara Mchedlidze, Joachim Spoerhase, and Alexander Wolff. Drawing graphs with vertices at specified positions and crossings at large angles. Lecture Notes in Computer Science, 7157:186–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28076-4_19/.

Bachmaier:2012:DUT

[670] Christian Bachmaier and Marco Matzeder. Drawing unordered trees on k-grids. Lecture Notes in Computer Science, 7157:198–210, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-28076-4_
20/.

Anonymous:2012:FMbd

[671] Anonymous. Front matter. Lecture Notes in Computer Science, 7157: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28076-4/1.

Unterschutz:2012:ORI

[672] Stefan Unterschütz, Christian Renner, and Volker Turau. Opportunistic, receiver-initiated data-collection protocol. Lecture Notes in Computer Science, 7158:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28169-3_1/.

Hurni:2012:TPO

[673] Philipp Hurni, Ulrich Bürgi, Markus Anwander, and Torsten Braun. TCP performance optimizations for wireless sensor networks. Lecture Notes in Computer Science, 7158:17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28169-3_2/.

Klonowski:2012:ISA

[674] Marek Klonowski and Kamil Wolny. Immune size approximation algorithms in Ad Hoc radio network. Lecture Notes in Computer Science, 7158:33–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28169-3_3/.

Wang:2012:FSK

[675] Peng Wang and Chinya Ravishankar. Foisting and stealing of keys in sensor networks. Lecture Notes in Computer Science, 7158:49–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28169-3_4/.

Leontiadis:2012:STS

[676] Ilias Leontiadis, Christos Efstratiou, Cecilia Mascolo, and Jon Crowcroft. SenShare: Transforming sensor networks into multi-application sensing infrastructures. Lecture Notes in Computer Science, 7158:65–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28169-3_5/.

Haas:2012:RSE

[677] Christian Haas, Joachim Wilke, and Viktor Stöhr. Realistic simulation of energy consumption in wireless sensor networks. Lecture Notes in Computer Science, 7158:82–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28169-3_6/.

Ko:2012:LPH

[678] JeongGil Ko, Kevin Klues, Christian Richter, Wanja Hofer, and Branislav Kusy. Low power or high performance? A tradeoff whose time has come (and nearly gone). Lecture Notes in Computer Science, 7158:98–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28169-3_7/.

Sitanayah:2012:FTR

[679] Lanny Sitanayah, Kenneth N. Brown, and Cormac J. Sreenan. Faulttolerant relay deployment based on length-constrained connectivity and rerouting centrality in wireless sensor networks. Lecture Notes in Computer Science, 7158:115–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28169-3_8/.

Fotouhi:2012:SHR

[680] Hossein Fotouhi, Marco Zuniga, Mário Alves, Anis Koubaa, and Pedro Marrón. Smart-HOP: a reliable handoff mechanism for mobile wireless sensor networks. Lecture Notes in Computer Science, 7158:131– 146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28169-3_9/.

Fabbri:2012:OBT

[681] Flavio Fabbri, Marco Zuniga, Daniele Puccinelli, and Pedro Marrón. On the optimal blacklisting threshold for link selection in wireless sensor networks. Lecture Notes in Computer Science, 7158:147–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28169-3_10/.

Pettinato:2012:MCT

[682] Paolo Pettinato, Niklas Wirström, Joakim Eriksson, and Thiemo Voigt. Multi-channel two-way time of flight sensor network ranging. Lecture Notes in Computer Science, 7158:163–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28169-3_11/.

Buchli:2012:GEW

[683] Bernhard Buchli, Felix Sutton, and Jan Beutel. GPS-equipped wireless sensor network node for high-accuracy positioning applications. Lecture Notes in Computer Science, 7158:179–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28169-3_12/.

Alessandrelli:2012:SSC

[684] Daniele Alessandrelli, Andrea Azzarà, Matteo Petracca, Christian Nastasi, and Paolo Pagano. ScanTraffic: Smart camera network for traffic information collection. Lecture Notes in Computer Science, 7158:196– 211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28169-3_13/.

Somov:2012:TES

[685] Andrey Somov, Christine C. Ho, Roberto Passerone, James W. Evans, and Paul K. Wright. Towards extending sensor node lifetime with printed supercapacitors. Lecture Notes in Computer Science, 7158:212– 227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28169-3_14/.

Hasenfratz:2012:FCL

[686] David Hasenfratz, Olga Saukh, and Lothar Thiele. On-the-fly calibration of low-cost gas sensors. Lecture Notes in Computer Science, 7158:228– 244, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28169-3_15/.

Somov:2012:EAG

[687] Andrey Somov, Alexander Baranov, Alexey Savkin, Mikhail Ivanov, and Lucia Calliari. Energy-aware gas sensing using wireless sensor networks. Lecture Notes in Computer Science, 7158:245–260, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28169-3_16/.

Anonymous:2012:BMn

[688] Anonymous. Back matter. Lecture Notes in Computer Science, 7158: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28169-3/1.

Anonymous:2012:FMbe

[689] Anonymous. Front matter. Lecture Notes in Computer Science, 7158: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28169-3/1.

Barbu:2012:ARA

[690] Guillaume Barbu and Philippe Hoogvorst. Application-replay attack on Java cards: When the garbage collector gets confused. Lecture Notes in Computer Science, 7159:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28166-2_1/.

Beckers:2012:SDD

[691] Kristian Beckers and Stephan Faßbender. Supporting the development and documentation of ISO 27001 information security management systems through security requirements engineering approaches. Lecture Notes in Computer Science, 7159:14–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28166-2_2/.

Breuer:2012:TAR

[692] Peter T. Breuer and Jonathan P. Bowen. Typed assembler for a RISC crypto-processor. Lecture Notes in Computer Science, 7159:22– 29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28166-2_3/.

Casalino:2012:TPC

[693] Matteo Maria Casalino and Henrik Plate. Transversal policy conflict detection. Lecture Notes in Computer Science, 7159:30–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28166-2_4/.

Gejibo:2012:CIE

[694] Samson Gejibo and Federico Mancini. Challenges in implementing an end-to-end secure protocol for Java ME-based mobile data collection in low-budget settings. Lecture Notes in Computer Science, 7159:38– 45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28166-2_5/.

Kovacs:2012:REI

[695] Máté Kovács and Helmut Seidl. Runtime enforcement of information flow security in tree manipulating processes. Lecture Notes in Computer Science, 7159:46–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28166-2_6/.

Masi:2012:FIX

[696] Massimiliano Masi and Rosario Pugliese. Formalisation and implementation of the XACML access control mechanism. Lecture Notes in Computer Science, 7159:60–74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28166-2_7/.

Moyano:2012:TOA

[697] Francisco Moyano and Carmen Fernandez-Gago. A task ordering approach for automatic trust establishment. Lecture Notes in Computer Science, 7159:75–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28166-2_8/.

Nguyen:2012:IIV

[698] Viet Hung Nguyen and Fabio Massacci. An idea of an independent validation of vulnerability discovery models. Lecture Notes in Computer Science, 7159:89–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28166-2_9/.

Ochoa:2012:SDP

[699] Martín Ochoa, Jan Jürjens, and Daniel Warzecha. A sound decision procedure for the compositionality of secrecy. Lecture Notes in Computer Science, 7159:97–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28166-2_10/.

Potharaju:2012:PSA

[700] Rahul Potharaju and Andrew Newell. Plagiarizing Smartphone applications: Attack strategies and defense techniques. Lecture Notes in Computer Science, 7159:106–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28166-2_11/.

Saadatmand:2012:DAS

[701] Mehrdad Saadatmand and Antonio Cicchetti. Design of adaptive security mechanisms for real-time embedded systems. Lecture Notes in Computer Science, 7159:121–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28166-2_12/.

Stergiopoulos:2012:HAL

[702] George Stergiopoulos and Bill Tsoumas. Hunting application-level logical errors. Lecture Notes in Computer Science, 7159:135–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28166-2_13/.

Zhu:2012:OTM

[703] Huafei Zhu and Hong Xiao. Optimal trust mining and computing on keyed MapReduce. Lecture Notes in Computer Science, 7159:143–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28166-2_14/.

Anonymous:2012:BMo

[704] Anonymous. Back matter. Lecture Notes in Computer Science, 7159: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28166-2/1.

Anonymous:2012:FMbf

[705] Anonymous. Front matter. Lecture Notes in Computer Science, 7159: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28166-2/1.

Marcus:2012:ARA

[706] Solomon Marcus. The art of reaching the age of sixty. Lecture Notes in Computer Science, 7160:1–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27654-5_1/.

Maurer:2012:CFO

[707] Hermann Maurer. Calude as father of one of the computer science journals. Lecture Notes in Computer Science, 7160:20–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27654-5_2/.

Bienvenu:2012:RSR

[708] Laurent Bienvenu and Alexander Shen. Random semicomputable reals revisited. Lecture Notes in Computer Science, 7160:31–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27654-5_3/.

Bridges:2012:CIT

[709] Douglas S. Bridges and Luminita S. Vîţă. Constructing the infimum of two projections. Lecture Notes in Computer Science, 7160:46–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27654-5_4/.

Brodhead:2012:BR

[710] Paul Brodhead, Rod Downey, and Keng Meng Ng. Bounded randomness. Lecture Notes in Computer Science, 7160:59–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27654-5_5/.

Campeanu:2012:NBS

[711] Cezar Câmpeanu. A note on Blum static complexity measures. Lecture Notes in Computer Science, 7160:71–80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-27654-5_6/.

Dinneen:2012:PSC

[712] Michael J. Dinneen. A program-size complexity measure for mathematical problems and conjectures. Lecture Notes in Computer Science, 7160: 81–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27654-5_7/.

Dumitrescu:2012:DRG

[713] Monica Dumitrescu. On degrees of randomness and genetic randomness. Lecture Notes in Computer Science, 7160:94–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27654-5_8/.

Freivalds:2012:HSC

[714] Rūsiņš Freivalds. Hartmanis–Stearns Conjecture on real time and transcendence. Lecture Notes in Computer Science, 7160:105–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27654-5_9/.

Gao:2012:LFC

[715] Ziyuan Gao, Frank Stephan, Guohua Wu, and Akihiro Yamamoto. Learning families of closed sets in matroids. Lecture Notes in Computer Science, 7160:120–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-27654-5_10/.

Jurgensen:2012:IUC

[716] Helmut Jürgensen. Invariance and universality of complexity. Lecture Notes in Computer Science, 7160:140–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27654-5_11/.

Kucera:2012:DPR

[717] Antonín Kučera and André Nies. Demuth’s path to randomness. Lecture Notes in Computer Science, 7160:159–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27654-5_12/.

Manin:2012:CCA

[718] Yuri I. Manin. A computability challenge: Asymptotic bounds for errorcorrecting codes. Lecture Notes in Computer Science, 7160:174–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27654-5_13/.

Patarin:2012:STG

[719] Jacques Patarin. Some transfinite generalisations of Gödel’s Incompleteness Theorem. Lecture Notes in Computer Science, 7160:183–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27654-5_14/.

Staiger:2012:OFC

[720] Ludwig Staiger. On oscillation-free Chaitin h-random sequences. Lecture Notes in Computer Science, 7160:194–202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27654-5_15/.

Tadaki:2012:PTB

[721] Kohtaro Tadaki. Phase transition between unidirectionality and bidirectionality. Lecture Notes in Computer Science, 7160:203–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-27654-5_16/.

Zenil:2012:CRL

[722] Hector Zenil. Computer runtimes and the length of proofs. Lecture Notes in Computer Science, 7160:224–240, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27654-5_17/.

Zimand:2012:SIC

[723] Marius Zimand. Symmetry of information: a closer look. Lecture Notes in Computer Science, 7160:241–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-27654-5_18/.

Chaitin:2012:HMI

[724] Gregory Chaitin. How much information can there be in a real number? Lecture Notes in Computer Science, 7160:247–251, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-27654-5_19/.

Cooper:2012:MMM

[725] S. Barry Cooper. Mathematics, metaphysics and the multiverse. Lecture Notes in Computer Science, 7160:252–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-27654-5_20/.

Anonymous:2012:FMbg

[726] Anonymous. Front matter. Lecture Notes in Computer Science, 7160: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-27654-5/1.

Ansper:2012:DGB

[727] Arne Ansper. Designing a governmental backbone. Lecture Notes in Computer Science, 7161:1–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29615-4_1/.

Vinkel:2012:IVE

[728] Priit Vinkel. Internet voting in Estonia. Lecture Notes in Computer Science, 7161:4–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29615-4_2/.

Beitollahi:2012:RBO

[729] Hakem Beitollahi and Geert Deconinck. A ring based onion circuit for hidden services. Lecture Notes in Computer Science, 7161:13–30, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29615-4_3/.

Boda:2012:UTW

[730] Károly Boda, Ádám Máté Földes, Gábor György Gulyás, and Sándor Imre. User tracking on the Web via cross-browser fingerprinting. Lecture Notes in Computer Science, 7161:31–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29615-4_4/.

Claes:2012:CSF

[731] Mathias Claes, Vincent van der Leest, and An Braeken. Comparison of SRAM and FF PUF in 65nm technology. Lecture Notes in Computer Science, 7161:47–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29615-4_5/.

Cucurull:2012:MAD

[732] Jordi Cucurull, Simin Nadjm-Tehrani, and Massimiliano Raciti. Modular anomaly detection for Smartphone ad hoc communication. Lecture Notes in Computer Science, 7161:65–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29615-4_6/.

Dossogne:2012:MVB

[733] Jérôme Dossogne and Frédéric Lafitte. Mental voting booths. Lecture Notes in Computer Science, 7161:82–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29615-4_7/.

Hamamoto:2012:MPP

[734] Kazutomo Hamamoto, Yasuyuki Tahara, and Akihiko Ohsuga. Methods for privacy protection considering status of service provider and user community. Lecture Notes in Computer Science, 7161:98–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29615-4_8/.

Helkala:2012:SMP

[735] Kirsi Helkala and Nils Kalstad Svendsen. The security and memorability of passwords generated by using an association element and a personal factor. Lecture Notes in Computer Science, 7161:114–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29615-4_9/.

Kani-Zabihi:2012:ISU

[736] Elahe Kani-Zabihi and Martin Helmhout. Increasing service users’ privacy awareness by introducing on-line interactive privacy features. Lecture Notes in Computer Science, 7161:131–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29615-4_10/.

Lemay:2012:OIR

[737] Frédérick Lemay, Raphaël Khoury, and Nadia Tawbi. Optimized inlining of runtime monitors. Lecture Notes in Computer Science, 7161:149– 161, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29615-4_11/.

Lupu:2012:IBK

[738] Radu Lupu, Eugen Borcoci, and Tinku Rasheed. Identity-based key derivation method for low delay inter-domain handover re-authentication service. Lecture Notes in Computer Science, 7161:162–175, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-29615-4_
12/.

Moonsamy:2012:FRS

[739] Veelasha Moonsamy, Ronghua Tian, and Lynn Batten. Feature reduction to speed up malware classification. Lecture Notes in Computer Science, 7161:176–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29615-4_13/.

Nazar:2012:RAE

[740] Assem Nazar, Mark M. Seeger, and Harald Baier. Rooting Android — extending the ADB by an auto-connecting WiFi-accessible service. Lecture Notes in Computer Science, 7161:189–204, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29615-4_14/.

Sen:2012:APP

[741] Jaydip Sen and Subhamoy Maitra. An attack on privacy preserving data aggregation protocol for wireless sensor networks. Lecture Notes in Computer Science, 7161:205–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29615-4_15/.

Stefan:2012:DCL

[742] Deian Stefan, Alejandro Russo, David Mazières, and John C. Mitchell. Disjunction category labels. Lecture Notes in Computer Science, 7161:223–239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29615-4_16/.

Tondel:2012:VCE

[743] Inger Anne Tøndel. Visualization control for event-based public display systems used in a hospital setting. Lecture Notes in Computer Science, 7161:240–255, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29615-4_17/.

Wenger:2012:EDS

[744] Erich Wenger and Michael Hutter. Exploring the design space of prime field vs. binary field ECC-hardware implementations. Lecture Notes in Computer Science, 7161:256–271, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29615-4_18/.

Anonymous:2012:BMp

[745] Anonymous. Back matter. Lecture Notes in Computer Science, 7161: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29615-4/1.

Anonymous:2012:FMbh

[746] Anonymous. Front matter. Lecture Notes in Computer Science, 7161: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29615-4/1.

Best:2012:PND

[747] Eike Best and Philippe Darondeau. Petri net distributability. Lecture Notes in Computer Science, 7162:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29709-0_1/.

Bruni:2012:CAP

[748] Roberto Bruni, Hernán Melgratti, and Ugo Montanari. Connector algebras, Petri nets, and BIP. Lecture Notes in Computer Science, 7162: 19–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29709-0_2/.

Buneman:2012:MP

[749] Peter Buneman. Models of provenance. Lecture Notes in Computer Science, 7162:39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29709-0_3.

Majumdar:2012:EEG

[750] Rupak Majumdar. End-to-end guarantees in embedded control systems. Lecture Notes in Computer Science, 7162:40, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-29709-0_4.

Zeller:2012:MPS

[751] Andreas Zeller. Mining precise specifications. Lecture Notes in Computer Science, 7162:41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29709-0_5.

Baar:2012:DEP

[752] Thomas Baar and Philipp Kumar. Detecting entry points in Java libraries. Lecture Notes in Computer Science, 7162:42–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29709-0_6/.

Beyer:2012:SAR

[753] Michael Beyer and Sabine Glesner. Static analysis of run-time modes in synchronous process network. Lecture Notes in Computer Science, 7162: 55–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29709-0_7/.

Bozhenkova:2012:CMC

[754] Elena Bozhenkova. Compositional methods in characterization of timed event structures. Lecture Notes in Computer Science, 7162:68–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29709-0_8/.

Caballero:2012:ADS

[755] Rafael Caballero and Yolanda García-Ruiz. Algorithmic debugging of SQL views. Lecture Notes in Computer Science, 7162:77–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29709-0_9/.

Dubtsov:2012:TTS

[756] Roman Dubtsov. Timed transition systems with independence and marked Scott domains: An adjunction. Lecture Notes in Computer Science, 7162:86–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29709-0_10/.

Fried:2012:ESE

[757] Michael A. H. Fried and Anna Fensel. An extensible system for enhancing social conference experience. Lecture Notes in Computer Science, 7162:95–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29709-0_11/.

Garanina:2012:EAM

[758] Natalia O. Garanina. Exponential acceleration of model checking for perfect recall systems. Lecture Notes in Computer Science, 7162:111– 124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29709-0_12/.

Gluck:2012:BCG

[759] Robert Glück. Bootstrapping compiler generators from partial evaluators. Lecture Notes in Computer Science, 7162:125–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29709-0_13/.

Gribovskaya:2012:LCT

[760] Natalya S. Gribovskaya. A logic characteristic for timed extensions of partial order based equivalences. Lecture Notes in Computer Science, 7162:142–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29709-0_14/.

Hamilton:2012:PCU

[761] Geoff W. Hamilton and Neil D. Jones. Proving the correctness of unfold/fold program transformations using bisimulation. Lecture Notes in Computer Science, 7162:153–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29709-0_15/.

Jaskelioff:2012:SME

[762] Mauro Jaskelioff and Alejandro Russo. Secure multi-execution in Haskell. Lecture Notes in Computer Science, 7162:170–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29709-0_16/.

Khoroshilov:2012:TOF

[763] Alexey Khoroshilov, Vadim Mutilin, and Eugene Novikov. Towards an open framework for C verification tools benchmarking. Lecture Notes in Computer Science, 7162:179–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29709-0_17/.

Klimov:2012:SCP

[764] Andrei V. Klimov. Solving coverability problem for monotonic counter systems by supercompilation. Lecture Notes in Computer Science, 7162:193–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29709-0_18/.

Klyuchnikov:2012:MRS

[765] Ilya Klyuchnikov and Sergei A. Romanenko. Multi-result supercompilation as branching growth of the penultimate level in metasystem transitions. Lecture Notes in Computer Science, 7162:210–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29709-0_19/.

Knoop:2012:SLB

[766] Jens Knoop, Laura Kovács, and Jakob Zwirchmayr. Symbolic loop bound computation for WCET analysis. Lecture Notes in Computer Science, 7162:227–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29709-0_20/.

Anonymous:2012:FMbi

[767] Anonymous. Front matter. Lecture Notes in Computer Science, 7162: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29709-0/1.

Kawai:2012:SHS

[768] Yutaka Kawai and Noboru Kunihiro. Secret handshake scheme with request-based-revealing. Lecture Notes in Computer Science, 7163:1– 16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29804-2_1/.

Gjosteen:2012:PBS

[769] Kristian Gjøsteen and Øystein Thuen. Password-based signatures. Lecture Notes in Computer Science, 7163:17–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29804-2_2/.

Lancrenon:2012:IPI

[770] Jean Lancrenon and Roland Gillard. Isolating partial information of indistinguishable encryptions. Lecture Notes in Computer Science, 7163: 34–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29804-2_3/.

Al-Sinani:2012:UCB

[771] Haitham S. Al-Sinani and Chris J. Mitchell. A universal client-based identity management tool. Lecture Notes in Computer Science, 7163:49– 74, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29804-2_4/.

Stegelmann:2012:DEP

[772] Mark Stegelmann and Dogan Kesdogan. Design and evaluation of a privacy-preserving architecture for vehicle-to-grid interaction. Lecture Notes in Computer Science, 7163:75–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29804-2_5/.

vanDeursen:2012:IAP

[773] Ton van Deursen and Saša Radomirović. Insider attacks and privacy of RFID protocols. Lecture Notes in Computer Science, 7163:91–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29804-2_6/.

Garcia:2012:CBR

[774] Flavio D. Garcia, Eric R. Verheul, and Bart Jacobs. Cell-based roadpricing. Lecture Notes in Computer Science, 7163:106–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29804-2_7/.

Pereira:2012:BAM

[775] Olivier Pereira. Ballot aggregation and Mixnet based open-audit elections. Lecture Notes in Computer Science, 7163:123–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29804-2_8/.

Gutmann:2012:PPI

[776] Peter Gutmann. PKI as part of an integrated risk management strategy for Web security. Lecture Notes in Computer Science, 7163:128– 146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29804-2_9/.

VanDamme:2012:PBM

[777] Gauthier Van Damme, Nicolas Luyckx, and Karel Wouters. A PKIbased mobile banking demonstrator. Lecture Notes in Computer Science, 7163:147–158, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29804-2_10/.

MBarka:2012:CVB

[778] Moez Ben M’Barka and Julien P. Stern. Certification validation: Back to the past. Lecture Notes in Computer Science, 7163:159–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29804-2_11/.

Winter:2012:HGL

[779] Johannes Winter and Kurt Dietrich. A hijacker’s guide to the LPC bus. Lecture Notes in Computer Science, 7163:176–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29804-2_12/.

Braeken:2012:SEL

[780] An Braeken, Antonio De La Piedro, and Karel Wouters. Secure event logging in sensor networks. Lecture Notes in Computer Science, 7163:194– 208, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29804-2_13/.

Anonymous:2012:BMq

[781] Anonymous. Back matter. Lecture Notes in Computer Science, 7163: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29804-2/1.

Anonymous:2012:FMbj

[782] Anonymous. Front matter. Lecture Notes in Computer Science, 7163: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29804-2/1.

Jansen:2012:AAS

[783] Klaus Jansen. Approximation algorithms for scheduling and packing problems. Lecture Notes in Computer Science, 7164:1–8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29116-6_1/.

Nutov:2012:ASC

[784] Zeev Nutov. Approximating subset k-connectivity problems. Lecture Notes in Computer Science, 7164:9–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29116-6_2/.

Marban:2012:LSM

[785] Sebastián Marbán, Cyriel Rutten, and Tjark Vredeveld. Learning in stochastic machine scheduling. Lecture Notes in Computer Science, 7164: 21–34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29116-6_3/.

Bein:2012:OAO

[786] Wolfgang Bein, Naoki Hatta, and Nelson Hernandez-Cons. An online algorithm optimally self-tuning to congestion for power management problems. Lecture Notes in Computer Science, 7164:35–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29116-6_4/.

Bazgan:2012:SAB

[787] Cristina Bazgan, Laurent Gourvès, and Jérôme Monnot. Single approximation for Biobjective Max TSP. Lecture Notes in Computer Science, 7164:49–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29116-6_5/.

Brankovic:2012:PAA

[788] Ljiljana Brankovic and Henning Fernau. Parameterized approximation algorithms for hitting set. Lecture Notes in Computer Science, 7164:63– 76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29116-6_6/.

Schwartges:2012:AAM

[789] Nadine Schwartges and Joachim Spoerhase. Approximation algorithms for the maximum leaf spanning tree problem on acyclic digraphs. Lecture Notes in Computer Science, 7164:77–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29116-6_7/.

Goetzmann:2012:OIR

[790] Kai-Simon Goetzmann and Sebastian Stiller. Optimization over integers with robustness in cost and few constraints. Lecture Notes in Computer Science, 7164:89–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29116-6_8/.

Ebenlendr:2012:LBD

[791] Tomáš Ebenlendr and Jiří Sgall. A lower bound on deterministic online algorithms for scheduling on related machines without preemption. Lecture Notes in Computer Science, 7164:102–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29116-6_9/.

Jansen:2012:SJI

[792] Klaus Jansen and Christina Robenek. Scheduling jobs on identical and uniform processors revisited. Lecture Notes in Computer Science, 7164:109–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29116-6_10/.

Shmoys:2012:AAF

[793] David B. Shmoys and Gwen Spencer. Approximation algorithms for fragmenting a graph against a stochastically-located threat. Lecture Notes in Computer Science, 7164:123–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29116-6_11/.

Zhu:2012:NCW

[794] Jianqiao Zhu, Ho-Leung Chan, and Tak-Wah Lam. Non-clairvoyant weighted flow time scheduling on different multi-processor models. Lecture Notes in Computer Science, 7164:137–149, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29116-6_12/.

Dorrigiv:2012:NPL

[795] Reza Dorrigiv and Alejandro López-Ortiz. A new perspective on list update: Probabilistic locality and working set. Lecture Notes in Computer Science, 7164:150–163, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29116-6_13/.

Brodal:2012:OFS

[796] Gerth Stølting Brodal and Gabriel Moruz. OnlineMin: a fast strongly competitive randomized paging algorithm. Lecture Notes in Computer Science, 7164:164–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29116-6_14/.

Paluch:2012:FSA

[797] Katarzyna Paluch. Faster and simpler approximation of stable matchings. Lecture Notes in Computer Science, 7164:176–187, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29116-6_15/.

vanZuylen:2012:SAA

[798] Anke van Zuylen. Simpler 3/4-approximation algorithms for MAX SAT. Lecture Notes in Computer Science, 7164:188–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29116-6_16/.

Renault:2012:OAA

[799] Marc P. Renault and Adi Rosén. On online algorithms with advice for the k-server problem. Lecture Notes in Computer Science, 7164:198– 210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29116-6_17/.

Harren:2012:ILB

[800] Rolf Harren and Walter Kern. Improved lower bound for online strip packing. Lecture Notes in Computer Science, 7164:211–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29116-6_18/.

Mansour:2012:CRS

[801] Yishay Mansour, Boaz Patt-Shamir, and Dror Rawitz. Competitive router scheduling with structured data. Lecture Notes in Computer Science, 7164:219–232, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29116-6_19/.

Bazgan:2012:AFN

[802] Cristina Bazgan, Laurent Gourvès, and Jérôme Monnot. Approximation with a fixed number of solutions of some biobjective maximization problems. Lecture Notes in Computer Science, 7164:233–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29116-6_20/.

Anonymous:2012:FMbk

[803] Anonymous. Front matter. Lecture Notes in Computer Science, 7164: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29116-6/1.

Axelsen:2012:TCT

[804] Holger Bock Axelsen. Time complexity of tape reduction for reversible Turing machines. Lecture Notes in Computer Science, 7165: 1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29517-1_1/.

Yokoyama:2012:TRF

[805] Tetsuo Yokoyama, Holger Bock Axelsen, and Robert Glück. Towards a reversible functional language. Lecture Notes in Computer Science, 7165: 14–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29517-1_2/.

Thomsen:2012:RPA

[806] Michael Kirkedal Thomsen, Holger Bock Axelsen, and Robert Glück. A reversible processor architecture and its reversible logic design. Lecture Notes in Computer Science, 7165:30–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29517-1_3/.

Rahman:2012:ORC

[807] Md. Mazder Rahman, Gerhard W. Dueck, and Anindita Banerjee. Optimization of reversible circuits using reconfigured templates. Lecture Notes in Computer Science, 7165:43–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29517-1_4/.

Moraga:2012:HBE

[808] Claudio Moraga. Hybrid GF(2)–Boolean expressions … for quantum computing circuits. Lecture Notes in Computer Science, 7165:54–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29517-1_5/.

Soeken:2012:ROS

[809] Mathias Soeken, Stefan Frehse, Robert Wille, and Rolf Drechsler. RevKit: An open source toolkit for the design of reversible circuits. Lecture Notes in Computer Science, 7165:64–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29517-1_6/.

Sasanian:2012:TMC

[810] Zahra Sasanian and D. Michael Miller. Transforming MCT circuits to NCVW circuits. Lecture Notes in Computer Science, 7165:77–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29517-1_7/.

Matsuo:2012:CGO

[811] Atsushi Matsuo and Shigeru Yamashita. Changing the gate order for optimal LNN conversion. Lecture Notes in Computer Science, 7165:89– 101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29517-1_8/.

Burignat:2012:TLC

[812] Stéphane Burignat, Mariusz Olczak, Michał Klimczak, and Alexis De Vos. Towards the limits of cascaded reversible (quantum-inspired) circuits. Lecture Notes in Computer Science, 7165:102–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29517-1_9/.

Burignat:2012:IRP

[813] Stéphane Burignat, Michael Kirkedal Thomsen, and Michał Klimczak. Interfacing reversible pass-transistor CMOS chips with conventional restoring CMOS circuits. Lecture Notes in Computer Science, 7165:112– 122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29517-1_10/.

Anonymous:2012:BMr

[814] Anonymous. Back matter. Lecture Notes in Computer Science, 7165: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29517-1/1.

Anonymous:2012:FMbl

[815] Anonymous. Front matter. Lecture Notes in Computer Science, 7165: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29517-1/1.

Negeri:2012:DSM

[816] Ebisa Negeri and Nico Baken. Distributed storage management using dynamic pricing in a self-organized energy community. Lecture Notes in Computer Science, 7166:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28583-7_1/.

Paquereau:2012:ABS

[817] Laurent Paquereau and Bjarne E. Helvik. Ant-based systems for wireless networks: Retrospect and prospects. Lecture Notes in Computer Science, 7166:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28583-7_2/.

Squartini:2012:TMD

[818] Tiziano Squartini and Diego Garlaschelli. Triadic motifs and dyadic selforganization in the world trade network. Lecture Notes in Computer Science, 7166:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28583-7_3/.

Stys:2012:MIV

[819] Dalibor Štys, Petr Jizba, Štěpán Papáček, Tomáš Náhlík, and Petr Císař. On measurement of internal variables of complex self-organized systems and their relation to multifractal spectra. Lecture Notes in Computer Science, 7166:36–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28583-7_4/.

Prothmann:2012:SOR

[820] Holger Prothmann, Sven Tomforde, Johannes Lyda, Jürgen Branke, and Jörg Hähner. Self-organised routing for road networks. Lecture Notes in Computer Science, 7166:48–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28583-7_5/.

Kavalionak:2012:PCM

[821] Hanna Kavalionak and Alberto Montresor. P2P and cloud: a marriage of convenience for replica management. Lecture Notes in Computer Science, 7166:60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28583-7_6/.

Punithan:2012:SOS

[822] Dharani Punithan and R. I. (Bob) McKay. Self-organizing spatiotemporal pattern formation in two-dimensional daisyworld. Lecture Notes in Computer Science, 7166:72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28583-7_7/.

Al-Asfoor:2012:HRS

[823] Muntasir Al-Asfoor, Brendan Neville, and Maria Fasli. Heuristic resource search in a self-organised distributed multi agent system. Lecture Notes in Computer Science, 7166:84–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28583-7_8/.

Georgiev:2012:QMM

[824] Georgi Yordanov Georgiev. A quantitative measure, mechanism and attractor for self-organization in networked complex systems. Lecture Notes in Computer Science, 7166:90–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28583-7_9/.

Blumenfeld-Lieberthal:2012:MMS

[825] Efrat Blumenfeld-Lieberthal and Juval Portugali. MetroNet: a metropolitan simulation model based on commuting processes. Lecture Notes in Computer Science, 7166:96–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28583-7_10/.

Gogolev:2012:RSO

[826] Alexander Gogolev and Christian Bettstetter. Robustness of selforganizing consensus algorithms: Initial results from a simulation-based study. Lecture Notes in Computer Science, 7166:104–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28583-7_11/.

Majercik:2012:IEU

[827] Stephen M. Majercik. Initial experiments in using communication swarms to improve the performance of swarm systems. Lecture Notes in Computer Science, 7166:109–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28583-7_12/.

Anonymous:2012:BMs

[828] Anonymous. Back matter. Lecture Notes in Computer Science, 7166: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28583-7/1.

Anonymous:2012:FMbm

[829] Anonymous. Front matter. Lecture Notes in Computer Science, 7166: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28583-7/1.

Evermann:2012:DSM

[830] Joerg Evermann and Ivan Porres. Doctoral symposium at MoDELS 2011. Lecture Notes in Computer Science, 7167:1–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29645-1_1/.

Agt:2012:SSL

[831] Henning Agt. Supporting software language engineering by automated domain knowledge acquisition. Lecture Notes in Computer Science, 7167: 4–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29645-1_2/.

Qureshi:2012:ISE

[832] Muhammad Atif Qureshi. Interoperability of software engineering metamodels. Lecture Notes in Computer Science, 7167:12–19, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29645-1_3/.

Brandsteidl:2012:SME

[833] Marion Brandsteidl and Andreas Winter. Software modeling in education: The 7th educators’ symposium at MoDELS 2011. Lecture Notes in Computer Science, 7167:20–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29645-1_4/.

Chiorean:2012:UCT

[834] Dan Chiorean, Vladiela Petraşcu, and Ileana Ober. Using constraints in teaching software modeling. Lecture Notes in Computer Science, 7167: 25–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29645-1_5/.

Whittle:2012:MBI

[835] Jon Whittle and John Hutchinson. Mismatches between industry practice and teaching of model-driven software development. Lecture Notes in Computer Science, 7167:40–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29645-1_6/.

Haugen:2012:VVY

[836] Øystein Haugen. VARY — variability for you. Lecture Notes in Computer Science, 7167:48–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29645-1_7/.

Svendsen:2012:TEG

[837] Andreas Svendsen, Xiaorui Zhang, Øystein Haugen, and Birger Møller-Pedersen. Towards evolution of generic variability models. Lecture Notes in Computer Science, 7167:53–67, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29645-1_8/.

Abu-Matar:2012:FBV

[838] Mohammad Abu-Matar and Hassan Gomaa. Feature-based variability meta-modeling for service-oriented product lines. Lecture Notes in Computer Science, 7167:68–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29645-1_9/.

Amaral:2012:SWM

[839] Vasco Amaral, Cécile Hardebolle, Hans Vangheluwe, László Lengyel, and Peter Bunus. Summary of the workshop on multi-paradigm modelling: Concepts and tools. Lecture Notes in Computer Science, 7167:83–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29645-1_10/.

Cicchetti:2012:SIS

[840] Antonio Cicchetti, Federico Ciccozzi, and Thomas Leveque. Supporting incremental synchronization in hybrid multi-view modelling. Lecture Notes in Computer Science, 7167:89–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29645-1_11/.

Boulanger:2012:SAU

[841] Frédéric Boulanger, Ayman Dogui, Cécile Hardebolle, and Christophe Jacquet. Semantic adaptation using CCSL clock constraints. Lecture Notes in Computer Science, 7167:104–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29645-1_12/.

Chaudron:2012:SFI

[842] Michel R. V. Chaudron, Marcela Genero, Silvia Abrahão, Parastoo Mohagheghi, and Lars Pareto. Summary of the first international workshop on experiences and empirical studies in software modelling. Lecture Notes in Computer Science, 7167:119–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29645-1_13/.

Zugal:2012:AIH

[843] Stefan Zugal, Jakob Pinggera, Barbara Weber, Jan Mendling, and Hajo A. Reijers. Assessing the impact of hierarchy on model understandability — a cognitive perspective. Lecture Notes in Computer Science, 7167:123–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29645-1_14/.

Fernandez-Saez:2012:DLD

[844] Ana M. Fernández-Sáez, Marcela Genero, and Michel R. V. Chaudron. Does the level of detail of UML models affect the maintainability of source code? Lecture Notes in Computer Science, 7167:134–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29645-1_15/.

Bencomo:2012:SIW

[845] Nelly Bencomo, Gordon Blair, Robert France, Betty H. C. Cheng, and Cédric Jeanneret. Summary of the 6^{th} International Workshop on Models@run.time. Lecture Notes in Computer Science, 7167:149–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29645-1_16/.

Ghezzi:2012:RMF

[846] Carlo Ghezzi, Andrea Mocci, and Mario Sangiorgio. Runtime monitoring of functional component changes with behavior models. Lecture Notes in Computer Science, 7167:152–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29645-1_17/.

Vogel:2012:RAL

[847] Thomas Vogel and Holger Giese. Requirements and assessment of languages and frameworks for adaptation models. Lecture Notes in Computer Science, 7167:167–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29645-1_18/.

Lucio:2012:MWS

[848] Levi Lúcio, Stephan Weißleder, Frédéric Fondement, and Harald Cichos. MoDeVVa 2011 Workshop Summary. Lecture Notes in Computer Science, 7167:183–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29645-1_19/.

Svendsen:2012:STO

[849] Andreas Svendsen, Øystein Haugen, and Birger Møller-Pedersen. Specifying a testing oracle for train stations — going beyond with product line technology. Lecture Notes in Computer Science, 7167:187–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29645-1_20/.

Anonymous:2012:FMbn

[850] Anonymous. Front matter. Lecture Notes in Computer Science, 7167: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29645-1/1.

Hashimoto:2012:AUA

[851] Junichi Hashimoto, Akihiro Kishimoto, and Kazuki Yoshizoe. Accelerated UCT and its application to two-player games. Lecture Notes in Computer Science, 7168:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31866-5_1/.

VanEyck:2012:RMG

[852] Gabriel Van Eyck and Martin Müller. Revisiting move groups in Monte– Carlo tree search. Lecture Notes in Computer Science, 7168:13–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31866-5_2/.

Baudis:2012:PSA

[853] Petr Baudiš and Jean loup Gailly. PACHI: State of the art open source go program. Lecture Notes in Computer Science, 7168:24–38, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31866-5_3/.

Baier:2012:TMM

[854] Hendrik Baier and Mark H. M. Winands. Time management for Monte– Carlo tree search in go. Lecture Notes in Computer Science, 7168:39– 51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31866-5_4/.

Lorentz:2012:MPP

[855] Richard J. Lorentz. An MCTS program to play EinStein Würfelt Nicht! Lecture Notes in Computer Science, 7168:52–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31866-5_5/.

Stankiewicz:2012:MCT

[856] Jan A. Stankiewicz and Mark H. M. Winands. Monte–Carlo tree search enhancements for Havannah. Lecture Notes in Computer Science, 7168: 60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31866-5_6/.

Nijssen:2012:PSM

[857] J. (Pim) A. M. Nijssen and Mark H. M. Winands. Playout search for Monte–Carlo tree search in multi-player games. Lecture Notes in Computer Science, 7168:72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31866-5_7/.

Chou:2012:TSG

[858] Cheng-Wei Chou, Ping-Chiang Chou, and Hassen Doghmen. Towards a solution of 7x7 go with meta-MCTS. Lecture Notes in Computer Science, 7168:84–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31866-5_8/.

Bouzy:2012:MEV

[859] Bruno Bouzy, Marc Métivier, and Damien Pellier. MCTS experiments on the Voronoi game. Lecture Notes in Computer Science, 7168:96– 107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31866-5_9/.

Wang:2012:PBL

[860] Jiao Wang, Shiyuan Li, Jitong Chen, Xin Wei, and Huizhan Lv. 4*4pattern and Bayesian learning in Monte–Carlo go. Lecture Notes in Computer Science, 7168:108–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31866-5_10/.

Wu:2012:TDL

[861] I-Chen Wu, Hsin-Ti Tsai, Hung-Hsuan Lin, and Yi-Shan Lin. Temporal difference learning for connect6. Lecture Notes in Computer Science, 7168:121–133, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31866-5_11/.

Papahristou:2012:ITD

[862] Nikolaos Papahristou and Ioannis Refanidis. Improving temporal difference learning performance in backgammon variants. Lecture Notes in Computer Science, 7168:134–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31866-5_12/.

Coulom:2012:CCL

[863] Rémi Coulom. CLOP: Confident local optimization for noisy black-box parameter tuning. Lecture Notes in Computer Science, 7168:146–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31866-5_13/.

Kaneko:2012:AEF

[864] Tomoyuki Kaneko and Kunihito Hoki. Analysis of evaluation-function learning by comparison of sibling nodes. Lecture Notes in Computer Science, 7168:158–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31866-5_14/.

Neller:2012:AOD

[865] Todd W. Neller and Steven Hnath. Approximating optimal dudo play with fixed-strategy iteration counterfactual regret minimization. Lecture Notes in Computer Science, 7168:170–183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31866-5_15/.

Hoki:2012:GLO

[866] Kunihito Hoki and Tomoyuki Kaneko. The global landscape of objective functions for the optimization of shogi piece values with a game-tree search. Lecture Notes in Computer Science, 7168:184–195, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31866-5_16/.

Saffidine:2012:SBR

[867] Abdallah Saffidine, Nicolas Jouandeau, and Tristan Cazenave. Solving breakthrough with race patterns and job-level proof number search. Lecture Notes in Computer Science, 7168:196–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31866-5_17/.

Yamaguchi:2012:ICF

[868] Yoshiaki Yamaguchi, Kazunori Yamaguchi, and Tetsuro Tanaka. Infinite connect-four is solved: Draw. Lecture Notes in Computer Science, 7168:208–219, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31866-5_18/.

Brausen:2012:BCG

[869] Henry Brausen, Ryan B. Hayward, Martin Müller, and Abdul Qadir. Blunder cost in go and hex. Lecture Notes in Computer Science, 7168:220–229, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31866-5_19/.

Regan:2012:UDC

[870] Kenneth W. Regan, Bartlomiej Macieja, and Guy McC. Haworth. Understanding distributions of chess performances. Lecture Notes in Computer Science, 7168:230–243, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31866-5_20/.

Anonymous:2012:FMbo

[871] Anonymous. Front matter. Lecture Notes in Computer Science, 7168: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31866-5/1.

vanderHoek:2012:CD

[872] Wiebe van der Hoek. Control and delegation. Lecture Notes in Computer Science, 7169:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29113-5_1.

Dennis:2012:PIS

[873] Louise A. Dennis. Plan indexing for state-based plans. Lecture Notes in Computer Science, 7169:3–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29113-5_2/.

Winikoff:2012:IFF

[874] Michael Winikoff. An integrated formal framework for reasoning about goal interactions. Lecture Notes in Computer Science, 7169:16–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29113-5_3/.

Bijani:2012:PAM

[875] Shahriar Bijani, David Robertson, and David Aspinall. Probing attacks on multi-agent systems using electronic institutions. Lecture Notes in Computer Science, 7169:33–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29113-5_4/.

Gunay:2012:DCC

[876] Akın Günay and Pınar Yolum. Detecting conflicts in commitments. Lecture Notes in Computer Science, 7169:51–66, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29113-5_5/.

Son:2012:FCU

[877] Tran Cao Son, Enrico Pontelli, and Chiaki Sakama. Formalizing commitments using action languages. Lecture Notes in Computer Science, 7169: 67–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29113-5_6/.

Robertson:2012:LCC

[878] David Robertson. Lightweight coordination calculus for agent systems: Retrospective and prospective. Lecture Notes in Computer Science, 7169: 84–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29113-5_7/.

Chopra:2012:EI

[879] Amit K. Chopra and Munindar P. Singh. The evolution of interoperability. Lecture Notes in Computer Science, 7169:90–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29113-5_8/.

Mascardi:2012:YCB

[880] Viviana Mascardi and Davide Ancona. 1000 years of coo-BDI. Lecture Notes in Computer Science, 7169:95–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29113-5_9/.

Garcia-Camino:2012:DAN

[881] Andrés García-Camino, Juan-Antonio Rodríguez-Aguilar, and Carles Sierra. A distributed architecture for norm-aware agent societies: a retrospective. Lecture Notes in Computer Science, 7169:102–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29113-5_10/.

Moreira:2012:SAB

[882] Álvaro F. Moreira, Renata Vieira, and Rafael H. Bordini. Speech-act based communication: Progress in the formal semantics and in the implementation of multi-agent oriented programming languages. Lecture Notes in Computer Science, 7169:111–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29113-5_11/.

Fornara:2012:SEN

[883] Nicoletta Fornara and Marco Colombetti. Specifying and enforcing norms in artificial institutions: a retrospective review. Lecture Notes in Computer Science, 7169:117–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29113-5_12/.

Torroni:2012:RRE

[884] Paolo Torroni, Federico Chesani, Paola Mello, and Marco Montali. A retrospective on the reactive event calculus and commitment modeling language. Lecture Notes in Computer Science, 7169:120–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29113-5_13/.

Corsar:2012:WSC

[885] David Corsar, Alison Chorley, and Wamberto W. Vasconcelos. Web service composition via organisation-based (re)planning. Lecture Notes in Computer Science, 7169:128–148, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29113-5_14/.

Bordini:2012:AMA

[886] Rafael H. Bordini. Agent and multi-agent software engineering: Modelling, programming, and verification. Lecture Notes in Computer Science, 7169:149–151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29113-5_15/.

Anonymous:2012:BMt

[887] Anonymous. Back matter. Lecture Notes in Computer Science, 7169: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29113-5/1.

Anonymous:2012:FMbp

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29113-5/1.

Solhaug:2012:UST

[889] Bjørnar Solhaug and Ketil Stølen. Uncertainty, subjectivity, trust and risk: How it all fits together. Lecture Notes in Computer Science, 7170: 1–5, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29963-6_1/.

Josang:2012:TEI

[890] Audun Jøsang. Trust extortion on the Internet. Lecture Notes in Computer Science, 7170:6–21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29963-6_2/.

Rudolph:2012:TAS

[891] Carsten Rudolph. Trust areas: a security paradigm for the future Internet. Lecture Notes in Computer Science, 7170:22–27, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29963-6_3/.

Marsh:2012:NST

[892] Stephen Marsh, Sylvie Noël, Tim Storer, and Yao Wang. Non-standards for trust: Foreground trust and second thoughts for mobile security. Lecture Notes in Computer Science, 7170:28–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29963-6_4/.

Morgenstern:2012:PCF

[893] Jamie Morgenstern, Deepak Garg, and Frank Pfenning. A proof-carrying file system with revocable and use-once certificates. Lecture Notes in Computer Science, 7170:40–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29963-6_5/.

Genovese:2012:NMA

[894] Valerio Genovese and Deepak Garg. New modalities for access control logics: Permission, control and ratification. Lecture Notes in Computer Science, 7170:56–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29963-6_6/.

Sarier:2012:SNB

[895] Neyire Deniz Sarier. Security notions of biometric remote authentication revisited. Lecture Notes in Computer Science, 7170:72–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29963-6_7/.

Muller:2012:HPC

[896] Sascha Müller and Stefan Katzenbeisser. Hiding the policy in cryptographic access control. Lecture Notes in Computer Science, 7170:90– 105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29963-6_8/.

Rechert:2012:LPR

[897] Klaus Rechert and Benjamin Greschbach. Location privacy in relation to trusted peers. Lecture Notes in Computer Science, 7170:106–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29963-6_9/.

Jamroga:2012:FNR

[898] Wojciech Jamroga, Sjouke Mauw, and Matthijs Melissen. Fairness in non-repudiation protocols. Lecture Notes in Computer Science, 7170:122–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29963-6_10/.

Chen:2012:RAR

[899] Liang Chen and Jason Crampton. Risk-aware role-based access control. Lecture Notes in Computer Science, 7170:140–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29963-6_11/.

Armando:2012:AAI

[900] Alessandro Armando and Silvio Ranise. Automated analysis of infinite state workflows with access control policies. Lecture Notes in Computer Science, 7170:157–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29963-6_12/.

Pohls:2012:RDI

[901] Henrich C. Pöhls and Focke Höhne. The role of data integrity in EU digital signature legislation — achieving statutory trust for sanitizable signature schemes. Lecture Notes in Computer Science, 7170:175–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29963-6_13/.

Greveler:2012:MRA

[902] Ulrich Greveler, Benjamin Justus, and Dennis Löhr. Mutual remote attestation: Enabling system cloning for TPM based platforms. Lecture Notes in Computer Science, 7170:193–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29963-6_14/.

Najera:2012:SAI

[903] Pablo Najera, Rodrigo Roman, and Javier Lopez. Secure architecure for the integration of RFID and sensors in personal networks. Lecture Notes in Computer Science, 7170:207–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29963-6_15/.

Rahman:2012:AIP

[904] Syed Sadiqur Rahman, Sadie Creese, and Michael Goldsmith. Accepting information with a pinch of salt: Handling untrusted information sources. Lecture Notes in Computer Science, 7170:223–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29963-6_16/.

Anonymous:2012:BMu

[905] Anonymous. Back matter. Lecture Notes in Computer Science, 7170: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29963-6/1.

Anonymous:2012:FMbq

[906] Anonymous. Front matter. Lecture Notes in Computer Science, 7170: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29963-6/1.

Gupta:2012:OAO

[907] Chitralekha Gupta and Preeti Rao. Objective assessment of ornamentation in Indian classical singing. Lecture Notes in Computer Science, 7172: 1–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31980-8_1/.

Patra:2012:ONA

[908] Moujhuri Patra, Soubhik Chakraborty, and Dipak Ghosh. Optimized neural architecture for time series prediction using raga notes. Lecture Notes in Computer Science, 7172:26–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31980-8_2/.

Gulati:2012:MDA

[909] Sankalp Gulati, Vishweshwara Rao, and Preeti Rao. Meter detection from audio for Indian music. Lecture Notes in Computer Science, 7172: 34–43, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31980-8_3/.

Singh:2012:ALR

[910] Shashi Bhushan Singh, Soubhik Chakraborty, and Keashav Mohan Jha. Assessment of level of recovery of cognitive impairment in the cerebrovascular accident and head injuries cases: Therapeutic impact of North Indian ragas. Lecture Notes in Computer Science, 7172:44–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31980-8_4/.

Braeunig:2012:TDB

[911] Matthias Braeunig, Ranjan Sengupta, and Anirban Patranabis. On tanpura drone and brain electrical correlates. Lecture Notes in Computer Science, 7172:53–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31980-8_5/.

Wu:2012:MII

[912] Jun Wu and Shigeki Sagayama. Musical instrument identification based on new boosting algorithm with probabilistic decisions. Lecture Notes in Computer Science, 7172:66–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31980-8_6/.

Jensen:2012:MGC

[913] Kristoffer Jensen. Music genre classification using an auditory memory model. Lecture Notes in Computer Science, 7172:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31980-8_7/.

deQuay:2012:IME

[914] Yago de Quay. Interactive music 3.0: Empowering people to participate musically inside nightclubs. Lecture Notes in Computer Science, 7172: 89–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31980-8_8/.

Ross:2012:HCM

[915] Joe Cheri Ross and John Samuel. Hierarchical clustering of music database based on HMM and Markov chain for search efficiency. Lecture Notes in Computer Science, 7172:98–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-31980-8_9/.

Stables:2012:FFM

[916] Ryan Stables, Cham Athwal, and Jamie Bullock. Fundamental frequency modulation in singing voice synthesis. Lecture Notes in Computer Science, 7172:104–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31980-8_10/.

Nymoen:2012:SAA

[917] Kristian Nymoen, Jim Torresen, Rolf Inge Godøy, and Alexander Refsum Jensenius. A statistical approach to analyzing sound tracings. Lecture Notes in Computer Science, 7172:120–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31980-8_11/.

Necciari:2012:ATF

[918] Thibaud Necciari, Peter Balazs, Richard Kronland-Martinet, Sølvi Ystad, and Bernhard Laback. Auditory time-frequency masking: Psychoacoustical data and application to audio representations. Lecture Notes in Computer Science, 7172:146–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31980-8_12/.

Aramaki:2012:PCE

[919] Mitsuko Aramaki, Richard Kronland-Martinet, and Sølvi Ystad. Perceptual control of environmental sound synthesis. Lecture Notes in Computer Science, 7172:172–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31980-8_13/.

Bhattacharjee:2012:RAP

[920] Utpal Bhattacharjee. Recognition of assamese phonemes using RNN based recognizer. Lecture Notes in Computer Science, 7172:187–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31980-8_14/.

Datta:2012:SAL

[921] Asoke Kumar Datta and Arup Saha. A system for analysis of large scale speech data for the development of rules of intonation for speech synthesis. Lecture Notes in Computer Science, 7172:197–206, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31980-8_15/.

Mohanty:2012:AIW

[922] Sanghamitra Mohanty and Basanta Kumar Swain. Adaptive and iterative Wiener filter for Oriya speech processing applications. Lecture Notes in Computer Science, 7172:207–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31980-8_16/.

Datta:2012:RFC

[923] Asoke Kumar Datta and Bhaswati Mukherjee. On the role of formants in cognition of vowels and place of articulation of plosives. Lecture Notes in Computer Science, 7172:215–234, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31980-8_17/.

Anonymous:2012:BMv

[924] Anonymous. Back matter. Lecture Notes in Computer Science, 7172: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31980-8/1.

Anonymous:2012:FMbr

[925] Anonymous. Front matter. Lecture Notes in Computer Science, 7172: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31980-8/1.

Bono:2012:RSM

[926] Viviana Bono, Sara Capecchi, and Ilaria Castellani. A reputation system for multirole sessions. Lecture Notes in Computer Science, 7173:

1–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30065-3_1/.

Chen:2012:ADM

[927] Tzu-Chun Chen, Laura Bocchi, and Pierre-Malo Deniélou. Asynchronous distributed monitoring for multiparty session enforcement. Lecture Notes in Computer Science, 7173:25–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30065-3_2/.

Werther:2012:MCN

[928] Bernd Werther and Nicklas Hoch. E-mobility as a challenge for new ICT solutions in the car industry. Lecture Notes in Computer Science, 7173: 46–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30065-3_3/.

Alvarez:2012:ENE

[929] Carme Àlvarez, Amalia Duch, Maria Serna, and Dimitrios Thilikos. On the existence of Nash equilibria in strategic search games. Lecture Notes in Computer Science, 7173:58–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30065-3_4/.

Barthe:2012:SEI

[930] Gilles Barthe and Exequiel Rivas. Static enforcement of information flow policies for a concurrent JVM-like language. Lecture Notes in Computer Science, 7173:73–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30065-3_5/.

Bernardo:2012:WMB

[931] Marco Bernardo. Weak Markovian bisimulation congruences and exact CTMC-level aggregations for sequential processes. Lecture Notes in Computer Science, 7173:89–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30065-3_6/.

Buscemi:2012:CSC

[932] Maria Grazia Buscemi and Mario Coppo. Constraints for service contracts. Lecture Notes in Computer Science, 7173:104–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30065-3_7/.

Chen:2012:SET

[933] Liqun Chen, Ming-Feng Lee, and Bogdan Warinschi. Security of the enhanced TCG privacy-CA solution. Lecture Notes in Computer Science, 7173:121–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30065-3_8/.

DErrico:2012:CAS

[934] Liliana D’Errico and Michele Loreti. Context aware specification and verification of distributed systems. Lecture Notes in Computer Science, 7173:142–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30065-3_9/.

DeNicola:2012:OTB

[935] Rocco De Nicola, Andrea Margheri, and Francesco Tiezzi. Orchestrating tuple-based languages. Lecture Notes in Computer Science, 7173:160– 178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30065-3_10/.

Duggan:2012:TCS

[936] Dominic Duggan and Ye Wu. Transactional correctness for secure nested transactions. Lecture Notes in Computer Science, 7173:179–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30065-3_11/.

Gabarro:2012:OUS

[937] Joaquim Gabarro, Maria Serna, and Alan Stewart. Orchestrating unreliable services: Strategic and probabilistic approaches to reliability. Lecture Notes in Computer Science, 7173:197–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30065-3_12/.

Melgratti:2012:CSC

[938] Hernán Melgratti and Christian Roldán. On correlation sets and correlation exceptions in ActiveBPEL. Lecture Notes in Computer Science, 7173:212–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30065-3_13/.

Zhang:2012:CIF

[939] Chenyi Zhang. Conditional information flow policies and unwinding relations. Lecture Notes in Computer Science, 7173:227–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30065-3_14/.

Anonymous:2012:BMw

[940] Anonymous. Back matter. Lecture Notes in Computer Science, 7173: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30065-3/1.

Anonymous:2012:FMbs

[941] Anonymous. Front matter. Lecture Notes in Computer Science, 7173: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30065-3/1.

Reinders:2012:OFS

[942] James Reinders. Only the first steps of the parallel evolution have been taken thus far. Lecture Notes in Computer Science, 7174:1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30397-5_1/.

Garcia:2012:DLB

[943] Marta Garcia, Julita Corbalan, Rosa Maria Badia, and Jesus Labarta. A dynamic load balancing approach with SMPSuperscalar and MPI. Lecture Notes in Computer Science, 7174:10–23, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-30397-5_2/.

Christadler:2012:PPN

[944] Iris Christadler, Giovanni Erbacci, and Alan D. Simpson. Performance and productivity of new programming languages. Lecture Notes in Computer Science, 7174:24–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30397-5_3/.

Heinecke:2012:THP

[945] Alexander Heinecke, Michael Klemm, Hans Pabst, and Dirk Pflüger. Towards high-performance implementations of a custom HPC kernel using (R) array building blocks. Lecture Notes in Computer Science, 7174:36– 47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30397-5_4/.

Bechara:2012:AAH

[946] Charly Bechara, Nicolas Ventroux, and Daniel Etiemble. AHDAM: An asymmetric homogeneous with dynamic allocator manycore chip. Lecture Notes in Computer Science, 7174:48–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-30397-5_5/.

Fularz:2012:FIR

[947] Michał Fularz, Marek Kraft, Adam Schmidt, and Andrzej Kasiński. FPGA implementation of the robust essential matrix estimation with RANSAC and the 8-point and the 5-point method. Lecture Notes in Computer Science, 7174:60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30397-5_6/.

Bielecki:2012:UFS

[948] Wlodzimierz Bielecki and Marek Palkowski. Using free scheduling for programming graphic cards. Lecture Notes in Computer Science, 7174: 72–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30397-5_7/.

Kawanami:2012:GAC

[949] Katsuya Kawanami and Noriyuki Fujimoto. GPU accelerated computation of the longest common subsequence. Lecture Notes in Computer Science, 7174:84–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30397-5_8/.

Hernandez:2012:EHL

[950] Oscar Hernandez, Wei Ding, Barbara Chapman, Christos Kartsaklis, and Ramanan Sankaran. Experiences with high-level programming directives for porting applications to GPUs. Lecture Notes in Computer Science, 7174:96–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-30397-5_9/.

Auer:2012:GAG

[951] Bas O. Fagginger Auer and Rob H. Bisseling. A GPU algorithm for greedy graph matching. Lecture Notes in Computer Science, 7174:108– 119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30397-5_10/.

Krause:2012:HPL

[952] Dorian Krause, Mark Potse, Thomas Dickopf, Rolf Krause, Angelo Auricchio, and Frits Prinzen. Hybrid parallelization of a large-scale heart model. Lecture Notes in Computer Science, 7174:120–132, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30397-5_11/.

Kraus:2012:EAH

[953] Jiri Kraus and Malte Förster. Efficient AMG on heterogeneous systems. Lecture Notes in Computer Science, 7174:133–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-30397-5_12/.

Rupp:2012:GAP

[954] Karl Rupp, Ansgar Jüngel, and Tibor Grasser. A GPU-accelerated parallel preconditioner for the solution of the Boltzmann transport equation for semiconductors. Lecture Notes in Computer Science, 7174:147– 157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-30397-5_13/.

Heuveline:2012:PSM

[955] Vincent Heuveline, Dimitar Lukarski, Nico Trost, and Jan-Philipp Weiss. Parallel smoothers for matrix-based geometric multigrid methods on locally refined meshes using multicore CPUs and GPUs. Lecture Notes in Computer Science, 7174:158–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-30397-5_14/.

Anonymous:2012:BMx

[956] Anonymous. Back matter. Lecture Notes in Computer Science, 7174: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-30397-5/1.

Anonymous:2012:FMbt

[957] Anonymous. Front matter. Lecture Notes in Computer Science, 7174: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-30397-5/1.

Angelov:2012:TOD

[958] Krasimir Angelov and Ramona Enache. Typeful ontologies with direct multilingual verbalization. Lecture Notes in Computer Science, 7175: 1–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31175-8_1/.

Bunzli:2012:CAL

[959] Alexandra Bünzli and Stefan Höfler. Controlling ambiguities in legislative language. Lecture Notes in Computer Science, 7175:21–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31175-8_2/.

Cramer:2012:IPN

[960] Marcos Cramer and Bernhard Schröder. Interpreting plurals in the Naproche CNL. Lecture Notes in Computer Science, 7175:43–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31175-8_3/.

Dantuluri:2012:ECN

[961] Pradeep Dantuluri, Brian Davis, and Pierre Ludwick. Engineering a controlled natural language into Semantic MediaWiki. Lecture Notes in Computer Science, 7175:53–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-31175-8_4/.

Fuchs:2012:FOR

[962] Norbert E. Fuchs. First-order reasoning for Attempto Controlled English. Lecture Notes in Computer Science, 7175:73–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31175-8_5/.

Kuhn:2012:CPN

[963] Tobias Kuhn. Codeco: a practical notation for controlled English grammars in predictive Editors. Lecture Notes in Computer Science, 7175:95– 114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-31175-8_6/.

Ranta:2012:CLE

[964] Aarne Ranta, Ramona Enache, and Grégoire Détrez. Controlled language for everyday use: The MOLTO phrasebook. Lecture Notes in Computer Science, 7175:115–136, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31175-8_7/.

Camilleri:2012:CNL

[965] John J. Camilleri, Gordon J. Pace, and Michael Rosner. Controlled natural language in a game for legal assistance. Lecture Notes in Computer Science, 7175:137–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-31175-8_8/.

Schwitter:2012:WES

[966] Rolf Schwitter. Working with events and states in PENG light. Lecture Notes in Computer Science, 7175:154–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-31175-8_9/.

Spreeuwenberg:2012:UCT

[967] Silvie Spreeuwenberg, Jeroen van Grondelle, and Ronald Heller. Using CNL techniques and pattern sentences to involve domain experts in modeling. Lecture Notes in Computer Science, 7175:175–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-31175-8_10/.

Anonymous:2012:BMy

[968] Anonymous. Back matter. Lecture Notes in Computer Science, 7175: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-31175-8/1.

Anonymous:2012:FMbu

[969] Anonymous. Front matter. Lecture Notes in Computer Science, 7175: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-31175-8/1.

Honda:2012:UDS

[970] Kohei Honda. Understanding distributed services through the π-calculus. Lecture Notes in Computer Science, 7176:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/accesspage/chapter/10.1007/978-3-642-29834-9_1.

Lozes:2012:RCU

[971] Étienne Lozes and Jules Villard. Reliable contracts for unreliable halfduplex communications. Lecture Notes in Computer Science, 7176:2– 16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29834-9_2/.

Chang:2012:BBS

[972] Fangzhe Chang, Pavithra Prabhakar, and Ramesh Viswanathan. Behavior based service composition. Lecture Notes in Computer Science, 7176: 17–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29834-9_3/.

Masson:2012:CDC

[973] Benoît Masson, Loïc Hélouët, and Albert Benveniste. Compatibility of data-centric Web services. Lecture Notes in Computer Science, 7176:32– 47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29834-9_4/.

Lopez:2012:TEB

[974] Hugo A. López and Jorge A. Pérez. Time and exceptional behavior in multiparty structured interactions. Lecture Notes in Computer Science, 7176:48–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29834-9_5/.

Mazzara:2012:TDM

[975] Manuel Mazzara, Faisal Abouzaid, and Nicola Dragoni. Toward design, modelling and analysis of dynamic workflow reconfigurations. Lecture Notes in Computer Science, 7176:64–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29834-9_6/.

Mateo:2012:OSB

[976] José Antonio Mateo, Valentín Valero, and Gregorio Díaz. An operational semantics of BPEL orchestrations integrating Web services resource framework. Lecture Notes in Computer Science, 7176:79–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29834-9_7/.

Fares:2012:DBV

[977] Elie Fares, Jean-Paul Bodeveix, and Mamoun Filali. Design of a BPEL verification tool. Lecture Notes in Computer Science, 7176:95–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29834-9_8/.

Bruni:2012:APA

[978] Roberto Bruni, Andrea Corradini, and Gianluigi Ferrari. Applying process analysis to the Italian eGovernment Enterprise Architecture. Lecture Notes in Computer Science, 7176:111–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29834-9_9/.

Gallino:2012:DSM

[979] Juan Pedro Silva Gallino and Miguel de Miguel. Domain-specific multimodeling of security concerns in service-oriented architectures. Lecture Notes in Computer Science, 7176:128–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29834-9_10/.

Anonymous:2012:BMz

[980] Anonymous. Back matter. Lecture Notes in Computer Science, 7176: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29834-9/1.

Anonymous:2012:FMbv

[981] Anonymous. Front matter. Lecture Notes in Computer Science, 7176: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29834-9/1.

DalLago:2012:HOC

[982] Ugo Dal Lago and Paolo Parisen Toldin. A higher-order characterization of probabilistic polynomial time. Lecture Notes in Computer Science, 7177:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32495-6_1/.

Cesena:2012:TCB

[983] Emanuele Cesena, Marco Pedicini, and Luca Roversi. Typing a core binary-field arithmetic in a light logic. Lecture Notes in Computer Science, 7177:19–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32495-6_2/.

Montenegro:2012:IBH

[984] Manuel Montenegro, Olha Shkaravska, Marko van Eekelen, and Ricardo Peña. Interpolation-based height analysis for improving a recurrence solver. Lecture Notes in Computer Science, 7177:36–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32495-6_3/.

Lopez-Garcia:2012:IBR

[985] Pedro Lopez-Garcia, Luthfi Darmawan, Francisco Bueno, and Manuel Hermenegildo. Interval-based resource usage verification: Formalization and prototype. Lecture Notes in Computer Science, 7177:54–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32495-6_4/.

Amadio:2012:CRC

[986] Roberto M. Amadio and Yann Régis-Gianas. Certifying and reasoning on cost annotations of functional programs. Lecture Notes in Computer Science, 7177:72–89, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32495-6_5/.

Asavoae:2012:ATA

[987] Mihail Asăvoae, Irina Măriuca Asăvoae, and Dorel Lucanu. On abstractions for timing analysis in the K framework. Lecture Notes in Computer Science, 7177:90–107, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32495-6_6/.

Rodrigues:2012:CET

[988] Vítor Rodrigues, João Pedro Pedroso, Mário Florido, and Simão Melo de Sousa. Certifying execution time. Lecture Notes in Computer Science, 7177:108–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32495-6_7/.

Herrmann:2012:CAP

[989] Christoph A. Herrmann and Kevin Hammond. Counter automata for parameterised timing analysis of box-based systems. Lecture Notes in Computer Science, 7177:126–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32495-6_8/.

Anonymous:2012:BMaa

[990] Anonymous. Back matter. Lecture Notes in Computer Science, 7177: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32495-6/1.

Anonymous:2012:FMbw

[991] Anonymous. Front matter. Lecture Notes in Computer Science, 7177: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32495-6/1.

Heinig:2012:CBI

[992] Andreas Heinig, Vincent J. Mooney, Florian Schmoll, Peter Marwedel, and Krishna Palem. Classification-based improvement of application robustness and quality of service in probabilistic computer systems. Lecture Notes in Computer Science, 7179:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28293-5_1/.

Brehm:2012:CSE

[993] Christian Brehm, Matthias May, Christina Gimmler, and Norbert Wehn. A case study on error resilient architectures for wireless communication. Lecture Notes in Computer Science, 7179:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28293-5_2/.

Tahan:2012:UDT

[994] Oussama Tahan and Mohamed Shawky. Using dynamic task level redundancy for OpenMP fault tolerance. Lecture Notes in Computer Science, 7179:25–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28293-5_3/.

Xu:2012:VFQ

[995] Yang Xu, Rafael Rosales, Bo Wang, Martin Streubühr, Ralph Hasholzner, and Christian Haubelt. A very fast and quasi-accurate power-state-based system-level power modeling methodology. Lecture Notes in Computer Science, 7179:37–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28293-5_4/.

Nikitin:2012:STM

[996] Nikita Nikitin and Jordi Cortadella. Static task mapping for tiled chip multiprocessors with multiple voltage islands. Lecture Notes in Computer Science, 7179:50–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28293-5_5/.

Barthels:2012:APM

[997] Andreas Barthels, Joachim Fröschl, Hans-Ulrich Michel, and Uwe Baumgarten. An architecture for power management in automotive systems. Lecture Notes in Computer Science, 7179:63–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28293-5_6/.

Urena:2012:IMI

[998] Isaías A. Comprés Ureña, Michael Riepen, Michael Konow, and Michael Gerndt. Invasive MPI on Intel’s single-chip cloud computer. Lecture Notes in Computer Science, 7179:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28293-5_7/.

Zoni:2012:LOH

[999] Davide Zoni, Patrick Bellasi, and William Fornaciari. A low-overhead heuristic for mixed workload resource partitioning in cluster-based architectures. Lecture Notes in Computer Science, 7179:86–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28293-5_8/.

Boniol:2012:DEM

[1000] Frédéric Boniol, Hugues Cassé, Eric Noulard, and Claire Pagetti. Deterministic execution model on COTS hardware. Lecture Notes in Computer Science, 7179:98–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28293-5_9/.

Schleuniger:2012:DPS

[1001] Pascal Schleuniger, Sally A. McKee, and Sven Karlsson. Design principles for synthesizable processor cores. Lecture Notes in Computer Science, 7179:111–122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28293-5_10/.

Petrides:2012:HPD

[1002] Panayiotis Petrides, George Nicolaides, and Pedro Trancoso. HPC performance domains on multi-core processors with virtualization. Lecture Notes in Computer Science, 7179:123–134, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28293-5_11/.

Brandenburg:2012:GNI

[1003] Jens Brandenburg and Benno Stabernack. A generic and non-intrusive profiling methodology for SystemC multi-core platform simulation models. Lecture Notes in Computer Science, 7179:135–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28293-5_12/.

Membarth:2012:DTS

[1004] Richard Membarth, Jan-Hugo Lupp, Frank Hannig, Jürgen Teich, and Mario Körner. Dynamic task-scheduling and resource management for GPU accelerators in medical imaging. Lecture Notes in Computer Science, 7179:147–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28293-5_13/.

Hampel:2012:APE

[1005] Volker Hampel, Thilo Pionteck, and Erik Maehle. An approach for performance estimation of hybrid systems with FPGAs and GPUs as coprocessors. Lecture Notes in Computer Science, 7179:160–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28293-5_14/.

Mattheis:2012:WSS

[1006] Sebastian Mattheis, Tobias Schuele, Andreas Raabe, Thomas Henties, and Urs Gleim. Work stealing strategies for parallel stream processing in soft real-time systems. Lecture Notes in Computer Science, 7179:172– 183, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28293-5_15/.

Bielby:2012:DSE

[1007] Matthew Bielby, Miles Gould, and Nigel Topham. Design space exploration of hybrid ultra low power branch predictors. Lecture Notes in Computer Science, 7179:184–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28293-5_16/.

Fawibe:2012:NMO

[1008] Ademola Fawibe, Jared Sherman, Krishna Kavi, Mike Ignatowski, and David Mayhew. New memory organizations for 3D DRAM and PCMs. Lecture Notes in Computer Science, 7179:200–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28293-5_17/.

Zhang:2012:VLO

[1009] Hao Zhang, Hiroki Matsutani, Yasuhiro Take, Tadahiro Kuroda, and Hideharu Amano. Vertical link on/off control methods for wireless 3-D NoCs. Lecture Notes in Computer Science, 7179:212–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28293-5_18/.

Elsts:2012:SRC

[1010] Atis Elsts, Rihards Balass, Janis Judvaitis, Reinholds Zviedris, and Girts Strazdins. SADmote: a robust and cost-effective device for environmental monitoring. Lecture Notes in Computer Science, 7179:225–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28293-5_19/.

Oxman:2012:SNC

[1011] Gadi Oxman, Shlomo Weiss, and Yitzhak (Tsahi) Birk. Streamlined Network-on-Chip for multicore embedded architectures. Lecture Notes in Computer Science, 7179:238–249, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28293-5_20/.

Anonymous:2012:BMab

[1012] Anonymous. Back matter. Lecture Notes in Computer Science, 7179: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28293-5/1.

Anonymous:2012:FMbx

[1013] Anonymous. Front matter. Lecture Notes in Computer Science, 7179: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28293-5/1.

Albert:2012:AIR

[1014] Elvira Albert, Puri Arenas, and Samir Genaim. Automatic inference of resource consumption bounds. Lecture Notes in Computer Science, 7180: 1–11, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28717-6_1/.

Middeldorp:2012:MIP

[1015] Aart Middeldorp. Matrix interpretations for polynomial derivational complexity of rewrite systems. Lecture Notes in Computer Science, 7180: 12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-28717-6_2.

Motik:2012:PCF

[1016] Boris Motik. Parameterized complexity and fixed-parameter tractability of description logic reasoning. Lecture Notes in Computer Science, 7180: 13–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-28717-6_3.

Aavani:2012:ESM

[1017] Amir Aavani and Xiongnan (Newman) Wu. Enfragmo: a system for modelling and solving search problems with logic. Lecture Notes in Computer Science, 7180:15–22, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28717-6_4/.

Accattoli:2012:PC

[1018] Beniamino Accattoli and Delia Kesner. The permutative λ-calculus. Lecture Notes in Computer Science, 7180:23–36, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28717-6_5/.

Alama:2012:AHP

[1019] Jesse Alama, Daniel Kühlwein, and Josef Urban. Automated and human proofs in general mathematics: An initial comparison. Lecture Notes in Computer Science, 7180:37–45, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28717-6_6/.

Alberti:2012:LAI

[1020] Francesco Alberti and Roberto Bruttomesso. Lazy abstraction with interpolants for arrays. Lecture Notes in Computer Science, 7180:46–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28717-6_7/.

Alpuente:2012:BTS

[1021] María Alpuente and Demis Ballis. Backward trace slicing for conditional rewrite theories. Lecture Notes in Computer Science, 7180:62– 76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28717-6_8/.

Antoniou:2012:FDL

[1022] Grigoris Antoniou, Thomas Eiter, and Kewen Wang. Forgetting for defeasible logic. Lecture Notes in Computer Science, 7180:77–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28717-6_9/.

Aspinall:2012:QP

[1023] David Aspinall, Ewen Denney, and Christoph Lüth. Querying proofs. Lecture Notes in Computer Science, 7180:92–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28717-6_10/.

Baader:2012:SLE

[1024] Franz Baader and Alexander Okhotin. Solving language equations and disequations with applications to disunification in description logics and monadic set constraints. Lecture Notes in Computer Science, 7180:107– 121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28717-6_11/.

Benes:2012:DPM

[1025] Nikola Beneš and Jan Křetínský. Dual-priced modal transition systems with time durations. Lecture Notes in Computer Science, 7180:122– 137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28717-6_12/.

Borgwardt:2012:FFH

[1026] Stefan Borgwardt and Barbara Morawska. Finding finite Herbrand models. Lecture Notes in Computer Science, 7180:138–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28717-6_13/.

Bulwahn:2012:STF

[1027] Lukas Bulwahn. Smart testing of functional programs in Isabelle. Lecture Notes in Computer Science, 7180:153–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28717-6_14/.

Bulychev:2012:MBS

[1028] Peter Bulychev and Alexandre David. Monitor-based statistical model checking for weighted metric temporal logic. Lecture Notes in Computer Science, 7180:168–182, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28717-6_15/.

Chacon:2012:DBM

[1029] José Luis Chacón and Ramón Pino Pérez. Duality between merging operators and social contraction operators. Lecture Notes in Computer Science, 7180:183–196, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28717-6_16/.

Fietzke:2012:AGI

[1030] Arnaud Fietzke and Evgeny Kruglov. Automatic generation of invariants for circular derivations in SUP(LA). Lecture Notes in Computer Science, 7180:197–211, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28717-6_17/.

Han:2012:MRU

[1031] The Anh Han and Ari Saptawijaya. Moral reasoning under uncertainty. Lecture Notes in Computer Science, 7180:212–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28717-6_18/.

Hetzl:2012:TAC

[1032] Stefan Hetzl and Alexander Leitsch. Towards algorithmic cutintroduction. Lecture Notes in Computer Science, 7180:228–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-28717-6_
19/.

Katebi:2012:CAS

[1033] Hadi Katebi and Karem A. Sakallah. Conflict anticipation in the search for graph automorphisms. Lecture Notes in Computer Science, 7180:243– 257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28717-6_20/.

Anonymous:2012:FMby

[1034] Anonymous. Front matter. Lecture Notes in Computer Science, 7180: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28717-6/1.

Bangalore:2012:TOB

[1035] Srinivas Bangalore. Thinking outside the box for natural language processing. Lecture Notes in Computer Science, 7181:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28604-9_1/.

Gonzalez:2012:GBM

[1036] Aitor González, German Rigau, and Mauro Castillo. A graph-based method to improve WordNet domains. Lecture Notes in Computer Science, 7181:17–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28604-9_2/.

Yildirim:2012:CDH

[1037] Savaş Yıldırım and Tuğba Yıldız. Corpus-driven hyponym acquisition for Turkish language. Lecture Notes in Computer Science, 7181:29– 41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28604-9_3/.

Garcia:2012:ATE

[1038] Renato Domínguez García, Sebastian Schmidt, Christoph Rensing, and Ralf Steinmetz. Automatic taxonomy extraction in different languages using Wikipedia and minimal language-specific information. Lecture Notes in Computer Science, 7181:42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28604-9_4/.

Tan:2012:ODC

[1039] He Tan, Rajaram Kaliyaperumal, and Nirupama Benis. Ontology-driven construction of domain corpus with frame semantics annotations. Lecture Notes in Computer Science, 7181:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28604-9_5/.

Abbas:2012:BHA

[1040] Qaiser Abbas. Building a hierarchical annotated corpus of Urdu: The URDU.KON-TB Treebank. Lecture Notes in Computer Science, 7181: 66–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28604-9_6/.

Carreras-Riudavets:2012:MAU

[1041] Francisco J. Carreras-Riudavets, Juan C. Rodríguez del Pino, and Zenón Hernández-Figueroa. A morphological analyzer using hash tables in main memory (MAHT) and a lexical knowledge base. Lecture Notes in Computer Science, 7181:80–91, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28604-9_7/.

Vasudevan:2012:OSI

[1042] N. Vasudevan and Pushpak Bhattacharyya. Optimal stem identification in presence of suffix list. Lecture Notes in Computer Science, 7181:92– 103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28604-9_8/.

Alfared:2012:ATP

[1043] Ramadan Alfared and Denis Béchet. On the adequacy of three POS taggers and a dependency parser. Lecture Notes in Computer Science, 7181:104–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28604-9_9/.

Nongmeikapam:2012:WIR

[1044] Kishorjit Nongmeikapam, Aribam Umananda Sharma, Laishram Martina Devi, and Nepoleon Keisam. Will the identification of reduplicated multiword expression (RMWE) improve the performance of SVM based Manipuri POS tagging? Lecture Notes in Computer Science, 7181:117– 129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28604-9_10/.

Kubon:2012:FWO

[1045] Vladislav Kuboň, Markéta Lopatková, and Martin Plátek. On formalization of word order properties. Lecture Notes in Computer Science, 7181:130–141, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28604-9_11/.

Ashraf:2012:CPO

[1046] Md. Izhar Ashraf and Sitabhra Sinha. Core-periphery organization of graphemes in written sequences: Decreasing positional rigidity with increasing core order. Lecture Notes in Computer Science, 7181:142–153, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28604-9_12/.

Bechet:2012:DLP

[1047] Nicolas Béchet, Peggy Cellier, Thierry Charnois, and Bruno Crémilleux. Discovering linguistic patterns using sequence mining. Lecture Notes in Computer Science, 7181:154–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28604-9_13/.

Quiniou:2012:WAS

[1048] Solen Quiniou, Peggy Cellier, Thierry Charnois, and Dominique Legallois. What about sequential data mining techniques to identify linguistic patterns for stylistics? Lecture Notes in Computer Science, 7181:166– 177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28604-9_14/.

Bajwa:2012:RSA

[1049] Imran Sarwar Bajwa, Mark Lee, and Behzad Bordbar. Resolving syntactic ambiguities in natural language specification of constraints. Lecture Notes in Computer Science, 7181:178–187, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28604-9_15/.

Liyanage:2012:CGS

[1050] Chamila Liyanage, Randil Pushpananda, Dulip Lakmal Herath, and Ruvan Weerasinghe. A computational grammar of Sinhala. Lecture Notes in Computer Science, 7181:188–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28604-9_16/.

Salehi:2012:AIP

[1051] Bahar Salehi, Narjes Askarian, and Afsaneh Fazly. Automatic identification of Persian light verb constructions. Lecture Notes in Computer Science, 7181:201–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28604-9_17/.

Dutta:2012:CAW

[1052] Sudakshina Dutta and Anupam Basu. A cognitive approach to word sense disambiguation. Lecture Notes in Computer Science, 7181:211– 224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28604-9_18/.

Gutierrez:2012:GBA

[1053] Yoan Gutiérrez, Sonia Vázquez, and Andrés Montoyo. A graph-based approach to WSD using relevant semantic trees and N-cliques model. Lecture Notes in Computer Science, 7181:225–237, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28604-9_19/.

Nguyen:2012:UWI

[1054] Kiem-Hieu Nguyen and Cheol-Young Ock. Using wiktionary to improve lexical disambiguation in multiple languages. Lecture Notes in Computer Science, 7181:238–248, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28604-9_20/.

Anonymous:2012:FMbz

[1055] Anonymous. Front matter. Lecture Notes in Computer Science, 7181: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28604-9/1.

Paukkeri:2012:EEL

[1056] Mari-Sanna Paukkeri, Jaakko Väyrynen, and Antti Arppe. Exploring extensive linguistic feature sets in near-synonym lexical choice. Lecture Notes in Computer Science, 7182:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28601-8_1/.

Klabunde:2012:AGF

[1057] Ralf Klabunde, Sebastian Reuße, and Björn Schlünder. Abduction in games for a flexible approach to discourse planning. Lecture Notes in Computer Science, 7182:13–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28601-8_2/.

Roukos:2012:DSS

[1058] Salim Roukos, Abraham Ittycheriah, and Jian-Ming Xu. Documentspecific statistical machine translation for improving human translation productivity. Lecture Notes in Computer Science, 7182:25–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28601-8_3/.

Okita:2012:MBR

[1059] Tsuyoshi Okita and Josef van Genabith. Minimum Bayes risk decoding with enlarged hypothesis space in system combination. Lecture Notes in Computer Science, 7182:40–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28601-8_4/.

Cao:2012:PSC

[1060] Hailong Cao, Eiichiro Sumita, Tiejun Zhao, and Sheng Li. Phrasal syntactic category sequence model for phrase-based MT. Lecture Notes in Computer Science, 7182:52–59, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28601-8_5/.

Mathur:2012:INC

[1061] Prashant Mathur and Soma Paul. Integration of a noun compound translator tool with Moses for English–Hindi machine translation and evaluation. Lecture Notes in Computer Science, 7182:60–71, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28601-8_6/.

Harastani:2012:NCA

[1062] Rima Harastani, Béatrice Daille, and Emmanuel Morin. Neoclassical compound alignments from comparable corpora. Lecture Notes in Computer Science, 7182:72–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28601-8_7/.

Hazem:2012:QNM

[1063] Amir Hazem and Emmanuel Morin. QAlign: a new method for bilingual lexicon extraction from comparable corpora. Lecture Notes in Computer Science, 7182:83–96, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28601-8_8/.

Petran:2012:AAP

[1064] Florian Petran. Aligning the un-alignable — a pilot study using a noisy corpus of nonstandardized, semi-parallel texts. Lecture Notes in Computer Science, 7182:97–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28601-8_9/.

Oliver:2012:PCW

[1065] Antoni Oliver and Salvador Climent. Parallel corpora for WordNet construction: Machine translation vs. automatic sense tagging. Lecture Notes in Computer Science, 7182:110–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28601-8_10/.

Yasuda:2012:MBB

[1066] Keiji Yasuda, Andrew Finch, and Eiichiro Sumita. Method to build a bilingual lexicon for speech-to-speech translation systems. Lecture Notes in Computer Science, 7182:122–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28601-8_11/.

Tripathi:2012:FST

[1067] Nandita Tripathi, Michael Oakes, and Stefan Wermter. A fast subspace text categorization method using parallel classifiers. Lecture Notes in Computer Science, 7182:132–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28601-8_12/.

Zheng:2012:RTC

[1068] Dequan Zheng, Chenghe Zhang, Geli Fei, and Tiejun Zhao. Research on text categorization based on a weakly-supervised transfer learning method. Lecture Notes in Computer Science, 7182:144–156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28601-8_13/.

Garcia-Plaza:2012:FCC

[1069] Alberto Pérez García-Plaza, Víctor Fresno, and Raquel Martínez. Fuzzy combinations of criteria: An application to Web page representation for clustering. Lecture Notes in Computer Science, 7182:157–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28601-8_14/.

Shrestha:2012:CST

[1070] Prajol Shrestha, Christine Jacquin, and Béatrice Daille. Clustering short text and its evaluation. Lecture Notes in Computer Science, 7182:169– 180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28601-8_15/.

Castillo:2012:IEW

[1071] Carlos Castillo, Héctor Valero, José Guadalupe Ramos, and Josep Silva. Information extraction from Webpages based on DOM distances. Lecture Notes in Computer Science, 7182:181–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28601-8_16/.

Li:2012:CFS

[1072] Qi Li, Javier Artiles, Taylor Cassidy, and Heng Ji. Combining flat and structured approaches for temporal slot filling or: How much to compress? Lecture Notes in Computer Science, 7182:194–205, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28601-8_17/.

Wonsever:2012:EAS

[1073] Dina Wonsever, Aiala Rosá, Marisa Malcuori, Guillermo Moncecchi, and Alan Descoins. Event annotation schemes and event recognition in Spanish texts. Lecture Notes in Computer Science, 7182:206–218, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28601-8_18/.

Arnulphy:2012:AGN

[1074] Béatrice Arnulphy, Xavier Tannier, and Anne Vilnat. Automatically generated noun lexicons for event extraction. Lecture Notes in Computer Science, 7182:219–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28601-8_19/.

Carroll:2012:LAC

[1075] John Carroll, Rob Koeling, and Shivani Puri. Lexical acquisition for clinical text mining using distributional similarity. Lecture Notes in Computer Science, 7182:232–246, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28601-8_20/.

Anonymous:2012:FMca

[1076] Anonymous. Front matter. Lecture Notes in Computer Science, 7182: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28601-8/1.

Asarin:2012:MIT

[1077] Eugene Asarin. Measuring information in timed languages. Lecture Notes in Computer Science, 7183:1–2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-28332-1_1.

Boigelot:2012:ABS

[1078] Bernard Boigelot, Julien Brusten, and Jean-François Degbomont. Automata-based symbolic representations of polyhedra. Lecture Notes in Computer Science, 7183:3–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28332-1_2/.

Dowek:2012:APC

[1079] Gilles Dowek. Around the physical Church–Turing thesis: Cellular automata, formal languages, and the principles of quantum theory. Lecture Notes in Computer Science, 7183:21–37, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28332-1_3/.

Downey:2012:PCT

[1080] Rod Downey. A parameterized complexity tutorial. Lecture Notes in Computer Science, 7183:38–56, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28332-1_4/.

Lutz:2012:CSD

[1081] Jack H. Lutz. The computer science of DNA Nanotechnology. Lecture Notes in Computer Science, 7183:57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/accesspage/chapter/10.1007/978-3-642-28332-1_5.

Abdulla:2012:MCR

[1082] Parosh Aziz Abdulla, Mohamed Faouzi Atig, and Jari Stenman. The minimal cost reachability problem in priced timed pushdown systems. Lecture Notes in Computer Science, 7183:58–69, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28332-1_6/.

Anantharaman:2012:UMC

[1083] Siva Anantharaman, Christopher Bouchard, Paliath Narendran, and Michael Rusinowitch. Unification modulo chaining. Lecture Notes in Computer Science, 7183:70–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28332-1_7/.

Arvind:2012:ITB

[1084] Vikraman Arvind and Yadu Vasudev. Isomorphism testing of Boolean functions computable by constant-depth circuits. Lecture Notes in Computer Science, 7183:83–94, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28332-1_8/.

Axelsen:2012:RMH

[1085] Holger Bock Axelsen. Reversible multi-head finite automata characterize reversible logarithmic space. Lecture Notes in Computer Science, 7183:95–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28332-1_9/.

Barash:2012:DCC

[1086] Mikhail Barash and Alexander Okhotin. Defining contexts in contextfree grammars. Lecture Notes in Computer Science, 7183:106–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28332-1_10/.

Bille:2012:LCE

[1087] Philip Bille, Inge Li Gørtz, and Jesper Kristensen. Longest common extensions via fingerprinting. Lecture Notes in Computer Science, 7183:119–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28332-1_11/.

Bille:2012:FCO

[1088] Philip Bille and Morten Stöckel. Fast and cache-oblivious dynamic programming with local dependencies. Lecture Notes in Computer Science, 7183:131–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28332-1_12/.

Bollig:2012:EIO

[1089] Beate Bollig and Tobias Pröger. An efficient implicit OBDD-based algorithm for maximal matchings. Lecture Notes in Computer Science, 7183:143–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28332-1_13/.

Bozzelli:2012:STG

[1090] Laura Bozzelli. Strong termination for gap-order constraint abstractions of counter systems. Lecture Notes in Computer Science, 7183:155–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28332-1_14/.

Cervelle:2012:CSB

[1091] Julien Cervelle. Covering space in the Besicovitch topology. Lecture Notes in Computer Science, 7183:169–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28332-1_15/.

Champarnaud:2012:ARE

[1092] Jean-Marc Champarnaud, Hadrien Jeanne, and Ludovic Mignot. Approximate regular expressions and their derivatives. Lecture Notes in Computer Science, 7183:179–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28332-1_16/.

Ciobanu:2012:CPN

[1093] Gabriel Ciobanu and G. Michele Pinna. Catalytic Petri nets are Turing complete. Lecture Notes in Computer Science, 7183:192–203, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28332-1_17/.

Dennunzio:2012:CCR

[1094] Alberto Dennunzio, Enrico Formenti, and Julien Provillard. Computational complexity of rule distributions of non-uniform cellular automata. Lecture Notes in Computer Science, 7183:204–215, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28332-1_18/.

Dube:2012:CGR

[1095] Danny Dubé, Mario Latendresse, and Pascal Tesson. Conservative groupoids recognize only regular languages. Lecture Notes in Computer Science, 7183:216–227, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28332-1_19/.

Forisek:2012:ACO

[1096] Michal Forišek, Lucia Keller, and Monika Steinová. Advice complexity of online coloring for paths. Lecture Notes in Computer Science, 7183:228– 239, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28332-1_20/.

Anonymous:2012:FMcb

[1097] Anonymous. Front matter. Lecture Notes in Computer Science, 7183: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28332-1/1.

Alhazov:2012:PMS

[1098] Artiom Alhazov. Properties of membrane systems. Lecture Notes in Computer Science, 7184:1–13, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28024-5_1/.

Beal:2012:BBE

[1099] Jacob Beal. Bridging biology and engineering together with spatial computing. Lecture Notes in Computer Science, 7184:14–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28024-5_2/.

Giavitto:2012:MSF

[1100] Jean-Louis Giavitto. The modeling and the simulation of the fluid machines of synthetic biology. Lecture Notes in Computer Science, 7184:19– 34, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28024-5_3/.

Nicolescu:2012:PDA

[1101] Radu Nicolescu. Parallel and distributed algorithms in P Systems. Lecture Notes in Computer Science, 7184:35–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28024-5_4/.

Vaszil:2012:VDP

[1102] György Vaszil. Variants of distributed P automata and the efficient parallelizability of languages. Lecture Notes in Computer Science, 7184:51– 61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28024-5_5/.

Agrigoroaiei:2012:QCM

[1103] Oana Agrigoroaiei and Gabriel Ciobanu. Quantitative causality in membrane systems. Lecture Notes in Computer Science, 7184:62–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28024-5_6/.

Alhazov:2012:CPM

[1104] Artiom Alhazov, Alexander Krassovitskiy, and Yurii Rogozhin. Circular post machines and P systems with exo-insertion and deletion. Lecture Notes in Computer Science, 7184:73–86, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28024-5_7/.

Cabarle:2012:SNP

[1105] Francis George C. Cabarle, Henry Adorna, and Miguel A. Martínez. A spiking neural P system simulator based on CUDA. Lecture Notes in Computer Science, 7184:87–103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28024-5_8/.

Cienciala:2012:MPC

[1106] Luděk Cienciala, Lucie Ciencialová, and Miroslav Langer. Modularity in P colonies with checking rules. Lecture Notes in Computer Science, 7184:104–119, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28024-5_9/.

Csuhaj-Varju:2012:FDA

[1107] Erzsébet Csuhaj-Varjú and György Vaszil. Finite dP automata versus multi-head finite automata. Lecture Notes in Computer Science, 7184:120–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28024-5_10/.

Dinneen:2012:AAP

[1108] Michael J. Dinneen, Yun-Bum Kim, and Radu Nicolescu. An adaptive algorithm for P System synchronization. Lecture Notes in Computer Science, 7184:139–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28024-5_11/.

Frisco:2012:PSA

[1109] Pierluigi Frisco and Gordon Govan. P systems with active membranes operating under minimal parallelism. Lecture Notes in Computer Science, 7184:165–181, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28024-5_12/.

Hinze:2012:CAC

[1110] Thomas Hinze, Christian Bodenstein, Benedict Schau, Ines Heiland, and Stefan Schuster. Chemical analog computers for clock frequency control based on P modules. Lecture Notes in Computer Science, 7184:182– 202, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28024-5_13/.

Huang:2012:EDS

[1111] Xiaoli Huang, Gexiang Zhang, Haina Rong, and Florentin Ipate. Evolutionary design of a simple membrane system. Lecture Notes in Computer Science, 7184:203–214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28024-5_14/.

Ipate:2012:FVP

[1112] Florentin Ipate, Raluca Lefticaru, Ignacio Pérez-Hurtado, and Mario J. Pérez-Jiménez. Formal verification of P systems with active membranes through model checking. Lecture Notes in Computer Science, 7184:215– 225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28024-5_15/.

Ivanov:2012:BCR

[1113] Sergiu Ivanov. Basic concurrency resolution in clock-free P systems. Lecture Notes in Computer Science, 7184:226–242, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28024-5_16/.

Pan:2012:AES

[1114] Linqiang Pan, Jun Wang, and Hendrik Jan Hoogeboom. Asynchronous extended spiking neural P systems with astrocytes. Lecture Notes in Computer Science, 7184:243–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28024-5_17/.

Macias–Ramos:2012:PLB

[1115] Luis F. Macías-Ramos, Ignacio Pérez-Hurtado, and Manuel García-Quismondo. A P-Lingua based simulator for spiking neural P systems. Lecture Notes in Computer Science, 7184:257–281, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28024-5_18/.

Manca:2012:CMM

[1116] Vincenzo Manca and Rosario Lombardo. Computing with multimembranes. Lecture Notes in Computer Science, 7184:282–299, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-28024-5_
19/.

Marchetti:2012:MBM

[1117] Luca Marchetti and Vincenzo Manca. A methodology based on MP theory for gene expression analysis. Lecture Notes in Computer Science, 7184:300–313, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28024-5_20/.

Anonymous:2012:FMcc

[1118] Anonymous. Front matter. Lecture Notes in Computer Science, 7184: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28024-5/1.

Fang:2012:MCR

[1119] Jun Fang, Zhisheng Huang, and Frank van Harmelen. A method of contrastive reasoning with inconsistent ontologies. Lecture Notes in Computer Science, 7185:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29923-0_1/.

Ren:2012:PAR

[1120] Yuan Ren, Jeff Z. Pan, and Kevin Lee. Parallel ABox reasoning of EL ontologies. Lecture Notes in Computer Science, 7185:17–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29923-0_2/.

Kim:2012:RFP

[1121] Kisung Kim, Bongki Moon, and Hyoung-Joo Kim. RP-filter: a pathbased triple filtering method for efficient SPARQL query processing. Lecture Notes in Computer Science, 7185:33–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29923-0_3/.

Zhang:2012:CVD

[1122] Hang Zhang, Wei Hu, and Yuzhong Qu. Constructing virtual documents for ontology matching using MapReduce. Lecture Notes in Computer Science, 7185:48–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29923-0_4/.

Fionda:2012:SFN

[1123] Valeria Fionda and Giuseppe Pirró. Semantic flow networks: Semantic interoperability in networks of ontologies. Lecture Notes in Computer Science, 7185:64–79, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29923-0_5/.

Wang:2012:BLS

[1124] Zhichun Wang, Zhigang Wang, Juanzi Li, and Jeff Z. Pan. Building a large scale knowledge base from Chinese Wiki encyclopedia. Lecture Notes in Computer Science, 7185:80–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29923-0_6/.

Kozaki:2012:DHG

[1125] Kouji Kozaki, Keisuke Hihara, and Riiciro Mizoguchi. Dynamic Is — a hierarchy generation system based on user’s viewpoint. Lecture Notes in Computer Science, 7185:96–111, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29923-0_7/.

Zhao:2012:MOL

[1126] Lihua Zhao and Ryutaro Ichise. Mid-ontology learning from linked data. Lecture Notes in Computer Science, 7185:112–127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29923-0_8/.

Celino:2012:OFO

[1127] Irene Celino and Daniele Dell’Aglio. An ontological formulation and an OPM profile for causality in planning applications. Lecture Notes in Computer Science, 7185:128–143, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29923-0_9/.

Du:2012:NMA

[1128] Jianfeng Du, Shuai Wang, Guilin Qi, Jeff Z. Pan, and Yong Hu. A new matchmaking approach based on abductive conjunctive query answering. Lecture Notes in Computer Science, 7185:144–159, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29923-0_10/.

Gao:2012:GGU

[1129] Qi Gao, Fabian Abel, and Geert-Jan Houben. GeniUS: Generic user modeling library for the social semantic Web. Lecture Notes in Computer Science, 7185:160–175, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29923-0_11/.

Tian:2012:ESS

[1130] Yuan Tian, Jürgen Umbrich, and Yong Yu. Enhancing source selection for live queries over linked data via query log mining. Lecture Notes in Computer Science, 7185:176–191, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29923-0_12/.

Yang:2012:SCS

[1131] Mengdong Yang and Gang Wu. Semantic caching for semantic Web applications. Lecture Notes in Computer Science, 7185:192–209, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29923-0_13/.

Wang:2012:EGT

[1132] Xin Wang, Thanassis Tiropanis, and Hugh C. Davis. Evaluating graph traversal algorithms for distributed SPARQL query optimization. Lecture Notes in Computer Science, 7185:210–225, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29923-0_14/.

Cheng:2012:BRS

[1133] Gong Cheng, Feng Ji, Shengmei Luo, Weiyi Ge, and Yuzhong Qu. BipRank: Ranking and summarizing RDF vocabulary descriptions. Lecture Notes in Computer Science, 7185:226–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29923-0_15/.

Horne:2012:OSS

[1134] Ross Horne, Vladimiro Sassone, and Nicholas Gibbins. Operational semantics for SPARQL update. Lecture Notes in Computer Science, 7185:242–257, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29923-0_16/.

Gu:2012:KDD

[1135] Peiqin Gu and Huajun Chen. Knowledge-driven diagnostic system for traditional Chinese medicine. Lecture Notes in Computer Science, 7185:258–267, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29923-0_17/.

Zhou:2012:LUL

[1136] Wenlei Zhou, Haofen Wang, Jiansong Chao, Weinan Zhang, and Yong Yu. LODDO: Using linked open data description overlap to measure semantic relatedness between named entities. Lecture Notes in Computer Science, 7185:268–283, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29923-0_18/.

Nikolov:2012:WSL

[1137] Andriy Nikolov, Mathieu d’Aquin, and Enrico Motta. What should I link to? Identifying relevant sources and classes for data linking. Lecture Notes in Computer Science, 7185:284–299, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29923-0_19/.

Haller:2012:ILD

[1138] Armin Haller, Tudor Groza, and Florian Rosenberg. Interacting with linked data via semantically annotated widgets. Lecture Notes in Computer Science, 7185:300–317, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29923-0_20/.

Anonymous:2012:FMcd

[1139] Anonymous. Front matter. Lecture Notes in Computer Science, 7185: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29923-0/1.

Barringer:2012:IVE

[1140] Howard Barringer and Klaus Havelund. Internal versus external DSLs for trace analysis. Lecture Notes in Computer Science, 7186:1–3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29860-8_1/.

Wang:2012:PCF

[1141] Chao Wang and Malay Ganai. Predicting concurrency failures in the generalized execution traces of x86 executables. Lecture Notes in Computer Science, 7186:4–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29860-8_2/.

Bonakdarpour:2012:RMT

[1142] Borzoo Bonakdarpour and Sebastian Fischmeister. Runtime monitoring of time-sensitive systems. Lecture Notes in Computer Science, 7186:19– 33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29860-8_3/.

Leucker:2012:TRV

[1143] Martin Leucker. Teaching runtime verification. Lecture Notes in Computer Science, 7186:34–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29860-8_4/.

Malik:2012:RVC

[1144] Sharad Malik. Runtime verification: a computer architecture perspective. Lecture Notes in Computer Science, 7186:49–62, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29860-8_5/.

Ziarek:2012:IDM

[1145] Lukasz Ziarek, Siddharth Tiwary, and Suresh Jagannathan. Isolating determinism in multi-threaded programs. Lecture Notes in Computer Science, 7186:63–77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29860-8_6/.

Jula:2012:EOI

[1146] Horatiu Jula, Silviu Andrica, and George Candea. Efficiency optimizations for implementations of deadlock immunity. Lecture Notes in Computer Science, 7186:78–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29860-8_7/.

Westbrook:2012:PRR

[1147] Edwin Westbrook, Jisheng Zhao, Zoran Budimlić, and Vivek Sarkar. Permission regions for race-free parallelism. Lecture Notes in Computer Science, 7186:94–109, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29860-8_8/.

Serebryany:2012:DRD

[1148] Konstantin Serebryany, Alexander Potapenko, Timur Iskhodzhanov, and Dmitriy Vyukov. Dynamic race detection with LLVM compiler. Lecture Notes in Computer Science, 7186:110–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29860-8_9/.

Milea:2012:NRA

[1149] Narcisa Andreea Milea, Siau Cheng Khoo, David Lo, and Cristian Pop. NORT: Runtime anomaly-based monitoring of malicious behavior for Windows. Lecture Notes in Computer Science, 7186:115–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29860-8_10/.

Maggi:2012:RVL

[1150] Fabrizio Maria Maggi, Michael Westergaard, Marco Montali, and Wil M. P. van der Aalst. Runtime verification of LTL-based declarative process models. Lecture Notes in Computer Science, 7186:131–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29860-8_11/.

Asarin:2012:PIT

[1151] Eugene Asarin, Alexandre Donzé, Oded Maler, and Dejan Nickovic. Parametric identification of temporal properties. Lecture Notes in Computer Science, 7186:147–160, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29860-8_12/.

Sumner:2012:MDA

[1152] William N. Sumner, Christian Hammer, and Julian Dolby. Marathon: Detecting atomic-set serializability violations with conflict graphs. Lecture Notes in Computer Science, 7186:161–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29860-8_13/.

Krena:2012:CMS

[1153] Bohuslav Křena, Zdeněk Letko, and Tomáš Vojnar. Coverage metrics for saturation-based and search-based testing of concurrent software. Lecture Notes in Computer Science, 7186:177–192, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29860-8_14/.

Stoller:2012:RVS

[1154] Scott D. Stoller, Ezio Bartocci, Justin Seyster, Radu Grosu, and Klaus Havelund. Runtime verification with state estimation. Lecture Notes in Computer Science, 7186:193–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29860-8_15/.

Navabpour:2012:ETN

[1155] Samaneh Navabpour, Chun Wah Wallace Wu, Borzoo Bonakdarpour, and Sebastian Fischmeister. Efficient techniques for near-optimal instrumentation in time-triggered runtime verification. Lecture Notes in Computer Science, 7186:208–222, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29860-8_16/.

Arcaini:2012:CCM

[1156] Paolo Arcaini, Angelo Gargantini, and Elvinia Riccobene. CoMA: Conformance monitoring of Java programs by abstract state machines. Lecture Notes in Computer Science, 7186:223–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29860-8_17/.

Reinbacher:2012:ATT

[1157] Thomas Reinbacher, Jörg Brauer, Daniel Schachinger, and Andreas Steininger. Automated test-trace inspection for microcontroller binary code. Lecture Notes in Computer Science, 7186:239–244, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29860-8_18/.

Reiss:2012:WMP

[1158] Steven P. Reiss and Alexander Tarvo. What is my program doing? Program dynamics in programmer’s terms. Lecture Notes in Computer Science, 7186:245–259, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29860-8_19/.

Basin:2012:AMR

[1159] David Basin, Felix Klaedtke, and Eugen Zălinescu. Algorithms for monitoring real-time properties. Lecture Notes in Computer Science, 7186:260–275, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29860-8_20/.

Anonymous:2012:FMce

[1160] Anonymous. Front matter. Lecture Notes in Computer Science, 7186: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29860-8/1.

Gjosteen:2012:NIV

[1161] Kristian Gjøsteen. The Norwegian Internet Voting Protocol. Lecture Notes in Computer Science, 7187:1–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32747-6_1/.

Spycher:2012:TTM

[1162] Oliver Spycher, Melanie Volkamer, and Reto Koenig. Transparency and technical measures to establish trust in Norwegian Internet voting. Lecture Notes in Computer Science, 7187:19–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32747-6_2/.

Allepuz:2012:IVS

[1163] Jordi Puiggalí Allepuz and Sandra Guasch Castelló. Internet voting system with cast as intended verification. Lecture Notes in Computer Science, 7187:36–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32747-6_3/.

DeYoung:2012:LLV

[1164] Henry DeYoung and Carsten Schürmann. Linear logical voting protocols. Lecture Notes in Computer Science, 7187:53–70, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32747-6_4/.

Schlapfer:2012:EVA

[1165] Michael Schläpfer, Rolf Haenni, Reto Koenig, and Oliver Spycher. Efficient vote authorization in coercion-resistant Internet voting. Lecture Notes in Computer Science, 7187:71–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32747-6_5/.

Heiderich:2012:BMM

[1166] Mario Heiderich, Tilman Frosch, Marcus Niemietz, and Jörg Schwenk. The bug that made me President: a browser- and Web-security case study on Helios voting. Lecture Notes in Computer Science, 7187:89– 103, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32747-6_6/.

Joaquim:2012:EHS

[1167] Rui Joaquim and Carlos Ribeiro. An efficient and highly sound voter verification technique and its implementation. Lecture Notes in Computer Science, 7187:104–121, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32747-6_7/.

Essex:2012:SLO

[1168] Aleksander Essex, Christian Henrich, and Urs Hengartner. Single layer optical-scan voting with fully distributed trust. Lecture Notes in Computer Science, 7187:122–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32747-6_8/.

Chaum:2012:PIV

[1169] David Chaum, Alex Florescu, Mridul Nandi, Stefan Popoveniuc, Jan Rubio, and Poorvi L. Vora. Paperless independently-verifiable voting. Lecture Notes in Computer Science, 7187:140–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-32747-6_9/.

Demirel:2012:FAP

[1170] Denise Demirel, Maria Henning, Peter Y. A. Ryan, Steve Schneider, and Melanie Volkamer. Feasibility analysis of Prêt à Voter for German federal elections. Lecture Notes in Computer Science, 7187:158– 173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32747-6_10/.

Schneider:2012:PVW

[1171] Steve Schneider, Sriramkrishnan Srinivasan, Chris Culnane, James Heather, and Zhe Xia. Prêt á Voter with write-ins. Lecture Notes in Computer Science, 7187:174–189, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32747-6_11/.

Bursuc:2012:TVD

[1172] Sergiu Bursuc, Gurchetan S. Grewal, and Mark D. Ryan. Trivitas: Voters directly verifying votes. Lecture Notes in Computer Science, 7187:190– 207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32747-6_12/.

Heiberg:2012:AVE

[1173] Sven Heiberg, Peeter Laud, and Jan Willemson. The application of Ivoting for Estonian parliamentary elections of 2011. Lecture Notes in Computer Science, 7187:208–223, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32747-6_13/.

Buckland:2012:TBP

[1174] Richard Buckland, Vanessa Teague, and Roland Wen. Towards best practice for E-election systems. Lecture Notes in Computer Science, 7187:224–241, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32747-6_14/.

Heather:2012:SEI

[1175] James Heather, Morgan Llewellyn, Vanessa Teague, and Roland Wen. On the side-effects of introducing E-voting. Lecture Notes in Computer Science, 7187:242–256, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32747-6_15/.

Anonymous:2012:BMac

[1176] Anonymous. Back matter. Lecture Notes in Computer Science, 7187: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32747-6/1.

Anonymous:2012:FMcf

[1177] Anonymous. Front matter. Lecture Notes in Computer Science, 7187: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32747-6/1.

Auer:2012:ITU

[1178] Peter Auer. Invited talk: UCRL and autonomous exploration. Lecture Notes in Computer Science, 7188:1, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

accesspage/chapter/10.1007/978-3-642-29946-9_1.

Kersting:2012:ITI

[1179] Kristian Kersting. Invited talk: Increasing representational power and scaling inference in reinforcement learning. Lecture Notes in Computer Science, 7188:2, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/accesspage/
chapter/10.1007/978-3-642-29946-9_2.

Stone:2012:ITP

[1180] Peter Stone. Invited talk: PRISM — practical RL: Representation, interaction, synthesis, and mortality. Lecture Notes in Computer Science, 7188:3, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349

(electronic). URL http://link.springer.com/accesspage/chapter/
10.1007/978-3-642-29946-9_3.

Szepesvari:2012:ITT

[1181] Csaba Szepesvári. Invited talk: Towards robust reinforcement learning algorithms. Lecture Notes in Computer Science, 7188:4, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/accesspage/chapter/10.
1007/978-3-642-29946-9_4.

Maes:2012:ADR

[1182] Francis Maes, Louis Wehenkel, and Damien Ernst. Automatic discovery of ranking formulas for playing with multi-armed bandits. Lecture Notes in Computer Science, 7188:5–17, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29946-9_5/.

Ong:2012:GDO

[1183] Sylvie C. W. Ong, Yuri Grinberg, and Joelle Pineau. Goal-directed online learning of predictive models. Lecture Notes in Computer Science, 7188: 18–29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29946-9_6/.

Robards:2012:GBA

[1184] Matthew Robards and Peter Sunehag. Gradient based algorithms with loss functions and kernels for improved on-policy control. Lecture Notes in Computer Science, 7188:30–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29946-9_7/.

Araya-Lopez:2012:ALM

[1185] Mauricio Araya-López, Olivier Buffet, Vincent Thomas, and François Charpillet. Active learning of MDP models. Lecture Notes in Computer Science, 7188:42–53, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29946-9_8/.

Lesner:2012:HAE

[1186] Boris Lesner and Bruno Zanuttini. Handling ambiguous effects in action learning. Lecture Notes in Computer Science, 7188:54–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29946-9_9/.

Nguyen:2012:FRL

[1187] Phuong Nguyen, Peter Sunehag, and Marcus Hutter. Feature reinforcement learning in practice. Lecture Notes in Computer Science, 7188:66– 77, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29946-9_10/.

Elkan:2012:RLB

[1188] Charles Elkan. Reinforcement learning with a bilinear Q function. Lecture Notes in Computer Science, 7188:78–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-29946-9_11/.

Geist:2012:PPB

[1189] Matthieu Geist and Bruno Scherrer. ℓ_{1}-penalized projected Bellman residual. Lecture Notes in Computer Science, 7188:89–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29946-9_12/.

Hoffman:2012:RLS

[1190] Matthew W. Hoffman, Alessandro Lazaric, Mohammad Ghavamzadeh, and Rémi Munos. Regularized least squares temporal difference learning with nested ℓ_{2} and ℓ_{1} penalization. Lecture Notes in Computer Science, 7188:102–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29946-9_13/.

Scherrer:2012:RLS

[1191] Bruno Scherrer and Matthieu Geist. Recursive least-squares learning with eligibility traces. Lecture Notes in Computer Science, 7188:115– 127, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29946-9_14/.

Tziortziotis:2012:VFA

[1192] Nikolaos Tziortziotis and Konstantinos Blekas. Value function approximation through sparse Bayesian modeling. Lecture Notes in Computer Science, 7188:128–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29946-9_15/.

Castro:2012:ACT

[1193] Pablo Samuel Castro and Doina Precup. Automatic construction of temporally extended actions for MDPs using bisimulation metrics. Lecture Notes in Computer Science, 7188:140–152, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29946-9_16/.

Levy:2012:UII

[1194] Kfir Y. Levy and Nahum Shimkin. Unified inter and intra options learning using policy gradient methods. Lecture Notes in Computer Science, 7188:153–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29946-9_17/.

Sairamesh:2012:OE

[1195] Munu Sairamesh and Balaraman Ravindran. Options with exceptions. Lecture Notes in Computer Science, 7188:165–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29946-9_18/.

Dimitrakakis:2012:RBR

[1196] Christos Dimitrakakis. Robust Bayesian reinforcement learning through tight lower bounds. Lecture Notes in Computer Science, 7188:177–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29946-9_19/.

Maes:2012:OLA

[1197] Francis Maes, Louis Wehenkel, and Damien Ernst. Optimized look-ahead tree search policies. Lecture Notes in Computer Science, 7188:189–200, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29946-9_20/.

Anonymous:2012:FMcg

[1198] Anonymous. Front matter. Lecture Notes in Computer Science, 7188: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29946-9/1.

Berger:2012:ARW

[1199] Andreas Berger and Eduard Natale. Assessing the real-world dynamics of DNS. Lecture Notes in Computer Science, 7189:1–14, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28534-9_1/.

Gehlen:2012:UBP

[1200] Vinicius Gehlen, Alessandro Finamore, Marco Mellia, and Maurizio M. Munafò. Uncovering the big players of the Web. Lecture Notes in Computer Science, 7189:15–28, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28534-9_2/.

Gebert:2012:IAT

[1201] Steffen Gebert, Rastin Pries, Daniel Schlosser, and Klaus Heck. Internet access traffic measurement and analysis. Lecture Notes in Computer Science, 7189:29–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28534-9_3/.

Chen:2012:IAA

[1202] Liang Chen, Xingang Shi, and Dah Ming Chiu. From ISP address announcement patterns to routing scalability. Lecture Notes in Computer Science, 7189:43–47, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28534-9_4/.

Timpanaro:2012:IUC

[1203] Juan Pablo Timpanaro, Isabelle Chrisment, and Olivier Festor. I2P’s usage characterization. Lecture Notes in Computer Science, 7189:48– 51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28534-9_5/.

Testa:2012:EAB

[1204] Claudio Testa, Dario Rossi, Ashwin Rao, and Arnaud Legout. Experimental assessment of BitTorrent completion time in heterogeneous TCP/uTP swarms. Lecture Notes in Computer Science, 7189:52–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28534-9_6/.

Fiadino:2012:STE

[1205] Pierdomenico Fiadino, Danilo Valerio, Fabio Ricciato, and Karin Anna Hummel. Steps towards the extraction of vehicular mobility patterns from 3G signaling data. Lecture Notes in Computer Science, 7189:66– 80, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28534-9_7/.

Jusko:2012:ISN

[1206] Jan Jusko and Martin Rehak. Identifying Skype nodes in the network exploiting mutual contacts. Lecture Notes in Computer Science, 7189: 81–84, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28534-9_8/.

Iacovazzi:2012:PFM

[1207] Alfonso Iacovazzi and Andrea Baiocchi. Padding and fragmentation for masking packet length statistics. Lecture Notes in Computer Science, 7189:85–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28534-9_9/.

Besiktas:2012:RTT

[1208] Cihangir Beşiktaş and Hacı Ali Mantar. Real-time traffic classification based on cosine similarity using sub-application vectors. Lecture Notes in Computer Science, 7189:89–92, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28534-9_10/.

Bartos:2012:TEF

[1209] Karel Bartos and Martin Rehak. Towards efficient flow sampling technique for anomaly detection. Lecture Notes in Computer Science, 7189:93–106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28534-9_11/.

Hafsaoui:2012:DPT

[1210] Aymen Hafsaoui, Guillaume Urvoy-Keller, Matti Siekkinen, and Denis Collange. Detecting and profiling TCP connections experiencing abnormal performance. Lecture Notes in Computer Science, 7189:107–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28534-9_12/.

Shavitt:2012:GIP

[1211] Yuval Shavitt and Noa Zilberman. Geographical Internet PoP level maps. Lecture Notes in Computer Science, 7189:121–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28534-9_13/.

Cui:2012:DTW

[1212] Heng Cui and Ernst Biersack. Distributed troubleshooting of Web sessions using clustering. Lecture Notes in Computer Science, 7189:125– 128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28534-9_14/.

Soderman:2012:UMI

[1213] Pehr Söderman, Markus Hidell, and Peter Sjödin. Using metadata to improve experiment reliability in shared environments. Lecture Notes in Computer Science, 7189:129–142, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28534-9_15/.

Deri:2012:TCD

[1214] Luca Deri, Simone Mainardi, and Francesco Fusco. tsdb: a compressed database for time series. Lecture Notes in Computer Science, 7189:143– 156, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28534-9_16/.

Bonelli:2012:FHP

[1215] Nicola Bonelli, Andrea Di Pietro, Stefano Giordano, and Gregorio Procissi. Flexible high performance traffic generation on commodity multi– core platforms. Lecture Notes in Computer Science, 7189:157–170, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28534-9_17/.

Silva:2012:INM

[1216] João Marco C. Silva and Solange Rito Lima. Improving network measurement efficiency through multiadaptive sampling. Lecture Notes in Computer Science, 7189:171–174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28534-9_18/.

Anonymous:2012:BMad

[1217] Anonymous. Back matter. Lecture Notes in Computer Science, 7189: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-28534-9/1.

Anonymous:2012:FMch

[1218] Anonymous. Front matter. Lecture Notes in Computer Science, 7189: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28534-9/1.

Bellinger:2012:PRC

[1219] Colin Bellinger and B. John Oommen. On the pattern recognition and classification of stochastically episodic events. Lecture Notes in Computer Science, 7190:1–35, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29356-6_1/.

Nguyen:2012:PRS

[1220] Linh Anh Nguyen and Andrzej Szałas. Paraconsistent reasoning for semantic Web agents. Lecture Notes in Computer Science, 7190:36–55, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29356-6_2/.

Memon:2012:AMC

[1221] Zulfiqar A. Memon and Jan Treur. An agent model for cognitive and affective empathic understanding of other agents. Lecture Notes in Computer Science, 7190:56–83, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29356-6_3/.

Jakiela:2012:MBS

[1222] Jacek Jakieła, Paweł Litwin, and Marcin Olech. Multiagent based simulation as a supply chain analysis workbench. Lecture Notes in Computer Science, 7190:84–104, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29356-6_4/.

Kotulski:2012:EDM

[1223] Leszek Kotulski and Adam Sedziwy. On the effective distribution and maintenance of knowledge represented by complementary graphs. Lecture Notes in Computer Science, 7190:105–120, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-29356-6_5/.

Godlewska:2012:ASM

[1224] Magdalena Godlewska. Agent system for managing distributed mobile interactive documents in knowledge-based organizations. Lecture Notes in Computer Science, 7190:121–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29356-6_6/.

Dzienkowski:2012:ACW

[1225] Bartłomiej Józef Dzieńkowski. Agent cooperation within adversarial teams in dynamic environment — key issues and development trends. Lecture Notes in Computer Science, 7190:146–169, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-29356-6_7/.

Kaminski:2012:PSB

[1226] Bogumil Kaminski and Maciej Latek. On pricing strategies of boundedly rational telecommunication operators. Lecture Notes in Computer Science, 7190:170–180, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-29356-6_8/.

Kamide:2012:RAT

[1227] Norihiro Kamide. Reasoning about time-dependent multi-agents: Foundations of theorem proving and model checking. Lecture Notes in Computer Science, 7190:181–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-29356-6_9/.

Neri:2012:LPM

[1228] Filippo Neri. Learning predictive models for financial time series by using agent based simulations. Lecture Notes in Computer Science, 7190:202– 221, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-29356-6_10/.

Anonymous:2012:BMae

[1229] Anonymous. Back matter. Lecture Notes in Computer Science, 7190: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-29356-6/1.

Anonymous:2012:FMci

[1230] Anonymous. Front matter. Lecture Notes in Computer Science, 7190: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-29356-6/1.

DeLathauwer:2012:BCA

[1231] Lieven De Lathauwer. Block component analysis, a new concept for blind source separation. Lecture Notes in Computer Science, 7191:1– 8, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28551-6_1/.

Michaeli:2012:PLE

[1232] Tomer Michaeli, Daniel Sigalov, and Yonina C. Eldar. Partially linear estimation with application to image deblurring using blurred/noisy image pairs. Lecture Notes in Computer Science, 7191:9–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28551-6_2/.

Chen:2012:CDL

[1233] Zhitang Chen and Laiwan Chan. Causal discovery for linear non-Gaussian acyclic models in the presence of latent Gaussian confounders. Lecture Notes in Computer Science, 7191:17–24, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28551-6_3/.

Haufe:2012:AIW

[1234] Stefan Haufe, Vadim V. Nikulin, and Guido Nolte. Alleviating the influence of weak data asymmetries on granger-causal analyses. Lecture Notes in Computer Science, 7191:25–33, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28551-6_4/.

Duan:2012:OPR

[1235] Zhiyao Duan, Gautham J. Mysore, and Paris Smaragdis. Online PLCA for real-time semi-supervised source separation. Lecture Notes in Computer Science, 7191:34–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28551-6_5/.

Loesch:2012:CRB

[1236] Benedikt Loesch and Bin Yang. Cramér–Rao bound for circular complex independent component analysis. Lecture Notes in Computer Science, 7191:42–49, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28551-6_6/.

Wang:2012:CNO

[1237] Ke Wang, Xiao-Feng Gong, and Qiu-Hua Lin. Complex non-orthogonal joint diagonalization based on LU and LQ decompositions. Lecture Notes in Computer Science, 7191:50–57, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28551-6_7/.

Deville:2012:EAQ

[1238] Yannick Deville and Alain Deville. Exact and approximate quantum independent component analysis for qubit uncoupling. Lecture Notes in Computer Science, 7191:58–65, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28551-6_8/.

Shen:2012:MJD

[1239] Hao Shen and Martin Kleinsteuber. A matrix joint diagonalization approach for complex independent vector analysis. Lecture Notes in Computer Science, 7191:66–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28551-6_9/.

Shen:2012:ASC

[1240] Hao Shen and Martin Kleinsteuber. Algebraic solutions to complex blind source separation. Lecture Notes in Computer Science, 7191:74– 81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28551-6_10/.

Yeredor:2012:SPS

[1241] Arie Yeredor. On the separation performance of the strong uncorrelating transformation when applied to generalized covariance and pseudocovariance matrices. Lecture Notes in Computer Science, 7191:82–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28551-6_11/.

Karhunen:2012:CCA

[1242] Juha Karhunen, Tele Hao, and Jarkko Ylipaavalniemi. A canonical correlation analysis based method for improving BSS of two related data sets. Lecture Notes in Computer Science, 7191:91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28551-6_12/.

Becker:2012:PBC

[1243] Julian Mathias Becker, Martin Spiertz, and Volker Gnann. A probabilitybased combination method for unsupervised clustering with application to blind source separation. Lecture Notes in Computer Science, 7191:99– 106, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28551-6_13/.

Slapak:2012:CMB

[1244] Alon Slapak and Arie Yeredor. Charrelation matrix based ICA. Lecture Notes in Computer Science, 7191:107–114, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28551-6_14/.

Palmer:2012:CFI

[1245] Jason A. Palmer and Scott Makeig. Contrast functions for independent subspace analysis. Lecture Notes in Computer Science, 7191:115– 122, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28551-6_15/.

Ylipaavalniemi:2012:DCS

[1246] Jarkko Ylipaavalniemi, Nima Reyhani, and Ricardo Vigário. Distributional convergence of subspace estimates in FastICA: a bootstrap study. Lecture Notes in Computer Science, 7191:123–130, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28551-6_16/.

LeCorff:2012:NOE

[1247] Sylvain Le Corff, Gersende Fort, and Eric Moulines. New online EM algorithms for general hidden Markov models. application to the SLAM problem. Lecture Notes in Computer Science, 7191:131–138, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28551-6_17/.

Almeida:2012:RWS

[1248] Miguel Almeida, Ricardo Vigário, and José Bioucas-Dias. The role of whitening for separation of synchronous sources. Lecture Notes in Computer Science, 7191:139–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28551-6_18/.

Meinecke:2012:SDS

[1249] Frank C. Meinecke. Simultaneous diagonalization of skew-symmetric matrices in the symplectic group. Lecture Notes in Computer Science, 7191:147–154, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28551-6_19/.

Lahat:2012:JBD

[1250] Dana Lahat, Jean-François Cardoso, and Hagit Messer. Joint block diagonalization algorithms for optimal separation of multidimensional components. Lecture Notes in Computer Science, 7191:155–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28551-6_20/.

Anonymous:2012:FMcj

[1251] Anonymous. Front matter. Lecture Notes in Computer Science, 7191: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28551-6/1.

Lee:2012:UGU

[1252] Changhyun Lee, D. K. Lee, and Sue Moon. Unmasking the growing UDP traffic in a campus network. Lecture Notes in Computer Science, 7192: 1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28537-0_1/.

Sarrar:2012:IIT

[1253] Nadi Sarrar, Gregor Maier, Bernhard Ager, Robin Sommer, and Steve Uhlig. Investigating IPv6 traffic. Lecture Notes in Computer Science, 7192:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28537-0_2/.

Reggani:2012:EHV

[1254] Ahlem Reggani, Fabian Schneider, and Renata Teixeira. An end-host view on local traffic at home and work. Lecture Notes in Computer Science, 7192:21–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28537-0_3/.

Heikkinen:2012:CUT

[1255] Mikko V. J. Heikkinen and Arthur W. Berger. Comparison of user traffic characteristics on mobile-access versus fixed-access networks. Lecture Notes in Computer Science, 7192:32–41, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28537-0_4/.

Lyra:2012:BDS

[1256] Christian Lyra, Carmem S. Hara, and Elias P. Duarte, Jr. Back-StreamDB: a distributed system for backbone traffic monitoring providing arbitrary measurements in real-time. Lecture Notes in Computer Science, 7192:42–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28537-0_5/.

Golab:2012:SOS

[1257] Lukasz Golab, Theodore Johnson, Subhabrata Sen, and Jennifer Yates. A sequence-oriented stream warehouse paradigm for network monitoring applications. Lecture Notes in Computer Science, 7192:53–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28537-0_6/.

Bonelli:2012:MGP

[1258] Nicola Bonelli, Andrea Di Pietro, Stefano Giordano, and Gregorio Procissi. On multi–gigabit packet capturing with multi–core commodity hardware. Lecture Notes in Computer Science, 7192:64–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28537-0_7/.

Beyene:2012:SSA

[1259] Yordanos Beyene, Michalis Faloutsos, and Harsha V. Madhyastha. SyFi: a systematic approach for estimating stateful firewall performance. Lecture Notes in Computer Science, 7192:74–84, 2012. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28537-0_8/.

Rotsos:2012:OOF

[1260] Charalampos Rotsos, Nadi Sarrar, Steve Uhlig, Rob Sherwood, and Andrew W. Moore. OFLOPS: An open framework for OpenFlow switch evaluation. Lecture Notes in Computer Science, 7192:85–95, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28537-0_9/.

DiCioccio:2012:PPU

[1261] Lucas DiCioccio, Renata Teixeira, Martin May, and Christian Kreibich. Probe and pray: Using UPnP for home network measurements. Lecture Notes in Computer Science, 7192:96–105, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28537-0_10/.

Rutgaizer:2012:DPA

[1262] Moshe Rutgaizer, Yuval Shavitt, Omer Vertman, and Noa Zilberman. Detecting pedophile activity in BitTorrent networks. Lecture Notes in Computer Science, 7192:106–115, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28537-0_11/.

Konte:2012:RWA

[1263] Maria Konte and Nick Feamster. Re-wiring activity of malicious networks. Lecture Notes in Computer Science, 7192:116–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28537-0_12/.

Barbosa:2012:DMS

[1264] Rafael R. R. Barbosa, Ramin Sadre, and Aiko Pras. Difficulties in modeling SCADA traffic: a comparative analysis. Lecture Notes in Computer Science, 7192:126–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28537-0_13/.

Elmokashfi:2012:CDN

[1265] Ahmed Elmokashfi, Amund Kvalbein, Jie Xiang, and Kristian R. Evensen. Characterizing delays in Norwegian 3G networks. Lecture Notes in Computer Science, 7192:136–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28537-0_14/.

Tie:2012:GWL

[1266] Xiaozheng Tie, Kishore Ramachandran, and Rajesh Mahindra. On 60 GHz wireless link performance in indoor environments. Lecture Notes in Computer Science, 7192:147–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28537-0_15/.

Triukose:2012:GIA

[1267] Sipat Triukose, Sebastien Ardon, Anirban Mahanti, and Aaditeshwar Seth. Geolocating IP addresses in cellular data networks. Lecture Notes in Computer Science, 7192:158–167, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28537-0_16/.

Goga:2012:SMR

[1268] Oana Goga and Renata Teixeira. Speed measurements of residential Internet access. Lecture Notes in Computer Science, 7192:168–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28537-0_17/.

Brownlee:2012:OWT

[1269] Nevil Brownlee. One-way traffic monitoring with iatmon. Lecture Notes in Computer Science, 7192:179–188, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28537-0_18/.

deDonato:2012:HLA

[1270] Walter de Donato, Pietro Marchetta, and Antonio Pescapé. A hands-on look at active probing using the IP prespecified timestamp option. Lecture Notes in Computer Science, 7192:189–199, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28537-0_19/.

Wahlisch:2012:ENC

[1271] Matthias Wählisch, Thomas C. Schmidt, Markus de Brün, and Thomas Häberlen. Exposing a nation-centric view on the German Internet — a change in perspective on AS-level. Lecture Notes in Computer Science, 7192:200–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28537-0_20/.

Anonymous:2012:FMck

[1272] Anonymous. Front matter. Lecture Notes in Computer Science, 7192: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28537-0/1.

Schilling:2012:CFT

[1273] Thomas Schilling. Constraint-free type error slicing. Lecture Notes in Computer Science, 7193:1–16, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32037-8_1/.

Komendantsky:2012:SFI

[1274] Vladimir Komendantsky. Subtyping by folding an inductive relation into a coinductive one. Lecture Notes in Computer Science, 7193:17– 32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32037-8_2/.

Brady:2012:ELG

[1275] Edwin Brady. Epic — a library for generating compilers. Lecture Notes in Computer Science, 7193:33–48, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32037-8_3/.

Day:2012:TMC

[1276] Laurence E. Day and Graham Hutton. Towards modular compilers for effects. Lecture Notes in Computer Science, 7193:49–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-32037-8_4/.

Filliatre:2012:FDC

[1277] Jean-Christophe Filliâtre and K. Kalyanasundaram. Functory: a distributed computing library for objective caml. Lecture Notes in Computer Science, 7193:65–81, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32037-8_5/.

Brown:2012:PFP

[1278] Christopher Brown, Hans-Wolfgang Loidl, and Kevin Hammond. ParaForming: Forming parallel Haskell programs using novel refactoring techniques. Lecture Notes in Computer Science, 7193:82–97, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32037-8_6/.

Berthold:2012:FHP

[1279] Jost Berthold, Andrzej Filinski, Fritz Henglein, Ken Friis Larsen, and Mogens Steffensen. Functional high performance financial IT. Lecture Notes in Computer Science, 7193:98–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-32037-8_7/.

Lesniak:2012:TSP

[1280] Michael Lesniak. Thread-safe priority queues in Haskell based on skiplists. Lecture Notes in Computer Science, 7193:114–129, 2012. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer.com/chapter/10.1007/978-3-642-32037-8_
8/.

Straka:2012:ATR

[1281] Milan Straka. Adams’ trees revisited. Lecture Notes in Computer Science, 7193:130–145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32037-8_9/.

Morazan:2012:FVG

[1282] Marco T. Morazán. Functional video games in CS1 II. Lecture Notes in Computer Science, 7193:146–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-32037-8_10/.

Henrix:2012:GGL

[1283] Jeroen Henrix, Rinus Plasmeijer, and Peter Achten. GiN: a graphical language and tool for defining iTask workflows. Lecture Notes in Computer Science, 7193:163–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-32037-8_11/.

Delbianco:2012:ASF

[1284] Germán Andrés Delbianco, Mauro Jaskelioff, and Alberto Pardo. Applicative shortcut fusion. Lecture Notes in Computer Science, 7193:179– 194, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-32037-8_12/.

Anonymous:2012:BMaf

[1285] Anonymous. Back matter. Lecture Notes in Computer Science, 7193: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:978-3-
642-32037-8/1.

Anonymous:2012:FMcl

[1286] Anonymous. Front matter. Lecture Notes in Computer Science, 7193: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-32037-8/1.

Lauesen:2012:WEL

[1287] Soren Lauesen. Why the electronic land registry failed. Lecture Notes in Computer Science, 7195:1–15, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28714-5_1/.

Paech:2012:ARP

[1288] Barbara Paech, Robert Heinrich, Gabriele Zorn-Pauli, Andreas Jung, and Siamak Tadjiky. Answering a request for proposal — challenges and proposed solutions. Lecture Notes in Computer Science, 7195:16– 29, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28714-5_2/.

Nekvi:2012:IRC

[1289] Md. Rashed Iqbal Nekvi, Nazim H. Madhavji, Remo Ferrari, and Brian Berenbach. Impediments to requirements-compliance. Lecture Notes in Computer Science, 7195:30–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28714-5_3/.

Poort:2012:HAS

[1290] Eltjo R. Poort, Nick Martens, Inge van de Weerd, and Hans van Vliet. How architects see non-functional requirements: Beware of modifiability. Lecture Notes in Computer Science, 7195:37–51, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28714-5_4/.

Koziolek:2012:RPP

[1291] Anne Koziolek. Research preview: Prioritizing quality requirements based on software architecture evaluation feedback. Lecture Notes in Computer Science, 7195:52–58, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28714-5_5/.

Wang:2012:SAI

[1292] Junjie Wang, Juan Li, Qing Wang, He Zhang, and Haitao Wang. A simulation approach for impact analysis of requirement volatility considering dependency change. Lecture Notes in Computer Science, 7195: 59–76, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28714-5_6/.

Anh:2012:CRR

[1293] Nguyen Duc Anh, Daniela S. Cruzes, Reidar Conradi, Martin Höst, and Xavier Franch. Collaborative resolution of requirements mismatches when adopting open source components. Lecture Notes in Computer Science, 7195:77–93, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28714-5_7/.

Gulke:2012:HLR

[1294] Tim Gülke, Bernhard Rumpe, Martin Jansen, and Joachim Axmann. High-level requirements management and complexity costs in automotive development projects: a problem statement. Lecture Notes in Computer Science, 7195:94–100, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28714-5_8/.

Mahaux:2012:CYC

[1295] Martin Mahaux, Alistair Mavin, and Patrick Heymans. Choose your creativity: Why and how creativity in requirements engineering means different things to different people. Lecture Notes in Computer Science, 7195:101–116, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28714-5_9/.

Raspotnig:2012:SFM

[1296] Christian Raspotnig and Andreas Opdahl. Supporting failure mode and effect analysis: a case study with failure sequence diagrams. Lecture Notes in Computer Science, 7195:117–131, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28714-5_10/.

Chowdhury:2012:AMA

[1297] Mohammad Jabed Morshed Chowdhury, Raimundas Matulevičius, Guttorm Sindre, and Peter Karpati. Aligning mal-activity diagrams and security risk management for security requirements definitions. Lecture Notes in Computer Science, 7195:132–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28714-5_11/.

Genon:2012:TMS

[1298] Nicolas Genon, Patrice Caire, Hubert Toussaint, Patrick Heymans, and Daniel Moody. Towards a more semantically transparent i* visual syntax. Lecture Notes in Computer Science, 7195:140–146, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28714-5_12/.

Adam:2012:PSP

[1299] Sebastian Adam. Providing software product line knowledge to requirements engineers — a template for elicitation instructions. Lecture Notes in Computer Science, 7195:147–164, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28714-5_13/.

Knauss:2012:SLO

[1300] Eric Knauss and Kurt Schneider. Supporting learning organisations in writing better requirements documents based on heuristic critiques. Lecture Notes in Computer Science, 7195:165–171, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28714-5_14/.

Daramola:2012:MIR

[1301] Olawande Daramola, Thomas Moser, Guttorm Sindre, and Stefan Biffl. Managing implicit requirements using semantic case-based reasoning research preview. Lecture Notes in Computer Science, 7195:172–178, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28714-5_15/.

Cleland-Huang:2012:TQS

[1302] Jane Cleland-Huang, Mats Heimdahl, Jane Huffman Hayes, Robyn Lutz, and Patrick Maeder. Trace queries for safety requirements in high assurance systems. Lecture Notes in Computer Science, 7195:179–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28714-5_16/.

Li:2012:WTV

[1303] Yang Li and Walid Maalej. Which traceability visualization is suitable in this context? A comparative study. Lecture Notes in Computer Science, 7195:194–210, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28714-5_17/.

Berry:2012:CDR

[1304] Daniel Berry, Ricardo Gacitua, Pete Sawyer, and Sri Fatimah Tjong. The case for dumb requirements engineering tools. Lecture Notes in Computer Science, 7195:211–217, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28714-5_18/.

Bruni:2012:AAM

[1305] Elia Bruni, Alessio Ferrari, Norbert Seyff, and Gabriele Tolomei. Automatic analysis of multimodal requirements: a research preview. Lecture Notes in Computer Science, 7195:218–224, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28714-5_19/.

Hochmuller:2012:MSQ

[1306] Elke Hochmüller. 10 myths of software quality. Lecture Notes in Computer Science, 7195:225–231, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28714-5_20/.

Anonymous:2012:FMcm

[1307] Anonymous. Front matter. Lecture Notes in Computer Science, 7195: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28714-5/1.

Skorupa:2012:MAS

[1308] Grzegorz Skorupa, Wojciech Lorkiewicz, and Radosław Katarzyniak. A multi-agent strategy for integration of imprecise descriptions. Lecture Notes in Computer Science, 7196:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28487-8_1/.

Wilkosz:2012:PEM

[1309] Kazimierz Wilkosz, Zofia Kruczkiewicz, and Tomasz Babczyński. Performance evaluation of multiagent-system oriented models for efficient power system topology verification. Lecture Notes in Computer Science, 7196:11–20, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28487-8_2/.

Khue:2012:BMI

[1310] Nguyen Tran Minh Khue and Nhon Van Do. Building a model of an intelligent multi-agent system based on distributed knowledge bases for solving problems automatically. Lecture Notes in Computer Science, 7196: 21–32, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28487-8_3/.

Hsieh:2012:TRM

[1311] Fu-Shiung Hsieh and Jim-Bon Lin. Temporal reasoning in multi-agent workflow systems based on formal models. Lecture Notes in Computer Science, 7196:33–42, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28487-8_4/.

Nguyen:2012:ARA

[1312] Vinh Gia Nhi Nguyen, Hiep Xuan Huynh, and Alexis Drogoul. Assessing rice area infested by brown plant hopper using agent-based and dynamically upscaling approach. Lecture Notes in Computer Science, 7196:43– 52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28487-8_5/.

Mannava:2012:SAD

[1313] Vishnuvardhan Mannava and T. Ramesh. A service administration design pattern for dynamically configuring communication services in autonomic computing systems. Lecture Notes in Computer Science, 7196: 53–63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28487-8_6/.

Liu:2012:CPF

[1314] Jialin Liu. Chemical process fault diagnosis based on sensor validation approach. Lecture Notes in Computer Science, 7196:64–73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28487-8_7/.

Brzostowski:2012:SAT

[1315] Krzysztof Brzostowski, Jarosław Drapała, and Jerzy Światek. System analysis techniques in eHealth systems: a case study. Lecture Notes in Computer Science, 7196:74–85, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28487-8_8/.

Wang:2012:DFF

[1316] Hsueh-Wu Wang, Ying-Ming Wu, Yen-Ling Lu, and Ying-Tung Hsiao. Detection of facial features on color face images. Lecture Notes in Computer Science, 7196:86–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28487-8_9/.

Wu:2012:ALD

[1317] YuLung Wu. Adaptive learning diagnosis mechanisms for E-learning. Lecture Notes in Computer Science, 7196:102–110, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28487-8_10/.

Yeh:2012:NIT

[1318] Wei-Ming Yeh. New integration technology for video virtual reality. Lecture Notes in Computer Science, 7196:111–117, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28487-8_11/.

Chu:2012:ISW

[1319] Hai-Cheng Chu and Szu-Wei Yang. Innovative semantic Web services for next generation academic electronic library via Web 3.0 via distributed artificial intelligence. Lecture Notes in Computer Science, 7196:118–124, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28487-8_12/.

Chen:2012:UFR

[1320] Shyi-Ming Chen, Yun-Hou Huang, Rung-Ching Chen, Szu-Wei Yang, and Tian-Wei Sheu. Using fuzzy reasoning techniques and the domain ontology for anti-diabetic drugs recommendation. Lecture Notes in Computer Science, 7196:125–135, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28487-8_13/.

Sheu:2012:CAI

[1321] Jia-Shing Sheu, Yi-Ching Kao, and Hao Chu. Content-aware image resizing based on aesthetic. Lecture Notes in Computer Science, 7196:136– 145, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28487-8_14/.

Do:2012:FKM

[1322] Nhon Do and Thu-Le Pham. A functional knowledge model and application. Lecture Notes in Computer Science, 7196:146–155, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28487-8_15/.

Duong:2012:LNE

[1323] Trong Hai Duong, Hai Bang Truong, and Ngoc Thanh Nguyen. Local neighbor enrichment for ontology integration. Lecture Notes in Computer Science, 7196:156–166, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28487-8_16/.

Liu:2012:NCI

[1324] Hsiang-Chuan Liu, Shang-Ling Ou, Hsien-Chang Tsai, Yih-Chang Ou, and Yen-Kuei Yu. A novel Choquet integral composition forecasting model based on M-density. Lecture Notes in Computer Science, 7196:167–176, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28487-8_17/.

Sugiyama:2012:AMR

[1325] Shota Sugiyama, Hidemi Yamachi, Munehiro Takimoto, and Yasushi Kambayashi. Aggregating multiple robots with serialization. Lecture Notes in Computer Science, 7196:177–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28487-8_18/.

Truong:2012:MAM

[1326] Hai Bang Truong and Ngoc Thanh Nguyen. A multi-attribute and multivalued model for fuzzy ontology integrationon instance level. Lecture Notes in Computer Science, 7196:187–197, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28487-8_19/.

Satta:2012:MAR

[1327] Keisuke Satta, Munehiro Takimoto, and Yasushi Kambayashi. Making autonomous robots form lines. Lecture Notes in Computer Science, 7196:198–207, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28487-8_20/.

Anonymous:2012:FMcn

[1328] Anonymous. Front matter. Lecture Notes in Computer Science, 7196: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28487-8/1.

Ngo:2012:AIS

[1329] Long Thanh Ngo and Binh Huy Pham. Approach to image segmentation based on interval type-2 fuzzy subtractive clustering. Lecture Notes in Computer Science, 7197:1–10, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28490-8_1/.

Zomorodian:2012:INN

[1330] M. Javad Zomorodian, Ali Adeli, Mehrnoosh Sinaee, and Sattar Hashemi. Improving nearest neighbor classification by elimination of noisy irrelevant features. Lecture Notes in Computer Science, 7197:11– 21, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28490-8_2/.

Kumar:2012:LBA

[1331] Naveen Kumar, Anamika Gupta, and Vasudha Bhatnagar. Lattice based associative classifier. Lecture Notes in Computer Science, 7197:22–31, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28490-8_3/.

Shieh:2012:ECA

[1332] Shu-Ling Shieh, Tsu-Chun Lin, and Yu-Chin Szu. An efficient clustering algorithm based on histogram threshold. Lecture Notes in Computer Science, 7197:32–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28490-8_4/.

Wu:2012:RRM

[1333] Mary Wu, InTaek Leem, Jason J. Jung, and ChongGun Kim. A resource reuse method in cluster sensor networks in ad hoc networks. Lecture Notes in Computer Science, 7197:40–50, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28490-8_5/.

Kim:2012:GTB

[1334] Hyon Hee Kim, Jinnam Jo, and Donggeon Kim. Generation of tag-based user profiles for clustering users in a social music site. Lecture Notes in Computer Science, 7197:51–61, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28490-8_6/.

Chiu:2012:PIB

[1335] Tzu-Fu Chiu, Chao-Fu Hong, and Yu-Ting Chiu. A proposed IPC-based clustering and applied to technology strategy formulation. Lecture Notes in Computer Science, 7197:62–72, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28490-8_7/.

Thinakaran:2012:CCM

[1336] Rajermani Thinakaran and Elankovan Sundararajan. Cluster control management as cluster middleware. Lecture Notes in Computer Science, 7197:73–82, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28490-8_8/.

He:2012:NNA

[1337] Xin He, Huiyun Jing, Qi Han, and Xiamu Niu. A novel nonparametric approach for saliency detection using multiple features. Lecture Notes in Computer Science, 7197:83–90, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28490-8_9/.

Jing:2012:MVB

[1338] Huiyun Jing, Xin He, Qi Han, and Xiamu Niu. Motion vector based information hiding algorithm for H.264/AVC against motion vector steganalysis. Lecture Notes in Computer Science, 7197:91–98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28490-8_10/.

Niu:2012:NCM

[1339] Xiamu Niu, Zhongwei Shuai, Yongqiang Lin, and Xuehu Yan. A novel coding method for multiple system barcode based on QR code. Lecture Notes in Computer Science, 7197:99–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28490-8_11/.

Guo:2012:RBS

[1340] Xiaolv Guo, Shu-Chuan Chu, Lin-Lin Tang, John F. Roddick, and Jeng-Shyang Pan. A research on behavior of sleepy lizards based on KNN algorithm. Lecture Notes in Computer Science, 7197:109–118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28490-8_12/.

Yan:2012:DDA

[1341] Lijun Yan, Shu-Chuan Chu, John F. Roddick, and Jeng-Shyang Pan. Directional discriminant analysis based on nearest feature line. Lecture Notes in Computer Science, 7197:119–128, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28490-8_13/.

Kim:2012:SSS

[1342] Cheonshik Kim, Dongkyoo Shin, Dongil Shin, and Ching-Nung Yang. A (2, 2) secret sharing scheme based on Hamming code and AMBTC. Lecture Notes in Computer Science, 7197:129–139, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28490-8_14/.

Huang:2012:AII

[1343] Jen-Chi Huang and Wen-Shyong Hsieh. An automatic image inpainting method for rigid moving object. Lecture Notes in Computer Science, 7197:140–147, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28490-8_15/.

Hu:2012:AIM

[1344] Wu-Chih Hu and Jung-Fu Hsu. Automatic image matting using component-hue-difference-based spectral matting. Lecture Notes in Computer Science, 7197:148–157, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28490-8_16/.

Samani:2012:TRL

[1345] Hooman Aghaebrahimi Samani, Jeffrey Tzu Kwan Valino Koh, and Elham Saadatian. Towards robotics leadership: An analysis of leadership characteristics and the roles robots will inherit in future human society. Lecture Notes in Computer Science, 7197:158–165, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28490-8_17/.

Quan:2012:UIP

[1346] Meinu Quan, Xuan Hau Pham, Jason J. Jung, and Dosam Hwang. Understanding information propagation on online social tagging systems: a case study on Flickr. Lecture Notes in Computer Science, 7197:166– 174, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28490-8_18/.

Chung:2012:WPS

[1347] Namho Chung, Chulmo Koo, and Seung-Bae Park. Why people share information in social network sites? integrating with uses and gratification and social identity theories. Lecture Notes in Computer Science, 7197:175–184, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28490-8_19/.

Shen:2012:IDE

[1348] Chien wen Shen, Ping-Yu Hsu, and Yen-Ting Peng. The impact of data environment and profitability on business intelligence adoption. Lecture Notes in Computer Science, 7197:185–193, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28490-8_20/.

Anonymous:2012:FMco

[1349] Anonymous. Front matter. Lecture Notes in Computer Science, 7197: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28490-8/1.

Seo:2012:IHB

[1350] Young Wook Seo, Seong Wook Chae, and Kun Chang Lee. The impact of human brand image appeal on visual attention and purchase intentions at an E-commerce Website. Lecture Notes in Computer Science, 7198: 1–9, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28493-9_1/.

Chae:2012:EHB

[1351] Seong Wook Chae, Young Wook Seo, and Kun Chang Lee. Exploring human brands in online shopping: An eye-tracking approach. Lecture Notes in Computer Science, 7198:10–18, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28493-9_2/.

Jo:2012:TPU

[1352] Nam Yong Jo, Kun Chang Lee, and Dae Sung Lee. Task performance under stressed and non-stressed conditions: Emphasis on physiological approaches. Lecture Notes in Computer Science, 7198:19–26, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28493-9_3/.

Hahn:2012:CDM

[1353] Min Hee Hahn, Do Young Choi, and Kun Chang Lee. Characteristics of decision-making for different levels of product involvement depending on the degree of trust transfer: a comparison of cognitive decisionmaking criteria and physiological responses. Lecture Notes in Computer Science, 7198:27–36, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28493-9_4/.

Choi:2012:CBD

[1354] Do Young Choi, Min Hee Hahn, and Kun Chang Lee. A comparison of buying decision patterns by product involvement: An eye-tracking approach. Lecture Notes in Computer Science, 7198:37–46, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28493-9_5/.

Lee:2012:ITM

[1355] Dae Sung Lee, Kun Chang Lee, and Nam Yong Jo. The influence of team-member exchange on self-reported creativity in the Korean information and communication technology (ICT) industry. Lecture Notes in Computer Science, 7198:47–54, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28493-9_6/.

Wang:2012:SPS

[1356] Lingzhi Wang. Semi-parametric smoothing regression model based on GA for financial time series forecasting. Lecture Notes in Computer Science, 7198:55–64, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28493-9_7/.

Asaithambi:2012:CRA

[1357] Mythili Asaithambi, Sujatha C. Manoharan, and Srinivasan Subramanian. Classification of respiratory abnormalities using adaptive neuro fuzzy inference system. Lecture Notes in Computer Science, 7198:65– 73, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28493-9_8/.

Chen:2012:ACO

[1358] Nai-Hua Chen. An ant colony optimization and Bayesian network structure application for the asymmetric traveling salesman problem. Lecture Notes in Computer Science, 7198:74–78, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-642-28493-9_9/.

Jafari:2012:EPB

[1359] S. A. Jafari, S. Mashohor, Abd. R. Ramli, and M. Hamiruce Marhaban. Expert pruning based on genetic algorithm in regression problems. Lecture Notes in Computer Science, 7198:79–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28493-9_10/.

Ghodrati:2012:HCP

[1360] Amirhossein Ghodrati and Shahriar Lotfi. A hybrid CS/PSO algorithm for global optimization. Lecture Notes in Computer Science, 7198:89– 98, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28493-9_11/.

Ghodrati:2012:HIP

[1361] Amirhossein Ghodrati, Mohammad V. Malakooti, and Mansooreh Soleimani. A hybrid ICA/PSO algorithm by adding independent countries for large scale global optimization. Lecture Notes in Computer Science, 7198:99–108, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28493-9_12/.

Ramezani:2012:MDE

[1362] Fatemeh Ramezani and Shahriar Lotfi. The modified differential evolution algorithm (MDEA). Lecture Notes in Computer Science, 7198:109– 118, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28493-9_13/.

Soltani-Sarvestani:2012:QCA

[1363] M. A. Soltani-Sarvestani, Shahriar Lotfi, and Fatemeh Ramezani. Quad countries algorithm (QCA). Lecture Notes in Computer Science, 7198:119–129, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28493-9_14/.

Mannava:2012:AFM

[1364] Vishnuvardhan Mannava and T. Ramesh. An aspectual feature module based adaptive design pattern for autonomic computing systems. Lecture Notes in Computer Science, 7198:130–140, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28493-9_15/.

Noor:2012:MAA

[1365] Noorhuzaimi Karimah Mohd Noor, Shahrul Azman Noah, and Mohd Juzaiddin Ab Aziz. Malay anaphor and antecedent candidate identification: a proposed solution. Lecture Notes in Computer Science, 7198:141– 151, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28493-9_16/.

Viswanathan:2012:RSA

[1366] V. Viswanathan and Ilango Krishnamurthi. Ranking semantic associations between two entities — extended model. Lecture Notes in Computer Science, 7198:152–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28493-9_17/.

Machova:2012:QLM

[1367] Svetlana Machova and Jana Kleckova. The quantum of language: a metaphorical view of mind dimension. Lecture Notes in Computer Science, 7198:163–168, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28493-9_18/.

Kim:2012:ESE

[1368] Hee-Cheol Kim. An experimental study to explore usability problems of interactive voice response systems. Lecture Notes in Computer Science, 7198:169–177, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28493-9_19/.

Chynal:2012:UEM

[1369] Piotr Chynał, Jerzy M. Szymański, and Janusz Sobecki. Using eyetracking in a mobile applications usability testing. Lecture Notes in Computer Science, 7198:178–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28493-9_20/.

Anonymous:2012:FMcp

[1370] Anonymous. Front matter. Lecture Notes in Computer Science, 7198: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28493-9/1.

Heyse:2012:ARC

[1371] Karel Heyse, Brahim Al Farisi, Karel Bruneel, and Dirk Stroobandt. Automating reconfiguration chain generation for SRL-based run-time reconfiguration. Lecture Notes in Computer Science, 7199:1–12, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28365-9_1/.

Vipin:2012:AAR

[1372] Kizheppatt Vipin and Suhaib A. Fahmy. Architecture-aware reconfigurationcentric floorplanning for partial reconfiguration. Lecture Notes in Computer Science, 7199:13–25, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28365-9_2/.

Luzhou:2012:DSL

[1373] Wang Luzhou, Kentaro Sano, and Satoru Yamamoto. Domain-specific language and compiler for stencil computation on FPGA-based systolic computational-memory array. Lecture Notes in Computer Science, 7199: 26–39, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28365-9_3/.

Kim:2012:EBP

[1374] Yongjoo Kim, Jongeun Lee, Jinyong Lee, Toan X. Mai, Ingoo Heo, and Yunheung Paek. Exploiting both pipelining and data parallelism with SIMD reconfigurable architecture. Lecture Notes in Computer Science, 7199:40–52, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28365-9_4/.

deDinechin:2012:TBD

[1375] Florent de Dinechin and Laurent-Stéphane Didier. Table-based division by small integer constants. Lecture Notes in Computer Science, 7199:53– 63, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28365-9_5/.

Liu:2012:HSE

[1376] Qiang Liu and Wayne Luk. Heterogeneous systems for energy efficient scientific computing. Lecture Notes in Computer Science, 7199:64–75, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28365-9_6/.

Ostadzadeh:2012:PFD

[1377] S. Arash Ostadzadeh, Roel Meeuws, Imran Ashraf, Carlo Galuzzi, and Koen Bertels. The Q^{2} profiling framework: Driving application mapping for heterogeneous reconfigurable platforms. Lecture Notes in Computer Science, 7199:76–88, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28365-9_7/.

Hussain:2012:PPP

[1378] Tassadaq Hussain, Muhammad Shafiq, Miquel Pericàs, Nacho Navarro, and Eduard Ayguadé. PPMC: a programmable pattern based memory controller. Lecture Notes in Computer Science, 7199:89–101, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28365-9_8/.

Anjam:2012:RTT

[1379] Fakhar Anjam, Quan Kong, Roel Seedorf, and Stephan Wong. A runtime task migration scheme for an adjustable issue-slots multi-core processor. Lecture Notes in Computer Science, 7199:102–113, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-642-28365-9_9/.

Ndu:2012:BST

[1380] Geoffrey Ndu and Jim Garside. Boosting single thread performance in mobile processors via reconfigurable acceleration. Lecture Notes in Computer Science, 7199:114–125, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28365-9_10/.

Zhou:2012:CAF

[1381] Gang Zhou, Li Li, and Harald Michalik. Complexity analysis of finite field digit serial multipliers on FPGAs. Lecture Notes in Computer Science, 7199:126–137, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28365-9_11/.

Takamaeda-Yamazaki:2012:SSS

[1382] Shinya Takamaeda-Yamazaki, Shintaro Sano, Yoshito Sakaguchi, Naoki Fujieda, and Kenji Kise. ScalableCore system: a scalable many-core simulator by employing over 100 FPGAs. Lecture Notes in Computer Science, 7199:138–150, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28365-9_12/.

Ma:2012:SMH

[1383] Sen Ma, Miaoqing Huang, Eugene Cartwright, and David Andrews. Scalable memory hierarchies for embedded manycore systems. Lecture Notes in Computer Science, 7199:151–162, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28365-9_13/.

Watanabe:2012:TMR

[1384] Takahiro Watanabe and Minoru Watanabe. Triple module redundancy of a laser array driver circuit for optically reconfigurable gate arrays. Lecture Notes in Computer Science, 7199:163–173, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-642-28365-9_14/.

Loke:2012:RAF

[1385] Wei Ting Loke and Yajun Ha. A routing architecture for FPGAs with dual-VT switch box and logic clusters. Lecture Notes in Computer Science, 7199:174–186, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28365-9_15/.

Jin:2012:MLC

[1386] Qiwei Jin, Diwei Dong, Anson H. T. Tse, Gary C. T. Chow, David B. Thomas, and Wayne Luk. Multi-level customisation framework for curve based Monte Carlo financial simulations. Lecture Notes in Computer Science, 7199:187–201, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28365-9_16/.

Nakahara:2012:LCH

[1387] Hiroki Nakahara, Tsutomu Sasao, and Munehiro Matsuura. A low-cost and high-performance virus scanning engine using a binary CAM emulator and an MPU. Lecture Notes in Computer Science, 7199:202– 214, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/978-3-
642-28365-9_17/.

Talip:2012:CEI

[1388] Mohamad Sofian Abu Talip, Takayuki Akamine, Yasunori Osana, and Naoyuki Fujita. Cost effective implementation of flux limiter functions using partial reconfiguration. Lecture Notes in Computer Science, 7199:215–226, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28365-9_18/.

Mingas:2012:PTM

[1389] Grigorios Mingas and Christos-Savvas Bouganis. Parallel tempering MCMC acceleration using reconfigurable hardware. Lecture Notes in Computer Science, 7199:227–238, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.springer.com/

chapter/10.1007/978-3-642-28365-9_19/.

Rafique:2012:HTF

[1390] Abid Rafique, Nachiket Kapre, and George A. Constantinides. A high throughput FPGA-based implementation of the Lanczos method for the symmetric extremal eigenvalue problem. Lecture Notes in Computer Science, 7199:239–250, 2012. CODEN LNCSD9. ISSN 0302-9743 (print),

1611-3349 (electronic). URL http://link.springer.com/chapter/10.
1007/978-3-642-28365-9_20/.

Anonymous:2012:FMcq

[1391] Anonymous. Front matter. Lecture Notes in Computer Science, 7199: ??, 2012. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:978-3-
642-28365-9/1.

Adam:2012:MMC

[1392] Gheorghe Adam, Ján Buša, and Michal Hnatič, editors. Mathematical Modeling and Computational Science: International Conference, MMCP 2011, Stará Lesná, Slovakia, July 4–8, 2011, Revised Selected Papers, volume 7125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28211-3 (print), 3-642-28212-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28212-6.

Aura:2012:IST

[1393] Tuomas Aura, Kimmo Järvinen, and Kaisa Nyberg, editors. Information Security Technology for Applications: 15th Nordic Conference on Secure IT Systems, NordSec 2010, Espoo, Finland, October 27–29, 2010, Revised Selected Papers, volume 7127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27936-8 (print), 3-642-27937-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-27937-9.

Shi:2012:DFW

[1394] Yun Qing Shi, Hyoung-Joong Kim, and Fernando Perez-Gonzalez, editors. Digital Forensics and Watermarking: 10th International Workshop, IWDW 2011, Atlantic City, NY, October 23–26, 2011, Revised Selected Papers, volume 7128 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32204-2 (print), 3-642-32205-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32205-1.

Bononi:2012:DCN

[1395] Luciano Bononi, Ajoy K. Datta, Stéphane Devismes, and Archan Misra, editors. Distributed Computing and Networking: 13th International Conference, ICDCN 2012, Hong Kong, China, January 3–6, 2012. Proceedings, volume 7129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25958-8 (print), 3-642-25959-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-25959-3.

Lopez:2012:CIP

[1396] Javier Lopez, Roberto Setola, and Stephen D. Wolthusen, editors. Critical Infrastructure Protection: Information Infrastructure Models, Analysis, and Defense, volume 7130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28919-3 (print), 3-642-28920-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-28920-0.

Schoeffmann:2012:AMM

[1397] Klaus Schoeffmann, Bernard Merialdo, Alexander G. Hauptmann, Chong-Wah Ngo, Yiannis Andreopoulos, and Christian Breiteneder, editors. Advances in Multimedia Modeling: 18th International Conference, MMM 2012, Klagenfurt, Austria, January 4–6, 2012. Proceedings, volume 7131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27354-8 (print), 3-642-27355-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-27355-1.

Modgil:2012:TAF

[1398] Sanjay Modgil, Nir Oren, and Francesca Toni, editors. Theorie and Applications of Formal Argumentation: First International Workshop, TAFA 2011. Barcelona, Spain, July 16–17, 2011, Revised Selected Papers, volume 7132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29183-X (print), 3-642-29184-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-29184-5.

Jonasson:2012:APSa

[1399] Kristján Jónasson, editor. Applied Parallel and Scientific Computing: 10th International Conference, PARA 2010, Reykjavík, Iceland, June 6– 9, 2010, Revised Selected Papers, Part I, volume 7133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28150-8 (print), 3-642-28151-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-28151-8.

Jonasson:2012:APSb

[1400] Kristján Jónasson, editor. Applied Parallel and Scientific Computing: 10th International Conference, PARA 2010, Reykjavík, Iceland, June 6– 9, 2010, Revised Selected Papers, Part II, volume 7134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28144-3 (print), 3-642-28145-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-28145-7.

Thilagam:2012:ACN

[1401] P. Santhi Thilagam, Alwyn Roshan Pais, K. Chandrasekaran, and N. Balakrishnan, editors. Advanced Computing, Networking and Security: International Conference, ADCONS 2011, Surathkal, India, December 16– 18, 2011, Revised Selected Papers, volume 7135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29279-8 (print), 3-642-29280-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29280-4.

Bubak:2012:BND

[1402] Marian Bubak, Tomasz Szepieniec, and Kazimierz Wiatr, editors. Building a National Distributed e-Infrastructure–PL-Grid: Scientific and Technical Achievements, volume 7136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28266-0 (print), 3-642-28267-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-28267-6.

Mossakowski:2012:RTA

[1403] Till Mossakowski and Hans-Jörg Kreowski, editors. Recent Trends in Algebraic Development Techniques: 20th International Workshop, WADT 2010, Etelsen, Germany, July 1–4, 2010, Revised Selected Papers, volume 7137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28411-6 (print), 3-642-28412-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28412-0.

Ardissono:2012:AUM

[1404] Liliana Ardissono and Tsvi Kuflik, editors. Advances in User Modeling: UMAP 2011 Workshops, Girona, Spain, July 11–15, 2011, Revised Selected Papers, volume 7138 of Lecture Notes in Computer Science.

Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28508-2 (print), 3-642-28509-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-28509-7.

Iwamura:2012:CBD

[1405] Masakazu Iwamura and Faisal Shafait, editors. Camera-Based Document Analysis and Recognition: 4th International Workshop, CBDAR 2011, Beijing, China, September 22, 2011, Revised Selected Papers, volume 7139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29363-8 (print), 3-642-29364-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-29364-1.

Barthe:2012:FAS

[1406] Gilles Barthe, Anupam Datta, and Sandro Etalle, editors. Formal Aspects of Security and Trust: 8th International Workshop, FAST 2011, Leuven, Belgium, September 12–14, 2011. Revised Selected Papers, volume 7140 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29419-7 (print), 3-642-29420-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-29420-4.

Arbab:2012:FSE

[1407] Farhad Arbab and Marjan Sirjani, editors. Fundamentals of Software Engineering: 4th IPM International Conference, FSEN 2011, Tehran, Iran, April 20–22, 2011, Revised Selected Papers, volume 7141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29319-0 (print), 3-642-29320-4 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-29320-7.

Wang:2012:WAI

[1408] Liwei Wang, Jingjue Jiang, Jiaheng Lu, Liang Hong, and Bin Liu, editors. Web-Age Information Management: WAIM 2011 International Workshops: WGIM 2011, XMLDM 2011, SNA 2011, Wuhan, China, September 14–16, 2011, Revised Selected Papers, volume 7142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28634-8 (print), 3-642-28635-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-28635-3.

Kundu:2012:PMI

[1409] Malay K. Kundu, Sushmita Mitra, Debasis Mazumdar, and Sankar K. Pal, editors. Perception and Machine Intelligence: First Indo-Japan Conference, PerMIn 2012, Kolkata, India, January 12–13, 2012. Proceedings, volume 7143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27386-6 (print), 3-642-27387-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-27387-2.

Nambiar:2012:TPE

[1410] Raghunath Nambiar and Meikel Poess, editors. Topics in Performance Evaluation, Measurement and Characterization: Third TPC Technology Conference, TPCTC 2011, Seattle, WA, USA, August 29–September 3, 2011, Revised Selected Papers, volume 7144 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32626-9 (print), 3-642-32627-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-32627-1.

Pan:2012:TEV

[1411] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, Maiga Chang, and Mingmin Zhang, editors. Transactions on Edutainment VII, volume 7145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29049-3 (print), 3-642-29050-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-29050-3.

Kuncak:2012:VMC

[1412] Viktor Kuncak and Andrey Rybalchenko, editors. Verification, Model Checking, and Abstract Interpretation: 13th International Conference, VMCAI 2012, Philadelphia, PA, USA, January 22–24, 2012. Proceedings, volume 7148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27939-2 (print), 3-642-27940-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-27940-9.

Russo:2012:PAD

[1413] Claudio Russo and Neng-Fa Zhou, editors. Practical Aspects of Declarative Languages: 14th International Symposium, PADL 2012, Philadelphia, PA, USA, January 23–24, 2012. Proceedings, volume 7149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27693-8 (print), 3-642-27694-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-27694-1.

Vanmechelen:2012:EGC

[1414] Kurt Vanmechelen, Jörn Altmann, and Omer F. Rana, editors. Economics of Grids, Clouds, Systems, and Services: 8th International Workshop, GECON 2011, Paphos, Cyprus, December 5, 2011, Revised Selected Papers, volume 7150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28674-7 (print), 3-642-28675-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-28675-9.

Liu:2012:FHI

[1415] Zhiming Liu and Alan Wassyng, editors. Foundations of Health Informatics Engineering and Systems: First International Symposium, FHIES 2011, Johannesburg, South Africa, August 29–30, 2011. Revised Selected Papers, volume 7151 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32354-5 (print), 3-642-32355-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-32355-3.

Joshi:2012:VST

[1416] Rajeev Joshi, Peter Müller, and Andreas Podelski, editors. Verified Software: Theories, Tools, Experiments: 4th International Conference, VSTTE 2012, Philadelphia, PA, USA, January 28–29, 2012. Proceedings, volume 7152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27704-7 (print), 3-642-27705-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-27705-4.

Lukasiewicz:2012:FIK

[1417] Thomas Lukasiewicz and Attila Sali, editors. Foundations of Information and Knowledge Systems: 7th International Symposium, FoIKS 2012, Kiel, Germany, March 5–9, 2012. Proceedings, volume 7153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28471-X (print), 3-642-28472-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28472-4.

Ramanujam:2012:DCI

[1418] R. Ramanujam and Srini Ramaswamy, editors. Distributed Computing and Internet Technology: 8th International Conference, ICDCIT 2012, Bhubaneswar, India, February 2–4, 2012. Proceedings, volume 7154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28072-2 (print), 3-642-28073-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28073-3.

Alexander:2012:EPPa

[1419] Michael Alexander, Pasqua D’Ambra, Adam Belloum, George Bosilca, Mario Cannataro, Marco Danelutto, Beniamino Di Martino, Michael Gerndt, Emmanuel Jeannot, Raymond Namyst, Jean Roman, Stephen L. Scott, Jesper Larsson Traff, Geoffroy Vallée, and Josef Weidendorfer, editors. Euro-Par 2011: Parallel Processing Workshops: CCPI, CGWS, HeteroPar, HiBB, HPCVirt, HPPC, HPSS, MDGS, ProPer, Resilience, UCHPC, VHPC, Bordeaux, France, August 29 — September 2, 2011, Revised Selected Papers, Part I, volume 7155 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29736-6 (print), 3-642-29737-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29737-3.

Alexander:2012:EPPb

[1420] Michael Alexander, Pasqua D’Ambra, Adam Belloum, George Bosilca, Mario Cannataro, Marco Danelutto, Beniamino Di Martino, Michael Gerndt, Emmanuel Jeannot, Raymond Namyst, Jean Roman, Stephen L. Scott, Jesper Larsson Traff, Geoffroy Vallée, and Josef Weidendorfer, editors. Euro-Par 2011: Parallel Processing Workshops: CCPI, CGWS, HeteroPar, HiBB, HPCVirt, HPPC, HPSS, MDGS, ProPer, Resilience, UCHPC, VHPC, Bordeaux, France, August 29 — September 2, 2011, Revised Selected Papers, Part II, volume 7156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29739-0 (print), 3-642-29740-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29740-3.

Rahman:2012:WAC

[1421] Md. Saidur Rahman and Shin ichi Nakano, editors. WALCOM: Algorithms and Computation: 6th International Workshop, WALCOM 2012, Dhaka, Bangladesh, February 15–17, 2012. Proceedings, volume 7157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28075-7 (print), 3-642-28076-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28076-4.

Picco:2012:WSN

[1422] Gian Pietro Picco and Wendi Heinzelman, editors. Wireless Sensor Networks: 9th European Conference, EWSN 2012, Trento, Italy, February 15–17, 2012. Proceedings, volume 7158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28168-0 (print), 3-642-28169-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-28169-3.

Barthe:2012:ESS

[1423] Gilles Barthe, Benjamin Livshits, and Riccardo Scandariato, editors. Engineering Secure Software and Systems: 4th International Symposium, ESSoS 2012, Eindhoven, The Netherlands, February, 16–17, 2012. Proceedings, volume 7159 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28165-6 (print), 3-642-28166-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-28166-2.

Dinneen:2012:CPB

[1424] Michael J. Dinneen, Bakhadyr Khoussainov, and André Nies, editors. Computation, Physics and Beyond: International Workshop on Theoretical Computer Science, WTCS 2012, Dedicated to Cristian S. Calude on the Occasion of His 60th Birthday, Auckland, New Zealand, February 21–24, 2012, Revised Selected and Invited Papers, volume 7160 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27653-9 (print), 3-642-27654-7 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-27654-5.

Laud:2012:IST

[1425] Peeter Laud, editor. Information Security Technology for Applications: 16th Nordic Conference on Secure IT Systems, NordSec 2011, Tallinn, Estonia, October 26–28, 2011, Revised Selected Papers, volume 7161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29614-9 (print), 3-642-29615-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29615-4.

Clarke:2012:PSI

[1426] Edmund Clarke, Irina Virbitskaite, and Andrei Voronkov, editors. Perspectives of Systems Informatics: 8th International Andrei Ershov Memorial Conference, PSI 2011, Novosibirsk, Russia, June 27–July 1, 2011, Revised Selected Papers, volume 7162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29708-0 (print), 3-642-29709-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29709-0.

Petkova-Nikova:2012:PKI

[1427] Svetla Petkova-Nikova, Andreas Pashalidis, and Günther Pernul, editors. Public Key Infrastructures, Services and Applications: 8th European Workshop, EuroPKI 2011, Leuven, Belgium, September 15–16, 2011, Revised Selected Papers, volume 7163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29803-6 (print), 3-642-29804-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29804-2.

Solis-Oba:2012:AOA

[1428] Roberto Solis-Oba and Giuseppe Persiano, editors. Approximation and Online Algorithms: 9th International Workshop, WAOA 2011, Saarbrücken, Germany, September 8–9, 2011, Revised Selected Papers, volume 7164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29115-5 (print), 3-642-29116-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-29116-6.

DeVos:2012:RCT

[1429] Alexis De Vos and Robert Wille, editors. Reversible Computation: Third International Workshop, RC 2011, Gent, Belgium, July 4–5, 2011. Revised Papers, volume 7165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29516-9 (print), 3-642-29517-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-29517-1.

Kuipers:2012:SOS

[1430] Fernando A. Kuipers and Poul E. Heegaard, editors. Self-Organizing Systems: 6th IFIP TC 6 International Workshop, IWSOS 2012, Delft, The Netherlands, March 15–16, 2012. Proceedings, volume 7166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28582-1 (print), 3-642-28583-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28583-7.

Kienzle:2012:MSE

[1431] Jörg Kienzle, editor. Models in Software Engineering: Workshops and Symposia at MODELS 2011, Wellington, New Zealand, October 16–21, 2011, Reports and Revised Selected Papers, volume 7167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29644-0 (print), 3-642-29645-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-29645-1.

vandenHerik:2012:ACG

[1432] H. Jaap van den Herik and Aske Plaat, editors. Advances in Computer Games: 13th International Conference, ACG 2011, Tilburg, The Netherlands, November 20–22, 2011, Revised Selected Papers, volume 7168 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31865-7 (print), 3-642-31866-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31866-5.

Sakama:2012:DAL

[1433] Chiaki Sakama, Sebastian Sardina, Wamberto Vasconcelos, and Michael Winikoff, editors. Declarative Agent Languages and Technologies IX: 9th International Workshop, DALT 2011, Taipei, Taiwan, May 3, 2011, Revised Selected and Invited Papers, volume 7169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29112-0 (print), 3-642-29113-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-29113-5.

Meadows:2012:STM

[1434] Catherine Meadows and Carmen Fernandez-Gago, editors. Security and Trust Management: 7th International Workshop, STM 2011, Copenhagen, Denmark, June 27–28, 2011, Revised Selected Papers, volume 7170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29962-8 (print), 3-642-29963-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-29963-6.

Ystad:2012:SSM

[1435] Sølvi Ystad, Mitsuko Aramaki, Richard Kronland-Martinet, Kristoffer Jensen, and Sanghamitra Mohanty, editors. Speech, Sound and Music Processing: Embracing Research in India: 8th International Symposium, CMMR 2011, 20th International Symposium, FRSM 2011, Bhubaneswar, India, March 9–12, 2011, Revised Selected Papers, volume 7172 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31979-3 (print), 3-642-31980-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31980-8.

Bruni:2012:TGC

[1436] Roberto Bruni and Vladimiro Sassone, editors. Trustworthy Global Computing: 6th International Symposium, TGC 2011, Aachen, Germany, June 9–10, 2011. Revised Selected Papers, volume 7173 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30064-2 (print), 3-642-30065-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-30065-3.

Keller:2012:FMC

[1437] Rainer Keller, David Kramer, and Jan-Philipp Weiss, editors. Facing the Multicore — Challenge II: Aspects of New Paradigms and Technologies in Parallel Computing, volume 7174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30396-X (print), 3-642-30397-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-30397-5.

Rosner:2012:CNL

[1438] Michael Rosner and Norbert E. Fuchs, editors. Controlled Natural Language: Second International Workshop, CNL 2010, Marettimo Island, Italy, September 13–15, 2010. Revised Papers, volume 7175 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-31174-1 (print), 3-642-31175-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-31175-8.

Carbone:2012:WSF

[1439] Marco Carbone and Jean-Marc Petit, editors. Web Services and Formal Methods: 8th International Workshop, WS-FM 2011, Clermont-Ferrand, France, September 1–2, 2011, Revised Selected Papers, volume 7176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29833-8 (print), 3-642-29834-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29834-9.

Pena:2012:FPA

[1440] Ricardo Peña, Marko van Eekelen, and Olha Shkaravska, editors. Foundational and Practical Aspects of Resource Analysis: Second International Workshop, FOPARA 2011, Madrid, Spain, May 19, 2011, Revised Selected Papers, volume 7177 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32494-0 (print), 3-642-32495-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-32495-6.

Herkersdorf:2012:ACS

[1441] Andreas Herkersdorf, Kay Römer, and Uwe Brinkschulte, editors. Architecture of Computing Systems — ARCS 2012: 25th International Conference, Munich, Germany, February 28 — March 2, 2012. Proceedings, volume 7179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28292-X (print), 3-642-28293-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). LCCN ???? URL http://www.springerlink.com/content/
978-3-642-28293-5.

Bjorner:2012:LPA

[1442] Nikolaj Bjørner and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 18th International Conference, LPAR-18, Mérida, Venezuela, March 11–15, 2012. Proceedings, volume 7180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28716-6 (print), 3-642-28717-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-28717-6.

Gelbukh:2012:CLIa

[1443] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 13th International Conference, CICLing 2012, New Delhi, India, March 11–17, 2012, Proceedings, Part I, volume 7181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28603-8 (print), 3-642-28604-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28604-9.

Gelbukh:2012:CLIb

[1444] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 13th International Conference, CICLing 2012, New Delhi, India, March 11–17, 2012, Proceedings, Part II, volume 7182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28600-3 (print), 3-642-28601-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28601-8.

Dediu:2012:LAT

[1445] Adrian-Horia Dediu and Carlos Martín-Vide, editors. Language and Automata Theory and Applications: 6th International Conference, LATA 2012, A Coruña, Spain, March 5–9, 2012. Proceedings, volume 7183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28331-4 (print), 3-642-28332-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-28332-1.

Gheorghe:2012:MCI

[1446] Marian Gheorghe, Gheorghe Păun, Grzegorz Rozenberg, Arto Salomaa, and Sergey Verlan, editors. Membrane Computing: 12th International Conference, CMC 2011, Fontainebleau, France, August 23–26, 2011, Revised Selected Papers, volume 7184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28023-4 (print), 3-642-28024-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-28024-5.

Pan:2012:SWJ

[1447] Jeff Z. Pan, Huajun Chen, Hong-Gee Kim, Juanzi Li, Zhe Wu, Ian Horrocks, Riichiro Mizoguchi, and Zhaohui Wu, editors. The Semantic Web: Joint International Semantic Technology Conference, JIST 2011, Hangzhou, China, December 4–7, 2011. Proceedings, volume 7185 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29922-9 (print), 3-642-29923-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

URL http://www.springerlink.com/content/978-3-642-29923-0.

Khurshid:2012:RVS

[1448] Sarfraz Khurshid and Koushik Sen, editors. Runtime Verification: Second International Conference, RV 2011, San Francisco, CA, USA, September 27–30, 2011, Revised Selected Papers, volume 7186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29859-1 (print), 3-642-29860-5 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-29860-8.

Kiayias:2012:VIT

[1449] Aggelos Kiayias and Helger Lipmaa, editors. E-Voting and Identity: Third International Conference, VoteID 2011, Tallinn, Estonia, September 28–30, 2011, Revised Selected Papers, volume 7187 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32746-X (print), 3-642-32747-8 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/978-3-642-32747-6.

Sanner:2012:RAR

[1450] Scott Sanner and Marcus Hutter, editors. Recent Advances in Reinforcement Learning: 9th European Workshop, EWRL 2011, Athens, Greece, September 9–11, 2011, Revised Selected Papers, volume 7188 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29945-8 (print), 3-642-29946-6 (ebook). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL

http://www.springerlink.com/content/978-3-642-29946-9.

Pescape:2012:TMA

[1451] Antonio Pescapè, Luca Salgarelli, and Xenofontas Dimitropoulos, editors. Traffic Monitoring and Analysis: 4th International Workshop, TMA 2012, Vienna, Austria, March 12, 2012. Proceedings, volume 7189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28533-3 (print), 3-642-28534-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ???? URL http://www.springerlink.com/content/978-3-
642-28534-9.

Nguyen:2012:TCC

[1452] Ngoc Thanh Nguyen, editor. Transactions on Computational Collective Intelligence VI, volume 7190 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29355-7 (print), 3-642-29356-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-29356-6.

Theis:2012:LVA

[1453] Fabian Theis, Andrzej Cichocki, Arie Yeredor, and Michael Zibulevsky, editors. Latent Variable Analysis and Signal Separation: 10th International Conference, LVA/ICA 2012, Tel Aviv, Israel, March 12–15, 2012. Proceedings, volume 7191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28550-3 (print), 3-642-28551-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.

springerlink.com/content/978-3-642-28551-6.

Taft:2012:PAM

[1454] Nina Taft and Fabio Ricciato, editors. Passive and Active Measurement: 13th International Conference, PAM 2012, Vienna, Austria, March 12– 14th, 2012. Proceedings, volume 7192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28536-8 (print), 3-642-28537-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-28537-0.

Pena:2012:TFP

[1455] Ricardo Peña and Rex Page, editors. Trends in Functional Programming: 12th International Symposium, TFP 2011, Madrid, Spain, May 16–18, 2011, Revised Selected Papers, volume 7193 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-32036-8 (print), 3-642-32037-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-32037-8.

Regnell:2012:REF

[1456] Björn Regnell and Daniela Damian, editors. Requirements Engineering: Foundation for Software Quality: 18th International Working Conference, REFSQ 2012, Essen, Germany, March 19–22, 2012. Proceedings, volume 7195 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28713-1 (print), 3-642-28714-X (e-book). ISSN 0302-9743 (print), 1611-

3349 (electronic). LCCN ???? URL http://www.springerlink.com/
content/978-3-642-28714-5.

Pan:2012:IIDa

[1457] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACIIDS 2012, Kaohsiung, Taiwan, March 19–21, 2012, Proceedings, Part I, volume 7196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28486-8 (print), 3-642-28487-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-28487-8.

Pan:2012:IIDb

[1458] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACIIDS 2012, Kaohsiung, Taiwan, March 19–21, 2012, Proceedings, Part II, volume 7197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28489-2 (print), 3-642-28490-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-28490-8.

Pan:2012:IID

[1459] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACIIDS 2012, Kaohsiung, Taiwan, March 19–21, 2012, Proceedings, Part III, volume 7198 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28492-2 (print), 3-642-28493-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://www.springerlink.

com/content/978-3-642-28493-9.

Choy:2012:RCA

[1460] Oliver C. S. Choy, Ray C. C. Cheung, Peter Athanas, and Kentaro Sano, editors. Reconfigurable Computing: Architectures, Tools and Applications: 8th International Symposium, ARC 2012, Hong Kong, China, March 19–23, 2012. Proceedings, volume 7199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28364-0 (print), 3-642-28365-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http://

www.springerlink.com/content/978-3-642-28365-9.